A Complete Bibliography of Publications in Annals of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 1.53

Title word cross-reference

$104 [Jon13]. $124.95 [Edd13]. $135 [SS23]. $14.95 [Kid13]. $143.00 [Gue23, Nut21]. $149 [Hal18]. $149.00 [Car18]. $150.00 [Hig17a]. $189.00 [Wit13]. $19.95 [Mil13]. $21.95 [Qui13]. $25.75 [Kid13]. $28.95 [Coo13]. $29.95 [Sta21]. $32.50 [Wei19]. $35.00 [Fra13, Kin19b, Law13, Ren18]. $37.50 [Kri19]. $39.95 [Lau21, Mcv21, Yi13]. $40.00 [Lin22, Mil13, Row18, Sei13]. $41.00 [Hay18]. $45 [SS23]. $45.00 [Gie16, Lan13, Mar21a, Ver13]. $47.50 [Smi18]. $48.00 [Van19]. $49.95 [Mel21]. $50 [Gol19]. $50.00 [May21]. $54.95 [Gat18]. $55.00 [Pro18, Reg19, Wol13a, Sta19]. $65.00 [Sla19]. $74.00 [Sha19b]. $85.00 [Wil23]. $89.95 [Bri13, Par13]. $90.00 [Wei19]. $95.00 [Lov13]. $99.99 [Coz19, Sch18b]. ^{e} [LS15a]. ^{th} [Car18]. d [PB45, PB47, PB48]. l [PB45, PB47, PB48]. μ [Mon05].

-tartrate [PB45, PB47, PB48].

0 [Akt09, Bin12, Coo13, Hen00, How19, Kid13, Lai04, Lin22, May21, Sei13, Vic12, Wil11a, Jon13]. 0-19-878895-9 [How19]. 0-226-73238-I [Lin22]. 0-387-25961-9 [Bin12]. 0-520-21077-8 [Hen00]. 0-674-02667-5 [Wil11a]. 0-691-01042-0 [Swe03]. 0-691-09259-1 [Swe03]. 0-8166-4763-1 [Akt09]. 0-822-94562-2 [May21]. 0-85404-479-5 [Lai04].

1 [Akt09, Bar11, Fra13, Gol19, Luc19, Row18, Sch08a, Vic04, WN16]. 1-4020-0848-1 [Vic04]. 1-84893-415-7 [Sus16]. 1127 [Cha12a]. 1127-6045 [Cha12a]. 1385-0180 [Sch18b]. 14th [Big22]. 14th-century [Big22]. 1500-1630 [Egm16]. 1610 [Lat15]. 1624 [Kah01]. 1630 [Kus18]. 1639 [Gue23]. 1650s [CB22]. 1664/5 [Geo52]. 1680s [CB22]. 1690s [Car13]. 1694 [VvM92]. 1700 [Mos14a]. 1730s [Gre84]. 1760s [FB22]. 1770s [Mcc79]. 1777 [Bec13a]. 1783 [Sch54]. 1790s [Bro09]. 1793 [Swe63]. 1797 [Swe67]. 17th [Coh10a, Gau14, Jac52a, Jac52b, Old74, DS16]. 17th-century [Coh10a]. 1800 [Boa12b, Edd08]. 1820 [Joh15a]. 1830 [Mos14b]. 1850s [Bro95, Kvi22]. 1860s [Bar98, Cah90, Ku06, Ma21, SS13]. 1870s [Cah90, DeK83]. 1890 [Bow13, Wil13]. 1893 [Cah10]. 18th [Old74]. 19 [VvM92]. 1900 [CS12]. 1910s [Ku06, Ma21]. 1914 [Kid13, Sta13]. 1920s [Pav17]. 1922 [Hen92]. 1923 [Ste76]. 1930 [End17]. 1930s [Bon14, Hoc83, Sea79]. 1931 [Lyn14]. 1936 [Bro36a]. 1939 [Cha12b, Hal14b]. 1945 [Hen00, Coz15]. 1950 [Hil91]. 1950s [Min21, Str97]. 1957 [Ano57b, CCC57]. 1960 [Leg17]. 1964 [Hen09]. 1970 [Ree14a]. 1980s [Bad09, Hol17b, Sei13]. 1999 [Sla03]. 19ème [CP97]. 19th [Hen89, Joh15b, Kra82, Tho55, CP97]. 19th-century [Joh15b, Kra82]. 1st [Eme02].

2 [Bsc12, Bur11b, Car18, Hei19, HA18, Jam12, May21, Ren18, Sch08b, Wei19]. 2-271-13458-7 [Swe21]. 2-8031-0231-5 [Bur11b]. 2000 [Dun13, McC13]. 2002 [Kin04, Rus04]. 2010 [Nor12]. 2011 [Ano12f]. 2016} [IM17]. 2017 [Ano18]. 2019 [CECE21]. 20th [Bad11, DS16]. 236 [Tre78a]. 24 [Kah01]. 25 [Gay13a]. 26 [Sch54]. 28 [CS12]. 2nd [MS22].

3 [Boa19, Gat18, Jar18a, Lau21, Maz10, Qui12, Ray18, Sta21, Wol13a]. 31st [CCC57]. 33 [Tho16]. 3rd [Eme02, Eme13, Gov21, Woo15, Wes97].

4 [Boa12b, Edd13, Egm18, Gie16, Gue23, Hal18, Nut21, Pal12, Sak12]. 440-year-old [SHM13].

5 [Bur11b, Cha12b, Del11a, Hay18, Kin19b, Kri19, Lai04, Law13, Mil13, MS22, SS23, Sha19b, Swe21, Wil11a, Ver13]. 56 [Col52d]. 57 [But57]. 58 [Smi23a].

6 [Pro18, Qui13, Smi18]. 6000 [GG11b]. 6045 [Cha12a]. 60th [Pan09b]. 611rb [Ray15]. 611rb/915ra [Ray15].

7 [Bil18, Bil16, Bri13, Del11b, Del12, Jam12, Kus18, Lov13, Mar21a, Wit13, Yi13]. 715 [Sta19]. 751pp [Swe21]. 79 [Zai17].

8 [Bow13, Byf19, Coz19, Dom09, Hen00, Lan13, Mcv21, Mil12, Por16, Sla19, Van19, Wil23]. 8968 [Sha19a].

9 [Bin12, Bur11a, Hig17a, How19, Kra13, Lat15, Mel21, Nor12, Par13, Reg19, Rob10, Tho11]. 915ra [Ray15]. 976 [Sha19a]. 976-1-78023-8968 [Sha19a]. 978 [Bar11, Bil18, Bil16, Boa12b, Boa19, Bow13, Bri13, Bsc12, Bur11a, Byf19, Car18, Cha12b, Coo13, Coz19, Del11b, Del11a, Del12, Dom09, Edd13, Egm18, Fra13, Gat18, Gie16, Gol19, Gue23, Hal18, Hay18, Hei19, Hig17a, HA18, Jam12, Jar18a, Kid13, Kin19b, Kra13, Kri19, Kus18, Lan13, Lat15, Lau21, Law13, Lov13, Luc19, Mar21a, Maz10, Mcv21, Mel21, Mil12, Mil13, MS22, Nor12, Nut21, Pal12, Par13, Por16, Pro18, Qui12, Qui13, Ray18, Reg19, Ren18, Rob10, Row18, Sak12, SS23, Sei13, Sha19b, Sla19, Smi18, Sta21, Swe21, Tho11, Van19, Vic12, Wei19, Wil23, Wit13, Wol13a, Yi13, Jon13, Ver13]. 978-0-19-023480-5 [Sha19b]. 978-0-19-284677-8 [Wil23]. 978-0-19-878549-1 [Row18]. 978-0-19-956272-5 [Cha12b]. 978-0-226-09318-5 [Law13]. 978-0-226-26145-4 [Gie16]. 978-0-226-27650-2 [Jam12]. 978-0-226-31703-8 [Por16]. 978-0-226-41874-2 [Ren18]. 978-0-226-43690-6 [Smi18]. 978-0-226-44712-4 [Egm18]. 978-0-226-48119-7 [Jam12]. 978-0-226-49697-9 [Reg19]. 978-0-226-50929-7 [Sta19]. 978-0-226-51951-7 [Del11b]. 978-0-226-58366-2 [Wei19]. 978-0-226-59226-8 [Sla19]. 978-0-226-59579-5 [Kri19]. 978-0-226-61089-4 [Sak12]. 978-0-226-61798-5 [Kin19b]. 978-0-226-73177-3 [Qui12]. 978-0-226-73658-7 [Mar21a]. 978-0-226-77057-5 [Del11a]. 978-0-226-81994-5 [SS23]. 978-0-226-82038-5 [SS23]. 978-0-226-87020-5 [Ver13]. 978-0-226-89377-8 [Lan13]. 978-0-253-22051-6 [Qui13]. 978-0-262-01272-0 [Sei13]. 978-0-262-01286-7 [Mil13]. 978-0-271-03525-3 [Wol13a]. 978-0-300-15197-8 [Bow13]. 978-0-367-35723-8 [Mcv21]. 978-0-444-52728-8 [Dom09]. 978-0-521-49103-7 [Lov13]. 978-0-521-88346-7 [Del12]. 978-0-674-02667-4 [Wil11a]. 978-0-674-02828-9 [Tho11]. 978-0-674-03544-7 [Yi13]. 978-0-674-04661-0 [Coo13]. 978-0-674-08866-5 [Hay18]. 978-0-674-24174-9 [Mel21]. 978-0-691-16503-5 [Ren18]. 978-0-691-17737-3 [Sta21]. 978-0-7190-8022-7 [Bri13]. 978-0-7190-8114-9 [Par13]. 978-0-7294-0981-0 [Jon13]. 978-0-7456-3676-4 [Kid13]. 978-0-7546-6150-4 [Edd13]. 978-0-7546-6332-4 [Boa12b]. 978-0-801-89313-1 [Fra13]. 978-0-81225-089-3 [Lau21]. 978-0-8229-4441-6 [Pro18]. 978-0-8229-4558-1 [Gol19].

978-0-88135-375-4 [Pal12]. 978-0-88135-383-9 [Nor12]. 978-1-107-15886-3 [Ray18]. 978-1-118-62077-9 [Hig17a]. 978-1-316-60262-1 [Luc19]. 978-1-316-64971-8 [Van19]. 978-1-4214-2096-7 [Bil18]. 978-1-4214-2177-3 [Gat18]. 978-1-4773-0359-7 [Bil16]. 978-1-78023-640-7 [Kus18]. 978-1-78374-512-8 [Byf19]. 978-1-78735-005-2 [HA18]. 978-1-78914-504-5 [MS22]. 978-1-84893-030-8 [Mil12]. 978-1-85168-681-0 [Kid13]. 978-1-897174-42-5 [Mil13]. 978-2-271-13458-5 [Swe21]. 978-2-503-53078-9 [Bur11a]. 978-2-906-05392-9 [Rob10]. 978-3-0343-1294-3 [Boa19]. 978-3-05-005095-9 [Lat15]. 978-3-319-56528-6 [Sch18b]. 978-3-319-69717-8 [Coz19]. 978-3-319-95251-2 [Hei19]. 978-3-631-57110-1 [Bar11]. 978-3-8319-0371-9 [Kra13]. 978-88-464-8818-3 [Maz10]. 978-88-6227-118-9 [Cha12a]. 978-90-04-16723-0 [Vic12]. 978-90-04-26399-4 [Hal18]. 978-90-04-28670-2 [Car18]. 978-90-04-32492-3 [Jar18a]. 978-90-04-43918-4 [Nut21]. 978-90-04-46511-4 [Gue23]. 978-90-481-2361-2 [Bsc12]. 978-90-481-2815-0 [Wit13]. 978-90-481-2816-7 [Wit13].

âge [Cam09, Rom06].

A. [Har96, Low75, Per86, Sme55a, Sme55b, Sne76]. Abbacus [Fol11]. Abbe [Het76, McK57]. Aberdeen [Woo87]. aberration [Rud07]. Abraham [Osp80]. Abroad [RS09]. absence [Swi62]. Absolute [CY05, Nay82, Ari73b, Col95, DeK83, Hop57b, Per83]. abstraction [Ver16]. Academic [Joh15b]. académie [Geo38a, Gue21]. academies [Son75]. Academy [Cal68, Cro09a, Gál88, Dav98, Gue10, Ung02]. Accademia [Bos03, Gia23, She11]. accelerators [Min21]. acceptance [Che98, Coc83, Mau69, Tre80a]. Accepted [Del11a]. acclimatization [Wei85b]. accomplishment [Tur82]. according [Hir07, Thi73, Uch22]. account [Bry75, Bur90, Eyl37, GG75b, SS58]. Accounts [CCC57, HT23, Mad57]. Accum [Col51]. Accuracy [Bro09, Bal23, Cha83a, Cha83b]. Accurate [Ish14]. accurately [Akk23]. Achievements [Pet14]. achromatic [Dei91, Jae10]. acid [Le 74, Par53b, Pat38, Pat39, Woo71c]. acidity [Le 72, Le 74]. Ackermann [Kin04]. acknowledgement [STDH22]. Acosta [Mar21b]. Acoustical [Ku09, Wit13, Pan09a]. Acoustics [Ku06, Ku13, Gou80, Hac79b]. acque [Duc14, Maf10]. Across [Des08, Lon19]. Action [Ano09g, Ear61, RO83]. activities [Cha68]. activity [Ehr95]. actuality [McG78]. Adam [Edw72a, Goo10]. adaptation [Sch75]. add [Sch19]. Addenda [Ano60a, Lop60, Mad88a, Mad88b]. Addendum [Bur00, PB48]. addiction [Ber79]. Addressing [Mol96]. Adler [Bur12b]. Administration [Kle12c]. Adolfo [CK89]. adoption [Jor82]. adsorption [FGR74]. advance [Fli59].

Advancement [Gay00, Kni13, Wit10, Por81, Son75]. Advances [GG84b, Str97]. Advantages [Alt07]. advertising [Bry92]. Advocacy [Ash17]. advocates [Sta88b]. Æemula [Nor12]. Æmula [ABNW09]. Æneas [Rot68]. aequipondio [Lev21]. Aere [Hen23]. aeronaut [Sme55c]. aesthete [Old80]. aether [Ait69]. aether-stream [Ait69]. Aethere [Hen23]. Affair [Rap12, Bik71, De 05]. affairs [Gay10]. Affinities [Gra01]. affinity [Dun62a, Dun62b, Tay08]. Africa [Sta88a]. After [Ako21, Bru11, Alm22, Kre23, RS89, Ste10a, Ben09]. afterlife [Pia21]. again [Moo87]. Against [Fli08, Cas15, Kah01, McC94, Ped75, Sar69]. Agassiz [Dav68]. Age [Ano05j, Cah10, Cha14, Del13, McE10a, Aga17b, Bro36b, Cam09, Dav90b, Dea81, Fus58, Gas10, Ham82, Kli15, Mor90, SS23, Sok96, Wat82, Ano09i, Are10, Goo11, Hes17, Sta17]. ageing [RO82]. agenda [Gia23]. ages [Rya91, Bur09, Rom06]. aggregate [Tre80a]. ago [Bik75]. agonistes [Sch13c, Tho16]. Agrarian [Pop12]. Agricultural [Bra95, Hol17b, Sto13, Jon85, Olb91, Sch21, Sne81]. Agriculture [Kni76, Stu89]. agrochemical [Hal20]. Ahmed [BvK73]. Aileen [Jam12]. aimants [Qué90]. ‘Aina [Tak11]. Air [Ano09h, Bru11, Gen07, War02, Bar73, Osm58, Tom14, vH94]. air-pump [vH94]. Airborne [Cre66]. Airy [Ben80, Cot76, Cot77]. Aiton [GG91b]. Aitor [Cha12b]. Akada [Kin04]. Akademie [Lat15]. al [Gre87, Mar11b, Hil85, Bur09, Bur12a]. Al-Biruni [Hil85]. al-Farisi [Bur09]. al-Tilmid [Bur12a]. Alan [Bil18, Joh14, Sch18b, Bsc12, Jac15a, Joh15b, Mau15, Ram15, RN15, UB15, Wei15]. alba [Mul76]. Albareda [Mal09]. Albert [Cam55, Ill89, Mar11b, Har96, Ill89, Mil81a, Pal83]. Alberti [Bin12, Par13, And07]. Alberto [Bey14, Sha19a]. Albrecht [Ken52, Son75, Wei48]. Alchandreana [Bur12c]. Alchandreus [Bur00, Bur98]. Alchemical [S´ 20, New96]. alchemist [Hen21a]. Alchemists [Ano09g]. Alchemy [Ano08f, Ano08h, Boa12b, dV21, Kah01, Lau21, Pri13, Boa12a, Ben14, Eve15, Gru11]. Alchimie [Boa12a, Kah01]. Aldini [FC36]. Aldous [Cal80]. Aldrovandi [SHM13]. Alejandro [Nut21]. Alessandro [Osm58]. Alexander [Coo13, Dau13, Dau14b, Swe03, Cot08, Dau13, Gib52a, HB74, Kle12b, Mar06, McC10, Nic11b, Pau78, TS14, de 60, Wer11, Dau14a]. Alexis [Gre81, Gre82]. Alfonso [Ela13]. Alfred [Ku13]. algebra [Ste11a, Ver16, Rei14, GG97]. algebraical [Ver16]. Algorismi [Fol11]. Alhazen [Eas86]. Alice [Lov79, Lov79]. alien [Tan80a]. aligned [Ari81]. alive [Sha19a]. Alkali [Gar04, Git66, Mat76]. Allan [FF68]. Allbutt [Leu08]. Alleged [Dra81]. Allegiance [Gov21]. allemande [DS97]. Allergy [Bur10c]. alliance [Kri19]. alloy [Pri55]. alloying [Wil95a, Wil95b]. alloys [McC89]. alluvial [Uss80]. Almanac [Tan14]. almanacs [Nic39]. along [Vet04a, Vet04b]. Alpert [SP11]. alpha [Tre74, Tre74]. Alpha-Teilchen [Tre74]. Alsace [CP97]. alsaciens [CP97]. Altered [Pan09a, Wit13]. alternative [Ben74, Cob18, Ink89]. altimetry [de 56]. Altitude [Mil96]. Aluminium [McC89]. amalgamation [Mor86, Tei75]. amateur [Geo38b, Joh18, SR71b].

Amazon [McC91]. Ambassador [Hud03]. ambitious [Sch19]. Ambrogio [FGR74]. Amedeo [Col64]. America [Bry13, Wil13, Aug11, Bed97, Bri10, Cah10, Cra12, Dol77, DK54b, Hal20, Het76, KK02, Ku13, Luc08, Mil91b, Par84, Plo78, Rad18, Rei81, Sok94, MSD23, War88, Wol13b, Fit17, Sav15]. American [Adl17, Ape20, Dav90b, Des08, Dym11, Gro72a, Gro74a, Gro74b, Het83, Hun99, Kai10, Kin19a, Lar91, Min21, Mor01, Sch75, SR73c, Str97, Uch22, Joh18]. Americas [Nic11b]. Amicus [Ung02]. Amir [Coo13]. ammonium [PB45, PB47, PB48]. Among [Bel12, Bil16, Dun94, Joh15b, SM15]. anaesthesia [Far83, Tee84, Und67]. anagram [Ben74]. analemmatic [Arc05]. Analemmatiques [Arc05]. Analogies [Ren18]. analogy [Hua22, RO84]. Analyse [Hen92]. analysing [Mac82]. Analysis [Dra82, GG11a, San00, WH11, Beg21, BB13, De 17, Deb67, Hen92, Ink89, Lai77, Tan77, Tom01]. Analytical [Bai93, Rap15a, CK76, Col37b, Pul94, Wil23]. analyticus [Tan77]. analytischen [Pul94]. Anatolica [Rob10]. anatomical [MT72, Nut21, Und63]. Anatomies [Ano08e, Car21]. Anatomizing [Hua22]. Anatomy [Fra10, Gre77, Koo10, Sch13b, Ste10a, Mac10, Und63, Wil62, Fra14a]. Ancestors [Goo10]. ancestral [Mag98]. anciens [Bur12c]. Ancient [Duc09, Jon06, Nut21, Pom11, Tre15, Eff21, Gel21, Ham12, HT23, Mol96, NGDP75, PLC48, Gel21]. Andersen [Bin12, Nor12]. Anderson [Swe72a, Ban14]. André [Bel12]. André-Jacob [Bel12]. Andreas [Ano09g, Wei48]. Andrew [OM95, SR73a]. Androids [Vos13, Mar15]. Anduaga [Cha12b]. Angle [Orr14]. Anglesey [Old93]. Anglicanism [Mul73]. Anglo [Kin19a]. Anglo-American [Kin19a]. Animal [Ben13b, BS12, Wal37, Bla19, Coh36, Coh10b, Ekh10, FLW11, FC36, Kin19b, Kip87, Shu68]. animalium [Kus10]. Animals [Ben13b, Egm16, Egm18, Kus18, Rit11, dC09a, FLW11, Gue10, MG10, New52]. Animaux [New52, Gue10]. Animi} [Nut14, Cor11]. Ann [Plo80]. Anna [Lau21]. Annals [Ano02a, Ano03a, Ano12f, Ano19, IM15, IM17, Ano74j, Ano09b, Ano16b, Ano18, CECE21, GG75a, GG76, GG10]. Anne [Bur14a]. Anne-Marie [Bur14a]. Anniversary [Pan09b, Hje10]. Annotated [Gui15, Swe03, BJ00, DC09b, Dar13]. Annotations [Bar95]. Announcement [Ano69a, Ano70a, Ano70b, Ano71a, Ano91a, Ano91b, Ano92a, Ano92b, Ano96a, Ano13a]. Announcements [Ano93a, Ano94a, Ano94b]. anomalous [WO87]. Antebellum [War88]. Anthony [Col52a, Kin04]. Anthropology [Wei85a]. anti [Hom88, Nye81]. anti-atomism [Nye81]. anti-Newtonian [Hom88]. antimony [Don50]. Antiquarian [SR71b]. antique [Pri54]. antiquitate [PLC48]. Antiquity [Hen14, Kin04, Wal12, Dar12, Wie19]. antitoxin [OS97]. Antoine [DK54a, DK55a, DK55b, DK57, Eyl63, Kah01]. Anton [Kon12]. Antonio [Gre87, Bla19, DG09, Gre87, SM16]. Antony [ZA16]. Antwerp [Mes96]. août [Kah01]. Ape [Qui13, Cos09]. apes [Hje10]. Aphorism [Sch36].

Apologia [McC39a]. apothecary [Di 17]. Apparatus [Hop57a, War88, Che98, CBL81, Fle97, Fra41, McC86]. appareils [Ben09]. appartenant [Ben09]. Appeal [Pra10]. appearances [Hol07]. Application [Mil03, Ege68, For70b, Lee23]. appointment [Ano88i]. Appraisal [CP02]. Appreciation [Ano15, GG91b]. Apprehension [Kol99]. apprenticeship [Rus78]. Approach [Moo87, Ink88, McD72, Mel10, Sch08c, Spr42, SW67, SS23]. Appropriate [Ink89]. Appropriation [Hal18, Lig16, Maf10]. Appropriazione [Duc14, Maf10]. approved [BS93]. approximation [Nay82]. Après [Ben09]. April [Sch54]. aqueous [Dym37]. Arab [Bur12c, Zia94]. arabe [Bur12c]. Arabic [Bur00, Pat17, Bur98, GW11, Jus11, Jus12, Nor16, Por10, Por11]. Arbeiten [Thi73]. arbitrary [Erl92]. Arbor [Plo80]. Archaean [Old91, Old92, Old93, Old94]. archaeological [Gen22]. Archaeology [Bry10, Eff21]. Archetype [Slo03]. Archibald [Old80, Old97, Woo15, Eme02, Eme13]. Archimedes [Sch18b, Wit13]. Architect [SP11]. Architects [Tka14, Val09]. Archives [Goe18, ZA16, GG74]. Arctic [Lev86, Lev94]. Arcubus [BvK73]. Area [Swi12, OM95, SR72b]. areas [For70a]. argument [Dra86, Dra88, Hil84]. Argumente [Ped75]. arguments [Qua72, Ped75]. Argyll [Eme02, Eme13, Woo15]. Ariadne [Ste00a]. Arias [Sla19]. Aristarchus [Sid07]. aristocrat [Ost92]. Aristote [Kah01]. Aristoteles [Bur11a]. Aristotelian [Ari73b, Bla10]. Aristotelis [SBT10, Bur14b]. Aristotle [Bro81, Bro79b, Hut77, Kah01, Slo11, Val09]. Arithmetica [Eva78]. Arithmetick [Wal97]. Armchair [Bow16]. armillary [Not21, Pri54]. Army [Mos14b, War02, Kel11]. Aromaticity [Wei15]. Arrangement [Mau15]. Arsenal [Mac89]. Art [Ano07j, Bin12, Egm16, HA18, Jar98, JR12, Kus18, Rap15a, Rei14, Sle17, And07, Ano85j, Bou17, PLC48, Sal92, Sch76, Ste11a, Wil91a, Yap83, Egm18, Wad15]. arte [Yap83, S´ 20]. artefact [Gen22]. Artful [Ano06a, Fra14a, Mac10]. Arthur [Kei84, Pal83]. arti [Duc14, Maf10]. Artifacts [Ano08l]. Artificial [Fra09, Nic11a]. artis [PLC48]. Artisan [Can10, Jar23, Ost92, SR72c]. Artisanal [Kle12a, Leh12a]. Artisanal-Scientific [Leh12a, Kle12a]. Artisans [Dav98, Mar15, Vos13]. artistic [For70b, Gou83]. artists [Har71]. Arts [BS65, Lan13, Coc56, Dun94, Gib51f, Gou83, Han21, Maf10, Rob53b, Rom78, Sti37, Tom14, Wer10, Sme55a, Sme55b]. Arun [YT13]. Aryabhatiya [Bur10a]. Ascelinus [Kun00]. ascent [How19]. ascribed [BS65]. Ashgate [Edd13]. Asia [Low03]. Asimov [Luc19]. Askesian [Ink77]. asparagine [ST51]. Aspden [For71a]. aspect [SD68]. Aspects [Ano08k, Bsc12, Péo10, Ste11b, Wri92, Ano91o, CR96, Dun62a, Dun62b, Hen89, McC75, Shi89, Ver56]. Aspekten [Hen89]. assembly [Sme57b]. Asses [GH08]. assessing [Dun23, Sok94]. assessment [Woo71b]. Association [Gay00, Kni13, Pap09, Wit10, Hon84, Lai04, Ora75]. Associative [GG97]. Assur [Gel21]. Assyrian [Gel21]. Astbury [Dav90a]. Astrid [Nor12]. astrobiology [Dun21]. Astrolabe [Bur00, Kun00, MVK02, NNS^{+}06, Bur98, Dek95, For71c, TD93]. Astrolabes [Bur12b, Cha83b, Gor87, Lam97, Tur94]. Astrological [Van01, Ave20, Bur12c, Rib20, Win66]. astrologiques [Bur12c]. Astrology [Alm22, Bur00, Bur98, Cam23, Dun94, GW11, Tar20]. Astronomer [CS12, Mor68, Arm53, Hol96, Wil91a, Sza72]. Astronomers [TH74, RS89, Plo80, BS21]. Astronomia [McC39b, Swe09]. Astronómica [NG09]. Astronomical [BB13, Tho51a, WH11, Arm56a, Bed97, FB22, Cha95, Din50, Fru16, Ges23, Kre23, Not18, Not21, Tir23]. astronomicus [Roc81]. Astronomie [Bur09]. Astronomy [Bur09, Bur10b, Bur12b, Gow82, Not17, Por10, Ste07, Ste09b, Ste11b, Wal00, Arm51, Arm54, BG94, Bec11, Cas15, CZ19, Gro74a, Het83, Hol96, Ink78, Mac19, Mas53a, NS52, Nic39, Ogi75, OEMB21, Sus15]. astrophysicist [Hen18]. astrophysics [Plo78]. asymmetries [Byr81]. atelier [Nau19]. Athanasius [Hir07]. Athias [Ama06]. Atlantic [May21, Mil13, Chr20, Des08, Row09]. Atlanticus [Ray15]. Atmosphere [And09a, Mea66]. Atmospheric [Péo10, Edw10]. Atom [Bsc12, Cha09, KL07, KN13, RS01, Hei67b, Smi23a]. Atomic [CR02, Del13, Ham03, Hen00, Mau15, UB15, Cam09, Ger23, RG23, Mau69, Par39, Pau78, Ros98, Sne76, Sok96, Low90]. atomique [Cam09]. Atomism [BV99, RN15, Ree80, Kah01, Nye81, Pau78]. atomisme [Kah01]. Atoms [Cha09, Min21, Kra82, Bsc12, Smi23a]. attack [Bro79b]. Attacks [Stu09]. attempt [Bro95]. attempts [Tur55]. attention [Kvi22]. attitudes [Hen80, Shu68, SR73c, SR73b]. attraction [Bow74, Cha75, Dun67, Haw68]. Attractione [Dun67]. attributed [TD93]. aube [Coh11, Lyn13]. auction [Tur67]. Aude [Lov13]. Audience [CS12, Bro55]. audiences [Lig10, Jam12]. auditory [Kvi22]. Aufrecht [Hen09]. August [Dau14a, Ecc80, Hor87, Kah01, Wer11, Cah90, Ecc76, Ecc80, Wal39]. Augustan [Coo90]. Auguste [Pot53]. Augustin [Duv54]. Augustus [Bro22, Lai79]. Aurora [Cha21]. Ausdehnungslehre [Lew77]. Austin [Bil16]. Australia [Hil91, And19, HW87, HW90]. Australian [Mac89, Tod93]. authentic [Feu75, PLC48]. Author [Ano09a, Del11a, End16, Bry72, Kur97, Not18, Spr77]. authorities [GL11]. Authority [Fin12, Hed22, Hon95, Mol96, STDH22, Coz19]. Authors [Ano09e]. authorship [BS65, Gie19]. Autisms [Ash17]. autobiographical [Bla78, FF68]. Autobiography [Bez75]. autograph [Ari74, Fra47, VvM92]. automatism [Coh36]. autonomous [vH92]. autour [Muc98]. Auvergne [de 62]. Aux [Muc98, Tom10, Qué90, Sto13, Sim13]. avant [Duc13]. Aviation [Ari79]. avibus [Yap83]. Avicenna [Bur11b]. Aviva [Reg19]. Avogadro [Col64]. awareness [Hon87]. away [Ros21].

B [Bod56, Bro79b, Cop87, Hal18, Jay78, McC14, New52, Sme54a, Sta19, Tre65a, Vic12, Wol13a, Woo71b, Woo71c]. b&w [Gol19]. B.A. [Hop58]. Babylonian [BB13, BB13, Ste07, Ste11b, WH11]. Bachem [Hen92]. Bachems [Hen92]. Background [RS01, Sch57, Cas15, NM37a, Nor75]. backgrounds [Ecc80]. Bacon [Coc56, FG83, Jal23, Mar21b, Ree80, Ree81, Son74, Wat92]. Bacterial [vH92]. Bacteriology [Fra13]. bacteriophage [vH92]. Bad [Edw04, Wes10]. Badash [Sei13]. Baer [DD21]. Bailey [OJY03]. Bainbridge [VV22]. Bains [Ben09]. Baklarish [Kah10]. Bala [YT13]. Balance [Plo99, And19, Bry75]. Baldwin [Gie16]. ballistics [Seg83]. ballooning [Lan83]. Balloons [Kim06]. balsam [Mil91a]. Balthasar [Att85]. Baltimore [Bil18, Fra13, Gat18]. Bancalari [BM94]. Bangladesh [WN16]. Bankruptcy [McC94]. Banks [Law09b]. Baptista [Par36]. Baptiste [Klo85a, Klo85b, Wil55c]. Barbados [Ste36b]. Barbara [Sus16]. Barcelona [NG09]. Bargielowska [Hal14b]. Barhebraeus [Tak11]. Barlow [Mau15]. barometer [Bry75, Goo67, McC05]. Barrow [Nau14]. Bart. [PM50]. Bartholomew [Tho51b]. Bartoli [CK89]. Bartolomeo [Cha52]. Bascelli [Pal12]. basis [Bra79, Sie09]. Bate [Not18]. Bates [Ste69a, Ste69b]. Bateson [Coc83, Olb89, Rus22]. Batteries [Mil03, Ock37]. battery [Hop59b, Mer98]. battle [Hed22]. Bayen [Nea51b]. Be [Rus04, Del11a, GG77, Pou91]. Beale [Stu82, Stu89]. beam [For98]. beast [Coh36]. beast-machine [Coh36]. Beastly [Kre11]. beat [Tan80b]. beautiful [Wri16]. Beauty [Hal14b, ZB10]. Became [Har09b, MS22]. Becker [Sus16]. Becoming [Wal04]. Beddoes [Rob55b, LS68]. Bedeutung [Ber97]. Beeldsnyder [Dei91]. Been [Nie10, Fis36]. Beethoven [GG94, Hui11]. Before [Fru16, Uga06, WH11, AM76, Bar98, BB13, Bro15, Bro18, Fin12, For98, Hol07, How02, How03, Hun99, Jay78, Kat23, Lar61, Mau69, Nau19, OG68, PK87, Rus21, SS58, Wal37, Mac73, Sta19]. Beginner [Kid13, Joh09]. Beginning [Gra01, Flo96, Ols70]. Beginnings [Cro03, Dea81, Gar10, Mil01, Yi14, Cha48, Hug11, Jon50, Plo78]. Begriffes [Nit96]. Begründung [Ste76]. Béguelin [Hak82]. Beguin [KH60, Pat37]. behind [De 05]. being [Dun23, OJY96, Sha72]. Bekker [Att85]. Belief [Ano99e, Duc97]. beliefs [GG91a, Old97]. Bellarmine [Tar20]. belonging [Ben09]. Bends [Stu09]. Benedetto [Ari73a, Ari81]. Bénédict [Sme78]. Benefits [Ano06i]. Benito [Sla19]. Benjamin [Par51b, Par53a, Bar73, Dea89a, DK55a, DK55b, DK57, GG97, GG12, Law08, Mil83, Rob63, Wil55c]. Bent [Bur12d]. Bent-Equivalent [Bur12d]. Bentham [LS68]. Benzene [Roc85, Wei15, Wal39]. benzoic [Pat39]. Béraut [McK36a]. Berggren [Swe03]. Bergman [CK76, LS68, Sme54a]. Bergson [Her17, Can15, Her18]. Bergstaat [Kon12]. Beringer [Mal82]. Berlin [Lat15, Ste76, Cal68, Mei79, Ste76]. Bern [Boa19]. Bernard [Cha72b, Hig17a, Jam12, Jam12, Mil13, Tat97, Tho54, You37]. Bernhard [de 53a]. Bernoulli [Ait55a, Gui95, Sch11, SS99]. Bernoullischen [Sch11]. Berthelot [Nye81, Tre78b]. Berthollet [DK54c, Gra01, Key90, Nea51c]. Berufung [Hen89]. Berzelius [Bec16, Rus63a, Rus63b]. Best [CECE21]. Bestandaufnahme [Wal14]. Bestandsaufnahme [VW10]. Bestimmung [Bur09]. Between [Alv07, Dij13, Ela11, End17, GE19, HA15, Key90, KS10, Mar11a, Ots05, Pal13, Por11, Pre08, SM16, Val09, vD11, And14, Bun77, Cun04, Dau13, Ecc80, For71b, Hac79a, Hen96, Hul02, Ill89, Jac52a, Jac52b, JTR94, JTR95, Kah10, Kim95, Lai79, Leg17, MHO13, Mil81b, Mil02, Pis14b, Ste69a, Ste69b, Tee79, VPW16, Wei51, Win66, vH94, vH92, Kah01, Hei14]. Bewell [Bil18]. Beyond [Aga12, Qui13, Wei15, Cos09, And15]. Beziehungen [Ecc80]. Bhaskara [Bur10a]. Bias [Hul85]. Bible [She98]. biblical [Sla19]. bibliographical [DK54b, Far79, FC36, Gib53d, NS81]. bibliography [MT72, PP77, Win66, Col37b]. Bicycle [Bra09]. Biddell [Ben80, Cot76]. ’Big [Gil02, Bai93]. Bildungsbürgertum [Kle21]. biochemistry [Bea76, Mor90]. Biographical [Col54a, Gre57, JL79, Col51, Rus49, Smi87, Swe72a]. Biographies [Ano05a, Ano08i, Qui12]. Biography [GG12, Old09, Gra71, HB74, Kur97]. Biological [Rus04, Sak12, CSVK80, Kot78, Mor90, Nyh09]. biologie [Cam09]. Biologist [Pav17, Bay38b, Bay38a, Bay39a, Bay39b]. Biology [Bow17, Car10, Maz12, Qui12, Slo03, Bea76, Cam09, Rup09, Sha72, Wei85b, Yi13]. Biomedicine [Par16]. Biopolitics [Pal13]. biotech [Hug11, Yi14]. Bird [Cha95]. Birds [Far18, Har58, Mei54, Yap83]. Birmingham [Sch56]. Birth [Del11b, Ell00, Nit96, Tan14, YT13, Bal06, Chi08, Chr95, Mel09]. Biruni [Hil85]. bis [Hen92, Kan70, Kan71]. Bishop [dCA36, McC39a]. Bit [Ano07e, Ano07f]. Black [Del11b, Sla03, Luc19, Mel09, Boa16, Cro59, Eyl63, McK36b, McK59, MK60, McK60, McK62, McK65, McK67, Old72c, Per83]. Blackburn [Cre10]. Blackwell [Hig17a]. Blagden [DK54c]. Blair [Eis05]. Blaise [Pat39]. Blanc [Tou89, de 56]. Bleaching [Pag03]. Blessed [Ano08a]. Blind [Har57]. Blomstrand [Kau75]. Blondel [Wil92]. Blood [Pel97, Fra45, Lau21, Mas53a, Ric86, Whi86a]. bloodstains [BS15]. Bloomington [Qui13]. Bloomsbury [Bou17, HA18]. blow [Old72a]. blow-pipe [Old72a]. blue [Col39]. Blume [MS22]. Board [Ano96i, Ano14a, Ano74j]. Böckh [Dau14a, Wer11]. Bodies [Fra10, Arm47]. Body [Car17, Hal13, Hal14b, Hua22, Low96, ZB10]. Boerhaave [Duc17, Gib57, Hom79, Lov74]. Boethius [Eva78]. Bohemia [Sta21]. Bohr [KN13, Kra17, Sce94]. Bolantio [BB95]. Bologna [Di 17]. Bomb [Hen00, Low90, Ros98]. Bonaparte [Bra75, Bra94]. bond [Hen78, How02]. Bonds [Bur12d]. Bone [Car05]. Bonelli [Bed82]. bones [Ego19]. Bonnet [Wil23]. Bonnier [Hil86]. Book [Adl17, Aga17b, Aga17a, Akt09, And15, Ank15, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano88a, Ano88b]. Book [Ano88c, Ano88d, Ano88e, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96c, Ano96d, Ano96e, Ano96b, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02b, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05b, Ano05c, Ano06b, Ano06c]. Book [Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano08d, Ban14, Bar11, Bar14, Bas15, Bec13b, Bec13a, Bec14, Ben14, Ben13b, Bey14, Bil18, Bil16, Bin12, Boa12b, Boa19, Boe14, Bou13, Bow13, Bow17, BB13, Bri13, Bry13, Bsc12, Bsc14a, Bsc14b, Bur17, Bur11a, Bur11b, Bur14b, Bur14a, Bur14c, Bur15, Byf19, Car17, Car18, Cha12a, Cha13, Cha12b, Cha14, Chu13, Coo13, Coz15, Coz19, Dau14b, Dau14a, Del09a, Del11b, Dij13, Dom09, Duc13, Duc14, Dun13, Edd13, Egm18, End17, Eve15, Fit17, Fol14a, For71a, Fra13, Fra14a, Fra14b, Gat18, Gau14, Gie16, Gol19, Gos15, GG13b, GG13a, Gro14, Gue13, Gue23, Hah15, Hal18, Hal14b, Hay18, Hei14, Hei19, Hen17, Hen14, Her17, Hig17a, HA15, HA18, Hol17a, How19]. Book [Hun14, Jac14, Jam12, Jar18a, Joh14, Joh15a, Jon13, Kid14, Kin04, Kin19b, Kni13, Kni17, Köb17, Kra13, KN13, Kri19, Kus18, Lan13, Lat14, Lat15, Lau21, Law13, Lee13, Leg17, Lin22, Lov13, Low03, Luc19, Lus14, Lyn13, Lyn14, Lyn15, Mar15, Mar21a, May21, Maz10, McC13, McC14, Mcv21, Mil12, Mil16, Mil13, MS22, Mos14a, Mos14b, Nic13, Nic14, Nor12, Nut14, Nut21, Pal12, Par13, Pat17, Pis13, Pis14b, Pis14a, Por16, Pre16, Pro18, Qui12, Qui13, Ray18, Ree14b, Ree14a, Reg19, Rei14, Ren18, Rob10, Row18, Rus04, Sak12, Sav15, Sch13b, Sch18b, SS23, Sei13, Sha19a, Sim13, Sla19, Smi15, Smi18, Sta13, Sta21, Str14, Sus15, Sus16, Tat17, Tau14, Tho16, Tho11, Tre15, Van19, Ver13, Wad15, Wal14, War14, Wee16, Wei19]. Book [Wer14, Wil11a, Wil15, Wil13, Wil23, Wit13, Wol13a, Woo15, Yi13, Yi14, YT13, Zai17, Bik65, Gia23, PK87, Wil37, BvK73, Kah10]. Bookreviews [Ano88f]. Books [Ano68a, Ano68b, Ano69b, Ano69c, Ano69d, Ano69e, Ano70c, Ano70d, Ano70e, Ano70f, Ano71b, Ano71c, Ano71d, Ano71e, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano74e, Ano74f, Ano74g, Ano74h, Ano92i, Ano92j, Ano92k, Ano93h, Ano94i, Ano96h, Ano97g, Far18, Fra10, Kus18, MS22, Sha19a, Dea83, Gue10, Tir20]. Boole [Hes52, Lai77, Lai79]. border [Kan70, Kan71]. borealis [Cha21]. Borell [Arm50]. Borelli [Ela13]. Borlase [Oli69]. Born [Del09a, Moo87, Gre05, Tei75]. Born-again [Moo87]. Borrello [Bar14]. Bosc [Ber12]. Boston [Gue23, Nut21, Vic04]. Botanic [Rob54a]. Botanical [LMBM06, Nic11b, Spr77]. Botanist [Mil11c, Swe72b, de 54]. Botanists [Gib57, Nic06]. Botany [Die12, Die16b, Hyd19, Ree76, Hol17b]. Bottles [Ren11]. Bouguer [AM76]. Boulter [HA18]. Boulton [Rob53b, Rob54c]. Bound [Bec14, Ink88, Mar16a, ML^{+}12]. boundaries [Bec16]. boundary [Old72b]. bounded [Rib20]. Boussingault [McC75]. Bovelles [San84]. Bowdich [Orr14]. Bowler [Luc19, Dun13, Mac85a]. Boxes [Ren11]. Boycott [Pre08]. Boyer [Jay78]. Boyle [Mad59, Nut14, Pat36a, Sha19b, Boa09, Bru11, Cec11, Cle90, Cor11, HP03, Lev14b, Mad57, Mad61a, Mad61b, Mad62, Mad64, Mad88a, Mad88b, New96, O’B65a, O’B65b, Ost92, Pal23, Ran63, Row50a, Sar02, Wes56a, Wes56b]. Brachistochrone [Sch11]. Bradley [Dom09, Wil55c]. Bradwardine [Mol96]. Bragg [Bey10]. Brahe [Cha84, Fig72, Sím93, Swe09, Tir20]. Branch [Bon14]. Brande [Spi69]. Branding [Gau12a]. Brass [NNS^{+}06]. Bray [Low03]. Brazil [Min21]. bread [Cob18]. Breaks [Dij04]. breakthrough [Coh10a, Gau14]. Bredekamp [Lat15]. Breeding [Fit17]. Breeds [Alt07]. Brenna [Nor12]. Brenni [Jar18a]. brevitatis [Ray82]. Brewster [Bro77]. Brianta [Maz10]. Bridge [Ela11, Pis14b, Val09]. Brief [Ill89, HT23]. briefe [Sok94]. Briefen [Hor87, Thi73]. Briefwechsel [Wer11, Dau14a]. Brill [Car18, Gue23, Hal18, Jar18a, Nut21]. Bringing [LMBM06]. Brinkman [Bry13]. Bristol [Kei84]. Britain [Bow13, Hal14b, Kni13, Ree14a, Zai17, And19, Bea12, Ber79, Bow09b, Bra92, Bra95, Cha48, Dub63, DK54b, FRW11, Fos37, Gay13b, Git66, Har09a, Hyd19, Jon95, Kni76, Law08, Mar09, Mat76, Nye18, Olb91, Old91, Old92, Old93, Old94, Pag03, Pel97, Rad18, Rit11, Sea79, Smi76, Smi23b, Ste00b, Ver94, Wie23, Wit96, Wit10, Yeo79, Yos99, ZB10, Dun13, Orr14, Ank15, Hun14]. Britannica [Jon13, KL09]. British [Cha12b, Cha14, Hal18, Hay18, Kni13, Mos14b, Alb89, And09b, Bar09, Bow16, Bro93, Bru11, Cha53b, Cha54a, Cha54b, Cha70, Dav68, Des08, DE50, Ell83, End16, Eyl67, Gar04, Gas10, Gib53a, GG93, Jon86, Kel11, Kni75, Ku06, Lig16, Lin16, Mac89, Mac93, Mea70, Ora75, Ozd14, Pau78, PP77, Por78, RS09, Sme87, Sta88a, SR72c, SR72b, Tay08, Ver16, Wit10, Yav93]. Britta [Nor12]. Brock [Hor87]. Brokering [Dal14]. Bronn [Mat11]. Brora [Wat82]. brothers [Mat54]. Brougham [Hay64]. Brown [Bry72]. Browne [Jon95]. Brownrigg [Rus49, Rus50, Rus51a, Rus51b]. Brückle [Lat15]. Brun [Coh10b]. Brünnow [Plo80]. Bruno [Sha19a, Arm48, Bon14, Mar16b, Mar19, Mas73]. Bt [Eyl63]. Buch [DD21]. Buchanan [WN16]. Buchwald [Mar11b]. Buckland [Nor42]. Buffon [Wil56a, Wil56b, Wil56c, Woo87]. Bugge [Bec13a, PdC10]. Building [Suz09]. Buildings [San00, Ege68]. Bulgaria [Bec09]. bull [Tar20]. Bulldog [Qui13, Cos09]. Bullock [Swe70b]. Burchard [Don50, Don51]. Bureau [BS65, Str97]. bureaucrat [Jar98]. bureaucrats [Olb89]. Buridan [Bur14b, SBT10]. Burke [Arm56c]. Burlington [Edd13]. Burma [WN16]. Burned [Sha19a]. burning [Sme87]. Busard [Bur14c]. Business [Far18, Yi13, Car10]. Bute [Tur67]. Byzantines [FW85].

C [Akt09, Arm58a, Chu13, Dij13, Dom09, Fra14b, Per86, Plo78, RH11, Sla03, Van19, Ell00, Mae05, Mar09, Ohn23, Wan18a, Whi13]. c.} [Dek95, Hol96, OEMB21, TD93, Bez75, Nea51c, Pul94, Swe74, Cro05, Don50, Ste36b]. c.1775 [Yos99]. ca [Gou23, Kvi22]. cabaliste [Kah01]. cabalists [Kah01]. cabinet [Hil86, dC88]. Cable [Mil13, FY09, Row09, Mil13]. Cadrans [Arc05, Arc07]. Caernarvonshire [Old93]. Calandrini [Gui15]. calcination [McK36a]. calculating [For72b]. calculation [Hal57]. calculations [Tho51a]. Calculus [Bec01, Fer02, Dub66, Gre81, Gre82, Pal21, Nau14]. calendar [Hil85, Not17]. calendrical [FW85]. Calibration [BD00]. calibrator [Sch11]. Calico [NG97]. call [Aga17b, Kli15]. Callan [dVH65]. caloric [Dia96]. calorimeter [LS74]. calotype [Woo71c]. Calvin [Mas53a]. Calvinist [Fli08]. Cambridge [Byf19, Col52d, Coo13, Hay18, Kid13, Lov13, Luc19, Mel21, Mil13, Ray18, Rus04, Sei13, Van19, Yi13, BG99, Col52b, Col52c, Col53, JL14, Pan09b, Pri52, Rus22]. Came [Del11a, Gau14, Coh10a]. Camel [Bra09]. Camera [HZ07]. Cameralism [Ver13, Wak09]. campaign [Fox96]. Campanella [McC39a]. Campbell [Eme02, Eme13, Woo15]. Campos [Ren18, Rap15b]. Campus [Mea09, Min21]. Can [BB13, Far19, Sid07]. Canada [And19, Mil91a]. Canadian [Mil13, Lev86, McB91]. canal [Jon82]. Canales [Law13]. Canaries [Stu09]. Cancer [Coo11, Gat11]. cannot [Ros21]. Canterbury [Dek08]. Cantor [Gra71]. capacity [Sim92]. capillarity [Mil47]. Capitalism [Mea09]. Capt. [Ste36a]. captains [Rob54c]. Captives [Gar04]. Carbon [RS01, Wil07]. card [Whi51]. Cardano [Rei14, Ste11a, Tan80a]. cardio [Fra41]. cardio-vascular [Fra41]. cards [Cra85]. Career [Wal04, WN16, Gay10]. Careers [Ank15, RS09, Bra81, Gay13b, Lov79]. caricatures [Hig17b]. Carl [Cun04, For71b, Jay78, Uch22]. Carlini [TT99]. Carlisle [Col52a]. Carlo [Gia23]. Carlson [Yi13]. Carnets [Ray15]. Carol [Ben13b]. Caroline [Ogi75]. Carolingian [Ste09b]. carried [DK54c]. Cartesian [Ait59, Cha19]. Cas [Fau05, Tom01]. Casares [SM16]. Case [Bri00, Cec11, Kid14, Kim12, Mil11a, SK11, SHM13, Sid07, vD08, Bro87, CBL81, Dei91, Ehr95, Fau05, GG86, GG84a, Gro72b, Ink75, Jon95, Key90, Kit82, Mil81b, Son75, Str93, Tom01, War12, Wit96, Woo71a, YO89, Sch76]. cases [Cha21]. Casimir [Pop12]. Castelli [Ari73a, Ari81]. catalog [Ben09]. Catalogue [Gay13a, Kin04, Ben09, McC86, Ste10b]. catalogues [Jay78]. Cataloguing [Bec13b, Ste10b]. catalysis [FGR74]. Catherine [Mos14b]. Catholic [RS89]. Cauchy [Dub66]. causa [Ray82]. causal [Byr80]. Cause [Ram10]. causes [Bla19, McD72, Wil87]. cave [Nor42]. Cavendish [LA22, Bro78a, Bro90, Jac52a, Jac52b, Mil02, Rob95, Sha19b]. CE [Die21]. Celebrate [Pan09b]. Celebrating [Smi23a]. Celestial [Ste07, TS14, Ait60, Ait62, Ait64a, Ait65, Arm50, Bra82a, Bra82b, Dek87]. Cells [Har09b]. celtium [Hei67a]. censoring [Mal20b]. censorship [Ave20, Mal20a, Mar21a, Tir20]. census [FS21]. Centennial [CP02]. Central [Vet04a, Vet04b, Ait64b, Erl92, Gui95, Hal57, Hoc83, GWN^{+}11, Sch08c]. centre [Erl92, Nof14]. Centuries [Bra09, GG11a, Ari73b, Deb67, Ear63, Old74, VT17, War12, Kid14]. Century [And15, Ank15, Bad11, Bed93, BS12, Cro09b, Dun13, Eis05, Fit17, Gau14, Gay13b, Jam12, Koj02, Leh12a, Mil02, MVK02, Nic06, Ots05, Péo10, Ree14a, Row18, Ser12, Smi15, SR71b, Tom09, Vic04, Wit13, dC09a, ATE11, App90, Ari73a, Arm53, Bel99, Ben91, BD00, Ben37, BS15, Big22, Bow09b, Bre16, Bru79, Buc13, Bur12c, CR96, Caw77, Cha70, Cha83a, Cha21, CZ19, Cob18, Coc56, Coh10a, Cro16, Dav66, De 17, Di 17, Die16b, Dol77, Dun62a, Dun62b, Dun23, DS16, Dym37, Eva78, Fei03, Fru16, FL07, Gál88, Geo38b, Ges23, Gib51e, Gie19, Glu23b, Gor87, Gou83, Gra72b, Gro74a, Hac79a, Har71, Hen89, Het76, Het83, Hom79, Jac52a, Jac52b, Jae10, Joh15b, Jon74, Ked16, Kit82, Kle12a, Klo85a, Klo85b, Kra82, Lam97]. century [LS23, Li94, Lon19, Low96, Mac23, Mac20b, McC36b, Mea70, Mel10, Mes96, Mil83, Mil88a, Mil45, Mil47, Mul73, NG97, NF73, Old72b, Oli69, Ols70, Pan09a, Pau78, Pel97, Pop12, Rad18, Rob62, RS71, RS74, Sar69, Sch08c, Sha75a, Sha72, Sha73, Sme87, Ste69a, Ste69b, SR72c, SR72e, SR72b, SR73c, SR74a, SM16, Swe64, Tay52b, Tho55, Tre63, Tre64, Tre65b, Tre65a, Tre70, Vac00, Val96, Wal36, Wei48, Whi86b, Woo53, Wri92, Yeo79, dBP86, dC88, Aga12, Bea12, Bri10, Car21, CP97, Cav06, Cro09a, Dar12, Dij04, Fli08, Gat11, Gra01, Kah10, Ku13, Law08, Li12, Mar01, McC83, Mil11b, RC08, Sal09, Sch90, SC04, SR71a, Tan14, Thi71, Tom14, Wil01, Bry13, Car18, Hen14, Wer14]. Certain [Del09a, Fra41, Gre05]. Certitude [San00]. Cesare [Duc14]. Chair [Col54b, SP92]. Challenge [Nic06, Roe90]. Chalmers [Bsc12, Sch18b]. Cham [Hei19]. Chambers [Hes17]. chameleon [Bes68]. chance [Gre22]. Chandrasekhar [Wes10]. Change [How02, Nay77, Bad60]. Changed [Her17, Jar18a, MLSB17, Can15]. changes [Bes68, Fra39b]. Changing [Hol17b, Par03, SR73b, Cah90]. Channel [Des08]. Chaos} [Sve19]. Chapelain [Geo38b]. Chapman [And09a]. Chappe [Arm54]. Chaptal [Per86]. Chapter [Bur10a, Sme55a, Sme55b]. chapters [BJ00, Swe03]. character [dCA36, Bro93, Coo90, Har72]. charcoal [Sch75]. charge [Rob95]. charges [Wal36]. Charles [HA18, Jac52a, Jac52b, Kee52, Kni13, Kur99, Wil23, Alt07, BJ96, Coh10b, Cop87, DD21, End16, Gro72b, Ink72a, Kot78, Pap09, Rus80, San84, Sch80, SP92, Tho67, Tou89, de 58, Spu87, Hol17a]. Charles-Édouard [Pap09]. Charleton [Sha73]. Charlotte [Egm18]. Charter [Gov21]. Chartres [Bur00, Bur98]. charts [Edd13, dVH36]. Chatto [Sus16]. Cheapside [Rus78]. checklist [Jay78, JL79]. Cheer [Car05]. Cheesemakers [Ano08a]. Chemica [Dun67]. chemicæ [PLC48]. Chemical [Ade10, Alb89, Bas95, Boa08, BL21, Cec11, Don88, Fin09, Gra01, Hir07, Jac15a, Mil08, Mil12, Nye89, Ram15, Ber97, CK76, Col64, Cro59, Dob36, Dun67, Hen78, Lan83, Li94, Mau69, McE10b, McK36b, McK59, McK60, McK62, McK65, McK67, NG97, Por81, RS71, SD68, Sme54a, Sme78, Tay52a, Tay08, Tom01, Ber97, Wil37]. chemicus [Duc17]. Chemie [Thi73]. Chemische [Ber97]. chemischer [Ber97]. Chemist [Gay10, Kur97, McE10a, RS09, Eyl63, Kau75]. chemistries [Chr20]. Chemistry [Ano08f, BV99, BS15, Boa12b, Bur12d, Col52b, Col52c, Col52d, Cro03, Kni17, Kno64, Kon12, Kra00, Kur01, Lai04, Lin10, Mil08, Mil16, Nea50, Nea51a, Nea51b, Nea51c, Nea51d, Nea52, Nye89, Par03, Pet14, RN15, Rei04, Ste20, Tom09, Tre63, Tre64, Tre65b, Tre65a, Tre70, Ape20, Bai93, Bec16, Bik71, Bre16, Bro90, Bro16, Bro55, Bro95, Cle90, Deb62, DK55a, Edd08, Fau05, Fle74, For15, Geo52, Gib51b, Gib52b, Kau79, Klo85a, Klo85b, Kni76, Mac20a, Mor02, Pat37, PLC48, Pot53, Rus59b, Sce94, Sce08, Sme54b, Sne81, SM16, Swa89, Tay08, Thi73, Tom10, Tur53, Uss78, Wig55, Wig56, Wig61, Ano88i, Col53, Lee13]. Chemists [Joh15b, Kur01, RS09, Swi12, CP97, Cha91, Mac89, Mac93, Rus78, Sme87, Tom10]. Chenevix [Uss78]. Cherso [Hen79]. Chesser [Jon95]. Cheyne [Bow74]. chez [Hir07]. Chicago [Egm18, Gie16, Kin19b, Kri19, Lan13, Law13, Lin22, Mar21a, Por16, Reg19, Ren18, SS23, Sla19, Smi18, Sta19, Ver13, Wei19, Cot08]. Chichester [Hig17a]. Chile [Don36a, Don36b]. Chilmead [FG83]. chimie [BV83, Fau05, Sim13, Tom10, Ber01, DK54b]. chimique [Tom01, Sme67]. Chimistes [CP97, Sim13, Tom10]. China [Cha84, CZ19, CBL81, Hsi09, Ink88, Kim99, Li94, Li12, Ma21, OJY96, Wan18a, YO89, Chu13]. Chinese [CZ19, Flo96, Ink89, Lea18, NGDP75, OJY03, Wan18a]. ching [AS81]. Chirotherium [Tre91]. Chisholm [Gra72a, Gib52a]. Christiaan [Ari79, Plo99, VvM92, Zig80, vHvG99]. Christian [Kau75, Swe09, Whi12]. Christianity [Fan09]. Christine [Rob10]. Christoforo [BB95]. Christoph [Cha21, For72a, Lat14, Fra13]. Christopher [Ben73, DK54a, Qui13]. Chromatic [Rud07]. chromosome [Coc83, Rus22]. Chromosomes [Com21, Rus21]. chronographe [LS15a]. chronological [Tho51a]. chronology [OEMB21]. chronometer [Bal23, Dav78, For66a]. Chronometers [Lev94, Akk23, Bro87]. Church [Hol17a, Sar69, SS58]. churches [Jam78]. chyli [Fra39b]. Chymicum [Pat37]. Chymie [NS81]. chymique [Hir07]. chymistry [Gue21]. Científicas [Mal09]. Cimento [Bos03, She11, Gia23]. circa [Edw72b]. Circle [Ank15, Can10, Péo10, Sin49, Sin50a, Sin50b, Gay13b]. Circles [Tka14, Qua68]. Circolazione [Maz10, Bri07]. circuit [Win49]. circular [Bro92]. circularity [Bra82a, Bra82b]. circulation [Coz19, Fra41, Mas53a]. circumstances [For72b]. City [Gay00]. Civil [Car05, Ape20, Rob62]. civilian [Mac93]. Civilisation [Dym11]. civilization [NG09, Bur17]. Civilizations [Gau14, YT13, Bal06, Coh10a]. cladistics [Hul85]. Clairaut [Cha75]. clandestinité [Coh11, Lyn13]. Clapeyron [Bra81]. clarification [Gál88]. Clark [Gre57]. Clarke [Old72a]. class [Jar98, Low96, Mac79]. classes [SR72f]. classic [RD16, Por16]. Classical [CR02, Lai04, Wal01, GM94, Mor02, Thi12]. Classicism [Hal13, Hui11]. Classification [Arc07, Bak62, Big22, Hig53, Hol17b]. classique [Thi12]. classroom [War88]. Claude [Tat97, You37]. Claudine [Lyn13]. Clausius [Bru58a, Dia94, Hut73]. Clave [Kah01]. Clavis [Ste00a]. Clavius [Gam21]. Cleghorn [MdVH58]. Clement [Not17]. Cleomedes [Tod82]. Clercq [Bec13a]. Clerk [Cra54, Hop57a, JTR94, JTR95]. Clerk-Maxwell [JTR94, JTR95]. Cleveland [Het76]. Clifford [Leu08]. Climate [Lus14]. climatology [Oli69, Rya23]. clinic [Mau90]. Clinical [Tau10]. Clio [Gue23]. clock [Kat16, Lau66, Mil88b]. Clockmaker [Bro79a]. clocks [Glu23b, Tir23]. closed [Hon95]. Closing [The94]. Cloth [Sta19, Boa19, Lan13, Wei19]. clots [Pel97]. cloud [Jon86]. Clouds [Dek90]. Club [Bow17, Kaa17, Rob55a]. clusters [Ste10b, Bec13b]. Clyde [Jon82]. CNRS [Swe21]. Co [Die12]. Co-production [Die12]. Coal [Wil13, Far69, Luc08]. coal-gas [Far69]. Coalfield [Wat82]. Coalsack [Dek90]. coast [Mil91b]. Cochrane [McK36b]. Cocker [Wal97]. Cockshutt [Gib55]. Code [Del11b, Ras17, Mel09]. Codex [Ray15]. coelestium [Kre23]. Coeli [Tar20]. Coffey [Rot68]. Cohen [Gau14, Lyn13]. cohesion [Mil45, Mil47]. Coignet [Mes97, Tur06]. coil [dVH65]. Coiled [Cav06]. coinage [Wil95a, Wil95b]. Colbatch [Coo90]. Cold [Bos03, Sav15, Ben37, Mal94, Ger23, RG23, Gor16, Kri22, MN13, Wol13b]. Cole [Jar18b]. Coleridge [Lev78, Sys93]. Collaboration [Bey10, Gor17, Lov11, Caw77, Gol19]. Collaborative [Die12, Rap12]. Collapse [Gue09]. Collected [Mar11b, Sce08, Har96, Lee13]. collection [Eyl67, Gen22, Jon84, Pri54, STDH22, Swe70a, Swe70b, Dor76]. Collections [Jac99, Kre13, Ben09, Bod56, WN16]. collectors [Swe72a]. College [Bed93, Dor76, Pri52, Tho55, Sti37, Ano88i, Bed93, KO83]. Collingwood [Old99b]. collision [CBL81]. collisions [Dia94]. Colloid [Aug11]. Colloquies [Win16]. Colombian [Uss80]. Colonial [Bew16, Bil18, Bri13, Lee23, dA21, Mac09, Bed97]. Colonialism [Ozd14]. colore [Fra39b]. colossus [Sta19]. Colour [Bel99, Nic06, Bes68, Duc88, For70b, Har71, Luc19, Mil42, NG97]. colour-changes [Bes68]. colour-mixing [For70b]. Colourful [Cob02]. Colouring [Nic06]. colourmen [Har71]. Columbia [Bed93]. Columbus [DK54a]. Combe [de 50a]. Combination [Pan05, Fle74]. Combining [UB15, Cro61, Mau69]. comburentibus [Ray15]. combustion [Ful61, PM38b]. come [Bra92]. comet [Alm22, SH89, VV22, Kol99, Wal84]. cometary [Arm51]. comets [VV22]. Comfort [Cro96]. Comité [Sme56a]. commemorating [Hje10]. comment [Mac85a]. commentaires [Vla10, Bur14a]. commentary [Eva78, McC37]. Comments [Hol76, Ari73a, Hil91]. commerce [Hyd19, Ano09i]. Commercialization [Bey11]. commitment [Bra82a, Bra82b, Hul85]. Committee [Ano57b, Mor86, SBB^{+}75]. Committees [Hul02]. Commodities [War88]. commodity [Kin19a]. Common [Car05, Cha19, New38, Tay08]. commotion [Bik75]. Commune [Sme57b]. Communication [Hal18, Hol07, Lig16]. Communications [FY09, Mil13]. communities [Hil91, HW90]. Community [Hun99, GG91a, GM94, Kim95, OH97, Par84, Reg19]. Companies [Bro79a]. Companion [Hig17a, JL14]. Company [Arn09]. ’Comparative [Kim99, Chr10, GG13a, Leu08, Beg21, Mel10, Sch21]. Comparison [Lat15, Byf19, HW87]. compass [Cot76]. Compendium [BB13]. Competing [SS13, Wol13b, Sav15]. competition [SP92]. Compiling [Ano05j, Car97a]. complaint [Yeo18]. complètes [Lov12]. complexions [Big22]. complexity [Kra82]. Composition [SD68, Hug52, Le 74, Par53b].

Compounds [Jac15a]. comprise [Bla88]. Computation [Nof14]. Computer [Lus14, Nof14, Edw10]. Computers [Par03]. computing [LA22]. computors [LA22]. concavities [Wil65]. Concentrated [Tan77]. concentric [Qua68]. Concept [Ela11, Pis14b, Pom11, CK76, Col95, Hac79a, Hen79, Kot78, Lov74, McC21, Old74, Sch39, Wil56d]. conception [Ari73b, Cro48, Dek96, Hua22, Hum37, Mit79]. conceptions [De 17]. Concepts [Ber79, Byr80, Cha19, Dun67, Sch11]. Conceptual [And09a, Ela11, Gas07, Pis14b, Nay77, Rou22]. conceptualizations [Ma21]. Conceptualizing [Pot22, Yea11]. concernant [Ben09]. concerning [Ben37, Dym37, Edw72b, Nak84, Old72b, Wil65, Wil87]. Condamine [McC91]. condemnation [RS89]. Condorcet [Bak62, Whi82]. conduction [GG91a]. Conferences [Nye89]. confession [Reg19]. Confessionalization [VV22]. confirm [Wil60c, Wil61]. Conflict [Fli08, Maf13, LS23, Lus18, Ols71]. confusing [Gab78]. Congestor [Tan77]. Congress [Cah10]. congresses [Fra38]. congruity} [GE19]. conics [Pou91]. Conklin [Pav17]. connaissance [Duc13]. Connecting [Row09, Mil13]. connection [Hon84, Woo87]. connections [Str97]. Connor [McC13]. Conquest [Moz04]. Conrad [LGKW08, Vic12]. Cons [Old99b, Kau79]. Consejo [Mal09]. Consequences [Mon05, Cum84, Sta88a]. Conservation [Jur10, Can19, Goo80, Kat23, Ole23]. conservative [Hen78]. considerations [Duc88, GG86, GG84a, KH60, Mau69, PB45, PB47, PB48]. Considering [FLW11, Ben13b]. consimilibus [AVJ^{+}06, Bur11b]. constant [Kid81, Mer98]. Constantijn [vHvG99]. Constantine [MJ89]. ’Constants [RB01]. Constellation [Sie09]. constitution [Pri55]. constitutional [Wal76]. Constructing [Kra21, Par16]. Construction [Boe14, Zui04, Ait78, Bow84, Erl92, For71c, Hol07, Nye11, Sys93, vH94]. consultant [Aug04]. Consultation [BS65]. Consulting [Wil13, Luc08]. containers [Har71]. contemporaries [Cor38a, Cor38b]. Contemporary [Ben13b, End16, Kra09a, Cha68, FLW11, Rob54a, Tho47, Woo71b]. Content [Mil13, Pul94, Row09]. Contention [Bos03]. Contentious [Bar14, Bra10, Bor10]. Contested [Coo11]. Context [DK11, KL07, KN13, RN15, She11, Wes10, Win16, Ink77, Ink91, Lev21, Wei85a]. contexte [Muc98]. Contexts [HA15, Gas10, IR11, MHO13, Ole23, Cha14, Fol14a]. Continent [Jac52a, Jac52b, de 57]. continental [YO89]. Continents [Mil13, Row09]. Contradictions [GG02]. contrariée [LS15a]. contrary [Osp80]. contrast [Ecc80]. contrasts [Gos76]. contre [Kah01]. Contribution [Die12, Sim04, Dub66, Ell83, Jac52a, Jac52b, Mes97, Oli69, Whi86b, Wil55a]. Contributions [Des08, Ait55a, Kin04, Kox90, Ogi75, Pot53, Sme54a]. Contributors [Ade10]. Control [Edw04, Sto13, Kat16, Lee23, Mac20b]. controlled [Hen87]. Controversial [MS22]. controversies [Old72b, OH97]. Controversy [Kim12, Kra09a, Lal12, Mar06, Mil02, Cot77, Dea81, Hon95, IM89, Lai79, McC38, Nak84, Old91, Old92, Old93, Old94, Ols71, Uss78, vH92]. convection [Ege68, IM89]. convention [Hen78]. conventional [Pul94]. Convergence [Fer02]. conversion [Goo80]. Cooperation [Bon14, vH94]. Copernican [Dek96, Har72, Por10, RS89]. Copernicus [Hil84, Hen17]. Copié [Ray15]. Copies [McK65, FS21, McK59, McK60, McK62, McK67, Old72c, Lat15]. Copley [Sin50b]. Copper [Mil04a]. Copy [Lat15, Bar95]. corde [Fra39b]. Core [Wei15]. Cores [MN13]. Cornelis [Gib48]. Cornelius [Cha12a, Hir08]. Cornhill [McC94]. Cornish [Don50]. Corps [Ano05e]. corpus [STDH22]. corpusculaire [Hir07]. corpuscular [Cle90, Hir07, New96, Sch13c, Tho16]. corpuscular-mechanism [Sch13c, Tho16]. correct [Tre78a, Wil60c, Wil61]. correction [Cot76, Feu76]. corrections [Cra54]. Corrective [BG99]. Correspondence [Ano83g, Ano86g, Ano90g, Ano91i, Ano94j, Ano95g, Cun04, DD21, Har96, Sch09, SS15, Sus16, Boa19, Cha61, Cha62, Cha63a, Cha63b, Cha64a, Cha64b, Coh36, DG09, Edw72b, EE51, Feu75, Feu76, For71b, Hen86, Hen88, HB14, JTR94, JTR95, Lop60, Pap09, Seg83, Ste61, Ste69a, Ste69b]. correspondents [Bau88]. Corrigenda [Ano97h, Ano60a]. Corrigendum [Ano55a, Ano09c, Ano12a, Fel03]. corruptione [Bur14b, SBT10]. Cosans [Qui13]. Cosmic [Bon14]. Cosmochemistry [Kra00]. cosmographer [OEMB21]. Cosmology [Cha12a, God09, Kra09a, Ste09b, Arm48, Har72, Hir08, KO83, Kra11b, Lep06, Fol14b]. cosmopolitanism [Aga19]. Cosmopolitics [Ste10c, Bsc14a]. Cosmos [Wer14, ATE11, Reg19, Ste09a]. Costa [Dar13]. Cottage [GG93]. Council [But57]. Count [Ols70, Spa55]. Counter [Tre86]. Counting [GG92, Rus21]. coureur [Sme57a]. cours [Aug04]. course [Edm75, SL77, Ste36a]. courses [Mil83]. court [dC88]. courtier [Sch19]. Courtrooms [BS12]. Cover [Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13c, Ano13d, Ano13e, Ano14c, Ano14d, Ano14b]. covered [GG76]. Cozar [OEMB21]. Crack [Ras17]. Cracking [Mel09, Del11b]. Cradle [LPST11, McC14]. Craft [Di 17, Key90]. Crafting [Tau10]. Craftsman [Bel12, Ost92]. Craftsmen [NNS^{+}06]. Crawford [May21]. Craze [Att85]. created [Bik75]. Creation [Bil16, Fin12, Hir07, Lai77, Ole23, SM15, Hir07]. creationism [Cos09, Qui13]. Creative [Mil13]. Creativity [Ben13a, GG91a]. credit [Hon96]. Crelle [Ecc76, Ecc80, Ecc76, Ecc80]. Crémieu [IM89]. crime [Rus95]. Crisis [Jac15a, GM94]. Critic [Rob54b]. Critical [San17, Ari73a, SBT10, VT17]. criticism [Ait62, Kem68]. criticisms [FB22]. critics [LS74, Ora75]. critique [Bur14a, Can75a, Cha75, FG83, Sha72, SBB^{+}75, Thi73, Vla10]. critiques [Hut77]. Crockatt [Row18]. Croll [Ham82]. Crookes [Bey11, DeK83]. Crop [Fus58, Hol17b]. Cross [Wal00, Dek90]. Cross-currents [Wal00]. crossfire [Alm22]. Crossroads [GG14, Maz12, Tob12]. Crowdsourcing [Kär14]. crucial [Hak82]. Cruickshank [Cou59]. Crusaders [Gar04]. cryolite [Kra95]. Cryptozoology [Bry10]. crystal [PB45, PB47, PB48]. Crystallina [McC39b]. Crystals [Mau15]. CSI [Ada17]. cube [Ols69]. Cullen [Cro59, Dob36, Wig55, Wig56]. Cultivating [Hyd19]. Cultivators [Cro10].

Cultura [Nut14, Cor11]. Cultural [Bil16, Lyn15, Sht17, CBL81, GG91a, GG11b, HY87, McA12, SM15]. Culture [Bez75, Bow13, Car21, Coo09, Das09, Die12, Fol11, Hei14, Hen00, Lov11, Maz12, Par16, Par13, Wal12, Zui04, Alb09, Ego19, Har09a, Hig17b, Ink78, Lun90, Mar11a, Ros98, SP85]. Cultures [Ano08e, Bri13, Mar15, Nyh16, Pop12, Aub23, Fre11, Hyd19, Kin19a, Mac09, Rad18, Vos13, Bur15]. Culturing [Har09b]. cum [Yap83]. Cumulative [Ano09b, Lev09b]. Cup [Ano09i]. curators [BJ96]. cure [OS97]. curia [Not17]. Curie [Dav95, Mic13]. Curiosities [Mar13]. curiosity [Kra95]. curious [Uss78]. current [Hum37, Low03]. currents [Wal00, Wei85b]. curriculum [Kau79]. Curry [Van19]. Curtis [McC13]. curve [Erl92]. Cutting [Nut21]. Cuvier [The86]. cybernetics [Aga17b, DS97, Kli15]. cybernétique [DS97]. Czechoslovak [Smi23a]. Czechoslovakia [Urb80].

D [Boa12b, Bry13, Jar18a, Sha19b, SS23]. D. [Kid13]. D.S.I.R. [Ver94]. d’Agriculture [Sme56a]. Daguerre [Woo97]. Daguerreotype [Woo97]. D’Alembert [Sch13a]. d’alliance [Muc98]. Dalton [Gas07, Mau69, Tha66]. Daltonian [Fle74]. Danes [Hje10]. Danica [Swe09]. Daniel [Old72a, Goo67, Wil92]. Daniell [Mer98]. Danish [Ran63, Whi13]. dans [BV83, Gen07, LS15a, Ray15, Swe21]. Dante [Mei54]. d’Antic [Ber12]. d’Antoine [Kah01]. Daqing [Mil13]. dark [Dek90, Mor90, Win23]. Darwin [Bey14, Hol17a, Moo87, Qui13, Alt07, Bar75, Bou17, BJ96, Cos09, Cur87, DD21, End16, Eva17, Fan09, Feu75, Feu76, Her18, Hje10, Hoa55, Kot78, Mar11c, Nor16, Qui12, Rob54a, Rup09, Rus80, Sch80, Ste61, Ste69a, Ste69b, Tee79, Win23, de 58, HA18, Ray18, Smi18]. Darwinian [Bel99, Gay10, Hes21, Wal04, Yal22]. Darwinism [Qui13, Bow78, Cos09, Cro96, Mat11, Moo87, Shi81, Wal76]. Darwins [End18]. Daston [Bou13, Por16]. Data [Lus14, Dra82, Edw10, Mac82]. d’Auteroche [Arm54]. David [Ber96, Ber19, Gol19, Kaa17, Ano15, Ste00b, VvM92, Wit13, Wol13a]. David’s [Old91, OM95]. Davy [Tat17, Ful64, Mil03, Rus59a, Rus59b, Rus63c, Spi68, Wei50]. Dawn [Ale10, Coo13]. Dawson [SP92]. Day [MT72]. Dead [Fra14a, Mac10]. Deaf [Kvi22]. Death [Ham03, Koo10, Ste10a, Ste00b, Dav78, Sch75, SS58, Sch13b]. Debate [Bel99, Can75b, Her17, Nor01, Pan05, Alm22, FB22, Bro77, Can15, Can75a, Chi70, Ink72a, Jon74, Kip87, Pel97, Yav93]. debates [Bru79]. debating [Yal22]. debt [McC39a]. début [LS15a]. Decay [Mon05]. Dechales [Gam21]. Deciphering [Cre10]. Decision [Gar04]. Declination [How02]. Decline [Aug11, Att85, Pum87]. Decomposition [Kim06]. deduction [Kat06]. Dee [Van01]. Deep [Cha16]. Deepest [MN13]. defects [Mea70]. Defence [Bur12d, Bar75, Hil84]. Defending [Gou10, War14]. Defied [Koo10, Sch13b]. defining [Rib20]. definite [Fle74]. Definitely [Bec01]. definition [Dek93]. Definitionen [Gab78]. definitions [Gab78]. deflecting [Bur65, For98]. Degeneration [Law09a]. degli [Chi08].

Degradation [Lyn15, McA12]. Degrees [Gre84]. deified [Hak92]. del [Bos03, Gia23, Gre87, Mil88b, She11]. delayed [Dek96]. Delbrück [vH92]. delectat [Ges23]. Delft [ZA16]. Delgado [Gre87]. delivered [Spr77]. Delmedigo [Adl97]. Deluge [Nor42]. Delusions [Mea09]. Demands [Bur12d]. demolished [Wal97]. demonology [Att85]. demonstration [AM76]. demonstrators [Wes97]. denudation [Dav66]. deodorized [McB91]. Dépôt [Bre91]. Department [Jar98, Mac18]. departures [Dra81]. dependency [Tod93]. Derby [Ell00, Rob53a]. Derek [Tur84]. Derham [Atk52]. Dernier [Bre93]. Descartes [Ait55b, Bro81, Bro36b, Buc08, Cha19, Coh36, Gou22, Lau66, Mit79, Rog72, Sch13c, SB13, Sch19, Tho16]. Descartes-agonistes [Sch13c, Tho16]. Descent [Bas15, Hal14a, Dra81]. described [Lin63]. description [VT17, de 50b]. Design [Fan09, Cha83a, Hac79a]. Designing [Jam12, Lig10]. designs [Fou89]. Desperate [Ree14a]. Detail [Egm16, Kus18]. detection [Wal36]. Determination [Mau15, Hop57b, Hop58, Hop59a]. determine [Bur09]. Deutsche [VW10, Wal14]. Deutschlands [Hen89]. develop [Wil60c, Wil61]. developing [Het83]. Development [Ade10, Bon14, Bro92, Bru57a, GG13b, Haa09, Har71, Rus04, Wea09, Wil01, Bra95, Bru57b, Bru58a, Bru58b, Caw77, Clo54, CC56, Cro96, Dav84, Flo60, For66a, Gee93, Geo52, Hen80, Hum37, Ink91, Lan83, Mat76, Mes97, O’B65a, O’B65b, Pas51, Por78, Roc85, Sch39, Sch90, Ste80, Swi65, Tod93, Tur91, Ver16, YO89]. Developments [Cav06]. deviation [Arm47]. Deviatory [Tan80a]. diagonal [Cha53a]. Diagram [Sid07, Arc90]. diagrams [Cro59]. dial [Mil93]. dialectical [Qua72]. Dialektik [Lew77]. Dialogue [YT13, Bal06, Dra88]. Dials [Kin04]. Dialysis [Tan11]. Diamagnetism [Jac15b, BM94]. Diamond [HBP08]. Diana [Mar11b]. diaries [Cha68]. diary [SR73a]. Dibner [Mil13]. dictionary [Bre16]. Dictionnaire [NS81]. Did [Cam80, Hah15, DD21, Leh12b, Zig80]. Didacus [McC37]. Difference [SS13]. Different [Duc17, Orr14, BL21, EE38]. differential [Dub63, Gre82, Gre81]. Differentiation [Mae05]. Difficult [Cro03]. Diffraction [Bro15, Bro18, Bro22]. Diffusion [Zui04, Gre84]. Dig [Law15]. Digital [Par03]. Digne [Ben09]. Digne-les-Bains [Ben09]. dilemma [Dav66, Ric86]. Dilemmas [Joh15b]. dilettante [Hen18]. Dimension [Aug11]. dimensional [De 17]. dimensions [Nut21, Wri92]. Dimetiri [Big22]. Dinosaur [Bri10, Bry13]. Dionysius [Hay81]. diphtheria [OS97]. diplomacy [RG23, Kri19]. Diplomat [Dau13, Sch19]. diplopia [Far19]. Dirac [Wri16]. direction [Pav17]. Directions [Tir23]. disagreement [Kaa17]. Disc [Van01]. Disciplinary [SS13, Cot08, Tay08]. Discipline [Par03]. Disciplines [Ama06, Par13, Alb09, Dol77]. Discorsi [Rap12, Rap15a]. Discorso [Ari73a]. Discourse [Qui09, Ave20, Low96, Ari81]. discourses [Sin50b]. discover [Cam80]. Discovered [Cha16, MVK02, OM95]. discoverer [Tre72]. discoveries [Sne75, Sok94, Wil55b]. Discovering [Bow16]. Discovery [Cun04, KL07, Mil02, Mil11a, Slo03, Bes68, BM94, Cha91, Dar85, Dek90, Dob45, HS22, McC36a, Mea66, Owe55, Pat38, Raw74, RP98, ST51, Woo71b]. discussion [GG75b]. Disease [Bea12, Fra13, Joh10, Ree14a, Wex08, Mau90, Wil55c, Cha13]. Disks [GG11a]. Disorder [Bea12, Ree14a]. Disordered [Ver13, Wak09]. Dispensatory [Bur12a]. Dispersion [Ber96, Che98, CR96]. display [RG23, Rad18]. Dispute [Ela13, Péo10, TT99, Ecc80, Wil92]. disputes [McC75]. disquiet [Sla19]. Disrupting [Kai10]. Dissection [Ano08e]. dissemination [PK87]. Distance [BG94, FB22]. Distances [Sid07]. Distinguished [Clo55]. ’Distributing [Mil02]. distribution [Gow82]. District [Old99a]. Diverse [Ano06h]. Diversity [SK11]. dividing [Bro92]. divination [Gel21]. Divine [Bow74]. division [Coh87]. DNA [Bra10, Cre10, Dav90a]. Do [Ano05g, BM17, Lei20, Lin22]. Doctor [Mcv21, McB91, Bro71]. doctoral [GG75b]. doctrine [Ebe39, McC36b, Tay08]. Documentation [DK57]. documents [Lin63]. Does [Ano05g]. Domain [End17]. Domestic [Alt07, Ree14a]. Domesticating [Goo08, Sta13]. dominance [Lun90]. Dominicus [Bur90]. Donald [Dau14b]. Donata [Maz10]. Doncaster [Rus22]. Dongen [Del12]. Donn [Rob63]. Donna [San17]. Donny [Tre80b]. don’t [Lin22]. Doody [Lov13]. door [The94]. Doppler [Thi71]. Dopplerprinzips [Thi71]. Dordrecht [Sch18b, Vic04, Wit13]. Dörffel [Arm51]. d’origine [Bur12c]. ‘Dorlodot [De 05]. Dosage [Gen07]. Dossie [Gib51f, Gib53d]. Dostoevsky [Gor16]. double [Cha88]. Douglas [Ano68c]. Down [Dij04, McC91]. Dr [Bro55, Mal82, Mil11a, WN16]. Dr. [Sch09, Stu09]. drafts [Ree81]. Dragon [SHM13]. Drake [Fin02]. Drama [Tka14]. Draper [Plo78]. Drawing [Ric86, Swe03]. Dream [Sla03]. Drebbel [Gib48]. Dresser [Far18]. Dreyer [Ste10b]. Drift [Lal12, YO89]. Drilled [MN13]. Drittel [Hen89]. Dropsy [Tan11]. Drug [Mae05]. Drugs [Win16, Ear61, Ear63, Hed22, May21]. dry [Sim92]. dualistic [GG91a]. duality [Hen80]. Dubious [Reg09]. Dudley [Ols71]. Duel [Ale10, Coo13]. Dufay [Coh51]. Dufresne [Swe70a]. Duhem [Hen88, Hen88, Mar76]. Duke [Eme02, Eme13, Woo15]. Dumas [Klo85a, Klo85b]. Dumfries [Und67]. Duncan [Ver16]. d’une [Cam09]. Dunglass [Eyl63]. Dunn [Kin04]. Durham [Hug51a]. During [Adl17, Bon14, Gau12a, Pre08, Bra76, Cha70, Hol17b, Lan83, SR73c, Tom14]. Duruy [Fau05]. Dury [Fau05]. dust [Cah90]. Dutch [Ano09i, dBP86, dC88]. Dwarfs [Wes10]. Dyck [Car18]. Dyeing [Mar01, Key90]. Dyes [Tre70]. Dynamical [Cra54, Bra90, Sha75b, Tur55]. Dynamics [RS09, Bra88, Nye89]. dynamo [Car92]. Dyse [Dav57].

e-ISBN-13 [Wit13]. E. [Bow84, Har74]. Ear [War07a]. Earl [Eme02, Eme13, Wes97, Woo15, Tur67]. earlier [Fra39a, Jac52a, Jac52b]. Earliest [Bur00, Bur98, Mad61a, Not18, Row50a, Sme57a]. Early [Bad11, Cab73, Car17, Cob02, CSVK80, Cor36, Cro05, Cro09a, Dan10, Dek87, Dij13, Dun13, Dup11, Ear61, Fli08, Fol11, God09, GG11a, GWN^{+}11, dVH36, Jac15b, Jac14, Jon13, Kär14, Kra09b, Maf13, May21, McB91, Mil03, Mil04a, NNS^{+}06, Nut14, Old99a, Ots05, Pag03, Pal83, Par03, Pom11, Rab77, RC08, Sim66, SD13, Ste10a, Wad15, Wal01, Wil15, Zia94, dA21, Ako21, Ari73a, Arm56b, Bak64, Bar57, Ben91, Bow09b, Bra76, Byr80, Cam23, Car97a, Caw77, Cha52, Cha53b, Cha54a, Cha54b, Cha70, Cop87, Cor11, Cot77, Cro96, Dun21, Ebe39, FG83, Fru16, Gee93, Ges23, Glu23b, Hal20, dVH67, Hon96, Hug51a, Jar23, JR12, Jur10, Kaa17, KL09, Kit82, Kra95, Lan83, Law08, Mac18, Mal20a, Mal20b, Mar21a, MG10, Mel21, Mil74]. early [Mil81a, Mor88, Nay80, Nay83, NG97, NF73, OS97, Ora75, Par42, Par51a, Pas51, Pav17, Pri52, Pri55, Roc85, Rud07, Rus78, Rut39, STDH22, Sch08c, Sco81, Sha75a, Sha73, Shu68, Sim92, Sme54b, Sme55a, Sme55b, SR71a, SR72c, Ste19, Und63, Und67, VB11, Wei19, Wit96, Woo71b, Woo71c, Yeo14, Yeo18, Gib51c, Gib53a, Hun23, KS10, Mar13, Nap09, Tom14]. early-modern [Jar23]. early-seventeenth-century [Ben91]. Earth [Bec14, Old74, Ano06f, Bru79, Bur65, Cha16, Dav66, Dea81, Gre84, Gre87, Gue09, Hug55, Maz10, ML^{+}12, Por78, RO82, RO84, SW67, Swe03, Vac00, Dea75]. Earth-Bound [Bec14, ML^{+}12]. earthquake [Dea93]. earthquakes [Dea89a, RO83, Wil87]. East [Arn09, Ell83, Low03, YO89, Wat82]. Eastern [GWN^{+}11, Bur12b, Gib51a]. Eberhard [Sau76]. echo [DR54]. echo-sounding [DR54]. Eckert [Kra13]. Eclipse [CS12, For72b]. Eco [Wea09]. Eco-Development [Wea09]. ecole [Sme54b, Bra75, Bra76, CP97]. ecological [Lee23]. Ecology [Ank15, Ano06f, Gay13b]. economic [CC47, Ink91, Jar18b]. Economy [Aga17a, Mil11a, Ink88]. Ecuadorian [Gen22]. ed [MS22]. Eddy [Boa12b, Hen18]. Eden} [Sve19]. Edge [Nof14, Mol84]. Edinburgh [Boa12b, Edd08, Bry90, Can75a, Can75b, Chi70, Jen16, Sha75a, SP92, Swe70b, Wil91b]. Edison [Fin12, Fox96, Wil07]. Edited [Lai04, Bec13a, Ben13b, Boa19, Bou13, Bur14a, Bur14c, Car18, Dij13, Dun13, Fol14a, Gue13, Hal18, Low03, May21, McC13, Ray18, Tau14, Van19, Wal14]. Editing [Gui15]. Edition [Ben14, Dar13, Gui15, Lat15, LS15b, DC09b, FS21, Gru11, Mad62, McC36a, SBT10, Bur14a, Vla10]. Editions [Pal12, Fra39b, Sim66, Sme67, Swe21]. Editor [OH85, Tur99, Bar11, Cha12a, Hig17a, Hor87, Lai04, Lat15, Vic04, Lev09b]. Editorial [Ait88, Ano36a, Ano96i, Ano14a, IL12, Ano74j, Lev14a]. editors [Akt09, Dom09, Jam12, Lat15, Lev09a, Mar11b, Mil13, Nea50, Pal12]. Edmond [Ste36a]. Edmund [Arm56c, FG83, Hua22]. Édouard [Pap09]. Education [And14, Bri00, LMBM06, Nor12, RC08, SBB^{+}75, SR71a, Bra75, Bra79, Don79, Lar61, Lun90, Mes96, Rob54c, RS74, SJ78, SR72c, SR72a, SR72e, SR72b, SR73c, SR73b, Tur73, War88, Whi86b, Gay00]. educational [BS65, SR74a]. educationalist [HR89]. Edward [Dom09, Gos15, Lev94, de 49, Cro96, Ham03, Har12, Kur99, Lep06, Old72a, Plo78, Rus78, Ter37, Wal97]. Edwardian [Olb91]. Edwards [Lus14]. Edwin [Pav17]. effect [Kat06, WO87, Wil57]. Effects [Bos03, CSVK80, GG91a]. Efficiency [Hon95]. effort [Mac89, Mac93, dBP86]. Egaña [Cha12b]. egg [Foo69]. Egmond [Kus18]. Egypt [Det18]. Eichstadt [Kre23]. Eighteenth [App90, Bed93, Cro09b, Eis05, Kid14, Kle12a, Leh12a, Nic06, Ree14a, Ser12, Tom09, Vac00, War12, Woo53, Arm53, Bre16, Cha21, Coc56, Dav66, Die16b, Dun62a, Dun62b, Dun23, Dym37, Ear63, Gib51e, Glu23b, Gou83, Hac79a, Hom79, Jae10, Low96, Mac20b, Mea70, Mil83, Mil88a, Mil47, Oli69, Pop12, Rob62, Sme87, Swe64, Tre63, Tre64, Tre65b, Tre65a, Tre70, Val96, Wal36, dC88, Bea12, Car21, Mar01, McC83, Mil11b, SC04]. Eighteenth-Century [Cro09b, Eis05, Leh12a, Nic06, Ser12, Tom09, App90, Kle12a, Vac00, Woo53, Arm53, Bre16, Coc56, Dav66, Die16b, Dun62a, Dun62b, Glu23b, Hac79a, Hom79, Jae10, Mea70, Mil83, Oli69, Sme87, Swe64, dC88, Bea12, Car21]. Eileen [Wil11a]. einigen [Hen89]. Einstein [Bey14, Del11b, Hen86, Her17, Pal83, Ano12a, Asp69, Byr80, Byr81, Can15, Cro16, Del11a, Del12, Feu71a, Feu71b, Har72, Har96, Hen86, Ill89, JL14, Mar11b, Mar11c, Mel09, Mil81a, Pal83, Sau14, Slo11, Zia94, vD10, For71a, Mar11b, Tho11, Row18, Sta21]. Einsteins [Ill89]. Einteilchen [Zac72]. Einteilchen-Schalenmodells [Zac72]. Elaine [Wei19]. elastic [Wil56d]. Elazar [Pis14b]. Elborg [Fra13]. Elderton [Lov79]. Election [Bec09, Lus18]. electric [Gee93, Hum37, IM89, Rob95, Wal36, Mil11a]. Electrical [Cah10, Ish14, Pas51, Rus59b, Dav98, Haw68, Hop57b, Hop58, Hop59b, Lei73, Rut39, Yav93]. electrically [Kan70, Kan71]. electricities [Haw71]. Electricity [Fin12, Goo08, Sta13, FC36, Hom79, Hop59a, Kip87, Wal37, dVH67]. electrochemical [Rus59a, Rus59b, Rus63a, Rus63b, Rus63c]. electrodynamics [Sne75]. electrolysis [Rus59a]. Electromagnetic [Cav06, Cra54, CR96, Hop59a, Sne84, Gos77]. electromagnetism [Sne75]. electrometallurgy [Mer98]. electron [Cha91, Owe55, RP98]. Electronic [Par03]. electrophysiologists [Hof36]. electrostatic [Hop59a]. elektrisch [Kan70, Kan71]. Element [Kra09b, CK76, Hei67a, Lov74]. element-instrument [Lov74]. élémentaire [DK54b]. elemental [Uss78]. elementary [Ari81, Hug52]. Elements [MWO07, Deb67, HS22, LR00]. Elevation [Nor01]. Eleven [OG68]. Eliakim [Par84]. Elie [dBP86]. Eliot [PM50]. Elisabeth [Wal14]. élites [Bri07, Maz10]. elixier [Fig72]. Elizabeth [Ben13b, Cre10, Jon95]. Elizabethan [Deb62, Jar18b, Joh91, Mor68, She09]. Ellert [Kra13]. Elliot [Man93, PM50]. elliptical [Bra82a, Bra82b]. elsewhere [Sme54b]. Emanuel [Pia21]. embassy [CBL81]. Embodied [Car17]. embryology [Bla10]. Emergence [Ama06, Ano06l, Das09, Pal83, Rei04, Kim95, Mil81a, SP85, Smi76, Zac72, LS15a]. Emerging [Gui11b]. Emery [Bal23]. Emma [Dij13]. Empire [Bra02, Bri13, Cha12b, Det18, Win16, And09b, Mac09, Mau18, And09b, Lin16, Hay18]. Empires [Rap15b, See18]. Empiricism [Car17, Jur10]. Empirics [Pom11]. Employed [Swi12]. Encounters [HA15, MHO13]. encyclopædias [Hug51b, Hug52, Hug53, Hug55]. Encyclopedia [Lov13, KL09, Doo10, Jon13]. End [Del09a, Boa12a, CR96, Gre05, Lau21, dBP86]. Ending [CCC57]. Ends [Cre10]. energeticist [Hak92]. Energy [Jur10, vD08, HL91, Kid81, Kra17, Ole23]. engagement [Han21]. Engine [Mil08, Bro92, Gro72b, Gol19]. engineer [App96, Bre91, Rob62, Sau76, Sch19]. engineer-geographers [Bre91]. Engineering [Ano05i, Ber09, Cha12b, Dym11, Lun90, Ozd14, Smi80, Yea11, Yi13, Bra81, Car10, Die16b, Don88, GG84b, Hon96, Woo36, Wri92]. Engineers [Bra92, Maf13, Hop81, SR71a]. Engines [Bro15, Bro18, Bro22]. England [Bec13a, Edd13, Ano07f, Arm55, Arm56b, Ben91, Bra94, Bro92, Coc56, Dan10, Don88, Dun94, EE38, Foo52, Geo52, Hal14a, How03, Jac52a, Jac52b, Jar23, Joh91, Jon37, Jon19, LPST11, Mad55, Mul73, PdC10, Pum87, Roc81, Sha73, Sim66, Ste10a, SR73b, Yeo18, Wei19, Ano07e, Bas15]. English [Pis13, Wil15, Ada17, Arn09, Bar98, Bed97, Ell00, Hug51b, Hug52, Hug53, Hug55, Lau66, Mil85, Nic39, Old99a, RS01, Rog72, Sch08a, Sch08b, SC04, SR71a, SR71b, SR73c, Tur00, Wei51, Yeo14, dV21]. enigma [Gil76, Hen21a]. enjeux [Muc98]. Enlightened [Eme13, Jon99, Mil01, Woo15]. Enlightenment [Ano05j, Boa16, Cha14, Cra12, Joh15a, Lyn15, Mar15, Vos13, Eme88, Tom12, Wil09, Car22, Edd13, For15, Gas10, Gau12a, LS15b, McA12, Rat09, Rya23, Ste20, Sto13, Wol13a, Mil16]. Enquiries [Hal20]. Ensei [Thi12]. ensure [Bal23]. Entanglement [Yea11]. entanglements [Kri22]. Entente [Pre08]. Entering [Reg09]. Enterprise [Kur03, RS71, Uss80]. enterprises [See18]. entomologist [Gay10]. Entre [Kah01]. entrepreneurialism [Coo90]. entrepreneurs [Kei84, Mel21]. entropiefunktion [Hut73]. entropy [Hut73]. Entstehung [Zac72, Hon84]. enunciation [Col95]. Environment [Ree14a, Ako21]. Environmental [Gar04, Kri19]. Environments [Gue13, Vet10]. Eötvös [Ill89, Ill89]. Eötvös-Versuch [Ill89]. Epact [Dek93]. Epicurea [McC39c]. Epidemica [dVH67]. Epigenetics [Bow17]. Episode [Bel99, Uss78]. epistemological [GG91a]. epistemologist [Cur87]. epistemology [Wri16]. equal [Coh87, Mil96]. equation [Cro59]. equations [Gre82]. equatorial [Cha95]. Equivalent [Bur12d]. Era [Bro15, Dav90b, Yal22, Adl17]. Erasmus [App90, Hoa55, Rob54a]. Erhaltung [Jur10]. Eric [Lee13, GG91b, Mea85]. Erik [Hay18]. Ernest [Ver56]. Ernst [DD21, Bla78, Mat11]. Errata [Ano36b, Ano87g]. Erratum [Ano86h, Ano89g, Ano92l, Ano95h, Ano05d, Ano09d, Ano13b, Ano16a]. error [Ait65, Hon87, Mac82, Ohn23]. erzeugter [Kan70, Kan71]. especially [Bod56, GG92]. Esq [Pat36a]. essai [Ben09]. Essay [Ano74i, Ano75g, Ano75h, Ano75i, Ano75j, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano77f, Ano77g, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano79g, Ano79h, Ano79i, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano83h, Ano83i, Ano83j, Ano84g, Ano84h, Ano84i, Ano84j, Ano85g, Ano85h, Ano85i, Ano86i, Ano86j, Ano88g, Ano88h, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93i, Ano93j, Ano93k, Ano94k, Ano94l, Ano94m, Ano94n, Ano94o, Ano95i, Ano95j, Ano95k]. Essay [Ano95l, Ano95m, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano98e, Ano98f, Ano99e, Ano01e, Ano03e, Ano04e, Ano04f, Ano05f, Ano05e, Ano05g, Ano07e, Ano12a, Blu13, Can10, Cor75, Del11a, Del12, Gre87, Hen00, Hor87, Hud03, Jac15a, Joh15b, Jon99, Kid13, Lai04, Mar11b, Mau15, Moo87, Moz04, Nyh16, Pal83, Ram15, San00, San17, Sha19b, Sla03, Sta19, Swe03, Swe21, UB15, Vic00, Vic04, Vic12, Wei15, Bak62, Wil23]. Essays [Nic11a, RN15, Wan18b]. essentially [Tre78a, Tre80a]. Establishing [Gia23, Tay08]. establishment [Dub66, HL91, Olb89]. estimation [Wal36]. Ethel [Lov79]. Ether [Lal12]. Ether-Drift [Lal12]. ethic [Kem68]. Ethics [Gar04]. ethos [Yeo18]. Étienne [Kah01]. Études [Bur12c]. Etzler [Arm55]. Euclid [Sim66]. Euclidis [Bur14c, Bus10]. eudiometer [Osm58]. Eugenics [HA19, Kre11, Lov79, Ots05, Sea79, Bow84, Byf19, Far79, Jon95, Wei85a, Lov79, Wei85a]. Eugenics-Land’ [Lov79]. Euler [Dom09, Ait55a, BS07, Cha21, Hom88, Kno89]. Eulers [Kno89]. eum [Fra39b]. EUR [Car18, Hal18, Hay18, Hei19, Jon13, Kra13, Mel21, Nut21, Swe21, Wit13]. EUR118 [Vic04]. Europa [Bri07, Maz10]. Europe [Dij13, Ako21, Cop87, Dal14, Dol77, Gor17, GWN^{+}11, Hes17, Hun23, KS10, Lun90, Mal20a, Mar13, Nap09, Not21, Sch08c, See18, Sil18, Ste19, Aga17a]. European [Dun13, Lan13, Bri00, Cha83b, Cha84, CZ19, Des08, Hoc83, Kre13, Mac23, PNGP09, She11, Wer10]. Evaluation [Edw04, Mil03]. Evan [Gue23]. Evans [Ray18]. evaporation [Dym37]. Evelyn [Rus04, Tay52a]. evening [Mil83]. eventful [Ham82]. eventual [Coc83]. Ever [MN13]. everyday [Wei19]. Everywhere [Ano07i]. Evidence [Bry92, Cha68, Cra85, Sim92, Rus63c, Woo71b]. Evolution [Bas15, Bel99, Bur17, End16, Fan09, Goo11, Sch75, Chi08, Ebe39, Eyl63, Hal14a, Hje10, Mal82, Old74, Pav17, STDH22, Tay45, Wil55a, Wil56a, Wil56b, Wil56c, Fit17]. Evolutionary [Bor10, Har85, Her18, Sha72, Bar14]. Evolutionism [De 05]. evolutionist [Lig16, Ste61, Hal18]. Evolving [Bur12d, Fra09, Mil13, FY09]. Ex [Ano06e]. Exact [Plo11, Ben09, Dun23, dBP86]. exactes [Ben09]. exactness [Aub23]. examination [Clo55, Ink77, Lam97]. examining [CZ19]. Example [Pis13, Lan83, Rom06, Sch08a, Sch08b]. excellent [Har57]. Excerpts [Swe63, Swe67, Mel88]. Exchange [Bil16, Maf13, SD13, Arm58b, Die21, SM15]. Exchange/Conflict [Maf13]. Exchaquet [Tou89]. Exclusively [BG99]. exemple [Rom06]. Exercises [Can10]. exhibita [Ray82]. exhibition [Ger23, MSD23]. exhibitions [RG23]. Existence [McG78, Tre80b]. Exotic [dC09a]. Expanded [Old10, Sok96]. Expanding [Vet04a, Vet04b]. expansion [Sta88a]. expansive [Wil54]. expansivity [Spu87]. Expedicion [Gre87]. expedition [Hig17b]. expeditionary [Hig17b]. expeditions [Lev86, Lev94]. Experience [Sla03, Ink91]. Experiences [Cav06, Jam12, FL07]. Experiences/Victorian [Jam12]. experiencing [CB22]. Experiment [Car18, Hil99, Ill89, Kär14, Pro18, Bak21, DR54, DS16, Hak82, Hed22, Mas73, Nay80, Roc85, Tho55]. Experimental [Bad11, Buc08, Cro03, Cro05, End17, Mit12, Tat17, dV21, And19, App90, Beg21, Cah90, Che98, Dra82, Ear60, Gia23, Hac79a, Hon87, Jon50, Mac82, Mil83]. Experimentalism [Coo90]. Experimentalist [Lev14b]. Experiments [BS12, Cec11, Ear63, Fol14a, Ku09, Lal12, URRF05, DK54c, Hal55, IR11, Kvi22, Mac76, Mad62, Nay83, Pat36b, Shu68, Sne76, Sne84, Spu87, Wil54, Nut21]. Expert [Kon12, Leh12a, Spi69]. Expertise [Dij13, KS10, Bre93, Leg17, VPW16]. Experts [Kle12a]. explain [Ros21]. explained [Mor08, Sch36, Nic13]. Explaining [CR02, Rus04, Ste09a, Ked16]. explanation [Mil93, Nye89, Tur55]. Explanatory [Bur12d]. Exploit [Pop12]. Exploration [Ano06f, Jon83, Nic11b]. explorations [Dek87]. Explored [End17]. explorers [Swe74]. Expo [Smi23a]. Export [Val96]. Exposed [Edw09a]. Expounding [Bur10a]. extant [Win66]. extension [GG76]. extinct [Tre78a]. extinction [Win23]. Extraordinary [Bey10, Gos15, Ari79, Har12]. Extreme [Stu09]. Eye [Egm16, La 11, Eas86, Kus18]. eyepieces [Rud07]. eyes [Lar91]. Eyewitness [Edw09a].

F [Boa19, Dun13, Kid13, Low03, Mil91b, Sch18b, Sme55a, Sme55b, Sne81, Ver16]. F. [Sza72]. F.R.S [Atk52, Eyl63, Gib52c, Ter37]. F.R.S. [Cha72b, Col52a, Col54a, Eyl37, Mad59, Pat36a, Rob62, Row50b, Rus49, Rus50, Rus51a, Rus51b, Woo71b, Woo71c]. Fabri [Pis14b, Ela11, Ela13]. fabrication [Bre93]. Fabricius [Arm58a, Ekh10]. fabrique [Cam09]. face [Fox96]. Fachzeitschriften [Ecc76]. facilities [Bra76]. Facsimile [Dar13, DC09b, McC37]. Fact [SB13]. factorial [Nor75]. Factors [MWO07]. factory [Cam09]. Facts [Por01]. faculties [Wil23]. Faculty [Fli08, Dav95]. Failed [Bsc12, Cha09, Fol14b, Mac20a, Bro95, HY87, Kra11b]. Failing [Tan11]. Failure [Bra95, Par16, Wil07, Bry90, Kat06, Str93]. Faith [Lov11, Duc97]. Faiths [Kah10]. Fall [Ber12, Hay81, Nay80, Nay82, Smi23b]. falling [Arm47]. falsification [Tre78a]. Fame [Fra10]. family [Mad55]. famine [CC56]. Faraday [BM94, Can10, Car92, Goo80, Mil03, Pel36, SS15]. Farey [Wat82]. Farisi [Bur09]. Farkas [Big21]. farmer [Jon85]. Farnham [Edd13]. Farquhar [Ano59a]. Fashionable [Bea12, Ree14a]. Fashioned [Rus04]. fatalisms [Yal22]. fate [Byf19, Hon84]. Father [Bey10]. features [GG94]. Fechner [Win49]. Fecundity [Ano06j]. Federal [Mei79]. Feingold [Vic04]. Felice [Ear60]. Female [McA12, Lyn15]. Feminism [Lov79]. Femmes [Mic13]. fenêtre [Thi12]. Fengzuo [CZ19]. Ferdinand [LMBM06]. Ferguson [Mil83, Mil85, Pat39]. Fermat [Moz04]. Fermentation [Fin09]. Fermilab [Kra11a]. ferrimagnetic [Qué90]. ferrimagnétiques [Qué90]. Fertile [Smi23b]. Feynman [Ott11, Gro14]. Fiasco [Lus18]. Fibonacci [Big21]. Fiction [End16, Smi15, Buc13]. Fictions [ATE11, Wer14]. Field [Gue13, Sau14, Vet04a, Vet04b, Vet10, GG91a, Kin19b, Roe90, Adl17, Cra54].

Fields [Fra14b, How02, How03, RH11]. ‘Fieldworker [Kim12]. fifteenth [VT17]. Fifty [Por16, RD16]. fig [Lau21, Mar21a]. Fighter [Flo96]. Fights [Ste00b]. figs [Bri13]. Figure [Ano12c, Ano12d, Ano12e]. figures [Cah90, Kra13, Par13]. Figuring [Coo09]. filia [Spe16]. Filling [Rei14, Ste11a]. fin [Boa12a, LS15a, Law09a]. finance [SR74a]. Finances [Mar13]. financial [Bra79]. Finches [Bey14, Mar11c]. find [FB22]. Finding [GH08]. finds [Tre91]. fine [Han21, Wil60a, Wil60b, Wil60c, Wil61]. Fingerprinting [Bra10]. Finn [Mil13]. Fire [Deb67, dV21, Hen21a, Hom79, Lov74]. Firenze [Fol11]. Fireworks [Wer10, Lan13]. First [Dar13, Lat15, OS97, SP11, Val09, Ait78, CR96, Cha21, DC09b, Dav90b, DR54, Edm75, Eyl37, EE38, FS21, Het72, Hon96, Mad62, OJY96, Old97, Sch36, Sch18a, Sme55c, SR74b, Tre91, Tre78b, Tur53, Col52b, Hal13, Nye89, Sil18]. Fiscal [Clo54]. Fisch [GG97]. Fischer [Joh15b]. Fish [Orr14]. Fisher [Nor75]. Fishes [Orr14]. Fitness [Hal14b, ZB10]. Fitting [How02, How03]. Fix [Bec09]. flames [Par45]. Flamsteed [Wil87]. Fleeming [Smi80]. Fleming [Hon96]. Fleury [Bur00, Bur98]. floating [Sne76]. flood [Ako21, SC15]. Florence [Chu13]. Florike [Kus18]. Florinskii [Byf19]. Floris [Gau14]. flow [Fra45]. flows [SS23]. fluidorum [Lev21]. Fluxio [Gre81]. Fluxio-differential [Gre81]. Flying [NM37b]. foetal [Fra41, Fra45]. Fog [Mil08, BT55]. fond [GG97]. Fontaine [Gre81, Gre82]. Fontana [Ear60]. Food [Hul02]. Foot [Bra09, Ari72]. Forbes [Gow82, Kaa17, Kid81, Mar06, Mea85]. Forbidden [Mar21a]. Force [War02, Bur65, Can19, Erl92, Goo80, Kat23, Wil54]. Forced [Bik71]. Forces [BS12, Ait64b, Gui95, Hal57, Hen23]. forecasting [Dav84]. Foreign [Plo11]. Forestry [Wea09]. forests [Dau13]. Forfait [Bra94]. Forgotten [Whi12, Cam23, Mel21]. Fork [Ku09]. form [PB45, PB47, PB48]. Formal [Fer02, Byr81]. Formation [Mil04b, Por11, Bre91, Gra70, Kaa17, Mal09, vH92]. former [Mad57]. Forming [HW90, Hil91]. Forms [Ku09, McO00]. formula [Erl92]. Formulation [Swi65]. Forster [Fra13]. Forth [Jon82]. Fortune [Fra10, Wes10]. fortunes [Bra94, Lew76]. fossils [Mal82, New52]. Foucault [Old99b]. Foundation [Jon13]. Foundations [Bec01, Kim99, Fle74]. founding [Dea91b, Par84, Plo80, Str94]. Four [Bra09, Gau14, Coh10a, Lar91]. Fourcroy [Sme55a, Sme55b, Sme67]. Fourier [Ano75k]. Fourth [HBP08, Col52c, Gay13a]. Fowler [Ung02]. Fownes [Row50b]. Fox [Rus04]. Fragment [Ray15, FF68]. Fragmentary [Tha66]. Frameworks [Ela11, Lee23, Pis14b]. Framing [Hig17b, Mau90]. francais [LS15a]. francaise [BV83]. France [Lyn14, Lyn15, Arm53, BS12, Boa12a, Bre91, Bre16, Car22, Dau13, Dav98, GG84b, Hyd19, Kle12a, Lac11, Mar09, McA12, Muc98]. Francesca [Gre87]. Francesco [Dur10, Hen79, HZ07, Tod82]. Francis [Col52b, Wal04, WN16, Coc56, Jal23, Mad61b, Mar21b, Ree80, Ree81, Son74, Wat92]. Francisco [Dea93]. Franck [Hen09, Hen09]. Franco [Bra81]. Francois [Bro71, Duv54, McK57, Tou89, BS12, Sme55c, Tre80b]. Frankland [Rus78]. Franklin [Pro18, Dav90a, Dea89a, DK55a, DK55b, DK57, Lev86, Lop60, Rob55a]. Frankreich [Dau13]. Franz [Dal14, Plo80, Pop12, Thi73]. fraud [Gre77]. Frederic [Mer98]. Frederick [Cal68, Dav92, Tat97, Yap83]. Frederico [CS12]. Frederik [Sch13b, Koo10]. Fredrik [Pri55]. Free [vS15, Yal22]. Freeman [Ben13b]. freezing [Wil54]. Freiberg [Dym11]. Fremy [Fau05]. French [Sta19, Ape20, Arc05, Aug04, Bau88, Ben09, BV99, Boa12a, Bra79, Bra92, Bre91, Bre93, Bur12c, Cam09, Car22, CP97, Cro09a, DS97, Dur10, Fau05, Gál88, Gen07, Gra01, Hir07, Kah01, Lan83, Mit12, Qué90, Rom06, Ser12, Sme87, Sta19, Sto13, Thi12, Tom01, Tom10]. frequency [Kat16]. Fresh [Orr14]. Fresh-Water [Orr14]. Friedrich [Cun04, For71b, de 53a, Cha21, Col51, Pop12]. friends [Jac52a, Jac52b]. Frobisher [Don50]. Frodsham [Lev94]. Front [Ano10c, Ano10d, Ano11a, Ano11b, Ano11c]. Frontier [Dym11, Kra11a]. frontiers [Aga19]. Frost [Sok96]. frosty [Kaa17]. FRS [Kur99, Kur03, Wal00]. Fruitful [Maf13]. fuel [Str93, Str97]. full [NG09]. Fulton [Ano59a]. Fulvio [Del11b]. function [You37]. fundamental [NG09, Nau14]. funding [Lin22]. furnaces [Gib48]. Further [Hal55, de 58, Don51, Gib53d]. fusible [Wil59]. Fusinieri [FGR74]. fusion [Hen87]. Future [Gue09, Luc19]. Futures [Goe18]. Fuzzy [Mar13]. Fyfe [Jam12].

G [Bar11, Luc19, Mat11, Mil12, PW08, Pul94, Sla03, Sza72, Wil91b]. G. [Win49]. Gabriel [May21]. Gain [Bsc12, Cha09]. Galan [Dun13]. Galen [Nut21, Fra41, Mcv21]. Galilean [Vic04]. Galilée [Duc13]. Galilei [Lat14]. Galileo [Gat18, Hei14, Lat14, Pis13, Sha19a, Ait54, Ano85j, Ari74, Bau88, BBHN14, Bur65, Dra81, Dra82, Dra86, Erl94, Fin02, Gro74c, Hil84, Lat15, Lew76, Mac76, Mar11a, Nay76, Nay77, Nay80, Nay82, Nay83, Pal12, Qua68, Qua70, Qua72, Qua74, Rap12, Rap15a, Rap17, Ree08, Rob74, RS89, Rus95, Sch08a, Sch08b, SB09, Val09, Wil11a, Wil57, McC39a]. Gallic [Woo71c]. Galton [Byf19, Ano04f, Swi65, Wal04, Wil55a]. Galvani [FC36, Hof36, Kip87, Wal37]. Galvanian [Hof36]. Galvanic [Win49]. Gap [Rei14, Ste11a]. Garcia [Win16]. Garden [Law15, Sve19, Rob54a]. gardener [Stu82, Stu89]. garnets [Qué90]. Garzoni [Uga06]. Gas [Stu09, Bru79, Far69, Kox90, Spu87]. gas-expansivity [Spu87]. Gascoigne [Cha14]. Gases [Bru57a, Fle74, Bru57b, Bru58a, Bru58b, Cah90, Cro61]. Gaslight [Joh15a, Tom09, Tom12]. Gassandi [Ben09]. Gassendi [Ben09]. gathering [Tei75]. Gaukroger [Wil23]. Gauss [Ben74, Cun04, For71b, Hen07]. gave [Bik65]. Gay [Cro61, Spu87]. Gay-Lussac [Cro61, Spu87]. gaze [MSD23]. gearing [FW85]. Gears [FW85]. Gebetszeiten [Bur09]. Geburt [Nit96]. Gegensatz [Ecc80]. Geiger [Tre86]. Geikie [Old97, Ham82, Old80, OH97]. Gemma [Cha12a, Hir08]. gemmologist [Swe64]. Gender [Coo09, Goo08, Low96, Sta13]. Gendered [Bad11]. genealogy [Sch18a]. Genentech [Hug11, Yi14]. General [Bre91, DS97, Lat15, Old94, Ste10b, Bre91, Mil11a, Sch13a]. générale [Bla88, DS97]. generally [BS65]. generated [Kan70, Kan71]. Generation [Ano08e, Boe14, Bru11, Del11a, Bla19, Dur10, Ekh10, Foo69, Gál88, Nye11, Dur10]. generatione [Bur14b, SBT10]. Generations [Smi13]. Genesis [Fli08, Nic11a, Geo38a, Mar76]. Genetic [Cha13, Ras17, Wex08]. Geneticists [Har85, Rus22]. Genetics [Kre11, Mag98, Nor75]. Geneva [Gui15]. Génie [Ano05e]. genius [Wil91a]. Genome [Com21]. Genomic [Ash17]. Gentleman [dBP86]. Gentleman-scientist [dBP86]. gentlemen [Tur73]. Geo [Gam21]. Geo-heliocentric [Gam21]. geochemical [Lei76]. Geodesica [Gre87]. Geodesy [War02, Ohn23]. Geodetic [War02]. geographers [Bre91]. géographes [Bre91]. geographic [Kot78]. geographical [Tho54]. Geography [Bra02, Dri12, Mos14a, Swe03, Wit96, Wit10, Sha12, BJ00, Kni13]. Geohistorical [Ano06l]. Geohistory [Old10]. Geological [Alv07, Bra09, Car05, Jac99, Mor88, OJY03, DE50, Edm75, Eyl37, EE38, EE51, Eyl67, Jon84, OH97, Osp80, Sch80, Sta88a, Tho54, Urb80]. Geologist [Cha63a, Cha63b, Cha64a, Cha64b, Cha61, Cha62, Edw72a, Old80, OJY96]. Geology [Ano03f, Ano06l, Dea75, GWN^{+}11, Jac99, Old99b, OJY03, Buc13, Cha53b, Cha54a, Cha54b, Cha70, Dea79, Old80, Osp80, PP77, Por78, Sha72, Sle69, Tho47, Smi15]. Geomagnetic [How02, How03]. geometriam [Bur14c, Bus10]. Geometrical [Dij04, Nau14, Jon74, Ver16]. geometrico [Sch80]. Geometrizacion [Gre87]. Geometry [Ben91, Bin12, HZ07, And07, Ver16]. Geophysical [And14]. Geophysics [Gro74b]. Geopolitics [Cha12b, And09b]. Georg [Arm51, For72a, Gra71, Lam97]. George [Ben80, Bow74, Bry72, Cot76, JTR94, JTR95, MT72, Mil91b, Mor68, Por78, Row50b, Tir23]. Georgetown [Gor16]. Georgian [Jon19]. Georgii [Swe21]. Gerard [TD93, Tur94, Van01]. germ [Cre66, Wil55c]. German [Fra13, Ver13, Ber97, Bur09, BvK73, Cha21, Dau13, DS97, Dym11, Ecc76, Ecc80, Gab78, Gor87, GG11b, Hen89, Hen86, Hen88, Hen92, Hen00, Hen09, Hon84, Hor00, Hut73, Ill89, Joh15b, Kan70, Kan71, Kle21, Kno89, Kre38, Mas53b, Mat11, Mit78, Nit96, Ped75, Pul94, Ros98, Sau76, Sch21, Sch11, Sie09, Slo03, Ste76, Str97, Thi71, Thi73, Tre74, VW10, Wak09, Wie23, Win66, Zac72]. Germany [Bec13a, Eff21, Har85, Hen89, Kle12a, LS23, Mei79, Nyh09, PdC10, Pre08, Sak12, Lau21]. Gerrit [Sne84]. Geschichte [Sie09]. Gessner [Vic12, Kus10, LGKW08]. Gewicht} [Ste76]. ghetto [GG93]. ghost [Hen18]. Ghosts [Pia21]. Giesecke [Swe74]. Gieseler [Sau76, Sau76]. Giessen [HB74]. Gift [Pan09b]. Gilbert [Fra47, Uga06]. Gilded [Cah10, Dav90b]. Gilded-Age [Cah10]. Giordano [Sha19a, Arm48, Mar16b, Mar19, Mas73]. Giovanni [BB95, Ela13]. Glaciated [Old99a]. Glaciation [Old99a]. glaciers [de 53a]. Gladstone [Bel99]. Glasgow [Wil37]. Glass [End16, Jac15a, Tur00, Adl97, Tur82]. Glassblowing [Jac15a]. Glassmaker [Ber12]. Glassworks [Wil11a, Ree08]. Global [Cha14, Gue13, Lig16, Lus14, Vet10, Ako21, Die21, Edw10, Gas10, SS23, Hal18]. globe [Dek96, dVH50]. glory [Wal84]. Glow [Ren18]. glycogenic [Tat97, You37]. glycosidic [Hen78]. Gmelin [Bro78a]. go [Mac93]. goals [Low75]. God [Goo11, Tac96]. Goethe [Duc88, Duc97]. Goiffon [Wil55c]. Going [Nor01]. gold [Don50]. Golden [Ano09i]. Good [Edw04, Lov11, Pro18, Wes10]. Goodacre [Ink78]. Gorai [AS81]. Gordin [Sta21]. Gordon [Aga17b]. Gorilla [Fan09, Goo11]. Gospel [KN13]. Got [Del11a]. Gothic [Cro03]. Göttingen [Son75]. Gottlob [Osp80]. Gould [Ols71]. Goulding [War14]. Governance [Zai17]. Government [Car05, Tur00, Hul02]. governmental [SR73b]. governor [Gro72b]. Gower [HB74]. Graaff [Min21]. Grace [Gra72a]. Gradmann [Fra13]. graduations [Cha83b]. Gradus [Big22]. Graham [Spr77, Tir23]. Grain [Sim92]. Granada [OEMB21]. Grand [Fol14b, Kra11b, Law09b]. Grant [Pav17]. Grassmann [Lew77, Pet14]. Gratings [Bro15, Bro18, Bro22]. Grattan [MIL15]. Grattan-Guinness [MIL15]. grave [LPST11, McC14]. Graves [Dor76]. Gravesande [vH94]. gravitation [Ait69]. Gravitational [Ano06g, Hen92, Hen96]. gravitatione [Lev21]. Gravitations [Hen92]. Gravitations-Rotverschiebung [Hen92]. gravity [Ait59, Cha19, Wal84, Kol99]. Gray [Mea85, Fra10]. Great [Kni14, Lus18, Orr14, Rei14, SS99, Wee16, Ink72a, Ock37, Ste11a, Cal68, Dub63, DK54b, Fos37, Her18, Hol07, Mat76, Rya91, Smi23b, SC15, Wie23]. Greatest [Ras17, Har12]. Grebe [Hen92]. Grebe/Bachem [Hen92]. Grebe/Bachems [Hen92]. Greco [Wie19]. Greco-Roman [Wie19]. Greece [PK87]. Greek [Hen14, Dar12, Høy23, PK87, Ste11b, Wal12]. Greeks [BB13, WH11]. Green [Gar04]. Greening [Kri19]. Greenland [MN13]. Greenwich [Kin04, Mac18, Kin04]. Gregor [Tre72]. Gregory [Ver16, VvM92, de 57]. grenats [Qué90]. Grenze [Kan70, Kan71]. Grocer [Bro79a]. Groningen [Mac72b]. ground [Tay08]. Group [Ano57b, Bar14, Bor10, Cha83b, Rus80]. groups [Lam97]. Grove [CH82]. Grover [Reg09]. Growth [Jon13, Fra41, Gro72a, Har74, KL09]. Grund [Ben14]. Guanientá [OEMB21]. Guericke [Coh51, dVH50]. Guerre [Sto13, Bre91]. Guide [Pou01, Swe03, Joh09, Kid13]. Guild [Bro79a]. Guilds [Gay00, Cra87]. Guillaume [Not21, Pap09]. Guillelmi [Bur11a]. Guinness [MIL15]. Gundissalinus [Bur90]. gunpowder [Ful64, Lop60]. Guthrie [Swe64]. Guyton [Cha72b, Sme54a]. Gynaecology [Bar09].

H [Akt09, Boa19, Bsc14b, Bur14c, Gau14, Gra72b, GG97, Jar98, Law15, Luc19, Mac82, McC13, RS01, Sco81, Str94, Van19, Yi13, vH92, vdS06]. H. [Bar11, Bod56, Kox90, Lew77, Mat11, McC39b, Mil91b, PW08]. H4 [HBP08]. Haak [Bar57]. Håkon [Nor12]. Haddingtonshire [Eyl63]. Hadley [Col52c]. Haeckel [Mat11]. hafnium [Sce94]. Hahn [Lat15]. Haldane [Stu09]. half [Cha21]. halftones [Smi18, Wei19]. Hall [Eyl63, Cha68]. Halle [Whi13]. Haller [Son75, de 53b]. Halley [Kol99, Ste36a, Wal84]. Hamburg [Kra13].

Hamilton [WN16, Lai79, Sle69]. Hampshire [Tho55]. Hanbury [Gib51d]. Hanckwitz [Mad55]. Hand [Dup11, Nic06, Ham01]. Hand-Colouring [Nic06]. handlist [Wel83]. Hands [MLSB17, Jar18a]. Hannah [Mar21a]. Hanoverian [Hen07]. Hans [Cha72a]. Hansen [Bec84]. Haraway [San17]. Hardback [Byf19, Lau21, Mar21a, Nut21, SS23, Sta21, Gue23, Hei19, Kin19b, Lin22, Sla19, Wil23]. hardcover [Gol19]. Hardly [Tee84]. Harman [Bow13]. harmonic [Mac72a]. harmonicae [Kre23]. harmony [Reg19]. Harold [For71a]. Harriot [Pis13, CT69, Pep76, Sch08a, Sch08b, Sok74, Sok94, Tan77, Tan80b]. Harris [McC94]. Harrison [Bsc14b, HBP08]. Harry [Gre77, SP11]. Hartgill [Mor68]. Hartley [OO51]. Hartlib [O’B65a, O’B65b]. Hartmann [Lam97]. Harumi [Tan14]. Harvard [Coo13, Hay18, Mel21, Rus04, Yi13]. Harvest [Gou22]. Harvey [Bay38b, Bay38a, Bay39a, Bay39b, Ekh10, Foo69, Fra41, Sal92, Whi86a, Wil62]. Hasledine [Kur03]. Hassenfratz [Duv52]. Hastings [Par84]. Haüy [Gen22]. Haven [Bow13]. Hay [SR73a]. Hayflick [Par16]. hbk [Coz19, Reg19, Sha19a, Swe03]. healer [Gel21]. Healers [Mel21, Mel21]. Healing [May21]. Health [Ano08g, CC09, Coz15, Hal14b, BJ13, Bro71, Coz19, Flo96, Mel21, Sin50a, ZB10]. healthcare [Mel21]. Hearing [Ano05f, Kvi22]. Heart [GG12, Mil13, Tay08, Row09]. Heartland [KK02]. Heat [Boa08, Bos03, Hil99, Sim04, Ben37, Ber97, CK89, Cor36, Cor38a, Cor38b, Ols69, Ols70, PM38b, Sco81]. heating [Ege68]. Heavens [Ste09b, RO84, Rib20]. Heaviside [Gos76, Gos77, Yav93]. Hebrew [Jus11, Jus12]. Hebrides [OH97]. Heidelberg [Wit13, Ken52]. Heike [Pap09, vD08]. Heinrich [Kan70, Kan71, de 50b, Eck10, Kan70, Kan71, Kra13]. Heisenberg [Del13, Hen00, Ros98]. Helen [Akt09, Fra14a]. Helena [TH74]. Helge [Fol14b]. heliocentric [Gam21, Ped75]. heliostats [Mil86]. heliozentrische [Ped75]. Helium [Kra09b]. Hell [Mad64]. Helmholtz [Cah10, Can19, Hui11, Jur10, Kat23, Mil42, Ole23]. Helmont [Bro79b, Par36, Pat36b]. Henderson [Woo71a]. Henry [Ver56, Ano58a, Ber96, Ber19, Bro22, Car92, Coh36, Cop80, Far18, Gra72a, Har57, How02, Mol84, Mor01, Mor09, Not18, Plo78, Plo80, Tom14, Win72]. Hentschel [Hei19]. Herapath [Bru57a]. Herausgeber [Ecc76]. Herbert [CP02, Gay13a, Jac09, Sha76]. herculis} [Tre91]. Heredity [Com21, Maz12, Wei85a, Mag98, End17]. Herefordshire [Stu82, Stu89]. heresy [Mar16b]. heritage [Dia96, Din50, GG91a]. Herman [Duc17, Lov74]. Hermann [DS97, Hui11, Pet14]. Hermes [Duc09]. Hermeticism [Duc09]. Heroes [Ale10, Coo13]. Herschel [Cas15, Cor38a, Cor38b, Hen82, Hil99, Kid81, Mac19, Ogi75, Ste10b, Woo71a]. Hertz [Eck10, Kra13]. Hester [Kin04]. heterogeneous [FGR74]. hexagrams [AS81]. Hidden [De 05, Fox96]. High [Bry10, HL91]. Higher [KK02, Kra11b, SJ78, Fol14b]. highly [Smi77]. Higton [Kin04]. Hill [Ano86g, Beg21, McC39c]. him [Sme78]. Hindu [NS52]. Hintergründe [Ecc80]. Hippocrates [Sch36, Yeo18]. Hirai [Cha12a]. Hiro [Cha12a]. Hispano [Gre87]. Histoire [Duc13, Ben09, Swe21, Gue10]. Histophysiology [Ama06]. Historia [Kus10, de 53b, LS15b]. historian [Mar76, Old80, Sco36]. Historians [Jac99, Sno01, JL79, Lar91]. Historic [Bur12b, San00]. Historical [Ano06l, Ano07j, Cro09b, Dav57, Fol14a, Gib51a, Gib53a, Gib53b, Gib53c, Gue13, Kra00, KN13, Mil08, PM37, PM38a, PM38b, PM39, PB45, PB47, PB48, Vet10, Eff21, GG86, GG84a, GG11b, IR11, Ink91, Lev21, Thi71, Wil59]. Historicity [Ano05f]. ’histories [Ham01, Alt07, Bry10, DL11, Mar21b, Tau10, Kni77, LS23, Bou13]. historiographical [Cop87]. historiographies [Hen21b]. Historiography [Ano08h, Mil12, McE10b, Old80]. History [Ank15, Ano03f, Ano07g, Ano08f, Ano11d, Bar09, Bar14, Bew16, Bin12, Boa12b, Bra10, Bri13, Bur10c, Cap12, Cha16, Chr10, Dom09, Don36a, Don36b, Duc09, Ego19, Fra14b, Gie16, Gil02, GG77, GG93, Gue09, Hen14, Hig17a, Hor91, Jac15b, Jac12, Joh10, Joh09, Kau79, Kid14, Kim99, Kin04, Kni17, Kra09a, Kra09b, Kur00, La 11, Lal12, Lea18, LR00, Lev14b, Low03, McC21, Mil12, Mil42, Mil01, Mil04a, Mil04b, Mil13, Nap09, Nic11a, Old99b, Pal13, Pal83, Pan09b, Pis14a, Rei14, Sch18b, Sht17, SR71b, Tan11, Tau14, Tre15, Wit13, Woo36, And07, Ano91o, Bak64, Bak21, Bea76, Bec84, Ben09, Bik71, Bod56, Bor10, Bro69a, Bro69b, Bro16, Can09, Car97a, Cha52, Cha71, Chi70, CC47, Col39, Cot77, Cro48, Dar12, Deb62, Die16a]. history [D^{+}11, Dob45, Don79, Doo10, Far79, For61, Fra36, Fra38, Fra39a, GG88, Gay13b, Gib39, Gou10, Hak92, Hal20, Ham12, Hei19, Hen88, HS22, Høy23, Jac52a, Jac52b, Jon37, Kni75, Kra95, Law21, Lev09a, Luc19, Mac09, Mac23, McB91, McE10b, Mei79, Mil74, Mor10, Mul92, Mul76, New38, Old72a, Par42, Par51a, Pot22, Pul94, Ren18, RH11, Rob95, Rou22, Rut39, SM82, Sco81, SBB^{+}75, SL77, SP92, Sme54b, Smi77, SJ78, Ste11a, Sti37, Swa89, Tho51b, Tur82, Und67, Urb80, Uss78, War12, WN16, Wer10, Why54, Woo71b, Woo71c, de 56, Bik75, Bro36a, Kid13, Lan13, Law13, Lov13, Bil18, War14, Van19, GG13a, Pis14a, Ree14b]. Hobbes [Hor00]. Hochschulen [Hen89]. Hodson [Hor87]. Hoff [RS01, vdS06]. Hofmann [Hor87, Ano88i, Hor87, Joh15b]. Hoggart [Por78]. Holbein [Hud03]. Holcot [Tac96]. Hold [Ano05j]. Hole [Del11b, MN13, Mel09]. Holland [Bec13a, AC99, PdC10]. Homberg [Gue21]. Home [Boa19, Hil91]. homogeneity [De 17]. Hon. [Har84, Ran63]. Honestly [Bec09]. Honolulu [Hig17b]. Honoré [Ela11, Ela13, Pis14b]. Honour [Jac15a, Joh15b, Mau15, Ram15, RN15, UB15, Wei15]. honourable [Mad59]. Hooke [Bar95, Ehr95, GE19, Gou80, KO83, Mul92, Nak84, Sch18a, Wil56d]. Hooker [DD21]. Hopkins [Bil18, Fra13, Gat18]. Horace [Sme78]. Horary [Kin04]. horizons [Wil11b, Str14]. horizontal [Mil93]. Hornsby [Wal00]. Horology [Ish14, Ben80]. Horrifying [Hen00]. Horse [SP92]. Horst [Lat15]. horticultural [Olb89]. Horticulture [RS17]. hospital [Mau90, Tho51b]. hospitals [Gra70]. Hottinger [de 50b]. Hour [Tur03, Tur06, Mil93]. hours [Arc90, Mil96]. House [McC83, Mor01, She09]. household [Wei19]. Households [Ser12]. Housewives [Ree14a]. Howard [Kur99]. Hsia [Chu13]. Huggins [Sus15, Bec11, HB14, Sus16]. Hugo [Bow78, Dar85, Kot79, Mei85, The94, Zev00]. Human [Bel99, Ben13b, Bow74, Bri13, Com21, Goo10, GG12, Li12, SS13, Wei85a, Fle97, FLW11, Kin19b, Mac09, Pav17, Rus21, VT17]. Human-Animal [Ben13b, FLW11]. humanism [Ree76]. Humanist [Lev14b]. Humanitarian [Shu68]. Humans [Car18, DS16]. Humboldt [Dau13, Dau14a, Wer11, Bik65, Bro93, Dau13, Hon84, Kle12b, McC10, Nic11b, TS14, Dau14b]. Humboldts [Hon84]. Humfrey [Ben14, Gru11]. Humphrey [Jar18b]. Humphry [Rus59a, Rus59b, Rus63c, Spi68, Tat17, Ful64]. Hundred [Cha17, Sch18b, Fos37]. Hunnings [Tho67]. Hunt [Bow78, SS13]. Hunter [Sch09, Spr42]. hunting [Yap83]. Huntington [Cha13, Wex08]. Hurry [Gos15, Har12]. Hutton [All94, Cha54b, Dea73, Dea75, Dea81, EE51, Jon82, Jon83, Jon84, Jon85, JTR94, JTR95]. Huttonian [Chi70, Dav66]. Huxley [Bar75, Cal80, Her18, Jar98, Law15, SS13]. Huygens [Ari79, Bar95, Fou89, MJ89, Plo99, VvM92, Whi12, Zig80, vHvG99]. Hybrid [Kon12]. Hybridism [MWO07]. Hybrids [McO00]. Hyde [TS14]. hydraulic [App96]. hydrocarbon [Par45]. hydrodynamics [Yam98]. Hydrogen [Kim06, Lan83]. Hydrology [SC15]. Hydrostatics [Cha17, Sch18b]. hygroscope [Tay49]. Hypatia [War14, Gou10]. Hyperbola [Par03]. hypotheses [Bro69a, Bro69b]. hypothesi [Arm50]. Hypothesis [Ham03, Roc85, Ait69, Bec84, Cal80, McD72]. hypothetical [HY87].

I. [For98]. iatrochemistry [Mas53b]. Iberia [Tir20]. Iberian [Rib20]. Ibn [Bur09, BvK73, Bur09, Bur12a, Kah10]. IBSN [SS23]. Ice [MN13, Ham82, LS74, Pat36b, de 50b, Hes17]. ice-mountains [de 50b]. iconography [Cha84]. idea [Gre84, Pav17, Spe16, Wil56a, Wil56b, Wil56c]. Ideas [Le 74, Old99a, Dym37, Fra45, Hom79, Lei76, S´ 20, Sar69, Yam98, YO89]. Identification [Jac15a]. identities [Hig17b, Mac09, Bri13]. Identity [Jac15a, Li12, Par03, Plo11, Wit96]. idols [Ora75]. Ierodiakonou [Fol14a]. Igne [MdVH58]. igneous [OH97]. Ignited [Del09a, Gre05]. ihren [Hor87]. ihrer [Ecc80]. II [AC99, Ait58a, Bay38b, Bra82b, Bro69b, Bru57b, Bur60, Cha54a, Cha62, Col37b, Cor38b, Don36b, Dun62b, DK55b, Feu71b, GG86, Gib51c, Gib53a, Gou23, Hop58, How03, Hug52, JTR95, Klo85b, Low90, Mas53b, McK59, Nea51a, O’B65b, Old92, PM38a, Par51a, Par53a, PB47, Qua72, Rob57, Rus51a, Rus59b, Rus63b, Sin50a, Ste69b, Sto68, Str93, Stu89, Tre64, Wes56b, Wig56, Wil56b, Wil60b, Wil61, Wil95b, Woo71c, Yap83]. III [Ait58b, Bay38a, Bru58a, Cha54b, Cha63a, DK57, Gib51d, Gib53b, Hop59b, Hug53, McK60, Nea51b, Old93, PM38b, Rus51b, Rus63c, Sin50b, Tre65b, Tre65a, Wil56c, Wil61]. IIIa [Wil60c]. Ilay [Eme02, Eme13, Woo15]. ill [Sta19]. illumination [AM76]. illuminators [Mil86]. Illus [Law13, Gol19, Ver13]. Illustrated [Kin04, Ree14b, D^{+}11]. Illustration [Gas07, dC09a]. Illustrations [Ano08l, Nic06, Gue10, Sha19a, Yap83]. illustrative [Old74]. im [Bur09, Ecc80, Hen89, Hen09, Hon84, Sie09, Ste76, Thi71]. Image [Ano10a, Roc10, Ked16, Joh14]. Images [Egm16, Goo11, Kus18, Eas86, Top80]. imaginary [Ait65, Mac76]. imagination [Bla10, Roc10, Wil92, Joh14]. Imagining [Ako21]. immediate [Wil60b]. immigrant [Cot08]. Immunity [Tau10]. Immunization [MS22]. Immunology [Tau10]. immutability [Duc97]. impact [CZ19, Lau66, Thi71]. impacts [Pep76]. Imperial [Chu13, Hsi09, Mcv21, Wea09, Ste76]. imperialism [Olb91]. Impetus [Ela11, Pis14b]. implication [Mac85a]. implications [Bow84, FY09, Mil13]. Implicit [Gab78]. Implizite [Gab78]. Imponderable [Yos99]. importance [Ber97, Osp80]. imprimant [LS15a]. improved [BS93]. improvement [Gol19, Rob56, Rob57, SR74b, Stu89, Hop59b]. impulse [Erl92]. Impure [Lin10]. Ina [Hal14b]. Incandescent [Fox96]. Incas [Gen22]. Inchofer [Mar19]. Incidentally [Pat36a]. incomparable [Pal23]. incongruity} [GE19]. incorrect [Tre80a]. Independent [dA21, Joh18]. independently [Cam80]. Index [Ano09b, Ano09e, Ano09f, Col37b, Lev09b, Ano09a]. India [Arn09, And10, Bas95, HR89, Ink88, Lar61, WN16, Win16, Ban14]. Indian [Arn09, GW11, Plo11]. Indiana [Qui13]. indices [Rib20, Mal20b]. Indigenous [Nap09]. Induction [Cav06, dVH65, Jor82, Mil81b]. Industrial [Aug04, KK02, Wal09, Per86, SR74b, Uss80]. Industrialization [Dup11, Ink88]. industrielle [Aug04]. Industry [Bri00, Cha12b, Gar04, Joh15a, Pag03, Tur00, And09b, Cra85, Don36a, Don36b, Far69, Gib50, Gib51d, Gib51e, Gib52b, Gib55, GG93, GL11, Jon19, Kle21, Kra95, Mat76, New38, Pas51, Rob54c, Sch75, Spi68, Tob12, Tom12, Ver94, Far72, Gib51c, GG14]. Inedita [For72a]. Inference [Alt07]. Infinite [Ano06j, Ben14, Bun77, Gru11]. Infinitely [SS99]. Infinitesimal [Bec01]. Infinity [Ben13a, Qua70, Qua72]. inflammable [Osm58, Tom14]. inflammable-air [Osm58]. Influence [Cro16, Gas07, Ben09, Bra75, Bro87, Dun21, Hen78, Hen79, Lai77, Lep06, O’B65a, O’B65b, Sle69, Sne81, Tho47, Row18]. Influences [Lai79, Kem68]. Information [Ano05j, Yea11, Aga17b, Kli15, Mac20b, Whi82, ZA16]. infrared [Sch89, Kan70, Kan71]. Infrastructure [Aga17a]. Ingenieur [Sau76]. ingénieurs [Bre91]. ingénieurs-géographes [Bre91]. Ingenious [Boa08]. ingredients [Ger23]. inhabitation [Dun21]. inhabited [Man93]. Inhalt [Pul94]. inheritance [Mag98, Nor75, Rus22]. initiatives [Kri19]. Ink [Arn09]. Innes [Olb89, Ste61]. Innovation [Fit17, Pop12, Por11, Dav98]. innovations [Per86]. innovators [Mel21]. Inquiry [Dup11, Hah15, Leh12b]. Inquisition [Lei20, Ave20, Mal20a, Sha19a]. insect [Sch18a]. Insectes [Sto13]. insects [Mel10]. Inséré [Ray15]. insight [Pal21, Sha75b]. insights [Cha75]. Inspectorate [Gar04]. Instability [vS15]. Instant [Low90]. instantaneous [Erl92]. instauration [Wat92]. Institute [Gay00, Hol17b, Mil13, Mor86, Ste76, Wei85a]. Institutes [SR72f, Ink75]. Institution [Alv07, Spr77, Ock37, Olb89, CH82, Edw72b, Kur00, Kur01, Mol84, Mor01, Rus63c, Tho57]. institutional [Ink77]. institutionalization [Bro95]. Institutions [Mar09, Nor12, Bra79, GL11]. Instruction [Ram15, Bra76, GG84b, Sme54b]. Instructions [Swe72a, Tir23]. Instrument [Car17, Cra87, Gau12a, Van01, And19, Bro79a, Bry92, Cha84, Cra85, Dav78, Hac79a, Lov74, Mil88a, Not21, Sch89, Sch90]. instrument-making [Bro79a, Cha84, Mil88a, Sch90]. Instrumental [Bla19, Lai04, Mor02]. Instrumentalizing [Wil07]. instrumentation [Bai93]. Instruments [Bur12b, Cav06, Dal14, Hui11, Jar18a, Jar18b, Kin04, Kre13, Ku09, MLSB17, Pan05, Pan09b, RC08, Wal09, Wil11a, Bed93, Bed97, Ben09, Cha83a, Dek93, Dun23, Ges23, Glu23a, Jar23, Joh91, Lev86, McC86, McC89, Pri52, Pri54, Rob56, Rob57, Rya91, Tur67, Tur73, Tur91, Wil91a, ZA16, dC88]. Insulated [Mil04a]. insurances [Sta88b]. Integrated [Gue09]. integration [CZ19, Gre82]. intellectual [Law21, How19]. intellectuels [Muc98]. Intelligent [Fan09]. intensity [Big22]. interaction [Mil81b]. Interests [Eme02, Aga19, Ell83]. Interim [SBB^{+}75]. interlude [Lin63]. International [Ban14, Bon14, Hei19, Nof14, And10, FB22, Caw77, Fra38, Tei75, Cah10, Pap09]. interparticulate [Hen23]. Interpretation [Mil12, Pal83, Sch08a, Hir07, Ait64a, Can75a, Eff21, Hir07, McE10b, Mil81a, Pis13, Tod93, Tur55]. Interpretations [Pan09b, Por01, Rus59b]. Interpreter [Dau14b, McC10]. Interpreting [Nyh16]. Intertextuality [Ano08l]. Interwar [Bow16, Har85, LPST11, Ste00b]. Introduction [Die16a, RG23, Lev09b, Mal20a, RN15, Tur91, Zui04, Bro16, Dub63, DK54b, Fra39a, HR89, Hen23, HA19, Li94, OH85, SBT10, YO89, Hen17, Kni17, Köb17]. Introductions [Lev99]. introductory [Bea76, Wil23]. invade [Bra94]. Invented [Plo99]. Inventing [Gau12a, Nap09, dV21]. Invention [Dup11, GH08, How02, How03, Jac09, Smi15, Whi12, Wil07, Buc13, Gro72b, Kat16, Tay49]. inventor [dVH65]. inventory [DG09, VW10]. inverse [Ait64b, Ait69, Gui95, Pou91]. inverse-square [Pou91]. inversion [Eas86]. invertebrate [New52]. Investigaciones [Mal09]. Investigating [Lac11, Sta17, Lyn14]. Investigation [Hil99, Ear60, Jae10]. investigations [BT55, DK55b, de 53a]. Invisible [Buc08, Edd13, Rat09]. Involvement [Car05, Woo71a]. Ioannes [Ait70]. Ionian [Ste09a]. Ionosphere [And09a, Cha12b, And09b]. Ireland [Ano07g]. Irene [Lat15]. Iris [Tob12, GG14]. Irish [Ade10, Swe67]. iron [Sch75]. irregular [For70a]. Irving [Mil11a]. Isaac [Luc19, McD72, Col54b, Cop80, Fei19, Gou23, Hal55, Lev21, Plo99]. Isabelle [Bsc14a]. ISBN [Akt09, Bar11, Bil18, Bil16, Bin12, Boa12b, Boa19, Bow13, Bri13, Bsc12, Bur11a, Bur11b, Byf19, Car18, Cha12a, Cha12b, Coo13, Coz19, Del11b, Del11a, Del12, Dom09, Edd13, Egm18, Fra13, Gat18, Gie16, Gol19, Gue23, Hal18, Hay18, Hei19, Hen00, Hig17a, HA18, How19, Jam12, Jar18a, Jon13, Kid13, Kin19b, Kra13, Kri19, Kus18, Lai04, Lan13, Lat15, Lau21, Law13, Lin22, Lov13, Luc19, Mar21a, May21, Maz10, Mcv21, Mel21, Mil12, Mil13, MS22, Nor12, Nut21, Pal12, Par13, Por16, Pro18, Qui12, Qui13, Ray18, Reg19, Ren18, Rob10, Row18, Sak12, Sch18b, SS23, Sei13, Sha19a, Sha19b, Sla19, Smi18, Sta19, Sta21, Sus16, Swe03, Swe21, Tho11, Van19, Ver13, Vic04, Vic12, Wei19, Wil11a, Wil23, Wit13, Wol13a, Yi13]. ISBN-10 [Bin12, Swe21, Wil11a]. ISBN-13 [Bar11, Boa12b, Bow13, Bri13, Bsc12, Bur11a, Cha12a, Cha12b, Coo13, Del11b, Del11a, Dom09, Edd13, Fra13, Jam12, Jon13, Kid13, Kra13, Lan13, Law13, Lov13, Maz10, Mil12, Mil13, Nor12, Pal12, Par13, Qui12, Qui13, Rob10, Sak12, Sei13, Swe21, Tho11, Ver13, Vic12, Wil11a, Wit13, Wol13a, Yi13]. Islam [Det18, Yal22]. Islamic [Att09, Bur09]. islamischen [Bur09]. Island [NM37b]. Isles [Bar09, Dav68, DE50]. Isoclinic [How03]. isomorphism [PB48]. ISSN [Cha12a, Sch18b]. Issue [Die16a, RG23, Gál88, Mau69]. issues [EE38, Old94]. Italian [Big22, Bri00, Dal14, Fol11, GM94, Gui11a, Maf10, Tur03, Tur06, Val09]. Italian-Hour [Tur03, Tur06]. Italiani’ [Chi08]. Italy [Car21, GG91a, Ger23, Maf13, Mar11a, Mel21, Vac00, Hei14, Mar21a]. item [Fra39b]. Itinéraire [Qué90]. itinerant [Ink72b]. itself [Ham01]. IV [Bay39a, Bru58b, Cha63b, Gib51e, Gib53c, Hug55, McK62, Nea51c, Old94, PM39, Tre70]. Ivan [Bec09]. Ivor [MIL15]. ix [May21, Mel21, Sch18b, Vic04, Nea52].

J [Ben14, Boa19, Cha91, Dun13, Edd13, Fra14b, Gre87, Gro14, Har96, dVH65, Hig17b, Jac15a, Joh14, Joh15b, Lai04, Luc19, Mau15, May21, McC13, McC14, New52, Por16, Pul94, Ram15, RN15, Rob95, Stu09, Sus16, Tau14, Top80, UB15, Van19, Wei15, Woo97]. J. [Adl97, Arm58a, Arm55, Bro79b, Cha91, McC39b, Per86, RS01, RH11, Rob95, Sco81, Sme54a, Sta19, Swe03, Top80, Woo71b, Woo71c, vdS06]. Jack [Hen18]. Jackson [How19]. Jacksonian [Col54b]. Jacob [Bec16, vH94, Bel12]. Jacobi [Pul94]. Jacobis [Pul94]. Jacopo [Fol11]. Jacqueline [Rei14]. Jacques [de 54]. Jacquier [Gui15]. Jahre [GG11b, Ste76]. Jahren [Thi73]. Jahrhundert [Thi71]. Jahrhunderts [Hen89]. Jakob [Ecc80, Ecc80, de 49]. James [BT55, Cha68, Dar13, Eyl63, Gol19, Hen09, Lev94, Wil92, de 49, All94, Dea73, Dea75, EE51, Ham82, Hen09, Jon82, Jon83, Jon84, Jon85, JTR94, JTR95, Kaa17, Mil83, Mil85, Rob54c, Sch54, SS13, Sne81, Swe72a]. Jameson [Jen16, Swe63, SW67, Swe67, Swe72a, Swe74]. Jan [vH94]. Janin [Dym11]. Janis [Ano05e]. January [Ano15, VvM92]. Japan [Hil91, Cro96, Fru16, HL91, HW87, HW90, Ink88, Kim95, Low90, Low96, Mac73, Ots05, Shi81, Str93, Suz09, Tan14, WT84]. Japanese [Ink91, Kim95, Shi89]. japanning [Gib51a, Gib53a, Gib53b, Gib53c]. Jardine [Van19]. Jazz [Goo11]. Jean [Fra45, Geo38b, Klo85a, Klo85b, Pat37, Sch13a, Sme55c, Wil55c, de 54]. Jean-Baptiste [Wil55c]. Jean-Francois [Sme55c]. Jean-Jacques [de 54]. Jenkin [Smi80]. Jeremias [Par51b, Par53a]. Jeremy [Gue13]. Jeroen [Del12]. Jesuit [Vic04, Cha84, Fei03, Whi86b]. Jesuits [Chu13, And14, Hsi09]. Jesús [Coz19, Gam21]. jewel [NF73, She09]. Jewish [Bur15, Cop80, Fre11]. Jimena [Law13]. Jnr [McC86]. Joan [Par36]. Job [McC37]. Johann [Gui95, SS99, de 50b]. John [Ano86g, Boa12b, Bur14b, Cha14, Coo90, Edd08, How19, Hud03, Mac19, McC39a, Mil13, Ols70, PM50, Por16, Pre16, SS23, Sin49, Sin50a, Sin50b, Tur84, Woo71a, AC99, Ano59a, App96, Bar11, Beg21, Bri13, Cas15, Cha95, Col54a, Col52b, Col52c, Col52d, Cum84, Dea83, Dea91a, DE50, Ela13, Eyl37, Gas07, Gib51d, Gui95, Gui11a, Har66, Hen18, Hes21, HBP08, Jac15b, Kaa17, Lev94, Man93, Mar11b, Mer98, Mil12, NG97, Olb89, PM50, Pat39, Rit62, Sch76, Sco36, SP92, Spr42, SBT10, Stu82, Stu89, Swe72a, Swe74, Tay52a, Tha66, Tho57, Uss80, VV22, Wan18b, Wat82, Whi51, Wil87, Yos99, vH92, vM86]. Johns [Bil18, Fra13, Gat18]. Johnson [Kid13, Uss80]. Johnston [Sne81]. Joiner [Bel12]. joint [Lev09a, DK55b]. Jonas [Wil91a]. Jones [Swe03]. Jöns [Bec16]. Jorgensen [Ren18]. José [Mal09, Mar21b]. Joseph [Ano75k, Arm56c, Boa16, Cop80, Cro59, DD21, Eyl63, Leh12a, MK60, Mol84, Mor09, Per83, Sch54, Sch57, Ste80]. Josiah [Bal23]. Journal [Bec13a, Nea51d, Nea52, Gie19, GG10, Hug51a, Mcc79, Nea50, Nea51a, Nea51b, Nea51c, PdC10, Sme57a, Swe67, Gie16, Lil48, Swe63, Swe67]. journalist [HR89]. Journals [Des08, Lev09a, Bar98, Ecc76, GG77]. Journey [Buc08, GG11b, McC91]. Journeys [Dal14]. Judd [OH97]. Julian [Her18]. Julius [Ecc80, Ecc80, Kra82, Wie23]. junior [Swe74]. Jupiter [Rob74]. Jurassic [Bry13, Bri10]. Jussieu [Gen22]. Just [Bar98, Gen22]. Justus [Hor87, Mad88c]. Jusuf [BvK73].

Kaiser [Wei85a]. Kalibrator [Sch11]. Kalibrator-Konzept [Sch11]. Kamerlingh [Pap09, vD08]. kansho [Thi12]. Kant [RO82, RO83, RO84]. Karl [DD21, GG02, Mac79, Mag98, Nor75]. Keeping [Zev00]. Keill [Gui95]. Keir [Dun67]. KEK [HL91]. Kekulé [Joh14, Roc85, Roc10, Wal39]. Keller [Rus04, Vic12]. Kellert [Akt09]. Kelly [Mos14b]. kelp [CC47]. Kelvin [Gos76]. Kenneth [Akt09]. Kepler [Reg19, Ait78, BG94, Bra82a, Bra82b, Cif86, GE19, Hon87, Kre23, Mal03]. Keuper [Nit96]. Kidneys [Tan11]. Kim [Low03]. Kin [San17]. Kind [SD13, BL21]. Kindred [Yal22]. kinds [Nak84]. Kinetic [Bru57a, Bru57b, Bru58a, Bru58b, Cah90, Kox90]. kinetics [Ber97]. Kinetik [Ber97]. King [Bur00, Gue10, App90, Bed93, Bur98, Hua22]. Kingdom [GH08, OEMB21]. Kinzō [AS81]. Kircher [Hir07]. Kirsti [Bin12]. Kirwan [Boa08, Sco81]. Kisir [Gel21]. Kisir-Assur [Gel21]. Kiyoshi [Kin04]. Klaus [Hei19]. Klein [Dij13, Har96]. Kluwer [Vic04]. Knight [Kid13]. Know [Hah15, Leh12b, Lin22]. Knowing [Vet10, Gue13]. Knowledge [Ano05j, Ano08l, Bra09, Bsc12, CMM08, Car17, Cha09, Die21, Gor17, Jus11, Jus12, KK02, May21, Mil11b, Nap09, Pop12, Sta17, Vac00, vD11, Bow84, CZ19, Fra41, Gie19, Hal20, Har74, Lon19, Mar21a, NG97, RS17, Sha75a, SM15, Uch22, Wei19, Bil16, Nic14, SS23]. known [Ari73a, Old72c]. Knox [Jen16]. Knut [Kär14]. Koch [Fra13]. Koelliker [Cam55]. Koenig [Pan09a, Wit13, Pan05]. Kohler [Kin19b]. konventionalen [Pul94]. Konzept [Sch11]. Kooijmans [Sch13b]. Kopp [Joh14, Roc10]. Kormos [Mar11b]. Korrespondenz [Hen86, Hen88]. Kossel [Ken52]. Kox [Har96]. Kraft [Jur10]. Kragh [Fol14b]. Kranich [Don50, Don51]. Krantz [Reg09]. Krementsov [Byf19]. Krige [SS23]. Kritik [Thi73]. Kt [Eme88]. Kuhn [Pot22, RD16, de 53a, Por16]. kulturgeschichtliche [GG11b]. Kundt [Cah90]. Kurt [Bec13a].

L [Bou13, Bsc14b, Bur14c, Car18, Gro14, Nea51c, Per86, Sch13b, Sme55a, Sme55b, Swe74]. L. [Sme54a, Woo97]. Lab [Mic13]. laboratoire [Mic13]. laboratoires [Fau05]. Laboratories [Cro05, Fau05]. Laboratory [Cro03, KS10, LR00, Bre16, HL91, McC83, Sme54b, LA22, Fra13, Rob95, vD08, Dij13]. Lacaille [Arm56a, FB22]. Lachapelle [Lyn14]. lacquer [Gib51a]. laden [Hen88]. Lady [Nor42, Pal23]. Lafuente [Gre87]. Lagrange [Rei14, Ste11a]. Lahore [NNS^{+}06]. Lake [Old99a]. Lalande [Cha88]. Lamanon [Car97b]. Lamarck [New52]. Lamarckian [Bow84]. Lamé [Bra81]. lament [Bik75]. Laming [Far69]. lamp [Mor36, Wei50]. L’Analyse [Tom01]. Lancashire [NG97]. Lancaster [Rus78]. Land [Bro09, Cas15, Chu13, Hsi09, Nor01, Lov79]. Land-marks [Cas15]. lands [Ano91o]. Landscape [TS14]. Lang [Boa19]. Langins [Ano05e]. Langmuir [Mil11a]. Language [Boa12b, Kim99, Edd08, Wig61]. Lapidary [Wal01]. Laplace [LS74, TT99]. Laputa [NM37a, NM37b]. Lardner [Hay81]. large [DK54c, Sme87]. large-scale [DK54c]. Last [Ber01, Koj02, Moz04, Bre93, Edw72a, Hen89, Pul94, PM39]. l’astronomie [Swe21]. Late [Ank15, Bea12, BD00, Chu13, Hei14, Kol99, Péo10, Pop12, Ree14a, Dol77, Dun94, Eas86, Fus58, Ma21, Mar11a, McC83, Mel21, Mes96, SR71a, Tac96, Whi86b, Bla10, Det18, Dup11, Gay13b, Hsi09]. Late-eighteenth-century [Pop12]. late-medieval [Eas86]. late-nineteenth-century [Whi86b]. later [Hen79, SR72a, WN16]. laterum [Tan77]. Latham [Ape20]. Latin [Jus11, Jus12, Min21, Not18, Ray82, MSD23]. latins [Bur12c]. Latinus [Bur11a, Bur11b]. latitudinarianism [Mul73]. Laurent [DK54a, DK55a, DK55b, DK57, Pot53]. lauri [ABNW09, Nor12]. L’avant [Sme57a]. L’avant-coureur [Sme57a]. Lavoisier [Ano07h, BS65, Ber01, Duv52, DK54a, DK54b, DK54c, DK55a, DK55b, DK57, Eyl63, Har41, Kim06, Le 72, LS74, Lop60, Par53b, Per86, SS58, Sme55a, Sme55b, Sme56b, Sme56a, Sme56c, Sme57a, Sme57b, Sme78, Sto66, Sto68]. Law [Aga17b, Ait69, AM76, Cam80, Cro61, Hor00, Nay80, Nay82, Par51b, Raw74, Swi65, Wal76, Wil56d, Spu87, Dra81, Kol99, Par53a]. Lawrence [Bey10, Hen00, Sei13]. Laws [Pis14a, Cap12, Kem17, Kot79]. lawyer [Sch19]. lay [Joh18]. Layering [And09a]. Leane [Ben13b]. Leap [SS99]. Learn [BB13, Sid07]. Learned [Kur03, Bik75, Jac52a, Jac52b]. Learning [KK02, Aga19, Wat92]. Least [How02, How03]. leather [Duv52, Spi68, Spi69]. leavening [Cob18]. Lebenswerk [Kan70, Kan71]. LeConte [Ste80]. Lecture [Pre08, Zev00, Cro59, Dar85, Edm75, Mer98, Mil85, New52, Old72c, Pul94]. lecturer [Har57]. lecturers [Ink72b, Mil83]. lectures [Bro69b, Bro55, Ink78, McK36b, McK59, McK60, McK62, McK65, McK67, Spr77, Wil62]. Lee [Lov79]. Leeuwenhoek [Col37a, Col37b, ZA16]. Lefroy [Lev86]. Legacy [BS07, Pul94, Dom09]. legal [BS15]. Legend [BV99, Gol19]. Legends [Tre15, Ham12]. Leibniz [Ait60, Ait62, Ait64a, Ait65, AS81, Ait81, Ano06j, Mel88, Mit79, Nau14, Roe90, Smi13, SC15]. Leiden [Car18, Gue23, Hal18, Jar18a, Nut21, vD08]. leisure [SM16]. Leith [Swe63]. Leland [Cha53b]. Lemaître [KL07]. length [Cha88]. lengths [Why54]. Lennart [Swe03]. Lens [vHvG99]. L’Ensei [Thi12]. lenses [MJ89, NF73, Sme87]. Leonard [Par16, Rus22]. Leonardo [Uga06, Eas86, Wil57, Wil65, Ray15]. Leong [Wei19]. Leonhard [BS07, Cha21, Dom09, Hom88, Kno89]. Leopold [Ecc76, Ecc80, DD21]. Leslie [Ols69, Ols70]. Lessons [Koo10, Sch13b]. Let [Kra17]. Letter [SC15, Ait81, Ari74, Arm56c, Bre16, DK54c, Fra47, Ill89, Lev94, Low75, Mad88c, Sch54, Sme78, VvM92, Wei50, Wil87]. Letters [Nor12, Ano86g, Arm58a, Ben73, Bik75, Bur60, Edw72a, Eva17, Hei76, JTR94, JTR95, MK60, Pel36, Sch80, Ste61, Ste69a, Ste69b, Thi73, de 49, de 58, ABNW09, Rya23, Tha66, Ray18]. letzten [Hen89, Pul94]. Leu [Vic12]. Levene [McC14]. Leviathan [Bru11]. Levitation [Eir18]. levity [PM37, Per83]. Lewis [Gib52c, Bro18, Gib52a]. lexical [Rou22]. Leyden [Lin63]. Libavius [Ano09g, KH60]. Liber [AVJ^{+}06, Bur14a, Vla10, Bur11b, BvK73]. Liberal [Rom78, Sti37]. Liberalism [Are10, Joh15b]. libertinage [Coh11, Lyn13]. liberty [Bed93]. Libraries [Nor12, Wel83, Win66]. Library [Vic12, LGKW08, Mor68, Swa89, Vic12]. libros [Bur14b, SBT10, Mil88b]. Licensed [Bra02]. Lichte [Ecc80]. Lichtenberg [For72a]. Liebig [Hor87, Mad88c, Son74, URRF05]. Lien [Flo96]. Lien-teh [Flo96]. Lieutenant [Hig17b]. Life [Adl17, Ano05h, Ano06f, Are10, Bec84, Blu13, Bow09a, Bow17, BS07, Dau14b, Del09a, Dom09, End16, Fra14b, Gos15, GG14, Har09b, Jac09, Kur99, Kur03, McO00, Mon05, Mor08, Nic11a, Prü09, Ras17, Rus04, Sak12, Ste00a, Stu09, Woo15, vS15, Aga17b, Ber19, Bow74, CMM08, Car10, CB22, Dav78, Eme13, Gol19, Gre05, Gui11a, Har12, Hir07, Jar23, Jon85, Kan70, Kan71, Mad57, Mad88a, McC86, McC10, NG97, Pal23, Ric86, RH11, Row50b, Sha73, Sta88b, Tob12, Wei85b, Ren18, Sin49, Yi13, Nic13, Nic14]. life-blood [Ric86]. lifelong [Han21]. Light [HZ07, Old99b, Ait62, Che98, Dea89b, Dek90, Duc97, Ecc80, GG74, GG97, Hak82, Hei19, Hen80, Hen21b, Hom88, Mal94, Nak84, Por78, Zig80, PM38b, Eis05]. Lighthouse [Det18]. lighting [Fox96]. Lightman [Hal18, Hig17a, Jam12]. Lignum [Par55]. Limbs [Qui09]. Lime [Pag03]. limit [Wil56d]. Limitations [Hil99]. Limited [Ano06h, Edd13]. Limits [Ano08h, Mil16, Nof14, Sta17, Beg21, For15, Rib20].

Lindsay [Rit62]. line [Hen92]. Linear [Big21, Rei89, GG97]. linguistic [Bec16]. Linienprofile [Hen92]. link [Tee79]. links [Bra81]. Linnaean [Die12, MWO07]. Linnaeus [Die16b, MW13]. Linstrum [Hay18]. L’introuvable [Dur10]. Lion [Lau21]. Lippmann [Mit12]. liquid [Adl97, Bru79]. liquid-in-glass [Adl97]. liquids [Tre78b, Tre80b]. Lisbon [Ama06]. Listening [Tka14]. Lister [Mor86]. lists [Gia23]. literary [SR74a]. Literature [Dan10, Smi22, Ste10a, Str14, Wer14, ATE11, MHO13, McA12, VT17, Wil11b, HA15, Lyn15]. Little [Ano07e, Ano07f, Ari73a, Old72c]. liver [You37]. Liverpool [Ish14, Kit82, RS71, RS74, SR72d, SR72e, Whi86b]. lives [Sme55a, Sme55b]. living [GG77, Kat23, Wil60a, Wil60b, Wil60c, Wil61]. LL.D [Swe72a]. Loadstone [Uga06, Ari81]. Lobatto [Sta88b]. Local [Nap09]. Location [Cro05]. locations [Win66]. Lochner [Wal76]. Lock [Ben14, Gru11]. Locke [Nut14, Car13, Col95, Cor11]. Lockyer [LR00]. locomotion [Fle97]. Lodestone [Mil04b]. Loeser [Cha72a]. Logarithmic [War07a]. Logic [Bin12, Del09b, Dom09, Haa09, Kim99, Rob10, Wol13a, Ano75l, Hes52, Lai77, Lai79, Wes97, Wil09, Tac96, GG13b]. logicism [GG74]. logics [GG75b]. lois [The86]. lok [Dek08]. Lombardy [SC15]. Lomonosov [Lei73, Lei76]. London [Coo13, Coz19, Hay18, HA18, Kin04, Kin19b, Kri19, Kus18, Lai04, Lan13, Lin22, Mcv21, Mil13, MS22, SS23, Sha19a, Sla19, Smi18, Sus16, Swe63, Ver13, Vic04, Wit13, Yi13, Arm58a, BT55, Bry92, Can10, CH82, Cra87, Dor76, Edm75, Gay00, Gra70, How02, Ink77, Kur01, Mil83, Mil11b, MJ89, Ock37, She09, SP85, Spr77, Swe63, Tho51b, Tre63, Tre64, Tre65b, Tre65a, Tre70, Win66]. Longino [Akt09]. Longitude [HBP08, Plo99, FB22, Gre84]. Longolius [Har58]. Longomontanus [Swe09, vM86]. look [HT23, Rei89]. Looked [MW13]. Looking [End16, Vic00]. Lore [Wal01]. Lorentz [Kox90]. Lorenz [Gou83, Kre23]. Lorraine [Por16]. Lost [HP03, Whi12]. Lot [Ano07e, Ano07f]. Louis [Bro36a, Arm56a, Cha72b, Dav68, Dym11, Qué90, Swe70a, Wil55c]. Lourenco [Rap15b]. Louvain [Cha12a, Hir08, Van01]. Love [Rob54b, Bik71, Boa19]. Low [Jar18a, vD08, GG86, GG84a, SP85]. Low-Temperature [vD08]. Lowell [Str94]. Lower [Fra39b]. Ltd [MS22]. Lucier [Wil13]. Luigi [Kip87]. Luis [Nut21, Ren18]. Luisa [Bed82]. Lumières [Lyn13, LL37, Coh11]. Lunar [Swe09, FB22, For66b, Rob55a, Rob56, Rob57, Sch56]. lunar-distance [FB22]. Lunars [Sch14]. Lunbeck [Bou13]. Lundmark [Kär14]. Lussac [Cro61, Spu87]. Lutheranism [Mas53b]. Lux [Ano06e]. Lycée [Sme55a, Sme55b]. Lyell [SP92]. Lyngurium [Wal01]. Lynn [Sak12]. Lysenko [RH85]. Lysenkoism [Ros21].

M [Boa19, Bow13, Bri13, Kid13, Ku13, May21, McC14, Sla19, Sne76, Sta21, Woo97, Wri16, Gen22]. M. [Har96, Hil91, Lei73, Swe74]. M.B [Gib52c]. M.D [Rus49, Rus50, Rus51a, Rus51b, Swe72a, Ter37]. M.D. [Rob55b]. M.P [Pat36a]. M42 [Sie09]. MA [Coo13, Hay18, Mel21, Mil13, Rus04, Vic04, Yi13]. Macartney [CBL81].

MacBride [Bow84]. Macculloch [Eyl37, Cum84]. MacDonald [Fra14a]. Macedonio [Sch89]. Mach [Hen88, Bla78, Hen88]. Machine [Bra10, Lus14, Cha53a, Coh36, Edw10, Gee93, Gor16]. Machines [BD00, Rus04]. Machlinenses [Not18]. MacKenzie [Bri13, Smi87, Bow85]. Maclure [Dea89b]. MacMillan [Coz19]. Macquer [Leh12a, NS81, Sme54a]. macroscopic [GG88]. madame [Sme78]. Maddison [Ben95]. Made [Fit17, Dav68]. Madness [SK11]. Maffioli [Duc14]. Magazines [Bow16]. Magellan [Boa19, Sco81]. Magellanic [Dek90]. Magendie [BS12]. Magic [Eir18, Hun23, Gel21]. Magne [Nor12]. magnesia [Mul76]. Magnetic [Bro87, GG91a, Hon84, How02, Lev86, Pra10, Yos99, BM94, Cot76, For98, Pum87, Mac18]. Magnetischen [Hon84]. Magnetism [Uga06, Caw77, Cot77, Pum87]. magneto [Gee93]. magneto-electric [Gee93]. magnets [Ari81, For98, Qué90, Sne76]. Magnus [Bel99]. Mahameleth [Bur14a, Vla10]. major [Min21]. Makers [Car21, NF73, Cra87]. Makes [Lat15, Pro18]. Making [Ada17, Ank15, Ano07g, Arn09, Cha13, Fra10, Gar04, Gay13a, Joh10, Kid13, Mac19, Mil11b, Mor01, Old99a, San17, Smi18, Tat17, Van01, Wal09, Wex08, Bak21, Bro79a, Bry92, Cha84, End18, Gay10, Gay13b, Hes21, Kni09, Mil88a, Min21, Sch90, Gie16, Rus04, Gue23]. Malacrida [Cha71]. Malady [Bur10c]. Malaquias [Boa19]. Malaria [Joh10]. Maligned [Haa19]. Malleable [Car21]. Mallet [Dea91b]. Malpangotto [Swe21]. Malpighi [Cob02]. Malta [Hop81]. Malthus [Bas15, Hal14a, Mei21]. Malverns [Old92]. mammals [Fra41]. Man [Gos15, Tat17, Ano75k, Bik75, Gay10, Har12, Por78, Win72, Fra41, Li12, OO51]. Management [Pop12]. Managing [Akk23, Nof14, ZB10, Hal14b]. Mance [Hop59b]. Manchester [Bri13, Par13, Alb09, Swi12]. maner [Ano91o]. Manichean [Ano03f]. Manipulation [Fer02]. manpower [Low90]. manufactory [Ful64]. manufacture [Gib39, Gib51c, Gib55, Git66, Gor87, Kra95, Lam97]. Manufacturing [Bre93, Gau12a, Pas51]. Manuscript [Eis05, Swa89, Dob36, Gos77, Pep76, Ree81, WT84, Yap83]. Manuscripts [Boa09, Bla78, CT69, Jay78, Rus63c, Tan80b]. many [End18, LS23, Mar16b, Rab77, Tur82]. Map [Swe03, Eyl37, EE38]. Mapping [STDH22, Urb80]. Maps [How03, Lan13, DE50]. Marc [Edd13]. Marcelin [Tre78b]. March [CCC57]. Marci [Ait70]. Marck [Ama06]. Marconi [Hon96]. Marcus [Ait70, Mar21a]. Margaret [Bec11, Sha19b, Sus15]. Marginalization [Gra01]. María [Coz19, Mal09, Bed82]. Marie [Bur14a, Dav95]. marine [Dav78, For66a, McC05]. Mariotte [Wil57]. Maritime [Kin04]. Mark [Bar14]. Market [Dij13, KS10]. Marketplace [Jam12, FL07]. Marking [Jac12, Tay08]. marks [Cas15]. Marques [Rap15b]. Marquess [Arm58b]. Marr [Hei14]. marriage [Bik71]. Mars [Ait78]. Marten [Wil55c]. Martin [Mil83]. Martínez [Sha19a, Bey14]. Martius [Uch22]. Martyrs [Ale10, Coo13]. Marx [Feu75, Feu76, Tee79]. Marxism [Mit78]. Mary [Jon95]. Masculinity [Nye18]. Maskelyne [Cun04, For71b]. Masonry [San00]. Mass [SC04, Ste76, Wes10, RP98].

Massachusetts [Sei13]. mast [Ari72]. Master [App96, Arm51, HR89]. matching [SL77]. matematiche [Maf10, Duc14]. Material [Bla10, Car21, Hal20]. materialism [Foo52, Mit78]. Materials [Bad11, Cro09b, KS10, Bak21, Gre22, Dij13]. maternal [Wil92]. Mateu [BS12]. mathematica [Mel88]. mathematicae [Ste00a]. Mathematical [Ben13a, Bin12, Bur10a, Des08, Dor76, Hei14, Hun99, Joh91, Pri10, Tak11, Tur73, Ung02, War14, And07, Bry92, Dun94, Gou10, Gra72a, Kno89, Low75, Mar11a, Par84, Rom06]. Mathematician [SS99, Ano75k, HR89, Sta88b, Wil91a]. Mathematicians [Maf13, Sno01, Hen89, Rei81]. Mathematics [Ale10, Ano11d, BG99, Bin12, Del09b, Des08, Dom09, Duc18, FRW11, Fra14b, GG14, Gui11b, Kid14, Mes96, Rob10, Ung02, War07a, Wil95a, Wil95b, CB22, GE19, Gra72b, GG11b, Haa19, Høy23, Jar18b, Maf10, RH11, Rob63, Roe90, Sch13c, Sco36, Tan80a, Tan80b, Tho16, Tob12, War12, Coo13, GG93, Hun14]. mathematicus [Duc17]. Mathematik [GG11b]. Mathematikern [Hen89]. mathématique [Rom06]. mathématiques [Rob10, Duc13]. Mathematische [Kno89]. mathematization [Mag98]. Mathesis [Mit79, Gam21]. Mathew [Boa12b]. Matter [Ano07g, Mil16, BM94, Cha19, Far69, For15, Hug52, Nye81, Rus63b, Smi80, Sne76, Wil60a, Wil60b, Wil60c, Wil61, Win23]. Matters [Cha16, SB13]. Matthew [Rob53b, Rob54c, Swe64]. Matthias [Pis13]. maturation [McC21]. mature [Bra88]. Maunder [Hen18]. Maurolyco [HZ07]. Max [Del09a, GG97, Gre05, RB01, vH92]. Maximilian [BL21]. Maxwell [Bru58b, Cra54, Dia94, Hop57a, Hop57b, Hop58, Hop59a, Hop59b, JTR94, JTR95, Tur55]. May [CS12, Rus04]. Mayer [Cha21, For66b, For70a, For70b, For71c, For72a, For72b, Hen82, Ku13, Sne76]. McCrory [Dau14b]. McEvoy [Mil12]. McEwen [Mil91b]. McKie [Ano68c]. McLeod [Gay13a]. MD [Fra13]. meaning [dVH67]. meanings [Ges23]. means [Pel97, Top80]. Measurement [CY05, Gre87, Hop57a, Mit12, Ste07, Cha95, Goo80, Hop59b, Hum37, Jal23, McC21, Rob95, Rut39]. Measurements [Bro15, Hen96, Wil11a, For98, Het72, Sch89]. Measures [Ste76, Rei89, Sim92]. Measuring [Ano91o, Cra12, For70a]. Mechanic [SR72f, Ink75]. Mechanical [Smi76, Ber97, Coc56, Cot76, Hil85, Top80, Val09]. Mechanics [Cap12, God09, Mar15, Pis13, Pis14a, Sal09, Slo11, Sta19, Ait60, Ait62, Ait64a, Ait65, App96, Arm50, GG84b, Kra21, Pul94, Sch08a, Sch08b, Vos13]. Mechanik [Ber97, Pul94, Pul94]. Mechanik-Vorlesung [Pul94]. mechanischen [Ber97]. Mechanism [Ehr95, Foo52, Rou22, Ber97, Bes68, CB22, Sch13c, Tho16]. Mechanizing [Hol96, Pum87]. Medal [RH11, Fra14b]. médecine [Sme56b]. media [Hol07]. mediaeval [Mar76]. mediator [Dau13]. Medica [Gue23]. Medical [Ano07j, Boa12b, Bro78b, Cro09a, Edd08, Fra13, Kah10, Kre11, Law08, Low96, Ber79, Deb62, Fra36, Gra70, Mor86, Mor86, Ape20]. Medicinal [Mad62]. Medicine [Ama06, Ano08g, Att09, CC09, Cha12a, CC11, Coo11, Gel21, Kid13, Lyn15, Mos14b, Nut21, Pom11, Ree14b, Sil18, Win16, BS15, Big22, Coo90, Dav90b, D^{+}11, Hir08, Kel11, Kni09, LPST11, Lev09a, Mac23, Mar21a, McA12, Mei79, SM16, Swa89, Wei19, Mor86, McC13]. medicines [Jon19]. medics [Sil18]. Medicus [Pop12]. Medieval [AC99, Att09, Bur10b, Bur11b, Bur15, God09, Wal01, Ano75l, Ave20, Dek95, Eas86, Fre11, Not21, Tac96]. medium [Gre77]. meeting [Bro36a]. Meetings [Ano57a]. Meets [Smi22]. Megascience [Kra11a]. Meiji [Cro96]. Meinungsstreites [Ecc80]. Melchior [Mar19]. Meldola [Gay10]. Melia [Del11b]. Melinda [Gie16]. Melloni [Cor38a, Cor38b, Sch89]. members [Mad55]. Membership [Sch56, Sme56b, Sme56a, Sme57b]. memoir [Dav90a, Par53b]. Mémoires [Ber01]. Memoranda [Mul92]. Memorial [Plo78]. Memory [Jac12, Mul92, GG97, Jay78]. Men [Car05, Clo55, de 64]. Mendel [Fis36, Kot79, Mac85b, OG68]. Mendeleev [Raw74]. Mendelian [Mag98, Mei85, MWO07, Zev00]. Mendelism [The94]. Mengoli [War07a]. Mental [Can10, Hei19]. Mercator [TD93, Tur94, Van01]. Mercer [NG97]. Mercure [Gen07]. mercury [Gen07, Mil88b]. mercy [Di 17]. Mere [Alv07]. merger [Ste76]. meridian [Cha95]. Merseyside [New38]. Méry [Fra45]. Mesotrons [Mon05]. mesozoic [Old72b]. mesozoic/tertiary [Old72b]. message [SB09, Pal12]. Messrs. [Lev94]. mestizo} [OEMB21]. metal [Wil59]. Metales [S´ 20]. metallic [Cha72a, GG91a]. metallurgic [Tei75]. metallurgical [Wil55b]. Metallurgy [NNS^{+}06]. Metals [Tre65b, Tre65a]. metalworking [Gor87]. metaphor [Lau66]. Metaphors [Rus04, Wil55a]. metaphysical [Bra82a, Bra82b]. Metaphysics [Goo80, Nor75]. Metapsychics [Lyn14, Lac11]. Meteorologica [VDdM08]. Meteorologica. [Bur11a]. Meteorological [Mac18, Dav84]. meteorology [Gar76, Gro72a, Kno64, Oli69, Har06, Hug53]. Meteors [Kär14]. method [FB22, For70a, For72b, Hop59b, LL37, Lai77, Rog72, Sch13c, Tho16, Wes56a, Wes56b, Gre81]. Méthodes [Sim13, Tom10]. Methodological [Fol14a, GG86, GG84a, IR11, Lau66]. methodology [McD72]. methods [Bak62, FB22, Tom10]. Methodus [Ray82, Pre16]. Métiers [BS65]. Metrics [Mar13]. Metrology [Hen07]. metropolis [Ink77]. Meusnier [DK54c]. Mexico [Min21]. Meyerstein [Hen07]. Michael [Bsc14b, HA18, Pis14b, Can10, Car92, Kra13, Moo87, Nye11, SS15, Boe14]. Michel [Nic13]. Michela [Swe21]. Michell [Cha54a, Gow82, Har66]. Michiel [Mes97, Tur06]. Michurinism [Wei85b]. Mickleburgh [Col52d]. Micro [Wal04, Ver94]. Micro-organisms [Ver94]. Micro-politics [Wal04]. Microbes [Ver94]. Micrographia [Bar95, Sch18a]. microhistorical [Gel21]. Microphone [Wil07]. Microscope [Ano06g, Com21, GH08, La 11, Law09a, Fou89, Har84]. microscopes [ZA16]. Microscopy [Edd13, BS15, Rat09, Smi23b]. microwave [CSVK80]. Micy [Bur00, Bur98]. Mid [Het83, Gál88, Het76, Mil88a, Sha73, Tho55, Wri92, Yeo79]. mid-19th [Tho55]. mid-eighteenth [Mil88a]. mid-nineteenth [Gál88]. Mid-nineteenth-century [Het83, Het76, Wri92]. mid-seventeenth [Sha73].

Middle [Bur09, Ell83, GG11a, Wil01, Gro74a, Mac79, Rya91, Rom06]. Midlands [Gib53c, Mil85]. might [Sch19]. Mikhail [Lei76]. milestones [Par45]. Militaire [Ano05e]. militaires [Bre91]. Militarization [Mos14b, Kel11]. Military [Coo11, Hop81, Kai10, Bra75, Bra92, Bre91, Fle97, Lin22, Sch19, Wei48]. mill [Gib51e]. Miller [Gol19, Pal83, Lal12]. Millikan [Sok96]. ‘Millstone [Cum84]. Milne [Har74, Lep06]. Milner [Col54b]. Mind [Jac12, Nut14, Cor11, Mei21, Wil92, Zig80]. Mindful [Dup11]. Minds [BD00, Hay18, Lin16]. Miner [Stu09, Don50]. mineral [SM16, Wat82, Ma21]. mineralogical [Gen22, Old74]. mineralogist [FF68, Kau75]. Mineralogy [Boa12b, Bre16, Edd08, Por81, Ma21]. mineraria [Bri07, Maz10]. Minerva [Car22, Smi22]. mines [Mor36, Str97]. Minimum [Hen18]. Minimus [Ozd14]. Mining [Bri00, Kle12b, Kle12c, Kon12, Mil16, Vac00, Don50, For15, Por81, SR72b, Kon12]. ministère [Fau05]. Ministry [Bre91, Fau05]. Minnesota [Akt09]. minus [Tan80a, Tan80b]. Minute [Wil37]. Miri [Bil16]. mirror [Gen22, Ree08, Wil11a]. mis [Spu87]. mis-teaching [Spu87]. mislead [DD21]. misplaced [Sal92]. Missing [Cun04, Hei67a]. mission [Cha84, Lin22, Whi13]. Missions [Chu13, Hsi09]. Mistical [Ben14]. Misticall [Gru11]. mistreatment [BL21]. Mitchill [Ape20]. mite [Sch18a]. Mittelalter [Bur09]. Mittler [Dau13]. mixing [For70b]. mobile [RG23, MSD23]. mobilization [Mac93]. Mode [Rei04, Ear61]. model [Hen88, Ink89, Und63, Zac72]. Modellbeladenheit [Hen88]. Modelling [And09a]. Models [BM17, Bri00, Car21, Lus14, Rus04, vdS06, CR96, Edw10, Gam21, Hei19, Wil55a, Wri92, dC88, Ano06h]. Modern [Ale10, Bur10c, Bur12d, Car17, Cob02, Coo13, Dan10, Dij13, Duc09, Gau14, God09, GG13b, Haa09, Hun23, Jac09, Jac14, Kid13, KS10, Kra00, Lai04, Maf13, Mar13, May21, Mil03, Nap09, Nut14, Nyh09, Pom11, Por01, SD13, Ste10a, Tre15, Wad15, Wal01, Wil15, YT13, Ako21, Bal06, Cam23, Coh10a, Cop87, Cor11, Dun21, Ham12, Jar23, JR12, Kem68, Kin19a, Kni09, Lep06, Mal20a, Mal20b, Mar21a, MG10, Mel21, Mor02, STDH22, Ste19, VB11, Wei19, Wit96, Yeo14, Yeo18, Sak12]. modernisation [NGDP75]. Modernism [Hal13]. Modernity [Das09, Kid13, Kni09]. Modernizing [Wea09]. modification [Nak84]. Modifications [Boa08]. moléculaire [Cam09]. Molecular [For98, Rei04, Yam98, vdS06, Bea76, CR96, Dia94, Cam09]. molecularization [Hol17b]. molecules [Ego19]. Moll [Sne84]. Møller [Bec13a]. moment [Aga17b, Kli15]. moments [For98]. monde [Duc13]. Money [Car13, Mar13, Di 17]. moneyed [Bik75]. Monge [Bin12, And07]. monk [Arm53]. Monkey [Fan09]. Mont [Tou89, de 56]. Mont-Blanc [Tou89]. Montano [Sla19]. Monteiro [FB22]. Moon [Bec84, TT99, Sid07]. moone [McC36a]. moons [Rob74]. Moore [Aga17b, Law08, Par84, Wil91a]. Moral [Cro16, Row18]. Morality [Ber79, Bur17]. Morange [Nic13]. Morbeka [Bur11a]. Mordechai [Vic04]. Morgan [Bow78, Lai79, Wal97]. Moritzi [de 60]. Morland [Bry75]. Morley [Ham03]. Morris [Lai04, Mat54, Tau14]. Morrison [Jar18a]. Morrison-Low [Jar18a]. Morse [Cro96]. Morten [Swe72b]. Morveau [Cha72b, Sme54a]. Moseley [Hei76, Hei67a]. Mossensohn [Bil16]. Mosson [Hil86]. Most [Bey10, MN13, Bak21]. Mote [Sok96]. Motion [Pis13, Erl92, Gro74c, Hen82, Kem17, Nay77, Sch08a, Sch08b]. motions [Ait57, Ait58a, Ait58b]. motive [Bro36b]. motu [AVJ^{+}06, Bur11b, Fra39b]. motuum [Kre23]. mountaineer [How19]. mountains [de 50b]. Moustiers [Not21]. mouvement [Bla88]. movement [Geo52, Gra39, Law21]. Movements [Kai10]. moving [Ari72]. Moyen [Rom06]. Moyes [Har57]. Mozambique [Rap15b]. Mozart [GG92]. Mr [Dav57]. Mr. [Fra10]. MRI [Pra10]. Mrs [Har84, Orr14]. ms [McK36b, McK59, McK60, McK62, McK65]. MS. [McK67]. Muddy [dA21]. Mueller [LMBM06]. Muir [Dea91a]. MUL.APIN [BB13, WH11, BB13]. Müller [Tre86]. multi [Nor75]. multi-factorial [Nor75]. Multiple [Ash17, UB15, Ges23]. multiplication [vH92]. Multiverse [Kra09a]. Mundane [Gar10, Ren11]. Mundi [Ano06e]. mundo [Pia21]. municipal [LPST11]. munitions [Mac89]. Murchison [Sta88a]. Murder [Ada17, Ste10a]. Murdoch [Smi87]. muriatic [Le 74]. Murray [Ste00b]. Musée [Ben09]. Museum [Kin04, Lai04, Pan09b, VW10, Wal14, Chi70, Alb09, Ben09, Bur12b, Eyl67, McC86, Mor10, Par13, VW10, Tau14]. Museums [Bri10, Bry13, Jac99, Mac09, Nor12, Rad18, Bri13]. mushroom [Jon86]. mushroom-shaped [Jon86]. Music [Ano05g, Hui11, War07a, Gou80, KO83]. Musical [Hui11]. Musick [KO83]. Muspratts [SR72d]. Musschenbroek [vH94]. must [Pou91]. mutation [Bow78]. Mutiny [Lar61]. Mutio [Hei14, Mar11a]. Mutualism [Bas15, Hal14a]. my [Swe67]. Myanmar [WN16]. Mydorge [Gou22]. Myriad [Mar11b]. Myrick [Gro14]. Mystery [Pri10, SHM13]. Mysticism [Ben13a]. Myth [Eir18, Far83, Pra10, Dei91, Tee84]. mythologie [BV83]. myths [Mar11c, Bey14].

N [Byf19, Low75, Lus14, Van19]. N. [dVH65]. nach [Thi73]. Nachschriften [Pul94]. Nagaoka [Hei67b]. Nageli [Wil60c, Wil60a, Wil60b, Wil60c, Wil61]. Name [Whi13]. Names [Ben14, Gru11]. Naming [Ben13a]. Nanko [Thi12]. Naomi [Lin22]. Napoleon [Bra75, Dal14]. narcotic [Ber79]. Nares [Lev86]. narrative [LS23]. narratives [Car97a]. Nathaniel [CT69, Mel53, Tan77]. Nation [Ban14, Nor01, Tau14, And10, Het83, Mor10]. National [Joh15b, Kin04, Aga19, Bec16, GG91a, HY87, Swa89, Wit96, Hol17b, VW10]. Nationalsozialismus [VW10, Wal14]. NATO [Kri19]. natron [Gib38]. Natural [Bew16, Bil18, Bos03, Bri13, Bro81, Bro77, Cha12a, Col54b, Duc18, Ell00, Fra09, Lov13, Mac73, Mar21b, Mil01, Nap09, SR71b, Wol13a, Bla19, Bod56, Chi70, Chr95, CC47, Doo10, Gol19, Gou83, Hir08, Jon37, Kni77, Lar91, Mac09, Mul92, STDH22, Sha73, SP92, Sla19, Tac96, Tre78a, Van19, WN16, Why54, Wil87, Wil91b, Wil09, Yeo79]. Naturalis [LS15b]. naturalist [Arm58a, Car97b, Mei21]. Naturalistic [Moo87, WH11, BB13]. naturalists [Ste69a, Ste69b]. Naturalium [AVJ^{+}06, Bur11b]. naturalization [GE19]. Nature [Alb09, Ano09i, Bow13, Bsc14b, Edw09a, Hil99, Jam12, McC10, Qui09, RB01, Sak12, Wea09, Wol13a, Ben37, Cah90, Car97a, Cro48, Hak82, Ham01, Har09a, HNS11, Hen21a, Jen16, Ked16, Lig10, Mas53b, Nak84, Nyh09, Ols70, Sce94, Sch76, Smi80, Top80, Uch22, Uss78, Wil09, Bar11, Gie16, Par13, Aga17a, Bar98, Dau14b, PW08, Ser12]. nature-philosophy [Mas53b]. Natures [Bew16, Bil18]. Nautical [Cam23, dVH36]. Navigating [Ano11d, Lon19]. Navigation [Wal00, Rob63]. Navigation. [Köb17]. Navigational [See18, Dun23]. Navy [Hac79b]. Naylor [Mac82]. Nazi [Hen00, Ros98]. Near [Bry10]. Nebel [Sie09]. nebula [Sie09, Dek90]. Nebulae [Bec13b, Ste10b]. necessity [McG78]. Needham [Lat15]. Néel [Qué90]. Negative [Alt07, PM38a]. Neglected [Aug11, Mac85b, SD68, Wil55c]. Negotiating [Alb89, Bec16]. Neile [Kem17]. ‘Neither [Vic04]. Neo [Gel21]. Neo-Assyrian [Gel21]. Nepal [WN16]. nephriticum [Par55]. Nervous [Bea12, Ree14a]. Netherlands [Bec01, Fli08, JR12, Sne75, Sne81, Sta88b, Zui04, Wad15]. Network [Mil13, CP97, FY09]. Networks [Ban14, McO00, Tan14, And10, Edw10, ML^{+}12, Bec14]. neue [Sie09]. Neunzehnten [Thi71]. Neuroses [Ree14a]. Never [Gau12a, Nie10]. Nevil [Cun04, For71b]. Newcomen [Gre22]. newer [Har41]. Newminster [NS52, Tho51a]. Newton [Ait55a, Ait69, Bra88, Bra90, Bur60, Bur65, Cha17, Cop80, Dav57, Duc17, Duc88, Fei19, FS21, Fle74, GE19, Gou23, Gui15, Hal55, Hal57, Haw68, Haw71, Hen21a, Hen23, Kol99, Lev21, McD72, McG78, Nak84, Nau19, Pou91, Pou01, Sim04, Sno01, Sta19, Ste19, Sys93, Tre66, Wan18a, WT84, Sch18b]. Newtonian [Sta19, Ait62, Ano06i, Cha75, Hom88]. Newtonianism [Mac20b]. Niccolò [BS21]. Nicholas [McC39c]. Nicholson [Lil48]. Nicolaas [Qui12]. Nicolas [Arm56a, Hak82]. Nicolas-Louis [Arm56a]. Nicole [Bur14c, Bus10]. Niels [Kra17]. Nierenstein [BL21]. Nieto [Dun13]. Nieto-Galan [Dun13]. Ni‘matallah [Tak11]. Nineteenth [Bel99, BD00, BS12, Bru79, GG11a, Hen14, Jam12, Kid14, Mil02, Old72b, Smi15, SR72e, SR74a, Tom14, War12, Wit13, BS15, Buc13, CR96, Caw77, Cha70, Cob18, De 17, Dol77, Fru16, FL07, Gál88, Gie19, Gro74a, Har71, Het76, Het83, Kit82, Klo85a, Klo85b, LS23, Li94, NG97, NF73, Ols70, Pan09a, Pau78, Pel97, Rad18, RS71, RS74, Sch08c, Sha75a, Sha72, Ste69a, Ste69b, SR72c, SR72b, SR73c, SM16, Tay52b, Whi86b, Wri92, Yeo79, dBP86, Cav06, Cro09a, Dar12, Gra01, Ku13, SR71a, SR71b, Thi71]. Nineteenth-Century [BS12, Jam12, Mil02, Smi15, Bel99, BD00, Bru79, BS15, Buc13, Fru16, FL07, LS23, Li94, Pan09a, Pau78, Pel97, Rad18, Sha75a, Ste69a, Ste69b, SR72b, SR73c, Cav06, Ku13]. Nippur [Rob10]. nitrate [Don36a, Don36b]. nitre [Gib38]. nitric [Par53b]. NJ [Sta21]. Njåstad [Nor12]. NL [Mil13]. NMR [Rei04]. no [Mei85]. Nobel [Del09a, Gre05]. Noble [Hig17b]. Nocturnal [Tur03, Tur06]. Nocturnals [Kin04]. Nollet [Hom79]. nomenclature [Hen78, Sme54a]. nomological [Ked16]. Non [End16, Jus11, Jus12, Old99b, VT17]. Non-Fiction [End16]. non-human [VT17]. Non-transferable [Jus11, Jus12]. Non-written [Old99b]. Nonproliferation [Gor17]. nor [Vic04]. Nordic [Rya23]. norms [Lus18]. North [McC13, CC11, Hes17, Aug11, Dol77, Hud03, Mil91b, Sok94, Swe74]. North-West [Swe74]. Northrop [vH92]. Norwegian [ABNW09, Nor12, Rya23]. nostro [Pia21]. Notation [Chr10, Gas07, GG13a, Alb89, Dub63]. Note [Ano54a, Ano55b, Ano55c, Dav57, Kun00, Mac82, Pat39, Bro93, Col54a, Don51, Gib53d, Hom79, Ink72b, Kni77, Mor86, New52, OM95, Ols69, Rus49, Spr77, Tod82, Tur55]. notebook [Gib52a]. Notebooks [Yeo14, Kno89, Kot78, Wil15]. Notes [Ano37a, Ano37b, Ano38a, Ano38b, Ano39a, Ano39b, Ano45, Ano49a, Mad55, Pel36, Cro59, GG92, Lev94, McK36b, Old72c, Ree81, Smi87, Swe72a, Tur82, Tur91, Wil59]. noticed [Ano92i, Ano92j, Ano92k, Ano93h, Ano94i, Ano96h, Ano97g]. notices [JL79]. notion [Ari81]. Notizbücher [Kno89]. nouvelle [Cam09]. nova [Ray82, Pre16]. novae [Swe21]. Novel [Bec09, Buc13, Tur06, Smi15]. novels [Bez75]. November [Ano15]. NSF [SP11]. Nuclear [Sei13, Bad09, For98, Hei67b, Jon86]. nucleosides [Hen78]. Nucleus [And10, Ban14]. Number [UB15, Hop59a]. Numbers [Bsc14b]. Numerical [Chr10, GG13a]. numerological [GG94]. Numerology [GG92]. Nummedal [Lau21]. Nuncius [Pal12, Lat15, SB09]. nutrition [Fus58]. Nutton [Mcv21]. Nyhart [Sak12].

O [Sme54a, BBHN14, Lat15]. Obituary [Bed82, Ben95, Mea85, Nor82, Ter37, Tur84]. Object [Car17]. objective [Dei91]. Objects [Par13, Sal09, Alb09]. Obscurity [Alt07]. Observaciones [Gre87]. Observation [Mon05, Pom11, DL11, Hua22, Not21, Bou13]. Observational [Mea70]. Observations [BS21, Bro77, GG11a, Jon06, SC15, McK57, Nak84, Tho54, Wil57, Wil65, Woo53, OO51]. observatoires [LS15a]. Observatories [Bec13a, Glu23b, PdC10]. Observatory [Det18, SM82, Cha83a, Mel53, Pri52, Aub23, Bry90, Dav84, Mac18, Ols71, Rap15b, Str94, Sza72]. Observer [Bec13a, PdC10]. Observing [SS13, Ste10b, Bec13b]. obstetric [Far83, Tee84]. Obstetrics [Bar09]. Occam [Gro74c]. Occident [vD11]. Occultism [Lyn14, Lac11]. ocean [Lin22]. Oceanic [Mil01]. oceanography [Jon83, Mil91b]. octave [Coh87]. Ocular [Str14, Wil11b]. Oddi [Hei14, Mar11a]. Odling [TW56]. Oeuvres [Lov12]. Off [Tan80b]. Off-beat [Tan80b]. Offal [Edw04]. Offering [Gre22]. office [Mil88a]. Official [Kle12b, Kon12]. Officials [Kle12c]. Oil [Har71, Wil13, Ell83, Luc08]. Old [Rus04, Att85, Rei89, SHM13]. oldest [Bur12c]. Oldroyd [Ano15]. Oliver [Lat15, Yav93]. Omens [Bsc14b, HNS11]. omnimoda [Ges23]. Once [Dij04]. One [Bel99, Cha17, Edw72a, Fos37, Gau14, Coh10a, Hop59a, Rab77, Cha19, Sch18b]. Oneworld [Kid13]. Ongoingness [San17]. Online [Ano11e, Ano12b, Ano12c, Ano12d, Ano12e, Ano13c, Ano13d, Ano13e, Ano14c, Ano14d, Ano14b]. Onnes [Pap09, vD08]. Onomancy [Jus11, Jus12]. Ontology [Cro09b, Wri16]. Oom [CS12]. Open [Byf19, Hon95]. Opera [For72a]. Ophthalmologist [Bel99]. opinion [Ecc80, Osp80, Rob54a]. Opistophone [Kvi22]. Oppenheimer [Tho11]. opponents [Bay38b, Bay38a, Bay39a, Bay39b]. opposition [Far83, Jon86, Tat97, Tee84]. Optical [Gou23, Tur00, Wil60c, Cha53a, Hal55, Pat38]. Optician [McC94]. Opticks [Gou23]. optico [Duc17]. optico-chemicus [Duc17]. Optics [Bur10b, Dij04, Gou22, Hen14, Wil01, Dar12, BB95]. orbit [Ait78]. orbits [Bra82a, Bra82b, Pou91]. Orcades [Smi87]. Orchid [Sht17]. Order [Fit17, Fle97, Sve19, Tac96]. ordered [Sha19b, Tan80b]. Ordering [Ste09b]. ordinary [Gre82]. ordnance [Mil88a]. Oreskes [Lin22]. Oresme [Bur14c, Bus10]. Orfila [BS12]. Organ [Tre15, Ham12]. Organic [Bow17, Jac15a, RN15, Kvi22, Tom10]. organique [Sim13, Tom10]. Organisation [Fau05, Bro79a]. organisms [Ego19, Ver94]. Organization [Fau05]. Orient [vD11]. Oriente [Ano06e]. Origin [Dar13, DC09b, KL07, Ari81, Bla19, Bur12c, Dea91a, Flo60, For66a, Hir07, Hut73, Mil91b, Tay42, Hon84, Dar13, DC09b]. origine [Hir07, Sch13a]. origines [Muc98, Sim13, Tom10]. Origins [Boa08, Boe14, Del11a, Fer02, Goo10, Jac14, Joh15a, Mat11, McC05, McE10a, Rus63a, Sta19, Tom09, BS65, Byr81, Cro61, Eme88, Gre81, Gre82, Hen87, Mac18, Mar76, Nye11, Pal21, Par39, Spe16, Tom10, Tom12, VB11]. Orinoco [Dau13]. Orinoko [Dau13]. Orion [Sie09, Sie09]. Ornithology [Far18]. Ornithorhynchus [Jen16]. Orrery [Mil74]. Ørsted [Chr95]. Orta [Win16]. Orthodoxy [Vic04]. Osstwald [Hak92]. Ostentation [dC09a]. Other [Ano09g, Bey14, Hen00, Kin04, Ren11, de 64, Cha68, Eyl63, Fra41, Her18, Lat15, Mar11c, PB45, PB47, PB48, Pia21, Pri54, Ree81, TH74, ZA16]. Ottaviani [Gro14]. Ottoman [Det18]. Ottomans [Bil16, SM15]. Oughtred [Ste00a]. our [Aga17b, Can15, Far19, Her17, Kli15]. output [GG94]. Outstanding [Jon13, KL09]. Outward [Mar16a]. oval [Ait78]. overlooked [Ari74]. overseas [dBP86]. Owen [MIL15, Qui12, Cos09, Qui13, Rup09, Slo03]. Oxford [How19, Jon13, Kid13, Kin04, Row18, Sha19b, Wil23, Cro03, Fox01, Fra36, HY87, O’B65b, Rob55b, Sve19, Tac96, Tay52b, Tur53]. Oxygen [Ham03, Le 72, Le 74].

P [Bez75, Dun13, Lus14, Tau14]. P. [Bow13, Sme54a, Wil91b, Wri16]. P.-J. [Sme54a]. P.R.S.E [Eyl63]. PA [Gol19, Pro18, Wol13a]. Pacific [Mil91b]. pacifists [Sil18]. Packaging [Ren11]. packing [Sim92]. Padua [Und63]. Page [Ano10c, Ano10d, May21, Ano10b, Ano11a, Ano11b, Ano11c, Cra54]. pages [Ano55f]. painting [Ran63]. paintings [SH89]. pair [SL77]. palaeontological [The86]. Palaeontology [Bry10, Cha48, Ell83]. Paleontology [Bri10, Bry13]. Palestine [Bod56]. Palgrave [Coz19]. Palissy [Tho54]. palladium [Uss78]. Pallets [HBP08]. Pamphlet [RS01, Whi12]. Pancreas [Edw04]. Pantagruel [Bro71]. Pantalony [Wit13]. Paolo [Jar18a]. papal [Tar20]. Papanelopoulou [Dun13]. Paper [CECE21, Egm18, Law15, Sle17, Mac85b, Pal83, Wei19]. Paperback [Mcv21, MS22, SS23, Kri19]. Papers [Boa09, Gou23, HP03, Lee13, Mar11b, Mor09, Ano86g, Geo52, Har96, Hol76, Sce08, Sce09, Wes56a, Wes56b, Gay13a]. parabolic [Nay76]. Paracelse [Kah01]. Paracelsian [S´ 20]. Paracelsism [Boa12a]. paracelsisme [Boa12a]. Paracelsus [Kah01, Sch36]. Paradigmatic [Lat15]. paradigms [Pot22]. parallax [Het72]. parent [Tre78a]. Parentalia [Ben73]. Paris [Aub23, Bra79, Bro78b, Bro81, CP97, Cha68, Dav84, Dav95, Fau05, Gre84, Gue10, Mau90, Pan09a, San84, SP85, Sme54b, Sme57b, Wit13, Hal20]. Parisian [Bel12, Fox96]. Park [Wol13a]. Parkes [Kur97]. Parkinson [Lev94]. Parlour [Law15]. Parry [Lev94]. Parson [Mor68]. Part [Ade10, Pop12, Cha70, Jac52a, Jac52b, Bay38a, Bay39a, Bay39b, Bra82a, Bra82b, Bro69a, Bro69b, Col37a, Col37b, DK55a, DK55b, DK57, Feu71a, Feu71b, For98, Fra41, JTR94, JTR95, Klo85a, Klo85b, O’B65a, O’B65b, Old91, Old92, Old93, Old94, Par51a, Qua72, Rus59a, Rus59b, Rus63a, Rus63b, Rus63c, Sin49, Sin50a, Sin50b, Ste69a, Ste69b, Sto66, Sto68, Stu82, Stu89, Swe74, Tha66, WN16, Wil95a, Wil95b, Woo71b, Woo71c]. particle [Hen80, Zac72]. particles [Cre66, Tre74]. particular [SR73b]. particularly [Kem68]. Partington [Bar11]. partnership [Chr95]. parts [Fra41, Bay38b]. Passage [Swe74]. passion [Han21]. Past [Buc08]. Pasteur [Han21, Wil55c]. Pasts [Goe18, Eff21]. patent [Jon19, Woo71a, Woo71c]. Path [Kol99, Ait78, Maf10]. pathfinder [Arm54]. Pathology [Leu08, Spr42]. paths [Gou22]. Patriarch [Tak11]. Patrizi [Hen79, Tod82]. Patron [Cro09a, Rob53b]. Patterns [Mil12, Die21, McE10b]. Paul [Bry13, Car97b, Lat15, Wil13, Ber12, Hen00]. Paula [Pop12]. Pauling [Bur12d]. Paviland [Nor42]. Pavy [Tat97]. Payment [Bik75]. Payson [Gos15, Har12]. pbk [Mil13, Swe03]. pdf [Ano10b]. peace [Kra17, Smi23a]. Peacock [Rob54b]. Pearson [Mac79, Mag98, Nor75]. pedagogy [Tay08]. Pedersen [Bec13a]. Peirce [GG97, GG12]. Peiresc [Sie09]. Peirescs [Sie09]. Peking [Cha84, Li12]. Pell [vM86]. Pender [IM89]. Pendulum [Whi12, Hal78, Tir23]. pendulums [Erl94, Mac76]. Penicillin [Edw09b, Coz19]. Penn [Wol13a]. Pennant [Mat54]. Pennsylvania [Lau21]. pension [Woo97]. Pensions [Cro10]. Penzance [SR71b]. People [Gue09, Li12]. Peoples [Whi13]. Pepys [Kur03, Dav57]. Perception [VT17]. perceptions [Rus21]. Percival [Str94]. Perdiguero [Dun13]. perfection [BS93]. performance [Leg17, VPW16]. Peril [Yi13, Car10]. Perils [Mea09, San00]. Period [Alv07, God09, Dun21, Eyl63, GG76, Kre38, MG10, Nay77, O’B65a, Sta88b]. Periodic [Lee13, Raw74, Sce09]. periodical [Zia94]. periodicals [SP85]. Periphery [Dun13, PNGP09, Tod93]. perpetuation [Bro78a]. perseth [Dek08]. persistence [Bra82a, Bra82b]. personal [Bar95, Smi77]. personality [Han21]. Perspective [Bin12, Sak12, And07, App96, Byf19, Far19, Nyh09, RH85]. Perspectives [Car18, Gue13, Low03, Tau14, Vet10, DS16, Fei03, Mor10, Vic04]. perspicuitatis [Ray82]. pertaining [Coh36]. Peru [Gre87]. Pest [Sto13]. pestilence [Fin22]. Peter [Bec13a, Boa19, Bsc14b, Lai04, Luc19, Gib51b, Rya91, Tur53]. Petroleum [McB91]. Petrov [Bec09]. Petworth [McC83]. Peurbachii [Swe21]. PGS [Alv07]. pharmaceutical [Rus78]. pharmaciens [Sim13, Tom10]. pharmacists [Tom10]. Pharmacology [Por11]. Pharmacopoeias [May21]. phases [PM39]. Phenetics [Hul85]. phenomena [GG88]. phenomenon [Tre80a]. Philadelphia [Lau21, Ink78]. Philanthropy [Mor86, Wat92]. Philip [Gol19, Kin04]. Philippines [And14]. Phillips [DE50]. Philomathique [Duv54]. Philosopher [Bsc12, Bur00, Cha09, Her17, SS99, App90, Can15, Bur98, Mit79]. Philosophia [McC39c]. Philosophiae} [SB13]. Philosophical [Bos03, BM17, Bsc12, Gro72a, Gro74a, Rob53a, Ape20, Gro74b, SR74a, Stu82, Stu89, Geo52]. Philosophie [Hen86, Sme67]. philosophique [Muc98]. philosophus [Duc17]. Philosophy [Akt09, Ano57b, Bsc12, Car13, Cha09, Cha12a, Cha14, Dom09, Duc18, Ell00, Kre11, Lee13, Lev14b, Nic11a, Pal13, Pat17, Sch18b, Slo03, Ste09a, Wol13a, Att85, Bro81, Cha75, Chr95, Cle90, Cop87, Gas10, Gol19, Har57, Hen86, Hes52, Hir08, Mas53b, Mil83, New96, Pal83, Pum87, Rab77, Ree80, STDH22, Sce08, Sha19b, Sha73, Sla19, Smi76, Wil91b, Wil09, Yeo79, Col54b, Moo87]. phlebitis [Fra39a]. Phlogistic [Boa08, Old74]. phlogiston [All94, Nea51a, PM37, PM38a, PM38b, PM39, Per83, Boa08]. phosgene [Dob45]. photoelectric [Kat06]. photographer [MSD23]. photographic [Woo71a]. Photography [Edw09a, Hol96, Woo71b, Woo71c]. photometric [Hen92]. photometrische [Hen92]. Photons. [Hei19]. Phrenological [Sha75a]. Phrenology [Can75b, Can75a]. Physical [Ade10, Ano08j, Bin12, Bow74, Dij04, Duc88, Rei04, Ste76, Mac19, Mil91b, Nye89, Ohn23]. Physical-Technical [Ste76]. physician [Bay38b, Bay38a, Bay39a, Bay39b, Don50]. physicians [Gue23]. Physicist [Her17, Ano75k, Can15, Gre05, Hen09, Sau76, Sne84, Wie23]. physicists [Hoc83]. Physick [Sve19]. Physico [CB22, Tho16, Col64, Sch13c]. physico-chemical [Col64]. Physico-mathematics [CB22, Tho16, Sch13c]. Physics [Ano06a, Asp69, Bad60, Bon14, CR02, Fol14b, Gar10, HW87, Koj02, Kra11a, Mil04b, Mit12, vS15, Cah90, CR96, Cha19, GG86, GG84a, GM94, Goo80, HL91, Hen23, Hen86, Hil91, HW90, Kei84, Kim95, Kra11b, Mel09, Pal83, PK87, SL77, Smi76, Del11b, For71a]. Physik [Hen86]. Physikalisch [Ste76]. Physikalisch-Technischen [Ste76]. Physiker [Hen09, Sau76]. physiological [Duc88, MT72]. Physiologist [Ric86]. physiologists [Fra38]. Physiology [vS15, Beg21, Ric86, Ste80, Tho51b, Ken52]. physique [Ber01]. Piazzi [SH89]. Pickering [Sus16, Plo78, Str94]. Picture [Gil02, Top80]. pictures [Kus10]. Picturing [Rit11]. Pieces [Mar11b]. Pierre [Bra94, Leh12a, Mar76, Nea51b]. Pierre-Joseph [Leh12a]. Pietro [War07a]. Piezoelectricity [Gar10]. Pigott [Mel53]. Pilâtre [Sme55c]. pile [Rus59a]. pilgrimage [Arm53]. Pingré [Arm53]. Pinhole [HZ07]. pioneer [Ari79, Tre80b, Whi82]. pioneering [LA22]. pipe [Old72a]. Pisum [Mac85b]. Pitcairne [Lin63]. Pitfalls [Ano06i, Ohn23]. Pittsburgh [Gol19, May21, Pro18]. Place [Ano09h, Dun94, Jen16, Sch21, Whi13]. plagiary [Mad61b]. Plague [Flo96, Fin22]. Plaister [Hal20]. Plana [TT99]. Planck [RB01]. plane [Erl92]. planes [Erl94, Mil93]. Planetarium [Bur12b]. planetarum [Swe21]. Planetary [Jon06, Ait57, Ait58a, Ait58b]. planets [Pia21]. Planning [Law08]. plans [Bra94, Tou89]. Plant [Beg21, Fit17, Hol17b, Rit62]. Plantarum [Pre16, Ray82]. Plants [Egm16, Kus18, Whi13]. Plata [dA21]. Plate [Zev00, YO89]. plates [Dar85, Gie16]. platina [Uss80, Val96]. platinum [Cha71, Cha72b]. Plato [Ung02, Ung02]. platypus [Jen16]. Playfair [Cha54b, Dea83]. Pliny [Doo10, Lov13, LS15b]. plts [Bow13, Gue23, How19, Kid13, Kri19, Luc19, Mcv21, Qui13, Sla19, Van19]. Plücker [Ecc80, Ecc80]. Plücker’s [Wie23]. pluralism [KLW06, Akt09]. plurality [Bro77, McC36b, Pia21]. plus [Bur12c]. Pneumatic [Tom09]. Poet [Mil11c, Sok96, Her18, Lev78]. Poincaré [IM89]. Poinsinet [LS15b]. Point [vD08]. poison [Hed22]. Poisons [BS12, Ear61, Ear63]. Polanyi [Boe14, Nye11]. Polarity [Fra09]. polemical [Nut21]. Police [Ver13, Wak09]. policies [GG75a]. Policy [Joh15b, SP11, Clo54, HL91, Jar18b, Jon86]. Polish [SJ78]. Political [Hal14a, Mil11a, War02, GL11, Ink88, Mac85a, McC75, Bas15]. Politics [Bas15, Cro16, Goo10, Kai10, Lus14, Maz12, Row18, Sei13, Tan14, Bad09, Edw10, Ego19, Gie19, Hal14a, HL91, Sea79, Wal04]. politique [Muc98]. Polity [Kid13]. polls [Lus18]. polygons [For70a]. polyhedra [San84]. Polymath [Vic12]. Polytechnic [Mor88, Spr77]. polytechnics [Kni75]. polytechnique [Sme54b, Bra75, Bra76]. Ponce [Bla19]. Pondering [Yos99, Hen21b]. Pontifex [Ozd14]. Ponts [Mac20a]. Pontypool [Gib53b]. Pope [Not17]. Popper [GG02]. Popular [Bow16, Bur09, KN13, Li12, SP85, Zui04, Bar98, Smi23b]. Popularisation [Dun13]. Popularization [NG09, Bar98, Bow09b]. Popularizers [Jam12, Lig10]. Popularizing [Aub23, PNGP09, Dun13]. population [Nor75, Sok74]. Portable [Kin04, Mil86, FW85, Goo67]. Porter [Gro72b]. portrait [Di 17, Rob62]. Portraiture [Hen02, Mad59, Mad88b]. Portugal [Car05, CS12, Mal20b, RS17, dC09a]. Portuguese [Alv07, Ama06]. Portuondo [Sla19]. posed [Tre91]. Positano [Jam78]. position [Hen78]. positional [Cas15]. Positivism [BV99]. Possessing [Mac10, Fra14a]. possession [MJ89, Ran63]. possible [Gil76]. post [HL91]. post-war [HL91]. Postmodern [Por01]. potassium [PB45, PB47, PB48]. potassium- [PB48]. potassium-ammonium [PB45, PB47, PB48]. potentiometer [Rut39]. potter [Tho54]. Poudre [Bre93]. poudres [Sme56c]. Pouillet [Kid81]. Pound [SC04]. Powder [Bre93]. Powell [McC14]. Power [Ban14, Fin12, And10, OH97, Pat38, Wil92]. pp [Bil18, Boa19, Bow13, Bri13, Byf19, Car18, Coo13, Coz19, Edd13, Egm18, Fra13, Gat18, Gie16, Gol19, Gue23, Hal18, Hay18, Hei19, Hig17a, HA18, How19, Jar18a, Jon13, Kid13, Kin19b, Kra13, Kri19, Kus18, Lai04, Lan13, Lau21, Law13, Lin22, Lov13, Luc19, Mar21a, May21, Mcv21, Mel21, Mil13, MS22, Nut21, Par13, Pro18, Qui13, Ray18, Reg19, Ren18, Row18, Sch18b, SS23, Sei13, Sha19a, Sha19b, Sla19, Smi18, Sta19, Sta21, Van19, Ver13, Vic04, Wei19, Wil23, Wit13, Wol13a, Yi13]. Practical [Fin12, Bra76, Jar18b, Mac20a]. Practice [Gau12b, Gay00, Mar09, Mit12, Par03, Ham12, Mal20b, Not21, OH97, Rib20, Wak09, vH94, Ver13, Tre15]. Practices [Bad11, Hui11, Glu23b, Jal23, Tom10]. practitioners [Joh91]. Prague [Kvi22, Sím93]. pratiques [Sim13, Tom10]. prayer [Bur09]. pre [Hof36, Kre38, Pay18]. pre-Galvanian [Hof36]. pre-Raphaelites [Pay18]. pre-war [Kre38]. precisio [Ges23]. Precision [BD00, Akk23, Aub23, Ges23, Glu23a, Glu23b, HT23, Jar23, Kre23, Ohn23, Tir23]. Preclassical [Pis13, Sch08a, Sch08b]. preconceptions [Rus21]. precursors [Bay38b, Bay38a, Bay39a, Bay39b, Wil55c]. predecessors [Cop80]. Prediction [Sce94, Kre23, Sta88a]. Predictions [How02]. predilections [Rus21]. preface [Dea75]. preliminary [FS21, Ink77]. Prelude [Gib52b, Kid81, Stu82]. Preludes [Ave20]. premature [Tre78a, Tre80a]. preparation [GG97]. presence [Swi62]. presence-and-absence [Swi62]. present [Lom11]. Presents [Goe18]. Preserving [SK11]. presidential [Lus18]. Press [Bil18, Bil16, Bow13, Bri13, Coo13, Egm18, Fra13, Gat18, Gie16, Gol19, Hay18, HA18, How19, Kin04, Kin19b, Kri19, Lan13, Lau21, Law13, Lin22, Lov13, Luc19, Mar21a, May21, Mel21, Mil13, Par13, Por16, Pro18, Qui13, Ray18, Reg19, Ren18, Row18, Rus04, SS23, Sei13, Sha19b, Sla19, Smi18, Sta19, Sta21, Van19, Ver13, Wei19, Wil23, Wol13a, Yi13, Mcc79, Zia94]. Pressure [Cha17, Sch18b]. Prestige [Cro96]. Preston [WO87]. Preußen [Dau13]. Preventive [Mor86]. Previously [Eis05]. Prévost [Hen82]. prewar [Str93]. Price [SP11, Gre77, Tur84]. Priestley [Jon99, Ano07h, Arm56c, Har41, Sch54, Sch57]. Priestleys [Mac20a]. primacy [Whi86a]. primates [VT17]. Primer [BM17]. Primeval [KL07]. Primeval-Atom [KL07]. Primitifs [Bur12c]. primitive [Bur12c]. primus [AVJ^{+}06, Bur11b]. Princeton [Ren18, Sta21]. ‘Principe [Sch13a]. Principia [SB13, FS21, Gui15, Mel88, Nau19, Pou01, Sno01, Wan18a]. principle [Thi71]. Principles [CP02, Bra90, For71c, GE19]. Pringle [Sin49, Sin50a, Sin50b]. Print [Arn09]. Printed [Ber01, PK87, Wel83, Win66]. Printing [Rap12, NG97]. prior [Ink88, Mau69, Tod93]. Prisca [Mol96]. Prism [Buc08]. Private [Are10, RS71, Vic12, LGKW08, McC83]. Prize [Ano02a, Ano03a, Ano12f, Ano18, CECE21, IM17, Cha88, Ano19, IM15, Ano16b]. Prizes [BG99]. pro [McC39a]. probabilism [Lau66]. probability [Gow82]. probably [Sme56c]. Problem [Old99a, Ait64b, CK89, Cha71, For61, Gui95, Hon87, McD72, Old80, Qua68, Qua70, Qua72, Ros21, Sch80, Sok94, Wal39]. Problems [NGDP75, Par16, Tre91]. procédés [Bre93]. Procedure [Bec09].

Proceedings [Ano57a, McC94]. Process [Fin09, Dav98, Lan83, McB91, Raw74, Tei75]. Processes [Nay77, Bre93, Cha84]. Produced [Maz12]. Production [vHvG99, Die12, Lan83, Ste36b, Str93]. professeur [Aug04]. professeur-consultant [Aug04]. profession [Rob62]. Professional [Har06, Mac79, Ols71]. professions [Lun90]. Professor [Col52b, Col52c, Aug04, Col52d, Col53, Kon12]. professorial [Gra70]. profiles [Hen92]. progetto [Gui11a]. Program [Sau14, Aug04]. Programme [NG09, Str97]. programmes [Mor90]. Programming [Big21]. progress [BM94, Cha53b, Cha54a, Cha54b, Cha70, Luc19, Rob54b, Smi87, Spe16]. Progressive [Tom12, Joh15a]. Progymnasmata [Swe09]. Prohibition [Pag03]. Project [Gay13a, Hen00, Gui11a, Low90, Ros98]. projectile [Nay83]. projectiles [Bla88]. projection [Dra86, Hil84]. Prometheus [Ink88]. Promise [Yi13, Car10]. Promises [Glu23a, Ohn23]. Promotion [Far72, Por81]. prompt [Roe90]. Proof [Lat15, Moz04, Nau14, Pou91]. proofs [Cra54]. propagation [Ols70]. Proper [Vic04]. Properties [Bos03, McD72, BM94]. prophet [Dav92]. prophets [Luc19]. proportion [Fle74]. Proportions [UB15, Mau69, Par51b, Par53a]. proposals [BS65, FB22, Hop58]. proposed [Mor86]. Proposition [Bla88]. Pros [Old99b, Kau79]. Prospective [Pis14a]. Prospectivea [Cap12]. prostheses [Kin19a]. Protestant [Kem68, Mas53a, Mas53b]. Proto [LR00]. Proto-elements [LR00]. Protogaea [Ant10]. Proust [Rob10]. Prout [Bro69a, Bro69b, Ham03]. Providence [Bra02]. Provinces [Ell00]. provincial [Kit82]. provision [Cab73]. PRS [Old97]. Prussia [Dau13]. Prussian [Col39, Fle97, Kle12b, Kle12c]. Pseudodoxia [dVH67]. psyche [Coh10b]. Psychical [Lyn14, Lac11]. psychologie [Muc98]. Psychology [BD00, Hay18, Lin16, Ste80, VB11, Jac14]. psychophysical [Kvi22]. Ptolemy [Bur00, BJ00, Bur98, Jon06, Swe03]. Pub [Edd13]. Public [Del13, Ell00, End17, Kim06, LMBM06, Mor01, Ols71, Dea73, Flo96, How19, RS17, Tur53, Sin50a]. Publication [RS01, Bec16, GG97, Old97, Tre78b]. publications [Sme56c, WN16]. publish [Cha88]. published [BS65, Geo52, Jay78, Mad61a, Row50a]. publisher [Ecc76]. Publishers [Byf19, Mil13]. Pulkovo [Sza72]. pulse [Hua22]. pump [vH94]. Pumpelly [OJY96]. Pumps [Bru11]. Punnett [Rus22]. Pure [BG99]. purine [Hen78]. Puritan [Kem68]. Purkyne [Kvi22]. purpose [Lev21]. purposes [Bar98]. Pursuing [Kat16, Wei15]. Pursuit [BD00, Ish14, Mel21, Reg19]. Putting [Mal20b]. pyrometers [Cha72b]. pyrometry [Cha71]. Pyrotechnic [Lan13, Wer10]. Pythagoras [Pri10, Qua74].

Qibla [Bur09, Bur09]. Qing [Ma21]. quadrant [Dek95, Dek08]. Quadrants [AC99, Kin04]. quadrature [vM86]. Quaestiones [SBT10, Bur14b]. Qualitative [De 17]. Qualities [Cec11]. quanta [GM94, Hei19]. Quantification [HT23, Mae05]. quantitatibus [Cif86]. Quantitative [Bun77, De 17]. quantities [Cha19]. Quantum [Del09a, Par03, Yea11, GG88, Gre05, Hen80, Kat06, Kra21, Pot22, Sce94]. Quare [Goo67]. quartz [Kat16]. queen [Don50]. Queenwood [Tho55]. Quest [Ano12a, Del12, Edd13, Moz04, DeK83, Hen21a, Rat09]. question [Duc97, Rom06]. Questionable [Gay13a]. Questiones [Bus10, Bur14c]. questioning [Glu23a]. questions [LL37, Val09]. Quetelet [Pel36]. quiet [Aga17b]. quinquina [Tom01].

R [Boa19, Kin19b, Lee13, Pal12, Row18]. R. [Ben95, Har96, Hil91, Mac82, Pan09b, Rob55a]. rôle [Old72a]. Rabelais [Bro71]. Rabi [For98]. Race [Goo10, Ras17, Smi13, SP92, Uch22]. racemate [PB45, PB47, PB48]. Radiant [Hil99, CK89, Cor36, Cor38a, Cor38b, Ols69]. radiation [CSVK80, Kan70, Kan71, Ren18, Sch89, Ren18]. radiations [GM94]. Radio [Cha12b, And09b, Hon96]. Radioisotope [MSD23]. Radium [Bad11, Ren11, Ren18]. Radius [Wes10, Roc81]. Raffael [Vic12]. Ragland [Gue23]. Railroad [Adl17, Vet04a, Vet04b, Ink72a]. Rainbow [Buc08]. Ralegh [Sok74]. Ramchandra [HR89]. ramifications [Wil65]. Ramsay [OM95]. Ramus [War14, Gou10, Péo10]. random [Ohn23]. Ranelagh [Pal23]. Rankine [Hut73, Hut73]. Raphael [Gat18, Hei14, Gay10, Mar11a, OJY96]. Raphaelites [Pay18]. Raphson [Cop80]. Raspe [Rob55a]. Ratcliff [Edd13]. rational [Tan77]. Rationale [Ste76]. rationality [HR89]. rationalium [Tan77]. Ratios [UB15, Mac72a]. Raw [Edw04]. Ray [Pre16, Wil61]. Rayleigh [Ku09]. Rays [Bon14]. Re [CZ19, Clo55, Hop58, Woo71b]. re-assessment [Woo71b]. re-determination [Hop58]. re-examination [Clo55]. Re-examining [CZ19]. Reaction [Car05, BL21, Rei81]. reactionary [Bra88]. reactions [Ber97]. Read [Sar02]. Reade [Woo71b, Woo71c]. reader [PW08, Bar11]. Reading [Nor16, Rap17, Ste19, Fru16, Pot22, Gat18]. Reaktion [Kus18, MS22, Sha19a]. Reaktionen [Ber97]. Real [Mac76, dCA36]. Reality [Bea12, Joh14, Ree14a, Roc10]. realm [Tan80a]. Realms [Reg09]. reason [Top80]. reasoning [Ano85j, Rab77]. Reasons [Pro18]. Reassessing [Mor90]. Reassessment [Mil02, Dek90]. Réaumur [Gau12a]. rebellion [Sve19]. received [Ano68a, Ano68b, Ano69b, Ano69c, Ano69d, Ano69e, Ano70c, Ano70d, Ano70e, Ano70f, Ano71b, Ano71c, Ano71d, Ano71e, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano74e, Ano74f, Ano74g, Ano74h]. Recently [MVK02, OM95]. Reception [Bar11, Lal12, Lov13, Mar13, Ste11b, Wes10, Doo10, Ebe39, Hoc83, Kah01, Ole23, PW08, Rei81, Sne75, Wil60b, Woo87, Zia94, Kah01]. Receptor [Mae05]. recherche [Aug04]. Recipes [Wei19]. Reciprocal [Par53a, Par51b]. Reckoning [Kin04]. Recognition [Har06, Lus18]. recoil [Tre80a]. recollections [Gra72b, Ken52]. reconciliation [SS58]. Reconsideration [Kra09b, Fli59, Mor88, Win72]. Reconstructing [Hal13, Tar20]. Reconstruction [Ung02]. records [Cha68]. Recovery [Pom11]. red [For61, Nor42]. redefinition [Cle90]. Redi [Dur10]. rediscovered [Bar95, Fis36]. Rediscovery [War14, Gou10, Kot79]. redshift [Hen92, Hen96]. Redux [Nic11a]. réel [Duc13]. Reeves [Wil11a]. reference [Sch08c, SR73b, Und63, Wal76, Wil55a]. references [OG68]. Reflecting [Zui04]. Reflections [La 11, Por16, Rei14, Sla03, Cro48, Jon50, RD16, Ste11a]. Reform [Sto13, Not17, Rob55b, Sme54a]. reformation [Lau21, Cha12a, Mas53a, Mas53b]. reformer [Kur97]. Refraction [Péo10, Gou22, Hor00]. Refraktion [Hor00]. Refrigeration [Pap09]. Refugee [Rei81, Hoc83]. refutation [Ari73b, Qua68, Qua70, vM86]. Regarded [Del11a]. regarding [PB45, PB47, PB48]. Regency [Can10]. Regia [Gam21]. régie [Sme56c]. régime [Aug04]. Regimens [Cor11, Nut14]. regiment [Tan80b]. Reginald [Rus22]. Regiomontanus [Ray15]. Region [Nor01, Old99a]. regional [Osp80]. Regnault [Mit12]. Regress’ [Kim12]. regular [San84]. Regulation [Coz15, BJ13, DS97, GL11, DS97]. Rehuel [Sta88b]. Reichsanstalt [Ste76]. Reid [Duc18]. rejection [Ait78, Coc83]. Related [Gou23]. relating [DK54c, Wes56a, Wes56b]. relation [Mas53a, Mas53b]. Relations [Cah10, Hul02, Pre08, Bun77, Ecc80, Eyl63, FLW11, Jac52a, Jac52b, Per86, Ben13b]. Relationship [Wes10, Hac79a, Hen86, Sha73, Wei51]. relative [Wil57]. relatives [Pat36a]. Relativity [Del11b, Del11a, Pal83, Del11a, Feu71a, Feu71b, GM94, Gro74c, Lal12, Mel09, Mil81a, Zia94, Eis05]. relevance [Le 74, The86]. Relief [Tou89]. Religion [Cha14, Dea75, Fli08, Goo10, Gov21, Hun23, Mos14a, Ram10, Fei19, Gas10, LS23, Sha12, dA21]. Religious [Ben13a, Far83, Kem68, Tee84, Pav17]. Remains [Mar11b, Tha66]. Remarkable [Ano06j]. Remedies [Kah10]. Remote [McC13, CC11]. Renaissance [Cha12a, Hei14, Kin04, Mel21, War14, Hed22, Hir08, Ano75l, Blu13, Boa12a, Cha12a, Dup11, Fin22, Gou10, Hut77, Mac23, Mar11a, Ree76, Ste09b, Val09, Wei51, Hud03, Pat17]. Renate [GG14]. René [Gen22, Sch19]. René-Just [Gen22]. Renée [Gat18]. renown [WO87]. Repeal [Pag03]. repercussions [Roe90]. repertory [Bro90]. Repetitive [Fle97]. Replication [Cob02, URRF05]. Reply [Far84, Roe90]. Report [Ano57b, Bro36a, But57, Bea76, Duv52, OM95, SBB^{+}75, Sok94]. reported [Sme57a]. reports [SR73b, Sto66, Sto68, WN16]. Repository [Ape20]. Representation [Cob02, Joh18, MG10, Mel10]. representations [Ekh10, Low96]. representatives [Sme57b]. reprints [Edd13]. reproductions [For70b]. Republic [Mei79, Rya23]. repugnant [BJ96]. repulsion [Dun67]. reputation [Ben09, Ben09]. Research [Ama06, And14, Bad11, Car05, Coo11, Kur03, Lyn14, Mor86, PP77, Ram15, Ram10, Aug04, Bas95, Bra95, CSVK80, Dav95, Fle97, Jon85, Klo85a, Klo85b, Lac11, Mor90, Nye89, Olb91, Sme57a, Tom14]. researches [Col37a, Col37b]. réseau [CP97]. réseaux [Muc98]. Resh [Tak11]. Resh-‘Aina [Tak11]. reshape [Far19]. resistance [Bar73, Gam21, Hop57b, Hop58, Hop59b]. resonance [For98]. Resonator [Ku09]. Resources [Pop12, Swa89]. respect [Erl92]. respiration [Sch39]. Response [Bow85, Lev09a]. Restaging [URRF05]. restatements [AM76]. restless [Die16b]. restoration [Pum87]. Restraints [Bar14, Bor10]. restricted [Uss80]. Resultados [Gre87]. Resurrecting [Hen18]. Rethinking [Kin19a]. retrieval [Whi82]. Retrospect [MW13]. Retrospective [CP02]. return [CB22]. Rev. [Cun04, For71b]. revealed [Dei91]. Revenge [Pri10]. reverend [Har66]. Review [Adl17, Aga17b, Aga17a, Akt09, And15, Ank15, Ano51a, Ano53a, Ano54b, Ano55d, Ano58b, Ano59b, Ano63a, Ano63b, Ano64a, Ano75i, Ano84i, Ano99e, Ano01e, Ano02b, Ano03e, Ano04e, Ano04f, Ano05f, Ano05g, Ano07e, Ano12a, Ban14, Bar11, Bar14, Bas15, Bec13b, Bec13a, Bec14, Ben14, Ben13b, Bey14, Bil18, Bil16, Bin12, Blu13, Boa12b, Boa19, Boe14, Bou13, Bow13, Bow17, BB13, Bri13, Bry13, Bsc12, Bsc14a, Bsc14b, Bur17, Bur11a, Bur11b, Bur14b, Bur14a, Bur14c, Bur15, Byf19, Car17, Car18, Cha12a, Cha13, Cha12b, Cha14, Chu13, Coo13, Coz15, Coz19, Dau14b, Dau14a, Del09a, Del11b, Del11a, Del12, Dij13, Dom09, Duc13, Duc14, Dun13, Edd13, Egm18, End17, Eve15, Fit17, Fol14a, For71a, Fra13, Fra14a, Fra14b, Gat18, Gau14, Gie16]. Review [Gol19, Gos15, GG11a, GG13b, GG13a, Gro14, Gue13, Gue23, Hah15, Hal18, Hal14b, Hay18, Hei14, Hei19, Hen17, Hen00, Hen14, Her17, Hig17a, HA15, HA18, Hol17a, How19, Hud03, Hun14, Jac14, Jam12, Jar18a, Joh14, Joh15a, Jon99, Jon13, Kid13, Kid14, Kin19b, Kni13, Kni17, Köb17, Kra13, Kri19, Kus18, Lai04, Lan13, Lat14, Lat15, Lau21, Law13, Lee13, Leg17, Lin22, Lov13, Low03, Luc19, Lus14, Lyn13, Lyn14, Lyn15, Mar15, Mar21a, Mar11b, May21, Maz10, McC13, McC14, Mcv21, Mil12, Mil16, MS22, Mos14a, Mos14b, Moz04, Nic13, Nic14, Nor12, Nut14, Nut21, Nyh16, Pal12, Par13, Pat17, Pis13, Pis14b, Pis14a, Por16, Pre16, Pro18, Qui12, Qui13, Ray18, Ree14b, Ree14a, Reg19, Rei14, Rob10, Row18, Sak12, San00]. Review [San17, Sav15, Sch13b, Sch18b, SS23, Sei13, Sha19a, Sim13, Sla19, Sla03, Smi15, Smi18, Sta13, Sta21, Str14, Sus15, Sus16, Swe03, Swe21, Tat17, Tau14, Tho16, Tho11, Tre15, Van19, Ver13, Vic00, Vic04, Vic12, Wad15, Wal14, War14, Wee16, Wei19, Wer14, Wil11a, Wil15, Wil13, Wil23, Wit13, Wol13a, Woo15, Yi13, Yi14, YT13, Zai17, Ano74i, Ano75g, Ano75h, Ano75j, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano77f, Ano77g, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano79g, Ano79h, Ano79i, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano83h, Ano83i, Ano83j, Ano84g, Ano84h]. review [Ano84j, Ano85g, Ano85h, Ano85i, Ano86i, Ano86j, Ano88g, Ano88h, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, Ano90h, Ano90i, Ano90j, Ano90k, Ano90l, Ano91j, Ano91k, Ano91l, Ano91m, Ano91n, Ano92m, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93i, Ano93j, Ano93k, Ano94k, Ano94l, Ano94m, Ano94n, Ano94o, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano96b, Ano96j, Ano96k, Ano96l, Ano96m, Ano96n, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano97n, Ano98e, Ano98f, Ano05e, Cor75, Far79, Gre87, Hor87, Kem68, Moo87, Pal83, Ren18, Sta19, VT17]. Reviewed [Ano09e]. Reviews [Ano36c, Ano36d, Ano36e, Ano37c, Ano37d, Ano37e, Ano37f, Ano38c, Ano38d, Ano38e, Ano39c, Ano39d, Ano39e, Ano47, Ano48, Ano49b, Ano50a, Ano50b, Ano51b, Ano51c, Ano52a, Ano52b, Ano52c, Ano53b, Ano53c, Ano53d, Ano54c, Ano55e, Ano56a, Ano56b, Ano59c, Ano60b, Ano60c, Ano60d, Ano61a, Ano61b, Ano61c, Ano61d, Ano62a, Ano62b, Ano63c, Ano64b, Ano65a, Ano65b, Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68d, Ano68e, Ano68f, Ano68g, Ano69f, Ano69g, Ano69h, Ano69i, Ano70g, Ano70h, Ano70i, Ano70j, Ano71f, Ano71g, Ano71h, Ano71i, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72n, Ano73d, Ano73e, Ano73f, Ano73g, Ano74k, Ano74l, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a]. Reviews [Ano01b, Ano01c, Ano01d, Ano02c, Ano02d, Ano02e, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08b, Ano08c, Ano08d, dVHDB57, Kin04, Mil13, Rus04, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano76a, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano78f, Ano79a, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d]. reviews [Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano98a, Ano98b, Ano98c, Ano98d, Ano05b, Ano05c]. revision [Feu76]. Revisited [Ham03, Old10, Ost92]. Revival [LS15b, Gib51b]. Revolution [Ano06l, Boa08, Dri12, Hes21, Kni14, Lal12, Mil12, She09, Wal09, Wee16, Bai93, Cop80, Dur10, Ehr95, Gre05, Kat06, Mas53a, Mas53b, McE10b, PK87, Por81, SD68, SR74b, Dur10, Aug04, Del11a, Lan83, Aug04, Por10, Mor02, Lai04, Del09a]. Revolutionary [Bra88, Bre93, Aga17b]. révolutionnaire [BV83]. révolutionnaires [Bre93]. Revolutions [Fol14b, Har74, Kra11b, Ape20, Por16, RD16]. Revue [Muc98]. Rewarding [MN13]. Rewards [Mea09]. Rhetoric [Dan10, Edw04, Vic00, Mit12]. Rhetorical [Hil84]. Ribot [Muc98]. Rich [Jac99]. Richard [Kin04, AC99, Arm58b, Col53, Del11a, Far69, Fra39b, Lew76, Pat36a, Qui12, Rup09, Sco81, Slo03, Wil55c]. Richards [Por16, Smi18]. Richardson [Kat06]. Richter [Kra13, Par51b, Par53a].

Righini [Bed82]. Rijckevorsel [dBP86]. Rinaldini [Gia23]. Río [dA21]. Rise [Ale10, Aug11, Ber12, Coo13, Sak12, Sus15, Arm50, Arm51, Bec11, Cob18, Coc56, Gib50, Gib51c, Gib51d, Gib51e, Gib55, Gro74a, Hay81, Hen18, Jar18b, Kei84, Kem68, Lun90, Nyh09, Smi23b, Tho47, Wei51]. Ritter [Chr95]. River [McC91]. Robert [Ban14, CH82, Eme88, Fra13, Mar11b, Mei21, Por16, Ran63, War14, Wit96, BG99, Boa09, Car97b, Cec11, Dea91b, Dom09, Ehr95, Eis05, Gib51f, Gib53d, Gou80, Hes17, HP03, Ink78, Jen16, KO83, Lev14b, Mad57, Mad59, Mad61a, Mad62, Mad88a, Mad88b, Mul92, New96, O’B65a, O’B65b, Ost92, Pal23, Row50a, Sok96, Swe63, SW67, Swe67, Swe72a, Swe74, Tac96, Tar20, Yi13]. Robertus [PLC48]. Robins [Bar73]. Robinson [Rob54c, Ano58a]. Robison [Yos99]. Rocha [FB22]. Rocke [Jac15a, Joh14, Joh15b, Mau15, Ram15, RN15, UB15, Wei15]. Rockingham [Arm58b]. Rocks [Old91, Old93, OH97]. rodent [Lee23]. Roderick [Sta88a]. Rodrigues [Rap15b]. Roger [Ano15]. Rogers [Bro18]. Roland [How19]. Role [Boa08, Slo03, vdS06, BM94, Cro96, Dav98, Gál88, Gou80, Hol96, IM89, KO83, Nay80, RS17, Tay08]. Roles [Pro18]. Roman [Bur12c, Ste09b, Wal12, Wie19]. Romans [Hah15, Leh12b]. Romantic [Ano05h, Bik71, Gra39, Old80, Chr95]. Romanticism [Ano04e, Bew16, Bil18]. Rome [De 05, Gui15, Mac20b, Mcv21]. Rommevaux [Duc13]. Ronald [Bsc14b]. Rond [Sch13a]. Ronde [Bre93]. room [Mer98]. Roots [Bow17, Kra00, Feu71a, Feu71b, Wei85b]. Rose [Hen00, Sve19]. Rosie [Dav90a]. Ross [Lev94, McC38, Swe74]. Rossbank [SM82]. Rossi [Bon14]. rotation [Bur65]. rotatory [Pat38]. Rothman [Reg19]. Rotula [Bry72]. Rotverschiebung [Hen92]. Roubo [Bel12]. Round [Bre93]. Rousseau [de 54]. route [Qué90]. Routledge [Mcv21]. Roux [Fol14a]. Rowe [Mil13]. Rowland [Bro15, Bro18, Bro22, Win72]. Rowley [App96]. Royal [Ano05e, Gay13a, Lai04, Nor12, Tho57, ABNW09, Bad60, Bar57, BS93, Car97a, Edw72b, Eme88, Gov21, Hac79b, Hul02, Mil11b, MJ89, Rus63c, Shu68, Spr77, Stu82, Stu89, Tre63, Tre64, Tre65b, Tre65a, Tre70]. Royale [Gue21, Sme56b, Sme56a]. Rozier [McK57, Mcc79, Nea50, Nea51a, Nea51b, Nea51c, Nea51d, Nea52, Sme55c]. Rubens [Kan70, Kan71]. Rudberg [Pri55]. Rudi [Gre77]. Rudiments [Pat37]. Rudolf [Cam55]. Rudolph [Pan09a, Wit13, Pan05]. ruins [Uch22]. Ruling [Bro15, Bro18, Lin16, Bro22, Hay18]. Rumania [Flo60]. Rumford [Ols70, Spa55]. Runge [GG14, Tob12]. Rupert [Bro09]. Rupke [Qui12]. Ruprecht [Kon12]. Rural [McC13, CC11]. Ruse [Moo87]. Rush [Bri10, Bry13]. Russell [BT55, GG74, GG75b]. Russia [Hoa55, Hon84, Kre11, Rya91, Byf19]. Russian [Bik71, Bra81, Bro95, Byf19, Cha21, Wei85b]. Rußlandreise [Hon84]. Rutherford [Hei67b, Tre74]. Rutherfords [Tre74]. Rutherfurd [Bro18]. Ruysch [Koo10, Sch13b]. Rview [Sha19b].

S [Adl97, Ban14, Bar11, Fol14a, Kid13, Kri19, Low03, Pan09b, Stu09, Tho11, LS15a]. S. [Lev78]. sa [Ben09]. Saber [Mil88b]. Sabine [Duc13]. saccharin [GL11]. Sacha [Sim13]. Sacred [Sha12, Jam78, Mos14a]. safety [Mor36, Wei50]. safety-lamp [Mor36, Wei50]. Sahl [GW11]. sail [Mas73]. Salabashev [Bec09]. Salas [Nut21]. Sale [Mea09, Eyl67]. sales [Tur67]. salpêtres [Sme56c]. salt [New38]. Saltpetre [Lop60]. Salusbury [McC37]. Salutary [Cro16, Row18]. Sam [Par13]. Samantha [Ray18]. Samuel [Bry75, de 49, Ape20, Kur97, O’B65a, O’B65b]. San [Dea93]. Sánchez [OEMB21]. sand [Tur82]. sand-glass [Tur82]. Sandifer [Dom09]. Sandra [Vic12]. sanguinis [Fra39b]. sans [New52]. Santacruz [Bla19]. Santesmases [Coz19]. Sapientia [Mol96]. Sara [Jar18a]. Sarah [Orr14]. Sasquatch [Reg09]. Satellite [Bec14, ML^{+}12]. satire [Pia21]. saturation [Dun67]. Saussure [Sme78]. savage [Uch22]. Savant [Kle12c, Sch19]. Savants [Bel12, Dav98]. Savile [War14, Gou10]. savoirs [Duc13]. Saxon [Sch90]. Scaffold [Kra21]. Scale [Wri92, Cha83b, DK54c]. scales [Ges23]. Scanning [GH08]. Scarburgh [Kee52]. Scerri [Lee13]. Schalenmodells [Zac72]. Scharlieb [Jon95]. Schechner [Jar18a]. Scheele [CK76, LS68]. Scheiner [Lat14]. schemes [Old74]. Schemmel [Pis13]. Schicksal [Hon84]. Schiller [Bik65]. Schleiermacher [Lew77]. Schlick [Hen86, Hen86]. Schmidt [DS97]. Schmitt [Cop87]. Schneider [Gre77]. scholar [Ost92]. Scholarship [Ano06i, Mos14a, Sha12]. scholastic [Hal78]. School [Boa12b, Edd08, Plo80, CP97, Dav95, Fra36, Klo85a, Klo85b, Lun90, Woo36, Ama06]. schools [Nye89]. Schrank [Pop12]. Schröder [GG75b]. Schrödinger [HY87]. Schulmann [Mar11b, Har96]. Schweber [Tho11]. Schwediauer [LS68]. ’Science [Ste00b, Adl17, Aga12, Aga19, Akt09, And14, Ano57b, Ano02a, Ano03a, Ano04e, Ano05g, Ano05h, Ano06l, Ano07g, Ano07j, Ano07k, Ano08k, Ano12f, Ano19, Are10, Ash17, Aug11, Bey10, Bey11, Bow09b, Bow16, BM17, Bro81, Bro36a, Bsc12, Bur17, Bur11b, Cah10, Car22, Car05, Cha09, Cob02, Coh11, Coo09, Cop87, Cra12, Cro03, Cro05, Cro09b, Cro10, Dan10, Del09a, Del13, Det18, Dun13, Edw09a, Egm16, Eir18, Ell00, End17, Fli08, Fox01, Fra14a, Fre11, FL07, Gas10, Gau14, RG23, Gil02, Goe18, Gra39, GG14, Hah15, Hak92, HL91, Hay64, Het83, HA15, HA18, Hol07, Hug51b, Hug52, Hug53, Hug55, Hui11, Hun23, IM15, IM17, Ink75, Ink77, Ink91, Jam12, Jar98, Joh15b, Kai10, KK02, Kid13, Kim99, Kim06, Kit82]. Science [Kle21, Kni13, Kni14, Kus18, Lai04, Lar91, Lar61, Law08, Li12, Lin10, Low03, LMBM06, Mac72b, Mar11c, Mea09, Mil01, Mil13, Mor10, Mor01, Nye18, Nyh16, Por16, Rad18, Reg09, Ren11, RS74, Sar69, Sav15, Sch18b, Sei13, SM15, Sim04, Sle17, Slo03, SD13, Smi15, SP11, Sta17, SR71b, SR72f, SR74b, Ste00b, Str14, Suz09, Tan14, Tau14, Tod93, Tom14, Uga06, Val09, Ver13, Vet04a, Vet04b, Vic00, Vic04, Wad15, War07b, WH11, Wer14, Wie19, Wit13, Woo36, dC88, ATE11, Ape20, Arm56b, Bad09, Bak64, Bak21, Bal06, Bar98, Bed93, Ber96, Ber79, Boa19, Bou17, BB13, Bra76, Bro93, Buc13, Bur90, Cab73, Cam09, Cha75, Coh10a, Cro48, Don79, Ego19, Fei03, Fle97, Flo60, Foo52, Fos37, Geo38b, GG77, Gre05, Hac79a]. science [HNS11, Hed22, Hen21a, Hen88, Het76, Hor91, Hul85, Ink72b, JL79, Joh09, Jon50, JR12, Kem68, Key90, Kin19b, Kni75, Kni09, Kri19, Kri22, LS23, Leh12b, Lei20, Lev78, Lev09a, Lig10, Lus18, Mac10, Mac73, Mac20a, Mal20a, Mal20b, MHO13, Mar21a, McE10b, Mei79, Met56, Mil81b, Mol84, Mul73, Nye11, Old80, PNGP09, Pas51, Pay18, Por81, RD16, Rob55b, RS17, Rog72, SBB^{+}75, SP85, Shi89, Smi77, Son74, Son75, SJ78, SR73b, Sti37, Tay52b, Thi12, Tho55, Tob12, Wak09, War88, Wei51, Wei19, Whi86b, Wil11b, Wit96, Wit10, Wol13b, Wri92, Yal22, Yeo79, Yeo14, Das09, Jam12, Mil12, Mor10, Pan09b, Wit10, dA21, Die16a, Ano16b, Ano18, Bsc14b, CECE21, Hig17a, Kni13, Pro18, Vic04, YT13, Kin19b, And15]. Science [Ano74j, Ano09b, Bey14, Bik75, Bil16, Boe14, Bur15, Car17, Cha14, Clo55, Dun13, GG75a, GG76, GG10, Lin22, Lyn13, Tat17, Tau14, Wee16, Wil15]. Sciences [Ade10, ABNW09, Ano06f, Ano08j, Bow09a, Cro09a, Jac14, Lan13, LS15b, Maz10, Nor12, Pat17, Plo11, Prü09, Sak12, Tak11, Vet10, Ben09, CB22, Geo38a, Gou83, Rap17, VB11, Wei85b, Wer10, dBP86, Bin12, Cal68, Cro09a, Dav98, Gál88, Gue21, Gat18, Gue13, Old97]. Scientific [LA22, Ama06, Ank15, Ano05j, Ano08l, Bon14, Bou13, Bra75, Bur12b, Cha68, Chu13, CBL81, Das09, Des08, Eme02, Far72, Fin12, Gay13a, Gie16, Hsi09, Jar18a, Joh14, KLW06, Kre13, Kur03, Leh12a, Mar09, MLSB17, Par03, Pet14, Por16, Pre08, Ram10, RD16, RC08, Rya91, Sch57, She09, Smi22, Ste09a, Wal04, Wal09, Bai93, Bed93, Ben09, Bow84, Bra79, Bre91, Bro55, Cha48, Cop80, Cra85, DL11, Dav78, Die21, Dol77, Dur10, Ecc76, Ecc80, Ehr95, GG91a, Gay10, Gay13b, Geo52, Gie19, Gor16, Gou80, Gre84, GL11, HR89, Har74, Har66, Hen87, Ink78, Jac52a, Jac52b, Jay78, Kle12a, LL37, Lov79, Low90, Mac85a, Mal82, Mar16a, Mas53a, Mas53b, McC86, McC89, McC91, McC75, McD72, Mcc79, Mit78, NM37a, O’B65a, O’B65b, PK87]. scientific [Pri54, Rob54b, Rob56, Rob57, Roc10, Rus50, Rus51a, Rus51b, Sch90, Sha72, SP85, Spe16, SR72c, SR72a, SR72e, SR73c, Tod93, Wes56a, Wes56b, Yeo18, YO89, ZA16, dC88, Akt09, Joh18, Zai17]. scientifique [Bre91, Dur10]. scientifiques [Ben09]. scientism [Hak92, Sha76]. Scientist [Bsc12, Cha09, Del09a, VPW16, Wel83, Dav92, Gou83, Han21, How19, Jon86, Sha75b, dBP86, Ros62, Mil17, Leg17]. Scientists [Ban14, HA18, Kai10, Kei84, Luc08, Nie10, Olb89, And10, Bou17, Eyl63, Joh18, TH74, Wil13]. Scoresby [McC86, Cot77, Swe74]. Scotland [Eme88, Eyl37, NF73, SR73a, Wit96, Woo87]. Scottish [Wol13a, Cab73, D^{+}11, Eme88, OH97, Sim92, Ste20, Wil09, Ree14b]. Scrapping [Zev00]. scratch [Mil93]. Scribal [Gat18, Rap17]. Script [Arn09, KO83]. Sea [Cha13, Wex08, Akk23, FB22, Wri16]. Search [Bra09, UB15, Hak82, Hei67a, Hes17, Lai77, Mor36, Nay76, Swe74]. Searching [Coh10b, GH08, Kre23]. Seas [Ano08m, Kni14, Mil13, Wee16, FY09]. Seasonal [Mil93, Mil96]. Seasonal-hour [Mil93]. Sebokht [Tak11]. secoli [Bri07, Maz10]. Second [Bri10, Bru11, Bry13, Law13, Can09, McC36a, Aug04]. secondary [RS74]. Seconde [Aug04]. Secord [Van19]. Secrecy [dC09a]. Secret [Duc09, Hud03, Ras17, Ren18, Hen21a, Low90, Woo97]. Secrets [Bey14, Eve15, Mar11c, Pri13]. sector [Cha95, Mes97]. secundus [AVJ^{+}06, Bur11b]. Sedgwick [Edw72a]. Seed [Bur10a]. Seeds [Ano07i, Ber96]. Seeing [Mil08]. Seeking [Wil09, Wol13a]. seen [Geo52]. segregation [Cam80, Dar85]. seiner [Hon84, Pul94]. seismology [Dea91b]. Selected [HB14, Sce09, Lee13, Sus16]. Selection [Ray18, Bor10, Eva17, Rus80, Bar14, Smi18]. Selections [Pal12]. Self [Tat17, Ave20, Di 17, Jor82, SR74b, Vos13, Mar15]. self-censorship [Ave20]. self-induction [Jor82]. self-portrait [Di 17]. Selling [Ren11, vdS06]. senior [Smi87]. Sensations [Wit13, Pan09a]. Sense [Old99a, Rus04, Law21]. Sensible [Cec11]. sensitive [Rit62]. Sent [Des08]. Sergius [Tak11]. Series [Ano06j, Duc13, Fer02, Smi13, Gay13a]. Seriously [Moo87]. Sermons [Fan09]. Servetus [Mas53a]. ses [Mic13]. sets [Bun77]. Seur [Gui15]. Seventeenth [GG11a, Kid14, Vic04, Wer14, ATE11, Ari73a, Ari73b, Ben91, Ben37, Cha83a, CZ19, Deb67, Ear63, Fei03, Geo38b, Jon74, McC36b, Mel10, Mil45, Mul73, Sha73, Wei48, Dij04, Sal09, Tan14, War12, Wil01]. Seventeenth-Century [Vic04, CZ19, Geo38b, Jon74, McC36b, Wei48, Tan14]. Seventy [Ano06k]. several [Gre81, Gre82]. Severus [Tak11]. Sextants [Bro09, Sím93]. Sexual [Smi18]. Sexuality [Lyn15, McA12]. shades [Duc17]. Shadow [Alv07, HZ07]. Shank [Sta19, Bsc14b]. shape [Gre84, Ren18, Zig80]. shaped [Jon86, Lin22]. Shapin [Can75a]. Shaping [Das09, Gor17, Joh15b, RS17]. Share [Ano05j]. Sharing [Gor17]. Sharlin [Hol76]. Sharon [Mel21]. sharp [Dek08]. Shaw [Gib51b]. Shea [Pal12]. Shefer [Bil16]. Shefer-Mossensohn [Bil16]. Sheffield [Bro55, Ink75]. Sheldonian [Cal80]. shell [Zac72]. Shibukawa [Tan14]. shift [For61]. shifts [Per86]. Shigeo [Ots05]. shine [Kra17]. ship [Ari72, Cot77]. shoe [Duv52]. Shore [dA21]. Short [Hen17, Joh10, Kni17, Köb17, Tan11, Bec84, Bro16, Fra36, Fra38, HA19, Mac23, SS58]. shoulders [de 64]. Shropshire [Old92]. Si [Ges23]. Sibbald [Eme88, Wit96]. Sidereal [Pal12, SB09]. Sidereus [Pal12, Lat15, SB09]. sides [Tan77]. siècle [Bur12c, CP97, Law09a, LL37]. sight [Wil92]. sign [Tan80b]. Significance [Jam78, Cop87, Pri55, Sch21, Yav93]. Silicon [Aga17b]. Silke [Kin04]. Silkworm [Cob02]. Silvan [Tho11]. Silvestre [Duv54]. Silwood [Ank15, Gay13b]. similarly [Ari81]. Similibus [BvK73, BvK73]. similitude [Wri92]. Simon [Coh87, Lan13]. simple [Fou89]. Simples [Win16, Kah10]. simplicity [Bra90]. since [BJ13, Coz15, Eff21]. sine [Hor00]. Single [Car05, Zac72]. single-particle [Zac72]. Singular [Mil11b]. Sinusgesetz [Hor00]. Sir [Ano86g, Ben73, Bry75, Cha68, Col52a, Coo90, Eme88, Eyl63, Haa19, Jac52a, Jac52b, Kee52, Mac19, McD72, Old80, Old97, Ols70, PM50, Rus59a, Rus59b, Rus63c, Sin49, Sin50a, Sin50b, Sle69, Sok74, Spi68, Wal04, Wit96, Woo71a, de 49]. sister [Pal23]. Sisters [Car22]. Sites [Jam12, FL07, Jam78]. Sixteenth [Gor87, MVK02, Péo10, dC09a, Ari73b, Deb67, Di 17, Ges23, Lam97, Lon19, Mac23, Mes96, Sar69, VT17, Sch90]. Sixteenth-Century [MVK02, Péo10, dC09a, Gor87, Di 17, Lam97, Lon19, Mes96, Sar69]. size [Wil57]. Sizes [Sid07]. Skeleton [Slo03, Old09]. sketch [Fra36]. Skills [Bec14, ML^{+}12, Sch90]. Skleno [Tei75]. sky [Dek87]. slavery [Ste20]. Sledge [Bra09]. Sleep [Beg21]. Sleigh [Egm18]. Sloan [Jon95]. Sloane [Ait81]. slow [Sch75]. Small [Bel99, SS99]. smell [BS15]. Smith [BG99, Cot08, Eyl67, de 49]. Smithsonian [Mol84, Mor01]. smuggling [Val96]. Smyth [SH89]. Snell [Dij04]. Snow [Bez75]. soap [Gib39]. Social [Ano08k, Boe14, Cro96, Kai10, Moo87, Olb91, SP11, Wal76, Bak64, Bow84, Can75a, Feu71a, Feu71b, Ink77, Jar23, Kri22, Lus18, McC75, Nye11, Old94, Sha75a, Wei85a]. Socialism [Ste00b, VW10]. Sociedad [NG09]. ‘Società [Chi08]. Société [Duv54, Sme56b, Sme56a]. Societies [Kur03, Mcc79, SR74a]. Society [ABNW09, Bro36a, Car05, Eme88, Gay13a, GG93, Hul02, Lai04, Mil11b, MJ89, Nor12, Sch56, SBB^{+}75, Stu89, Tre63, Tre64, Tre65b, Tre65a, Tre70, Wil37, Bro78a, Gib51f, Gro74b, Hay64, Ink77, Kit82, Leg17, VPW16, Alb89, Bad60, Bar57, Bro78a, Bro90, BL21, BS93, Car97a, Far72, Gam21, Gov21, Gro72a, Gro74a, Mor88, Rob53a, Rob55a, Rob56, Rob57, Shu68, SR71b, Stu82]. Sociological [Ano08i]. Sociologist [SP11]. Sociology [CP02]. soda [Kra95]. Soddy [Dav92]. sodium [PB45, PB47, PB48]. sodium-ammonium [PB45, PB47, PB48]. Sofie [Lyn14]. Sojourners [Hsi09, Chu13]. Solaires [Arc05, Arc07]. Solar [CS12, Kra09b, LR00, For61, For72b, Hen82, Kid81, Mil86]. solari [Pia21]. soldiers [Sil18]. Solid [Bru79]. Solla [Tur84]. Solution [SHM13, Ben74, Gil76, Qua68, Qua70, Qua72, Sch80]. Solvay [Nye89]. Some [Ade10, Ano91o, Ben37, Cha61, Cha62, Cha63a, Cha63b, Cha64a, Cha64b, Cha75, Cro48, Dun62a, Dun62b, Dym37, Edw72b, EE51, Fra39b, GG86, GG84a, Gil02, GG74, GG75a, GG94, dVH67, Hil91, Hut77, Jon50, Ken52, KH60, LL37, Nak84, Old74, Old97, Par45, Pat36a, Shi89, Sme67, Sme87, Tur91, Ver56, Wil54, Wil57, Wil59, de 49, Bar98, CR96, Cha83a, Fra41, HT23, Hen89, Kni75, LS74, Mad55, McK59, MK60, McK60, McK62, McK65, McK67, Pel36, Sme57a, Tur55, Old94]. Son [Bey10, Ben09]. Sonata [Hui11]. sonne [Dek08]. sopra [Ari73a]. Soul [Jac14, Mar19, VB11, Wil23]. souls [Bla10, Bla19]. Sound [Ano05f, Pan05, Tka14, FG83]. sounding [DR54]. Soup [Mar06]. source [Bur90, Not17]. Sourcebook [Gui11b]. Sources [Bin12, Jac99, Old99b, Pet14, Sch08b, Bea76, Kus10, New96, Wel83, Pis13]. sous [Fau05]. South [Cra12, Lar91, Wat82]. South-East [Wat82]. southern [Dek87, Dek90, How03]. Soviet [Kre11]. Soviets [Sav15, Wol13b]. Space [Pet14, Cop80, Hen79, McG78]. Spacesuits [Stu09]. Spain [Coz19, Hor91, Kah10, Pre08, RC08, Ser12]. Spanish [Sla19, And14, Mal09, MVK02, Ser12, SM16, Val96]. Sparks [Mar06]. Spary [Van19, Dij13]. spas [SM16]. Special [Die16a, RG23, Pal83, Sch08c, Und63, Wal76, Wil55a, Mil81a]. speciation [Kot78]. Species [Car13, MWO07, SD13, Kot78, Dar13, DC09b]. specific [Sco81]. Spectacle [Ser12]. Spectacles [BS93]. Spectra [CR02, LR00, TS14, Mac72a]. spectrometric [Gen07]. Spectrométrique [Gen07]. Spectroscopic [Hen02]. Spectroscopy [Rei04]. spectrum [Cor38a, Cor38b]. Spectrums [Ash17]. speculation [Eas86]. Speculations [Fol14b, Kra82, Kra11b]. speculis [Ray15]. Spencer [CP02, Jac09, Sha76]. Spencerism [Hal18, Lig16]. Spengler [Gou83]. Spettroscopisti [Chi08]. Sphere [Del13]. Spheres [Gas07, Pri54]. spies [Bra92]. Spirit [BG99, Bow78]. Spiritism [Lyn14, Lac11]. Spoken [Tka14]. spontanée [Dur10]. Spontaneous [Foo69, Mon05, Dur10, Ful61, Gál88]. Spreading [KN13]. Spring [Plo99]. Springer [Hei19, Sch18b, Wit13]. Sprouting [Ano07i]. spy [Spa55]. square [Ait69, Pou91]. Squares [How02, How03]. St [Bro36a, Mil13, Old91, OM95]. St. [TH74, Tho51b]. Stachel [Mar11b]. stadholders [dC88]. Stahl [Tur53]. Stalbridge [O’B65a]. stamps [Don50]. stand [Lei20]. standard [Sim92]. standards [Kat16, Sim92]. Standing [Sta19]. Stanhope [Wes97]. Stanisław [GWN^{+}11]. Star [Bec13b, TS14, Ste10b]. Stark [Ste00b]. Starlight [Sus15, Bec11]. Starling [Ver56]. Stars [GG11a, Bur90, Gow82, Kra17, Mac19, Kin04, Kun00]. starting [Gre22]. Staszic [GWN^{+}11]. State [Ano07j, Sav15, Ver13, War07b, Wol13a, Bad60, Leg17, Mor86, NG09, Olb91, Uch22, VPW16, Wak09, Wol13b, Woo97]. statecraft [Jar18b]. Statement [AM76]. States [Hal20, Rei81, Ebe39, Gro72a, Gro74a, Pas51, Ste80]. Statesman [Bel99]. statics [Mar76]. Statistical [Byr80]. Statisticians [Hun99]. statistics [Dia94]. Status [Bon14, Lus18]. Stayley [Del11a]. Steam [McE10a, Mil08, Gol19, Gro72b]. Stedall [Rei14]. Stefan [Wal14]. Steiner [Ecc80, Ecc80]. Steinicke [Bec13b]. Stella [Jon95]. Stellar [GG11a, LR00, Het72, Kra17]. Stengers [Bsc14a]. Stephen [McC13, Wil23, Akt09]. Stereochemistry [vdS06]. Sternbild [Sie09]. Stetin [Kre23]. Stevin [Sch18b, Cha17, Coh87]. Stewart [Car18, McC14]. still [Tay45]. Stirpium [de 53b]. STM [GH08]. Stochastic [She98, Rab77]. Stone [Bsc12, Cha09]. Stop [Ano07k]. Stories [Kin19b, Mil11a]. storm [Hol07]. storm-warning [Hol07]. stormy [Alm22, Hen09]. Story [Alv07, Ben13a, Cre10, GH08, Ren18, Zev00, Bry90, Gab78, Mil17, Ros62, Sie09, Whi51]. Strachey [Cha54a]. strand [Mac89]. Strange [Ano08m, Chu13, Hsi09, Kni14, Sta17, Wee16, Ren18]. Strategies [De 05, SB13, SS13, Bec16, Muc98]. strategy [Hil84, Mor90]. stratigraphic [Nit96]. stratigraphischen [Nit96]. Stratigraphy [Sch08c]. stream [Ait69]. street [HB74]. Strocchia [Mel21]. strontium [Par42, Par51a]. Structural [San00]. structure [Far69, Ink78, Sha75a, Sta88a, Wil60a, Wil60b, Wil60c, Wil61, Por16, RD16]. Struggle [Har06]. Struggles [Fin12]. Strutt [Ege68]. Struve [Sza72]. Stuart [MS22, Fus58]. students [Klo85a, Klo85b]. Studies [Akt09, Ben13b, Bin12, Bro69a, Bro69b, Bur10b, Bur12c, Dom09, Kid14, Kim12, Mil47, Mil13, Nut21, Sch18b, Wil60c, Wit13, Ano75l, Col64, Cor36, FLW11, PM37, PM38a, PM38b, PM39, SR73b, Tay52a, War12, WN16, Wil61, Wil61]. Study [Ben14, BS12, Com21, Gar10, Gre57, Kni13, Old99b, URRF05, AS81, Ben73, Bra82a, Bra82b, Bra90, Bro87, Cha83b, CZ19, Col51, CBL81, DK54b, FC36, Gel21, Gos76, Gru11, HR89, Mal82, Mil81b, NS81, Old97, Ols69, Ols70, Par45, Pau78, Ros98, SBB^{+}75, Str93, SM16, Uss80, Wit10, YO89]. Stukeley [Mil74]. Stunica [McC37]. Sturm [Hen09]. Styles [Hon96]. Subject [Ano09f]. subjection [Tre78b]. subjectivity [Kvi22]. substance [Pel97]. substrate [Smi23b]. subtlety [Ree80]. Succeeded [Bsc12, Cha09]. Success [SP11, Bra95, Pat17]. successful [Ger23]. succession [Smi23b]. successive [BM94, Fra39b]. successors [Bay38b, Bay38a, Bay39a, Bay39b]. Such [Hen00]. Suffer [Stu09]. sufficiently [Dun23]. Sugar [War07b, GL11, Ste36b]. sulphur [dVH50]. summary [Mad57]. sumptuous [Kni77]. Sun [Kin04, Sid07, Man93]. sundial [FW85]. Sundials [Kin04, Arc05, Hig53, Mil93, Mil96, Kin04]. Sunspots [SB13, GS10, Lat14]. super [Bur14b, Bur14c, Bus10, SBT10]. superconductivity [GG84a]. superfluidity [GG86, GG88]. Superior [Mal09]. Supernatural [Lyn14, Lac11]. Superposition [Ela13]. Supplementary [SR72b]. support [Mor86, Olb91]. Suppression [Pat17]. Supremacy [Gov21]. Surface [Hop57a]. surgeon [Spr77]. surgical [Gra70, Und67]. Surrey [Kur00]. Survey [Alv07, Bea76, Car05, Eyl37, Fra41, Gib51a, Gib53a, Gib53b, Gib53c, PP77, Smi87, Wat82, Cum84]. surveying [Ano91o, Ben91, Tur91]. Surveyor [GWN^{+}11, Wei48, Wil91a]. survival [Sch75]. Survive [Stu09]. Suspension [Ano99e, Eir18]. Sussex [Hig17a, McC83]. Sustainability [Gue09]. Sutherland [Wat82]. Sverdrup [Mil91b]. Swammerdam [Cob02]. Sweden [Nor01]. Swedenborg [Pia21, Sch14]. Swedish [Hol96, Hol07, Kär14, Kau75]. Sweet [War07b]. sweetening [McB91]. Swift [NM37a, NM37b]. Swindlers [Wil13, Luc08]. Swineshead [Lew76]. Swiss [Har96]. Switzerland [Hei19, de 50b]. Sydney [And09a]. Sylva [FG83, Ree81]. Sylvarum [FG83, Ree81]. Sylvester [Cro96, Ink72a]. symbols [Alb89]. Symmons [Col54a]. Sympathy [Bur17]. Synaptic [Mar06]. Synergies [And14]. synoptice [Ray82]. synthesis [DK54c, Har85, Her18, HY87, vH92]. Synthetic [Str93, Mat76, Str97]. Syriac [Tak11]. System [War02, CZ19, Die16b, Flo96, Hil84, Hol07, RS89, Rut39]. systematic [Ohn23]. Systematizing [SB13]. Système [New52]. Systems [MW13, Cha21].

’t [RS01, vdS06, Dar13, Dav90a, Kin04, Lai04, Lev78, McC13, Tau14, Win49]. T. [Jar98, Law15, Sme54a]. Table [Bry10, Sce09, Lee13, Kun00]. tables [Dek93, Dun62a, Dun62b, Edd13, For66b, Lam97, Not18]. Tablettes [Rob10]. Tablot [Gro72b]. Tabulae [Kre23]. Tabule [Not18]. tabulis [Ray82]. Tait [Wil91b]. Tajika [GW11]. take [GG10, Zig80]. taken [New52]. Taking [Moo87]. Talbot [Woo71a, Woo71c]. Tale [Sei13, Bad09]. Tales [Hor87]. Talking [Ano07k]. Talmud [She98]. Taming [Hol07]. Tanganyika [Lee23]. Tappan [Plo80]. Tara [Lau21]. tartaric [Pat38]. tartrate [PB45, PB47, PB48]. Tasmania [SM82]. taste [Nye81]. Taylor [McC14]. te [Ges23]. Tea [Ano09i, Mau18]. teacher [Rob63, Tur53]. Teaching [Kni75, Mei79, Bro78b, Fos37, Gra70, Gue23, Jar98, Rob55b, SBB^{+}75, Spu87, SJ78, Sti37, Tac96, Tay52b, Tho55, Und63, Wig55, Wig56, Woo36]. team [Lev14a]. Technical [Gay00, Kur03, Ste76, Bak62, Hen89, NGDP75, SR72c, SR72a, SR73b, Sto66, Sto68, Whi86b]. technique [Woo97]. techniques [Hol17b]. Technischen [Hen89, Ste76]. Technological [Fit17, Gor17, Nof14, Smi22, Dav90b, Kre38, Str93, Tod93]. Technologies [Gat18, Har09b, Cob18, Kat16, Rap17]. Technology [Ano05i, Arm55, Ber09, Cah10, Cha12b, Dun13, Goo08, Ink88, Kid13, Mil13, Nof14, Sav15, Sch18b, Sta13, Tom09, Wal12, Wit13, Yi13, Aga21, Car10, Cha95, Clo54, CC56, Fli59, Fox96, Gor87, Hen87, Hor91, Ink89, Ink91, Kni09, Lev09a, Mei79, Mil81b, PNGP09, Rom78, SBB^{+}75, SJ78, Wol13b, Sla03]. tecnologico [Bri07, Maz10]. tectonics [YO89]. Ted [Mil13]. teh [Flo96]. Teilchen [Tre74]. telegraphy [Hol07, Hon96]. telephony [Jor82]. Telescope [Mal03, Wil11a, Wil01, Zui04, Jae10, Ree08]. Telescopes [Bec14, Mea70, ML^{+}12, Rud07]. Telescopic [GG11a]. Telioux [Cha52]. Telliamed [Coh11, Lyn13]. Telluribus [Pia21]. Telomeres [Cre10]. Temperature [CY05, vD08, Col95, GG86, GG84a, McC21]. tempests [Hol07]. Temple [Roe90]. temporis [Spe16]. tending [Wil60c, Wil61]. Tension [Hop57a, Tre78b, Tre80b]. Tenth [Law13, Bur12c, Can09]. term [Bak64, Nit96]. Terminology [Ink89]. terms [NGDP75]. terrae [Tar20]. Terrestrial [Caw77]. tertiary [Old72b]. testing [And19]. Tetrahedral [RS01]. Texas [Bil16]. text [Sme78]. Textbook [Gra01]. Textgrundlage [Sie09]. Texts [Ano08l, Bur00, Bur98, Fru16]. textual [Die16b, Fra39b, Sie09]. theatre [Ora75, Und63]. Their [Dek93, Hsi09, Bal23, Chu13, Coh36, Cop80, DD21, Ecc80, Gou22, HS22, Jac99, Lam97, Le 74, McC75, Sme87, Sne76, TT99]. Theistic [De 05]. Them [MW13]. thematic [PP77]. themselves [Gre22]. Theodore [Bar57]. Théodule [Muc98]. theological [Tre66]. théologie [Kah01]. theologies [Cop80]. theology [Bro77, Kah01, Mac72b, Yeo79]. Theophrastus [Wal01]. Theorem [Moz04, Nau14, Qua74]. Theoretical [Bow17, Mit12, Swe03, BJ00, Dun62a, Dun62b, Kim95, Old94, Thi73]. theoretischen [Thi73]. Theoricae [Swe21]. Theorie [Ber97, Bla88, DS97]. Theories [Blu13, Fol14b, Mil45, Pan05, Swe09, Ear61, Kra11b, Lei73, Li94, Mac85a, Osp80, Sne76, Hug52, Hug55]. Theorizer [Mol84]. Theorizing [OJY03]. Theory [Ano09g, Bin12, Boa08, Bru57a, CR96, Cra54, End16, Feu71b, Gau12b, Gro72b, Mae05, Mil81a, OH97, Pal83, Sau14, TT99, UB15, Yos99, vH94, vdS06, Ait54, Ait55a, Ait55b, Ait57, Ait58a, Ait58b, Ait59, And07, Ber97, Bla19, Bow78, Bru57b, Bru58a, Bru58b, Cah90, Che98, Coc83, Cre66, Dav66, Dav84, DS97, FGR74, For70b, GG91a, Gou80, Gou22, Hen80, Hom88, Hut77, KO83, Kat06, Kem68, Kot78, Kox90, Le 72, Le 74, Mas53a, Mau69, McK36a, Mil42, Nak84, Nay77, Nea51a, Nor75, PM37, PM38a, PM38b, PM39, Par39, Pau78, Pep76, Per86, Por78, RO83, RO84, Roc85, Rus22, Rus59a, Rus59b, Rus63a, Rus63b, Rus63c, Sce94, Sha75b, SW67, Swi62, Tat97, Tre78a, Tre80a, Wil55a, Wil55c, YO89, You37, Zig80, Feu71a, Dea75, Gos77]. Therapeutic [Edw04]. Therapy [Edw04]. there [Gau12b]. Thermodynamic [Kat06]. Thermodynamics [Gar76, Mal94]. Thermometer [Gau12a, Adl97, Ari74, Cha52, Cha71, Tay42]. thermometers [Cha72a, Gib48]. thermopile [Sch89]. theses [Kah01, Kah01]. thesis [GG75b, LS23]. Things [Hen00, Ren11]. Thinking [Mcv21, Sal09, Sha75b, She98]. third [Hen89]. thirteen [Rob55a]. thirteenth [Eva78, Ked16]. Thirty [Sch90]. Thomas [Bec13a, Haa19, Pis13, Spi69, Bow78, Cha61, Cha62, Cha63a, Cha63b, Cha64a, Cha64b, Din50, Duc18, Edw72b, FF68, Fox96, Gre57, Hug51a, Leu08, Mat54, Mau69, Mei21, Mol96, Par49, PdC10, Rob54b, Rob55b, Rob62, Sch08a, Sch08b, SS13, Sok74, Sok94, Spr77, Tan77, Wal00, WO87]. Thomaz [Boa19]. Thomist [Tar20]. Thomsen [Kra82]. Thomson [CY05, Cha91, Dia96, Mau69, Par49, Rob95, Sha75b, Smi80, Top80]. Thoreau [Ber96, Ber19, Met56]. thorium [Tre78a]. Thoruranium [Tre78a]. Thos. [McK36b]. those [Win66]. Thought [Ano08m, Hul02, IR11, Bow74, Byr80, CZ19, Lau66, Lon19, PK87, Tho47, Fol14a]. thoughts [RO82]. Thread [Ste00a]. threat [Lev09a]. Three [AC99, Bla78, Kah10, MJ89, Sch80, Tur82, Tur94]. Thrifty [Bak21]. Thuret [Plo99]. tide [Ked16]. tides [Ait54, Ait55a, Ait55b]. Tierra [Gre87]. Tilmid [Bur12a]. Timber [Fli59, CC56]. Time [Ano09h, Glu23b, Kin04, Rap15b, Ari73b, Can15, GG11b, Hut77, Lew76, McG78, SS58, VW10, Her17]. Timekeeper [HBP08, Plo99]. timekeepers [Bal23]. Timekeeping [Ish14]. Times [Dau14b, Duc09, Fra14b, Gos15, Ste00a, Bur09, Har12, Hen09, Lau21, McC10, RH11]. Timothy [Ren18]. tin [Lop60]. tinplate [Gib50, Gib51c, Gib51d, Gib51e, Gib55]. Tiny [Car05]. Tissue [Par16, Sch39]. titanium [Tre72]. Title [Ano55f]. Tiziana [Pal12]. Tobias [For66b, For70a, For70b, For71c, For72a, For72b]. Tobies [GG14]. Today [Law15, Hor91]. Together [RS01]. Tomic [Sim13]. Tommaso [McC39a]. Tones [Pan05]. too [BG99]. Tool [Rap15a]. Toolbox [Gas07]. tools [Ben09]. top [Ari72, Mas73]. top-sail [Mas73]. Tornioli [BS21]. Torporley [CT69, Tan77]. Torricell [Seg83]. torsion [And19]. Toscanelli [Ray15]. Total [CS12]. Toulmin [Por78]. Tour [Bec13a, Swe67, de 57, Dav68, Mil85, PdC10]. tower [Bed97, Dra88]. Toxicants [BJ13, Coz15]. Traces [dA21]. Tracings [Hes17]. Tractatus [Bur11b, AVJ^{+}06, Fol11, Fra39b]. Trade [And14, Win16, Bro79a, Cra85, Gib51a, Gib53a, Gib53b, Gib53c, Mil88a, Stu89, Whi51]. Tradition [Gra01, Gue23, Por11, Ste09a, Wal01, Cam23, Cor11, Nut14]. Traditional [Kim99, Spu87]. Traditions [Mar21b]. Trafficking [Bad11]. Tragedy [Edw09b]. Training [Bri00, Rob54c, Bra75, RS71, SR71a, SR71b]. Traité [DK54b]. traités [Bur12c]. trajectories [Nay83]. trajectory [Nay76]. Transactions [Geo52]. transatlantic [Hon96, Mac20a]. transcripts [Pul94]. transdisciplinary [Sch08c]. Transfer [Nof14, Fox96]. transferable [Jus11, Jus12]. transference [YO89]. Transformation [Ku06, Sal09, Val09, vD11, Dun21, Maf10]. transformations [GW11]. transformer [Hon95]. transformisme [Coh11, Lyn13]. Transformist [Wei85b]. transfusion [Pel97]. transhistorical [Byf19]. Transit [Lom11, Bed97, Hig17b]. Transition [Aga17a, Mcc79, Rad18]. transits [Woo53]. transitu [Fra39b]. Translated [Fra13, Wil23]. Translating [Wan18a, Bre16, Die16a]. Translatio [Bur11a]. Translation [Bew16, Bil18, Bur10a, Cec11, Pou01, RS01, Swe03, BJ00, Byf19, Coh36, Die16b, Die21, Gor16, McC37, NGDP75, Ari73a]. translations [Die16a, Ma21]. Translators [LS68, Mar16a]. Transmission [Ano08l, Mar06, Plo11, Ste11b, Dol77, Mac20a]. transmutation [Kot78]. transmutations [Gue21]. transnational [Byf19, Kri22, SS23]. Transparency [Pra10]. Transplant [Tan11]. Transplantation [Tre15, Ham12]. trasferimento [Bri07, Maz10]. trasformazione [Duc14, Maf10]. Tratado [OEMB21]. Travel [Bra09, Mar13, Car97a, Mar16a, VT17]. Travellers [Car97a, Ste69a, Ste69b]. travels [Wie23]. Treatise [Ben14, WH11, Ari73a, BB95, BB13, FG83, Fin22, Not21, San84, Gru11, Uga06]. treatises [Bur12c]. treatment [Wei51]. Trends [Ano06i]. Trial [Edw04]. Trials [Fan09, Gue23, Hed22, Jar23, Sch21]. triangle [Tan77]. triangulis [Tan77]. Trinity [KO83, Pri52]. Trip [Pre08]. Trissotetras [Haa19]. Tristram [Bod56]. Triumph [Edw09b]. Tropical [Joh10]. Trouble [San17]. troubles [Glu23b]. Troy [SC04]. True [Ben13a, Dun94, Sok94]. Truth [Bey14, Bra10, Mar11c, Nay82]. Truths [GG02]. trying [GG10]. tubular [Hua22]. Tudor [Dun94]. Tulip [Ano09i]. Tuning [Ku09]. Tunnelling [GH08]. Turbulent [RH11, Fra14b]. Turchetti [Kri19]. Turn [Bri10, Bry13, Gra72b]. Turner [Kin04, Ter37, Wil92, Gou83]. tutor [Sok74]. twelfth [Eva78, Kah10]. Twelfth-Century [Kah10]. Twentieth [And15, Ank15, Bry13, Cro16, Dun13, Fit17, Gay13b, Ots05, Row18, Bow09b, Aga12, Bri10, Fli08, Gat11, Law08, Li12, RC08, SR71a]. Twentieth-Century [Ank15, Fit17, Gay13b, Ots05, Row18, Cro16, Bow09b, Law08, Li12, RC08]. Two [Alv07, Bez75, Gat18, Nak84, Nye89, Old72c, Ots05, Sme56c, SH89, Wil55b, Wil65, Ben73, Bik75, Dol77, Gor87, Haw71, JTR94, JTR95, Kim95, Lam97, Min21, Rap17, Ste69a, Ste69b, Sus16]. Tycho [Cha84, Fig72, Sím93, Swe09, Tir20]. Tyndall [How19, Hes21, Jac15b, Kaa17, Kei84, Tho57]. Typewriter [Lea18]. Tyrocinium [Pat37].

U [Mil91b, Tre78a]. U-236 [Tre78a]. U.K. [Hoc83]. U.S.A [SBB^{+}75]. U.S.A. [Hoc83, Lun90]. U.S.S.R. [Hoc83]. UCL [HA18]. UK [Byf19, Van19]. Ultrarotstrahlung [Kan70, Kan71]. Unaffected [Wes10]. Uncertainty [Goo08, Sta13, Are10]. Uncommon [MHO13, HA15]. uncovering [Han21]. Undergraduate [SBB^{+}75, Don79]. undermining [Ari73b]. Understanding [Boa09, GG88, Her17, McE10a, BM94, Can15]. Underwater [Hac79b]. Undescribed [Van01]. unequal [Arc90]. UNESCO [Nof14]. unexcited [Haw68]. Unfulfilled [WO87]. Unhappy [Alv07]. unification [vD10, Del12]. Unified [Sau14]. union [Bik71, Gou83, Gra72a, Kre11]. Unipolar [Mil81b]. unit [Hop57b, Hop58, Hop59a]. United [Ebe39, Gro72a, Gro74a, Hal20, Pas51, Rei81, Ste80]. Units [Mil03, Gra70, Hop59a]. Unity [And09a, Ano12a, Del12]. universal [Lai77]. Universalis} [Mit79]. Universalism [DK11]. Universe [Smi80, Cas15, Sha19b, Kra00, KL07]. universitaire [Muc98]. universities [Hen89, Kni75, Min21]. University [Bil18, Bil16, Bow13, Bri13, Col52b, Col52c, Col52d, Coo13, Egm18, Fra13, Gat18, Gie16, Gol19, Gra72b, Gue23, Hay18, How19, Jon13, Kin04, Kin19b, Kri19, Lan13, Lau21, Law13, Lin22, Lov13, Luc19, Mar21a, May21, Mel21, Par13, Por16, Pro18, Qui13, Ray18, Reg19, Ren18, Row18, Rus04, SS23, Sha19b, Sla19, Smi18, Sta19, Sta21, Van19, Ver13, Wei19, Wil23, Wol13a, Yi13, BG99, Bro78b, Bro81, Chi70, Col53, Cot08, Dor76, Edm75, Mac72b, Pan09b, Swe70b]. Unknown [Eis05, Nea51c, Mac85b]. unlikely [SL77]. unlucky [Car97b]. Unofficial [Hun99]. Unpublished [SC15, Wes56a, Wes56b, Ait81, Arm56c, Arm58b, Bak62, Ben73, Bra90, Dea75, Duv52, Gra72b, Mad88c, OM95, Pel36, Ree81, Sch80, Thi73, VvM92, Wei50, de 58]. Unravelling [Bec11, Sus15]. unreachable [Dur10]. unrecorded [Dek95, Sme56c, Sme67]. unsolved [Cha71]. until [Geo52]. unveröffentlichten [Thi73]. upon [McC37]. Upright [Hen09]. Urdu [HR89]. Urquhart [Haa19]. Urs [Vic12]. ursprung [Hut73]. Ursula [Dij13]. US$113 [Vic04]. USA [Har06, Lal12]. Use [Bro09, Car18, KK02, Akk23, Byr81, Cro59, Dek93, Kat23, Sme87, Sne76, Tod82, Wil55a, Wil91a]. used [Gor87, Lam97, Tur91]. Useful [Vic04, GG77]. users [Bal23]. Uses [Ku09, DS16]. Usk [Gib53b]. utilitaire [Aug04]. utilitarian [Aug04, Bro36b]. utility [RS17, Son74]. utopian [Sve19]. Utopianism [Arm55]. utopias [Mac20a, Kre38]. Utopists [Arm56b]. Utrecht [Sne84]. utter [Win23].

V [Lei73, Cam55, Wal76, Woo71a, Bay39b, Cha64a, Gib55, McK65, Nea51c]. vacation [Arm58a]. Vaccines [MS22]. vacuum [DeK83, Wri16]. Valencia [GL11]. Vallensis [PLC48]. Valley [Aga17b, Dea91a]. Vallisneri [DG09]. values [Ols71]. vapour [Dym37]. Varia [Rob10]. variables [Gre81, Gre82]. Varignon [Bla88]. Varin [de 50a]. vascular [Fra41]. Vasilii [Byf19]. Vast [Lus14]. Vatican [Yap83]. Vaunting [Joh18]. Vaupel [Wal14]. vegetables [Tom01]. végétaux [Tom01]. velocity [BG94]. venandi [Yap83]. Venice [Lat15]. venom [Ear60]. venous [CB22]. ventures [SR72e]. Venus [Hig17b, Lom11, Mea66, Woo53]. Verein [Hon84]. Verhältnis [Hen86]. Veritas [Spe16]. veritate [PLC48]. Verlag [Kra13, Lat15]. Vermächtnis [Pul94]. Vermont [Edd13]. vers [Gen07]. Verschmelzung [Ste76]. Versuch [Ill89]. Versus [Ano07j, Bra75, Fle97, Goo80, Gow82, Hon95, Kat06, McC75]. Vertebrate [Slo03]. Vertèbres [New52]. vertical [Mil93]. Verwechselungsgeschichte [Gab78]. Very [Hen17, Kni17, Köb17, Bro16, HA19, Sch75]. Vetter [Gue13]. VI [Cha64b, McK67, Nea51d, Not17]. via [Duc14, Maf10]. vicar [Ste61]. Victoria [LMBM06]. Victorian [Bas15, Bur17, How19, Hun14, Jam12, Ano07k, Edw09a, Far18, FRW11, Hal14a, Ish14, Jam12, Lig10, LMBM06, Rus78, Sta17]. Victorians [SR72a]. vie [Hir07]. Vienna [Are10, Bad11]. View [Ano03f, Bur12d, Hak92, Het76, Lev78, Mal82, Ped75, Rus63b, Smi77]. Views [Old99b, Ben37, Far19, Far69, Har41, Tre66, Wil60c, Wil61]. Vigani [Col52b]. Vigenère [Pat39]. vii [Mil13, Reg19, Yi13, Nea51d]. viii [Hay18, Kid13, Lan13, Lov13, Nea52]. Ville [McC75]. Villon [Kah01]. Vinci [Wil65]. viper [Ear60]. Virreinato [Gre87]. Virtual [Pis14a, Cap12]. Virtuosi [Wil15, Yeo14]. Virtuoso [App96]. vis [Can19]. Vision [Bel99, Wil11b, Ari73a, Hol96, Mil42, Ste80, Str14]. Visionary [Jar98, Mol84]. visit [Hon84]. Visiting [Nau19]. visits [Wie23]. Vista [Ari73a]. Visual [Cob02, Coo09, Nyh16, Ste11b, HS22, Rou22]. vita [Gui11a]. Vital [BS12, Fin09, vS15]. viva} [Can19]. Vivian [Mcv21]. Viviani [Roe90]. Vivisection [Ric86]. Vlasschaert [Bur14a]. vocation [Hen89]. vocational [Cab73]. vogue [Jon37]. Vol [Akt09, Gue23, Har96, Sch18b]. volcanic [RO83]. volcanoes [de 62]. Volkstümliche [Bur09]. Vols [Bur60]. Volta [Osm58]. Voltaic [Mil03, Gil76, Rus59a]. Voltaire [Jon13, Sta19]. Volume [Gou23, Lat15, Pis13, Wit13, Gos77, Kra21, Sch08a, Sch08b]. Volumes [Ano09b, Lev09b, Boa19, Cro61, Sus16]. Vorgeschichte [Pul94]. Vorlesung [Pul94]. vortex [Ait57, Ait58a, Ait58b]. Voting [Bec09]. Voyage [Swe63, Lon19, NM37a, NM37b]. Voyaging [Ano08m, Kni14, Wee16]. Vries [Bow78, Cam80, Dar85, Kot79, Mei85, The94, Zev00]. vs [De 17, LS23, Mar06]. Vulcan [Smi22].

W [Ben95, Boa19, Bow84, Hig17b, Hil91, Kni13, Pau78, Sne81, Sza72]. W. [Dav90a, Gra72b, Str94, Swe74]. Wake [HA18, Bou17, Mac20b]. Wald [Thi73]. Wäldern [Dau13]. Walds [Thi73]. Wale [Nor12]. Wales [EE38, LPST11, SR73b]. Walked [Cha13, Wex08]. Walker [Boa12b, Edd08, Har12, Clo55, Swe72a]. walking [Fle97]. Wall [Hol76]. Wallis [Ela13, Gui11a, Sco36]. Walter [Sok74, Sha73]. Walton [Ver16]. War [Aga17a, Ger23, Kel11, Sav15, Sil18, Str93, Wol13b, HL91, Kre38, Low90, Mac93, Pal21, Sar69, Bre91, RG23, Gor16, Hal13, Hol07, Kri22, MN13, Sch90, Mos14b]. Ward [Har84]. Wärmetheorie [Ber97]. Warming [Lus14]. warning [Hol07]. Wars [Alv07, Kim95]. Warwick [Bro55]. Was [Cha16, Gau12a, Mac85b, Rus95, Sme57a]. Watanabe [Hil91]. Watches [Bro09]. ’Water [Mil02, Kim06, Orr14, DK54c, Maf10, Pat36b, Wil54]. waterproofing [Duv52]. Waters [Akt09, SM16]. Waterston [Bru57b]. Watson [Arm58b, Col53]. Watt [Gol19, Ben13b, JTR94, JTR95, Mil02, Mil08, Rob54c, Sch54, de 57]. wave [Che98, Hen80, Zig80]. wave-particle [Hen80]. Wavelength [Bro15]. Waves [Ano06g, Kan70, Kan71]. Way [Ano04f, Sar02]. wealth [Coz19]. weapons [Jon86]. Weather [Dav84, Rap15b, SR73a]. Webers [Fle97]. Webster [Cha61, Cha62, Cha63a, Cha63b, Cha64a, Cha64b, Edw72b]. Weighing [Hen21b]. Weight [Ham03, Cha19, Mau69, PM38a]. Weights [Ste76]. Weimann [Vic12]. Weimar [Wei85a]. well [Ape20, Sha19b]. well-ordered [Sha19b]. Wellen [Kan70, Kan71]. Wells [Bar11, Luc19, PW08]. Weltbild [Ped75]. Werner [Osp80]. Wernerian [Chi70, SW67]. Werrett [Lan13]. West [Adl17, Dym11, Hig17a, Mil91b, YO89, Mei79, Swe74, Vet04a, Vet04b]. Western [CZ19, Fru16, Gib51a, Jay78, OJY96]. Weston [Har12, Gos15]. Where [Gos77, Lei20]. Whewell [Bro77, Slo03, Yeo79]. which [HY87, Sme57a]. Whig [Old80]. Whipple [Pan09b]. Whiston [How03]. Whitby [McC86]. White [Wes10, Luc19, Fra47]. Whitehead [Low75]. Who [Cha13, Wex08, Del09a, Gre05]. Whole [UB15]. wider [Ped75]. Wien [Ill89]. Wiener [GG75b]. Wife [Bey14, Mar11c]. Wikipedia [Lov11]. Wilde [Car92]. Wiley [Hig17a]. Wilhelm [Hor87, Kau75, Wei85a, Gue21, Hak92]. Wilkingson [Tho67]. Wilkins [dCA36, McC38, McC39a, Wan18b]. Will [Law15, vS15, Yal22]. Willcocks [Ozd14]. Willem [vH94]. William [Ano58a, HB74, Lev94, Sch09, SP92, Sle69, Sus15, Sus16, Atk52, Bay38b, Bay38a, Bay39a, Bay39b, Bec11, Bey10, Bey11, BT55, Bro18, CY05, Clo55, Coc83, CH82, Cou59, Cro59, DeK83, Dea89b, Dia96, Dun67, Ege68, Eyl67, Gib52a, Gib52c, Gra72a, Hor87, How03, HB14, Kem17, Kur03, Lai79, Lev94, Mau15, McC86, McC94, MdVH58, Mil74, Nor42, Olb89, Oli69, Ozd14, Pal12, Rus22, Rus49, Rus50, Rus51a, Rus51b, Sal92, Sha75b, Smi80, Spi69, Swe70b, TS14, TW56, Tom14, Tre72, Uss80, Ver16, Whi86a, Wig55, Wig56, Yeo79]. Williams [Ham03, Sla03]. Williamson [HB74, Pau78]. Willis [OJY03]. Willy [Ill89, Ill89]. Wilson [Kin04, Wol13a]. win [Cha88]. windfall [Har84]. window [Thi12]. Winds [Mar21b]. Winner [Ano12f, Ano16b, Ano18, Ano19, CECE21, IM15, IM17]. Winter [Sei13, Bad09]. Wire [Mil04a]. Wireless [And09b, Hon96, Cha12b]. Wirkungsgeschichte [Thi71]. Wisan [Ano85j]. Wisdom [Bar09]. wissenschaftlichen [Ecc80]. wissenschaftlicher [Ecc76]. Wissenschaftsgeschichte [Hen88]. Witch [Att85]. Witch-Craze [Att85]. Witchcraft [Bar09]. Withers [Kni13]. within [CR02, Ego19]. without [Aga19, Asp69, Byf19, For71a, GG10, Qui12, Rup09]. Wolff [Wal14]. Wolfgang [Bec13b]. Wollaston [Cha71, TS14, Uss78, Uss80]. Wollaston/Chenevix [Uss78]. Woman [Cha13, Wex08]. Women [Car22, Jon95, Mic13, Ray18, LA22, Eva17, Mar16a, Mel21, Sti37]. wonder [Hed22]. wonderful [Bro90]. Woodward [Cha53b]. Woolwich [Cou59]. Word [Pom11, Dea79, Ros62, Mil17]. Wordes [Ben14, Gru11]. Words [Mos14a, Fru16, Nut21, Sha12]. Wordsworth [Hol17a]. Work [Ano06j, Ber01, BS07, Dom09, GG84b, Kur99, Mil03, Pis13, Pis14a, Vet04a, Vet04b, Arm56a, Bar95, Bro18, Cap12, CK76, Cop87, Cor38a, Cor38b, Fis36, Gou80, Har66, Har71, Hop57b, Hop58, Hop59b, Kan70, Kan71, Mad57, Nay80, OJY96, Pri55, Row50b, Rus50, Rus51a, Rus51b, Sch08a, Sch08b, Sle69, Sme78, The86, Tre80b, Ver56, Ver94, Wil60a, Wil60b, Wil60c, Wil61, Win49, Wit96]. workers [Gib50, GG84b]. working [Jar98, Ree81, SR72f, Tau10]. Works [Ano09g, Bel12, Sar02, GG92, Thi73, Win66]. Workshop [Pan05, Wit13, Mer98, Pan09a]. Workshops [Ser12]. World [Ano04f, Buc08, Del09a, Gau14, Hal13, Hud03, Nor01, Ren11, Sil18, War02, Aga19, Cha21, Coh10a, Edw10, Gre05, Hak92, Har12, Lon19, Mal82, Mar19, McC36a, Ped75, Alv07, Cra12, Kim95, Low90, Str93, VT17, May21]. world-view [Hak92]. worldmaking [Leh12b, Hah15]. Worlds [Mos14a, Ots05, Bro77, Mar16b, McC36b, Pia21, Sha12, Van19]. Wormskiold [Swe72b, Swe74]. Wren [Ben73]. Wrestling [HNS11, Bsc14b]. Wright [Din50, Hug51a]. Writing [Mau18, WH11, Mad61a, Mar16a, Row50a, Wie19, BB13]. writings [FC36, Sch21, Tan80a, Wil56a, Wil56b, Wil56c]. written [Ham01, Old99b, Sme56c, Sme78]. Wu [Flo96]. Wurtz [Fau05]. Wyttenbach [de 49].

X [Bur11a, Coz19, Luc19, Nut21, Sha19b, Sta19, Kaa17, Wil61]. X-Club [Kaa17]. X-ray [Wil61]. Xaver [Dal14]. Xe [Bur12c]. xi [Bow13, Lau21, Mar21a, Sta21]. xii [Car18, Edd13, Gat18, Kid13]. xiii [Hei19, Jon13, Mcv21, Par13]. xiv [Gue23, Hal18]. XIX [Akt09, Maz10, Gol19, Bri07]. Xue [CZ19]. xv [Bri13, Sla19]. xvi [Lai04, Wol13a]. xvii [Bil18]. XVIII [Maz10, Bri07, Rob10]. xx [How19]. xxi [Qui13]. xxv [Van19]. xxvi [Byf19]. xxviii [Ray18]. xxxii [Jar18a]. xxxiv [Smi18, Wil23]. xxxvi [Wit13].

Yale [Bow13]. Yamanouchi [Ots05]. Yang [Mil13]. Yarwell [Whi51]. Yavana [Plo11]. Year [Ano57b, But57, CCC57, Cha88, SHM13, Ste36a]. Years [Ano06k, Bow16, Cha17, Egm18, Par03, Sch90, Sch18b, Sle17, Bar57, Bik75, Bra76, Fos37, GG11b, Har96, Mac18, Sme55a, Sme55b, Thi73, Tur82]. Yeoman [Rob62]. Yogas [GW11]. York [Bri13, Lat15, Mcv21, Par13, Ray18, Sha19b, Wit13, Wal76, Bed93]. Yorkshire [Mor88]. Yosemite [Dea91a]. Yoshio [Thi12]. Young [RS01, Fei19, Sch19, Gra72a, Gra72b, Mil42]. younger [Rob54c]. youthful [Old97]. Yvette [Ben13b].

Zach [Dal14]. Zauberflöte [GG92]. Zealand [Old72b]. Zeeman [RP98, WO87]. Zeit [Hen09, VW10, Wal14]. Zeitreise [GG11b]. Zero [vD08, Col95]. Zero-Point [vD08]. Zieglerin [Lau21]. zograscope [Cha53a]. Zoo [Egm18, Sle17]. Zoological [SHM13, Col37a, Col37b]. zoologiques} [The86]. Zum [Hen86, Sch11]. Zur [Bur09, Hen88, Hen86, Kan70, Kan71, Thi71, Zac72]. Zürich [Ram15]. Zusammenhang [Hon84]. zusetsu [Thi12]. Zweiniger [Hal14b]. Zweiniger-Bargielowska [Hal14b]. zwischen [Dau13, Ecc80].

References

Andersen:2009:ALR

[ABNW09] Håkon With Andersen, Britta Brenna, Magne Njåstad, and Astrid Wale. Æmula lauri: the Royal Norwegian Society of Sciences and Letters, 1760–2010. Science History Publications, Sagamore Beach, MA, USA, 2009. ISBN 0-88135-383-3 (hardcover). xvi + 400 + 16 pp. LCCN AS283.T9 A33 2009.

Ackermann:1999:RIJ

[AC99] Silke Ackermann and John Cherry. Richard II, John Holland and three medieval quadrants. Annals of Science, 56(1):3–23, 1999. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Adam:2017:MME

[Ada17] Alison Adam. Murder and the making of English CSI. Annals of Science, 74(3):241–243, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Adelman:2010:PWW

[Ade10] Juliana Adelman. It’s part of what we are: Some Irish contributors to the development of the chemical and physical sciences. Annals of Science, 67(2):294–295, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Adler:1997:JDL

[Adl97] Jacob Adler. J. S. Delmedigo and the liquid-in-glass thermometer. Annals of Science, 54(3):293–299, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Adler:2017:BRF

[Adl17] Antony Adler. Book review: Field Life: Science in the American West During the Railroad Era. Annals of Science, 74(4):339–341, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Agar:2012:STC

[Aga12] Jon Agar. Science in the Twentieth Century and Beyond. History of science. Polity Press, Cambridge, UK, 2012. ISBN 0-7456-3469-9, 0-7456-3470-2, 0-7456-7579-4 (e-book). ix + 614 pp. LCCN Q125 .A36 2012.

Agar:2017:BRE

[Aga17a] Jon Agar. Book review: Europe’s Infrastructure Transition: Economy, War, Nature. Annals of Science, 74(2):173–174, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Agar:2017:BRC

[Aga17b] Jon Agar. Book review: The cybernetics moment: or why we call our age the information age / Moore’s Law: the life of Gordon Moore, Silicon Valley’s quiet revolutionary. Annals of Science, 74(1):88– 90, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Agar:2019:SFC

[Aga19] Jon Agar. Science without frontiers: cosmopolitanism and national interests in the world of learning, 1870–1940. Annals of Science, 76(1):103–105, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Agar:2021:WT

[Aga21] Jon Agar. What is technology? Annals of Science, 77(3):377–382, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1954:GTT

[Ait54] E. J. Aiton, M.Sc. Galileo’s theory of the tides. Annals of Science, 10(1):44–57, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1955:CNB

[Ait55a] E. J. Aiton, M.Sc. The contributions of Newton, Bernoulli and Euler to the theory of the tides. Annals of Science, 11(3):206–223, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1955:DTT

[Ait55b] E. J. Aiton, M.Sc. Descartes’s theory of the tides. Annals of Science, 11(4):337–348, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1957:VTP

[Ait57] E. J. Aiton, M.Sc., Ph.D. The vortex theory of the planetary motions. I. Annals of Science, 13(4):249–264, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1958:VTPa

[Ait58a] E. J. Aiton, M.Sc., Ph.D. The vortex theory of the planetary motions. II. Annals of Science, 14(2):132–147, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1958:VTPb

[Ait58b] E. J. Aiton, M.Sc., Ph.D. The vortex theory of the planetary motions. III. Annals of Science, 14(3):157–172, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1959:CTG

[Ait59] E. J. Aiton, M.Sc., Ph.D. The Cartesian theory of gravity. Annals of Science, 15(1):27–49, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1960:CML

[Ait60] E. J. Aiton, M.Sc., Ph.D. The celestial mechanics of Leibniz. Annals of Science, 16(2):65–82, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1962:CML

[Ait62] E. J. Aiton, M.Sc., Ph.D. The celestial mechanics of Leibniz in the light of Newtonian criticism. Annals of Science, 18(1):31–41, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1964:CML

[Ait64a] E. J. Aiton, M.Sc., Ph.D. The celestial mechanics of Leibniz: a new interpretation. Annals of Science, 20(2):111–123, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1964:IPC

[Ait64b] E. J. Aiton, M.Sc., Ph.D. The inverse problem of central forces. Annals of Science, 20(1):81–99, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1965:IEC

[Ait65] E. J. Aiton, M.Sc., Ph.D. An imaginary error in the celestial mechanics of Leibniz. Annals of Science, 21(3):169–173, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1969:NAS

[Ait69] E. J. Aiton, M.Sc., Ph.D. Newton’s aether-stream hypothesis and the inverse square law of gravitation. Annals of Science, 25(3):255– 260, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1970:IMM

[Ait70] E. J. Aiton, M.Sc., Ph.D. Ioannes Marcus Marci (1595–1667). Annals of Science, 26(2):153–164, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1978:KPC

[Ait78] E. J. Aiton. Kepler’s path to the construction and rejection of his first oval orbit for Mars. Annals of Science, 35(2):173–190, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1981:ULL

[Ait81] E. J. Aiton. An unpublished letter of Leibniz to Sloane. Annals of Science, 38(1):103–107, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1988:E

[Ait88] E. J. Aiton. Editorial. Annals of Science, 45(5):447–450, 1988. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Akkermans:2023:MPH

[Akk23] Emily Akkermans. Managing precision: how to use chronometers accurately at sea. Annals of Science, 81(1–2):235–257, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2288142.

Akopyan:2021:AFI

[Ako21] Ovanes Akopyan. After the flood. imagining the global environment in early modern Europe. Annals of Science, 78(1):126–128, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aktunc:2009:BRS

[Akt09] M. Emrah Aktunç. Book review: Stephen H. Kellert, Helen E. Longino, and C. Kenneth Waters, editors, Scientific Pluralism. Minnesota Studies in the Philosophy of Science. Vol. XIX. ISBN 0-8166-4763-1. Annals of Science, 66(2):299–302, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Alborn:1989:NNC

[Alb89] Timothy L. Alborn. Negotiating notation: Chemical symbols and British Society, 1831–1835. Annals of Science, 46(5):437–460, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Alberti:2009:NCO

[Alb09] Samuel J. M. M. Alberti. Nature and culture: objects, disciplines, and the Manchester Museum. Manchester University Press, Manchester, UK, 2009. ISBN 0-7190-8114-9 (hardcover). xi + 239 pp. LCCN AM43.M26 A53 2009.

Alexander:2010:DDH

[Ale10] Amir Alexander. Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Allchin:1994:JHP

[All94] Douglas Allchin. James Hutton and phlogiston. Annals of Science, 51(6):615–635, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Almasi:2022:ACS

[Alm22] Gábor Almási. Astrology in the crossfire: the stormy debate after the comet of 1577. Annals of Science, 79(2):137–163, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2030409.

Alter:2007:AOC

[Alt07] Stephen G. Alter. The advantages of obscurity: Charles Darwin’s negative inference from the histories of domestic breeds. Annals of Science, 64(2):235–250, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

AlvesDaMota:2007:MSI

[Alv07] Teresa Salomé Alves Da Mota. A mere shadow of an institution: the unhappy story of the Portuguese Geological Survey (PGS) in the period between the two World Wars. Annals of Science, 64(1):19– 40, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1976:LIB

[AM76] Piero E. Ariotti and Francis J. Marcolongo. The law of illumination before Bouguer (1729): Statement, restatements and demonstration. Annals of Science, 33(4):331–340, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Amaral:2006:ENS

[Ama06] Isabel Amaral. The emergence of new scientific disciplines in Portuguese medicine: Marck Athias’s Histophysiology Research School, Lisbon (1897–1946). Annals of Science, 63(1):85–110, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Andersen:2007:GAH

[And07] Kirsti Andersen. The geometry of an art: the history of the mathematical theory of perspective from Alberti to Monge. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-25961-9. xxxvii + 812 pp.

LCCN QA515 .A53 2007. URL http://www.loc.gov/catdir/
enhancements/fy0818/2005927076-d.html;   http://www.loc.
gov/catdir/enhancements/fy0818/2005927076-t.html.

Anduaga:2009:SCL

[And09a] Aitor Anduaga. Sydney Chapman on the layering of the atmosphere: Conceptual unity and the modelling of the ionosphere. Annals of Science, 66(3):333–344, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

AnduagaEgana:2009:WEG

[And09b] Aitor Anduaga Egaña. Wireless and empire: geopolitics, radio industry, and ionosphere in the British Empire, 1918– 1939. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-956272-5. xxv + 386 pp. LCCN

QC973.4.I6.  URL http://site.ebrary.com/lib/alltitles/
docDetail.action?docID=10282097;   http://www.gbv.eblib.
com/patron/FullRecord.aspx?p=746658.

Anderson:2010:NNS

[And10] Robert S. Anderson. Nucleus and nation: scientists, international networks, and power in India. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-01975-6 (hardcover). xxvi + 683 + 18 pp. LCCN Q127.I4 A69 2010.

Anduaga:2014:SJP

[And14] Aitor Anduaga. Spanish Jesuits in the Philippines: Geophysical research and synergies between science, education and trade, 1865– 1898. Annals of Science, 71(4):497–521, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anderson:2015:BRB

[And15] Warwick Anderson. Book review: Science in the Twentieth Century and Beyond. Annals of Science, 72(2):266–267, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anderson:2019:EIT

[And19] Katharine Anderson. An ‘experimental’ instrument: testing the torsion balance in Britain, Canada and Australia. Annals of Science, 76 (1):58–86, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anker:2015:BRB

[Ank15] Peder Anker. Book review: The Silwood Circle: A History of Ecology and the Making of Scientific Careers in Late Twentieth-Century Britain. Annals of Science, 72(2):264–266, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1936:Ea

[Ano36a] Anonymous. Editorial. Annals of Science, 1(1):1–3, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1936:Eb

[Ano36b] Anonymous. Errata. Annals of Science, 1(1):i–??, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1936:Ra

[Ano36c] Anonymous. Reviews. Annals of Science, 1(1):111–113, 1936. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1936:Rb

[Ano36d] Anonymous. Reviews. Annals of Science, 1(2):235–238, 1936. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1936:Rc

[Ano36e] Anonymous. Reviews. Annals of Science, 1(4):468–470, 1936. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Na

[Ano37a] Anonymous. Notes. Annals of Science, 2(2):236–238, April 1937.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Nb

[Ano37b] Anonymous. Notes. Annals of Science, 2(3):353–356, 1937. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Ra

[Ano37c] Anonymous. Reviews. Annals of Science, 2(1):130–136, 1937. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Rb

[Ano37d] Anonymous. Reviews. Annals of Science, 2(2):239–241, April 1937.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Rc

[Ano37e] Anonymous. Reviews. Annals of Science, 2(3):357–360, 1937. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1937:Rd

[Ano37f] Anonymous. Reviews. Annals of Science, 2(4):474–476, 1937. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1938:Na

[Ano38a] Anonymous. Notes. Annals of Science, 3(1):149, 1938. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1938:Nb

[Ano38b] Anonymous. Notes. Annals of Science, 3(2):237–238, 1938. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1938:Ra

[Ano38c] Anonymous. Reviews. Annals of Science, 3(1):150–151, 1938. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1938:Rb

[Ano38d] Anonymous. Reviews. Annals of Science, 3(2):239, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1938:Rc

[Ano38e] Anonymous. Reviews. Annals of Science, 3(4):457–460, 1938. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1939:Na

[Ano39a] Anonymous. Notes. Annals of Science, 4(1):107, 1939. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1939:Nb

[Ano39b] Anonymous. Notes. Annals of Science, 4(3):322–324, 1939. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1939:Ra

[Ano39c] Anonymous. Reviews. Annals of Science, 4(1):108–111, 1939. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1939:Rb

[Ano39d] Anonymous. Reviews. Annals of Science, 4(2):244, 1939. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1939:Rc

[Ano39e] Anonymous. Reviews. Annals of Science, 4(3):325–327, 1939. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1945:N

[Ano45] Anonymous. Notes. Annals of Science, 5(3):295–296, 1945. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1947:R

[Ano47] Anonymous. Reviews. Annals of Science, 5(4):373–377, 1947. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1948:R

[Ano48] Anonymous. Reviews. Annals of Science, 6(1):102–104, 1948. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1949:N

[Ano49a] Anonymous. Notes. Annals of Science, 6(2):206, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1949:R

[Ano49b] Anonymous. Reviews. Annals of Science, 6(2):207–214, 1949. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1950:Ra

[Ano50a] Anonymous. Reviews. Annals of Science, 6(3):323–326, 1950. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1950:Rb

[Ano50b] Anonymous. Reviews. Annals of Science, 6(4):448–450, 1950. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1951:Rb

[Ano51a] Anonymous. Review. Annals of Science, 7(2):210, 1951. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1951:Ra

[Ano51b] Anonymous. Reviews. Annals of Science, 7(1):109–112, 1951. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1951:Rc

[Ano51c] Anonymous. Reviews. Annals of Science, 7(3):300–302, 1951. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1952:Ra

[Ano52a] Anonymous. Reviews. Annals of Science, 8(2):192–194, 1952. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1952:Rb

[Ano52b] Anonymous. Reviews. Annals of Science, 8(3):282–284, 1952. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1952:Rc

[Ano52c] Anonymous. Reviews. Annals of Science, 8(4):398–401, 1952. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1953:Rd

[Ano53a] Anonymous. Review. Annals of Science, 9(4):377, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1953:Ra

[Ano53b] Anonymous. Reviews. Annals of Science, 9(1):98–100, 1953. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1953:Rb

[Ano53c] Anonymous. Reviews. Annals of Science, 9(2):194–195, 1953. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1953:Rc

[Ano53d] Anonymous. Reviews. Annals of Science, 9(3):287–288, 1953. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1954:N

[Ano54a] Anonymous. Note. Annals of Science, 10(3):274–275, 1954. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1954:Rb

[Ano54b] Anonymous. Review. Annals of Science, 10(3):276, 1954. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1954:Ra

[Ano54c] Anonymous. Reviews. Annals of Science, 10(1):78–86, 1954. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:C

[Ano55a] Anonymous. Corrigendum. Annals of Science, 11(4):356, 1955. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:Na

[Ano55b] Anonymous. Note. Annals of Science, 11(3):270, 1955. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:Nb

[Ano55c] Anonymous. Note. Annals of Science, 11(4):356, 1955. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:Rb

[Ano55d] Anonymous. Review. Annals of Science, 11(3):268–269, 1955. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:Ra

[Ano55e] Anonymous. Reviews. Annals of Science, 11(1):99–102, 1955. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1955:TP

[Ano55f] Anonymous. Title pages. Annals of Science, 11(4):303–308, 1955. ISSN 0003-3790 (print), 1464-505X (electronic). URL https://www.

tandfonline.com/doi/abs/10.1080/1464505X.2018.12016718.

Anonymous:1956:Ra

[Ano56a] Anonymous. Reviews. Annals of Science, 12(1):84, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1956:Rb

[Ano56b] Anonymous. Reviews. Annals of Science, 12(2):162, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1957:PM

[Ano57a] Anonymous. Proceedings at meetings. Annals of Science, 13(3):77– 81, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1957:RCP

[Ano57b] Anonymous. Report of the Committee of the Philosophy of Science Group for the Year 1956–1957. Annals of Science, 13(3):85, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1958:HWR

[Ano58a] Anonymous. Henry William Robinson 1888–1960. Annals of Science, 14(4):211–213, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1958:R

[Ano58b] Anonymous. Review. Annals of Science, 14(2):154, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1959:JFF

[Ano59a] Anonymous. John Farquhar Fulton 1899–1960. Annals of Science, 15(3–4):139–140, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1959:Rb

[Ano59b] Anonymous. Review. Annals of Science, 15(2):135–136, 1959. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1959:Ra

[Ano59c] Anonymous. Reviews. Annals of Science, 15(1):63–64, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1960:AC

[Ano60a] Anonymous. Addenda & corrigenda. Annals of Science, 16(4):276, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1960:Ra

[Ano60b] Anonymous. Reviews. Annals of Science, 16(1):59–64, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1960:Rb

[Ano60c] Anonymous. Reviews. Annals of Science, 16(2):125–128, 1960. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1960:Rc

[Ano60d] Anonymous. Reviews. Annals of Science, 16(4):269–275, 1960. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1961:Ra

[Ano61a] Anonymous. Reviews. Annals of Science, 17(1):63–64, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1961:Rb

[Ano61b] Anonymous. Reviews. Annals of Science, 17(2):121–128, 1961. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1961:Rc

[Ano61c] Anonymous. Reviews. Annals of Science, 17(3):197–200, 1961. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1961:Rd

[Ano61d] Anonymous. Reviews. Annals of Science, 17(4):269–272, 1961. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1962:Ra

[Ano62a] Anonymous. Reviews. Annals of Science, 18(1):63–64, 1962. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1962:Rb

[Ano62b] Anonymous. Reviews. Annals of Science, 18(2):125–130, 1962. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1963:Rb

[Ano63a] Anonymous. Review. Annals of Science, 19(3):239–240, 1963. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1963:Rc

[Ano63b] Anonymous. Review. Annals of Science, 19(4):299–300, 1963. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1963:Ra

[Ano63c] Anonymous. Reviews. Annals of Science, 19(2):149–152, 1963. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1964:Rb

[Ano64a] Anonymous. Review. Annals of Science, 20(4):323, 1964. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1964:Ra

[Ano64b] Anonymous. Reviews. Annals of Science, 20(3):243–244, 1964. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1965:Ra

[Ano65a] Anonymous. Reviews. Annals of Science, 21(2):139–144, 1965. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1965:Rb

[Ano65b] Anonymous. Reviews. Annals of Science, 21(3):203–208, 1965. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1966:Ra

[Ano66a] Anonymous. Reviews. Annals of Science, 22(1):69–72, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1966:Rb

[Ano66b] Anonymous. Reviews. Annals of Science, 22(2):139–144, 1966. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1966:Rc

[Ano66c] Anonymous. Reviews. Annals of Science, 22(3):221–224, 1966. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1966:Rd

[Ano66d] Anonymous. Reviews. Annals of Science, 22(4):295–301, 1966. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1967:Ra

[Ano67a] Anonymous. Reviews. Annals of Science, 23(1):77–80, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1967:Rb

[Ano67b] Anonymous. Reviews. Annals of Science, 23(2):175–176, 1967. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1967:Rc

[Ano67c] Anonymous. Reviews. Annals of Science, 23(3):243–248, 1967. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1967:Rd

[Ano67d] Anonymous. Reviews. Annals of Science, 23(4):305–314, 1967. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:BRa

[Ano68a] Anonymous. Books received. Annals of Science, 24(2):178, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:BRb

[Ano68b] Anonymous. Books received. Annals of Science, 24(4):346, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:DM

[Ano68c] Anonymous. Douglas McKie 1896–1967. Annals of Science, 24(1): 1–5, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:Ra

[Ano68d] Anonymous. Reviews. Annals of Science, 24(1):89–98, 1968. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:Rb

[Ano68e] Anonymous. Reviews. Annals of Science, 24(2):169–177, 1968. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:Rc

[Ano68f] Anonymous. Reviews. Annals of Science, 24(3):251–258, 1968. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1968:Rd

[Ano68g] Anonymous. Reviews. Annals of Science, 24(4):339–345, 1968. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:A

[Ano69a] Anonymous. Announcement. Annals of Science, 25(4):370, 1969.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:BRa

[Ano69b] Anonymous. Books received. Annals of Science, 25(1):93–94, 1969.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:BRb

[Ano69c] Anonymous. Books received. Annals of Science, 25(2):176–177, 1969.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:BRc

[Ano69d] Anonymous. Books received. Annals of Science, 25(3):272–273, 1969.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:BRd

[Ano69e] Anonymous. Books received. Annals of Science, 25(4):369, 1969.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:Ra

[Ano69f] Anonymous. Reviews. Annals of Science, 25(1):81–92, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:Rb

[Ano69g] Anonymous. Reviews. Annals of Science, 25(2):165–175, 1969. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:Rc

[Ano69h] Anonymous. Reviews. Annals of Science, 25(3):261–271, 1969. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1969:Rd

[Ano69i] Anonymous. Reviews. Annals of Science, 25(4):351–368, 1969. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Aa

[Ano70a] Anonymous. Announcement. Annals of Science, 26(1):84, 1970. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Ab

[Ano70b] Anonymous. Announcement. Annals of Science, 26(4):368, 1970.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:BRa

[Ano70c] Anonymous. Books received. Annals of Science, 26(1):82–83, 1970.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:BRb

[Ano70d] Anonymous. Books received. Annals of Science, 26(2):174–175, 1970.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:BRc

[Ano70e] Anonymous. Books received. Annals of Science, 26(3):271–272, 1970.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:BRd

[Ano70f] Anonymous. Books received. Annals of Science, 26(4):367, 1970.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Ra

[Ano70g] Anonymous. Reviews. Annals of Science, 26(1):73–82, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Rb

[Ano70h] Anonymous. Reviews. Annals of Science, 26(2):165–174, 1970. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Rc

[Ano70i] Anonymous. Reviews. Annals of Science, 26(3):261–271, 1970. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1970:Rd

[Ano70j] Anonymous. Reviews. Annals of Science, 26(4):355–366, 1970. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:A

[Ano71a] Anonymous. Announcement. Annals of Science, 27(2):212, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:BRa

[Ano71b] Anonymous. Books received. Annals of Science, 27(1):107–108, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:BRb

[Ano71c] Anonymous. Books received. Annals of Science, 27(2):210–211, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:BRc

[Ano71d] Anonymous. Books received. Annals of Science, 27(3):310–311, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:BRd

[Ano71e] Anonymous. Books received. Annals of Science, 27(4):422–424, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:Ra

[Ano71f] Anonymous. Reviews. Annals of Science, 27(1):105–107, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:Rb

[Ano71g] Anonymous. Reviews. Annals of Science, 27(2):205–209, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:Rc

[Ano71h] Anonymous. Reviews. Annals of Science, 27(3):299–309, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1971:Rd

[Ano71i] Anonymous. Reviews. Annals of Science, 27(4):413–421, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRa

[Ano72a] Anonymous. Books received. Annals of Science, 28(1):107–108, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRb

[Ano72b] Anonymous. Books received. Annals of Science, 28(2):213–215, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRc

[Ano72c] Anonymous. Books received. Annals of Science, 28(3):318, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRd

[Ano72d] Anonymous. Books received. Annals of Science, 29(1):103–104, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRe

[Ano72e] Anonymous. Books received. Annals of Science, 29(3):319–320, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:BRf

[Ano72f] Anonymous. Books received. Annals of Science, 29(4):424, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Ra

[Ano72g] Anonymous. Reviews. Annals of Science, 28(1):101–107, 1972. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rb

[Ano72h] Anonymous. Reviews. Annals of Science, 28(2):205–213, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rc

[Ano72i] Anonymous. Reviews. Annals of Science, 28(3):307–318, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rd

[Ano72j] Anonymous. Reviews. Annals of Science, 28(4):413–418, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Re

[Ano72k] Anonymous. Reviews. Annals of Science, 29(1):99–103, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rf

[Ano72l] Anonymous. Reviews. Annals of Science, 29(2):203–211, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rg

[Ano72m] Anonymous. Reviews. Annals of Science, 29(3):313–319, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1972:Rh

[Ano72n] Anonymous. Reviews. Annals of Science, 29(4):415–423, 1972. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:BRa

[Ano73a] Anonymous. Books received. Annals of Science, 30(2):253–256, 1973.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:BRb

[Ano73b] Anonymous. Books received. Annals of Science, 30(3):379, 1973.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:BRc

[Ano73c] Anonymous. Books received. Annals of Science, 30(4):464–468, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:Ra

[Ano73d] Anonymous. Reviews. Annals of Science, 30(1):123–128, 1973. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:Rb

[Ano73e] Anonymous. Reviews. Annals of Science, 30(2):221–252, 1973. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:Rc

[Ano73f] Anonymous. Reviews. Annals of Science, 30(3):365–378, 1973. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1973:Rd

[Ano73g] Anonymous. Reviews. Annals of Science, 30(4):457–464, 1973. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRb

[Ano74a] Anonymous. Book reviews. Annals of Science, 31(2):171–179, 1974.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRd

[Ano74b] Anonymous. Book reviews. Annals of Science, 31(3):263–268, 1974.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRf

[Ano74c] Anonymous. Book reviews. Annals of Science, 31(5):471–472, 1974.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRg

[Ano74d] Anonymous. Book reviews. Annals of Science, 31(6):579–584, 1974.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRa

[Ano74e] Anonymous. Books received. Annals of Science, 31(1):91–92, 1974.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRc

[Ano74f] Anonymous. Books received. Annals of Science, 31(2):180, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRe

[Ano74g] Anonymous. Books received. Annals of Science, 31(4):371–372, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:BRh

[Ano74h] Anonymous. Books received. Annals of Science, 31(6):585–587, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:ER

[Ano74i] Anonymous. Essay review. Annals of Science, 31(6):575–578, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:NEB

[Ano74j] Anonymous. A new editorial board for Annals of Science. Annals of Science, 31(2):93, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:Ra

[Ano74k] Anonymous. Reviews. Annals of Science, 31(1):73–91, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1974:Rb

[Ano74l] Anonymous. Reviews. Annals of Science, 31(4):365–370, 1974. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRa

[Ano75a] Anonymous. Book reviews. Annals of Science, 32(1):75–84, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRb

[Ano75b] Anonymous. Book reviews. Annals of Science, 32(2):179–186, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRc

[Ano75c] Anonymous. Book reviews. Annals of Science, 32(3):299–304, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRd

[Ano75d] Anonymous. Book reviews. Annals of Science, 32(4):393–414, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRe

[Ano75e] Anonymous. Book reviews. Annals of Science, 32(5):519–524, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:BRf

[Ano75f] Anonymous. Book reviews. Annals of Science, 32(6):587–608, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:ERa

[Ano75g] Anonymous. Essay review. Annals of Science, 32(1):71–73, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:ERb

[Ano75h] Anonymous. Essay review. Annals of Science, 32(2):175–178, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:ERc

[Ano75i] Anonymous. Essay review. Annals of Science, 32(3):295–298, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:ERd

[Ano75j] Anonymous. Essay review. Annals of Science, 32(6):579–582, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:JFM

[Ano75k] Anonymous. On Joseph Fourier: the man, the mathematician and the physicist. Annals of Science, 32(5):503–514, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1975:RSM

[Ano75l] Anonymous. Recent studies of medieval and Renaissance logic. Annals of Science, 32(5):515–518, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:BRa

[Ano76a] Anonymous. Book reviews. Annals of Science, 33(1):111–118, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:BRb

Anonymous:1976:BRc

[Ano76c] Anonymous. Book reviews. Annals of Science, 33(3):319–330, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:BRd

[Ano76d] Anonymous. Book reviews. Annals of Science, 33(4):407–424, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:BRe

[Ano76e] Anonymous. Book reviews. Annals of Science, 33(5):489–518, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:BRf

[Ano76f] Anonymous. Book reviews. Annals of Science, 33(6):595–618, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERa

[Ano76g] Anonymous. Essay review. Annals of Science, 33(1):105–109, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERb

[Ano76h] Anonymous. Essay review. Annals of Science, 33(2):201–204, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERc

[Ano76i] Anonymous. Essay review. Annals of Science, 33(2):205–210, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERd

[Ano76j] Anonymous. Essay review. Annals of Science, 33(3):311–318, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERe

[Ano76k] Anonymous. Essay review. Annals of Science, 33(4):399–406, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERf

[Ano76l] Anonymous. Essay review. Annals of Science, 33(5):481–484, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERg

[Ano76m] Anonymous. Essay review. Annals of Science, 33(5):485–488, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERh

[Ano76n] Anonymous. Essay review. Annals of Science, 33(6):585–588, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1976:ERi

[Ano76o] Anonymous. Essay review. Annals of Science, 33(6):589–593, 1976.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:BRa

[Ano77a] Anonymous. Book reviews. Annals of Science, 34(1):71–102, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:BRb

[Ano77b] Anonymous. Book reviews. Annals of Science, 34(2):203–220, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:BRc

[Ano77c] Anonymous. Book reviews. Annals of Science, 34(3):315–330, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:BRd

[Ano77d] Anonymous. Book reviews. Annals of Science, 34(4):429–448, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:BRe

[Ano77e] Anonymous. Book reviews. Annals of Science, 34(6):613–636, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:ERa

[Ano77f] Anonymous. Essay review. Annals of Science, 34(1):63–70, 1977.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1977:ERb

[Ano77g] Anonymous. Essay review. Annals of Science, 34(6):607–611, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRa

[Ano78a] Anonymous. Book reviews. Annals of Science, 35(1):81–102, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRb

[Ano78b] Anonymous. Book reviews. Annals of Science, 35(2):195–220, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRc

[Ano78c] Anonymous. Book reviews. Annals of Science, 35(3):325–338, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRd

[Ano78d] Anonymous. Book reviews. Annals of Science, 35(4):433–440, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRe

[Ano78e] Anonymous. Book reviews. Annals of Science, 35(5):543–550, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:BRf

[Ano78f] Anonymous. Book reviews. Annals of Science, 35(6):623–654, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:ERa

[Ano78g] Anonymous. Essay review. Annals of Science, 35(2):191–194, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:ERb

[Ano78h] Anonymous. Essay review. Annals of Science, 35(4):427–431, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:ERc

[Ano78i] Anonymous. Essay review. Annals of Science, 35(5):527–531, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:ERd

[Ano78j] Anonymous. Essay review. Annals of Science, 35(5):533–541, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1978:ERe

[Ano78k] Anonymous. Essay review. Annals of Science, 35(6):619–622, 1978.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRa

[Ano79a] Anonymous. Book reviews. Annals of Science, 36(1):87–110, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRb

[Ano79b] Anonymous. Book reviews. Annals of Science, 36(2):183–204, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRc

[Ano79c] Anonymous. Book reviews. Annals of Science, 36(3):301–314, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRd

[Ano79d] Anonymous. Book reviews. Annals of Science, 36(4):409–424, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRe

[Ano79e] Anonymous. Book reviews. Annals of Science, 36(5):543–548, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:BRf

[Ano79f] Anonymous. Book reviews. Annals of Science, 36(6):625–672, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:ERa

[Ano79g] Anonymous. Essay review. Annals of Science, 36(3):291–299, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:ERb

[Ano79h] Anonymous. Essay review. Annals of Science, 36(4):403–408, 1979.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1979:ERc

[Ano79i] Anonymous. Essay review. Annals of Science, 36(5):539–542, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRa

[Ano80a] Anonymous. Book reviews. Annals of Science, 37(1):107–126, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRb

[Ano80b] Anonymous. Book reviews. Annals of Science, 37(2):235–252, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRc

[Ano80c] Anonymous. Book reviews. Annals of Science, 37(3):357–362, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRd

[Ano80d] Anonymous. Book reviews. Annals of Science, 37(4):469–488, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRe

[Ano80e] Anonymous. Book reviews. Annals of Science, 37(5):607–614, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:BRf

[Ano80f] Anonymous. Book reviews. Annals of Science, 37(6):685–724, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERa

[Ano80g] Anonymous. Essay review. Annals of Science, 37(1):101–106, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERb

[Ano80h] Anonymous. Essay review. Annals of Science, 37(2):229–233, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERc

[Ano80i] Anonymous. Essay review. Annals of Science, 37(3):353–356, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERd

[Ano80j] Anonymous. Essay review. Annals of Science, 37(4):463–467, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERe

[Ano80k] Anonymous. Essay review. Annals of Science, 37(6):673–678, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1980:ERf

[Ano80l] Anonymous. Essay review. Annals of Science, 37(6):679–683, 1980.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRa

[Ano81a] Anonymous. Book reviews. Annals of Science, 38(1):123–124, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRb

[Ano81b] Anonymous. Book reviews. Annals of Science, 38(2):217–250, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRc

[Ano81c] Anonymous. Book reviews. Annals of Science, 38(3):353–376, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRd

[Ano81d] Anonymous. Book reviews. Annals of Science, 38(4):483–502, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRe

[Ano81e] Anonymous. Book reviews. Annals of Science, 38(5):597–612, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:BRf

[Ano81f] Anonymous. Book reviews. Annals of Science, 38(6):691–700, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERa

[Ano81g] Anonymous. Essay review. Annals of Science, 38(1):109–122, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERb

[Ano81h] Anonymous. Essay review. Annals of Science, 38(2):207–216, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERc

[Ano81i] Anonymous. Essay review. Annals of Science, 38(3):343–352, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERd

[Ano81j] Anonymous. Essay review. Annals of Science, 38(4):477–482, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERe

[Ano81k] Anonymous. Essay review. Annals of Science, 38(5):591–596, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1981:ERf

[Ano81l] Anonymous. Essay review. Annals of Science, 38(6):663–690, 1981.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRa

[Ano82a] Anonymous. Book reviews. Annals of Science, 39(1):77–114, 1982.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRb

[Ano82b] Anonymous. Book reviews. Annals of Science, 39(2):199–227, 1982.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRc

[Ano82c] Anonymous. Book reviews. Annals of Science, 39(3):315–325, 1982.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRd

[Ano82d] Anonymous. Book reviews. Annals of Science, 39(4):421–432, 1982.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRe

[Ano82e] Anonymous. Book reviews. Annals of Science, 39(5):505–526, 1982.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:BRf

[Ano82f] Anonymous. Book reviews. Annals of Science, 39(6):605–630, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERa

[Ano82g] Anonymous. Essay review. Annals of Science, 39(2):187–192, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERb

[Ano82h] Anonymous. Essay review. Annals of Science, 39(2):193–198, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERc

[Ano82i] Anonymous. Essay review. Annals of Science, 39(4):399–406, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERd

[Ano82j] Anonymous. Essay review. Annals of Science, 39(4):407–420, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERe

[Ano82k] Anonymous. Essay review. Annals of Science, 39(5):491–504, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1982:ERf

[Ano82l] Anonymous. Essay review. Annals of Science, 39(6):593–603, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRa

[Ano83a] Anonymous. Book reviews. Annals of Science, 40(1):95–108, January 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRb

[Ano83b] Anonymous. Book reviews. Annals of Science, 40(2):197–216, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRc

[Ano83c] Anonymous. Book reviews. Annals of Science, 40(3):297–316, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRd

[Ano83d] Anonymous. Book reviews. Annals of Science, 40(4):397–409, July 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRe

[Ano83e] Anonymous. Book reviews. Annals of Science, 40(5):505–529, September 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:BRf

[Ano83f] Anonymous. Book reviews. Annals of Science, 40(6):663–676, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:C

[Ano83g] Anonymous. Correspondence. Annals of Science, 40(4):396, July 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:ERa

[Ano83h] Anonymous. Essay review. Annals of Science, 40(3):281–288, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:ERb

[Ano83i] Anonymous. Essay review. Annals of Science, 40(3):289–295, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1983:ERc

[Ano83j] Anonymous. Essay review. Annals of Science, 40(5):501–504, September 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:BRa

[Ano84a] Anonymous. Book reviews. Annals of Science, 41(1):87–104, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:BRb

Anonymous:1984:BRc

[Ano84c] Anonymous. Book reviews. Annals of Science, 41(3):287–311, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:BRd

[Ano84d] Anonymous. Book reviews. Annals of Science, 41(4):395–415, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:BRe

[Ano84e] Anonymous. Book reviews. Annals of Science, 41(5):489–511, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:BRf

[Ano84f] Anonymous. Book reviews. Annals of Science, 41(6):593–606, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:ERa

[Ano84g] Anonymous. Essay review. Annals of Science, 41(2):165–169, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:ERb

[Ano84h] Anonymous. Essay review. Annals of Science, 41(2):171–177, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:ERc

[Ano84i] Anonymous. Essay review. Annals of Science, 41(4):383–393, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1984:ERd

[Ano84j] Anonymous. Essay review. Annals of Science, 41(5):483–487, 1984.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRa

[Ano85a] Anonymous. Book reviews. Annals of Science, 42(1):81–85, 1985.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRb

[Ano85b] Anonymous. Book reviews. Annals of Science, 42(2):165–185, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRc

[Ano85c] Anonymous. Book reviews. Annals of Science, 42(3):347–354, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRd

[Ano85d] Anonymous. Book reviews. Annals of Science, 42(4):437–451, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRe

[Ano85e] Anonymous. Book reviews. Annals of Science, 42(5):545–546, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:BRf

[Ano85f] Anonymous. Book reviews. Annals of Science, 42(6):617–630, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:ERa

[Ano85g] Anonymous. Essay review. Annals of Science, 42(4):421–429, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:ERb

[Ano85h] Anonymous. Essay review. Annals of Science, 42(4):431–435, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:ERc

[Ano85i] Anonymous. Essay review. Annals of Science, 42(6):603–611, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1985:WGA

[Ano85j] Anonymous. Wisan on Galileo and the art of reasoning. Annals of Science, 42(6):613–616, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:BRa

[Ano86a] Anonymous. Book reviews. Annals of Science, 43(1):87–109, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:BRb

Anonymous:1986:BRc

[Ano86c] Anonymous. Book reviews. Annals of Science, 43(3):291–314, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:BRd

[Ano86d] Anonymous. Book reviews. Annals of Science, 43(4):407–412, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:BRe

[Ano86e] Anonymous. Book reviews. Annals of Science, 43(5):489–510, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:BRf

[Ano86f] Anonymous. Book reviews. Annals of Science, 43(6):563–607, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:CLP

[Ano86g] Anonymous. Correspondence the letters and papers of Sir John Hill, 1714–75. Annals of Science, 43(2):175–178, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:E

[Ano86h] Anonymous. Erratum. Annals of Science, 43(1):110, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:ERa

[Ano86i] Anonymous. Essay review. Annals of Science, 43(1):81–86, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1986:ERb

[Ano86j] Anonymous. Essay review. Annals of Science, 43(6):557–561, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRa

[Ano87a] Anonymous. Book reviews. Annals of Science, 44(1):85–105, January 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRb

[Ano87b] Anonymous. Book reviews. Annals of Science, 44(2):191–213, March 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRc

[Ano87c] Anonymous. Book reviews. Annals of Science, 44(3):297–318, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRd

[Ano87d] Anonymous. Book reviews. Annals of Science, 44(4):419–438, July 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRe

[Ano87e] Anonymous. Book reviews. Annals of Science, 44(5):525–544, September 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:BRf

[Ano87f] Anonymous. Book reviews. Annals of Science, 44(6):645–656, November 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1987:E

[Ano87g] Anonymous. Errata. Annals of Science, 44(1):106, January 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:BRa

[Ano88a] Anonymous. Book reviews. Annals of Science, 45(1):93–112, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:BRb

[Ano88b] Anonymous. Book reviews. Annals of Science, 45(2):201–220, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:BRc

[Ano88c] Anonymous. Book reviews. Annals of Science, 45(3):309–327, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:BRd

[Ano88d] Anonymous. Book reviews. Annals of Science, 45(4):427–446, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:BRe

[Ano88e] Anonymous. Book reviews. Annals of Science, 45(5):535–554, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:B

[Ano88f] Anonymous. Bookreviews. Annals of Science, 45(6):647–670, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:ERa

[Ano88g] Anonymous. Essay review. Annals of Science, 45(3):303–307, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:ERb

[Ano88h] Anonymous. Essay review. Annals of Science, 45(6):643–646, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1988:HAC

[Ano88i] Anonymous. Hofmann’s appointment to the College of Chemistry. Annals of Science, 45(2):199, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRa

[Ano89a] Anonymous. Book reviews. Annals of Science, 46(1):97–115, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRb

[Ano89b] Anonymous. Book reviews. Annals of Science, 46(2):201–219, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRc

[Ano89c] Anonymous. Book reviews. Annals of Science, 46(3):309–322, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRd

[Ano89d] Anonymous. Book reviews. Annals of Science, 46(4):423–435, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRe

[Ano89e] Anonymous. Book reviews. Annals of Science, 46(5):527–548, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:BRf

[Ano89f] Anonymous. Book reviews. Annals of Science, 46(6):627–639, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:E

[Ano89g] Anonymous. Erratum. Annals of Science, 46(5):519, 1989. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:ERa

[Ano89h] Anonymous. Essay review. Annals of Science, 46(1):93–96, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:ERb

[Ano89i] Anonymous. Essay review. Annals of Science, 46(2):195–199, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:ERc

[Ano89j] Anonymous. Essay review. Annals of Science, 46(3):303–307, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:ERd

[Ano89k] Anonymous. Essay review. Annals of Science, 46(5):521–526, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1989:ERe

[Ano89l] Anonymous. Essay review. Annals of Science, 46(6):621–625, 1989.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRa

[Ano90a] Anonymous. Book reviews. Annals of Science, 47(1):89–109, 1990.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRb

[Ano90b] Anonymous. Book reviews. Annals of Science, 47(2):191–212, 1990.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRc

[Ano90c] Anonymous. Book reviews. Annals of Science, 47(3):313–316, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRd

[Ano90d] Anonymous. Book reviews. Annals of Science, 47(4):411–421, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRe

[Ano90e] Anonymous. Book reviews. Annals of Science, 47(5):513–528, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:BRf

[Ano90f] Anonymous. Book reviews. Annals of Science, 47(6):619–628, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:C

[Ano90g] Anonymous. Correspondence. Annals of Science, 47(4):407–409, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:ERa

[Ano90h] Anonymous. Essay review. Annals of Science, 47(1):81–87, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:ERb

[Ano90i] Anonymous. Essay review. Annals of Science, 47(3):291–300, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:ERc

[Ano90j] Anonymous. Essay review. Annals of Science, 47(3):301–312, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:ERd

[Ano90k] Anonymous. Essay review. Annals of Science, 47(4):393–406, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1990:ERe

[Ano90l] Anonymous. Essay review. Annals of Science, 47(5):507–512, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:Aa

[Ano91a] Anonymous. Announcement. Annals of Science, 48(4):312, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:Ab

[Ano91b] Anonymous. Announcement. Annals of Science, 48(5):508, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRa

[Ano91c] Anonymous. Book reviews. Annals of Science, 48(1):91–102, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRb

[Ano91d] Anonymous. Book reviews. Annals of Science, 48(2):191–206, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRc

[Ano91e] Anonymous. Book reviews. Annals of Science, 48(3):295–304, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRd

[Ano91f] Anonymous. Book reviews. Annals of Science, 48(4):391–406, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRe

[Ano91g] Anonymous. Book reviews. Annals of Science, 48(5):487–507, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:BRf

[Ano91h] Anonymous. Book reviews. Annals of Science, 48(6):597–606, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:C

[Ano91i] Anonymous. Correspondence. Annals of Science, 48(5):508, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:ERa

[Ano91j] Anonymous. Essay review. Annals of Science, 48(1):69–86, 1991.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:ERb

[Ano91k] Anonymous. Essay review. Annals of Science, 48(1):87–90, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:ERc

[Ano91l] Anonymous. Essay review. Annals of Science, 48(3):289–293, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:ERd

[Ano91m] Anonymous. Essay review. Annals of Science, 48(4):385–389, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:ERe

[Ano91n] Anonymous. Essay review. Annals of Science, 48(6):589–595, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1991:MAM

[Ano91o] Anonymous. ‘measuring of all maner of lands’: Some aspects of the history of surveying. Annals of Science, 48(4):311, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:Aa

[Ano92a] Anonymous. Announcement. Annals of Science, 49(4):400, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:Ab

[Ano92b] Anonymous. Announcement. Annals of Science, 49(5):506, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRa

[Ano92c] Anonymous. Book reviews. Annals of Science, 49(1):91–98, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRb

[Ano92d] Anonymous. Book reviews. Annals of Science, 49(2):191–200, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRc

[Ano92e] Anonymous. Book reviews. Annals of Science, 49(3):277–300, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRd

[Ano92f] Anonymous. Book reviews. Annals of Science, 49(4):385–399, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRe

[Ano92g] Anonymous. Book reviews. Annals of Science, 49(5):489–502, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BRf

[Ano92h] Anonymous. Book reviews. Annals of Science, 49(6):589–600, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BNa

[Ano92i] Anonymous. Books noticed. Annals of Science, 49(1):1, 1992. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BNb

[Ano92j] Anonymous. Books noticed. Annals of Science, 49(1):99–100, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:BNc

[Ano92k] Anonymous. Books noticed. Annals of Science, 49(5):503–504, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:E

[Ano92l] Anonymous. Erratum. Annals of Science, 49(5):505, 1992. CODEN

ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERa

[Ano92m] Anonymous. Essay review. Annals of Science, 49(1):87–90, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERb

[Ano92n] Anonymous. Essay review. Annals of Science, 49(2):175–184, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERc

[Ano92o] Anonymous. Essay review. Annals of Science, 49(2):185–189, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERd

[Ano92p] Anonymous. Essay review. Annals of Science, 49(4):377–384, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERe

[Ano92q] Anonymous. Essay review. Annals of Science, 49(6):577–583, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1992:ERf

[Ano92r] Anonymous. Essay review. Annals of Science, 49(6):585–588, 1992.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:A

[Ano93a] Anonymous. Announcements. Annals of Science, 50(6):600, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRa

[Ano93b] Anonymous. Book reviews. Annals of Science, 50(1):95–98, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRb

[Ano93c] Anonymous. Book reviews. Annals of Science, 50(2):183–200, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRc

[Ano93d] Anonymous. Book reviews. Annals of Science, 50(3):291–301, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRd

[Ano93e] Anonymous. Book reviews. Annals of Science, 50(4):389–402, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRe

[Ano93f] Anonymous. Book reviews. Annals of Science, 50(5):491–500, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BRf

[Ano93g] Anonymous. Book reviews. Annals of Science, 50(6):585–599, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:BN

[Ano93h] Anonymous. Books noticed. Annals of Science, 50(1):99–100, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:ERa

[Ano93i] Anonymous. Essay review. Annals of Science, 50(2):175–181, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:ERb

[Ano93j] Anonymous. Essay review. Annals of Science, 50(4):373–381, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1993:ERc

[Ano93k] Anonymous. Essay review. Annals of Science, 50(4):383–388, 1993.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:Aa

[Ano94a] Anonymous. Announcements. Annals of Science, 51(5):iv, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcements. Annals of Science, 51(6):iv, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRa

[Ano94c] Anonymous. Book reviews. Annals of Science, 51(1):75–102, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRb

[Ano94d] Anonymous. Book reviews. Annals of Science, 51(2):185–199, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRc

[Ano94e] Anonymous. Book reviews. Annals of Science, 51(3):291–328, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRd

[Ano94f] Anonymous. Book reviews. Annals of Science, 51(4):416–452, 1994.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRe

[Ano94g] Anonymous. Book reviews. Annals of Science, 51(5):541–570, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BRf

[Ano94h] Anonymous. Book reviews. Annals of Science, 51(6):655–687, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:BN

[Ano94i] Anonymous. Books noticed. Annals of Science, 51(2):199–202, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:C

[Ano94j] Anonymous. Correspondence. Annals of Science, 51(5):539–540, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:ERa

[Ano94k] Anonymous. Essay review. Annals of Science, 51(2):177–183, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:ERb

[Ano94l] Anonymous. Essay review. Annals of Science, 51(3):281–285, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:ERc

[Ano94m] Anonymous. Essay review. Annals of Science, 51(3):287–290, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:ERd

[Ano94n] Anonymous. Essay review. Annals of Science, 51(4):413–415, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1994:ERe

[Ano94o] Anonymous. Essay review. Annals of Science, 51(5):531–538, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:BRa

[Ano95a] Anonymous. Book reviews. Annals of Science, 52(1):85–102, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:BRb

Anonymous:1995:BRc

[Ano95c] Anonymous. Book reviews. Annals of Science, 52(3):307–320, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:BRd

[Ano95d] Anonymous. Book reviews. Annals of Science, 52(4):413–430, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:BRe

[Ano95e] Anonymous. Book reviews. Annals of Science, 52(5):519–536, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:BRf

[Ano95f] Anonymous. Book reviews. Annals of Science, 52(6):619–628, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:C

[Ano95g] Anonymous. Correspondence. Annals of Science, 52(3):303–305, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:E

[Ano95h] Anonymous. Erratum. Annals of Science, 52(2):212, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:ERa

[Ano95i] Anonymous. Essay review. Annals of Science, 52(1):77–84, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:ERb

[Ano95j] Anonymous. Essay review. Annals of Science, 52(2):187–192, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:ERc

[Ano95k] Anonymous. Essay review. Annals of Science, 52(5):503–507, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:ERd

[Ano95l] Anonymous. Essay review. Annals of Science, 52(5):509–517, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1995:ERe

[Ano95m] Anonymous. Essay review. Annals of Science, 52(6):615–618, 1995.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:A

[Ano96a] Anonymous. Announcement. Annals of Science, 53(5):488, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRd

[Ano96b] Anonymous. Book review. Annals of Science, 53(4):413–430, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRa

[Ano96c] Anonymous. Book reviews. Annals of Science, 53(1):89–106, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRb

[Ano96d] Anonymous. Book reviews. Annals of Science, 53(2):195–212, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRc

[Ano96e] Anonymous. Book reviews. Annals of Science, 53(3):301–318, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRe

[Ano96f] Anonymous. Book reviews. Annals of Science, 53(5):527–540, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BRf

[Ano96g] Anonymous. Book reviews. Annals of Science, 53(6):633–654, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:BN

[Ano96h] Anonymous. Books noticed. Annals of Science, 53(3):319–322, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:EB

[Ano96i] Anonymous. Editorial board. Annals of Science, 53(4):ebii, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:ERa

[Ano96j] Anonymous. Essay review. Annals of Science, 53(1):85–88, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:ERb

[Ano96k] Anonymous. Essay review. Annals of Science, 53(3):297–300, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:ERc

[Ano96l] Anonymous. Essay review. Annals of Science, 53(5):521–526, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:ERd

[Ano96m] Anonymous. Essay review. Annals of Science, 53(6):617–625, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1996:ERe

[Ano96n] Anonymous. Essay review. Annals of Science, 53(6):627–631, 1996.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRa

[Ano97a] Anonymous. Book reviews. Annals of Science, 54(1):93–109, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRb

[Ano97b] Anonymous. Book reviews. Annals of Science, 54(2):207–220, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRc

[Ano97c] Anonymous. Book reviews. Annals of Science, 54(3):311–326, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRd

[Ano97d] Anonymous. Book reviews. Annals of Science, 54(4):423–430, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRe

[Ano97e] Anonymous. Book reviews. Annals of Science, 54(5):523–531, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BRf

[Ano97f] Anonymous. Book reviews. Annals of Science, 54(6):623–634, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:BN

[Ano97g] Anonymous. Books noticed. Annals of Science, 54(3):327–329, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:C

[Ano97h] Anonymous. Corrigenda. Annals of Science, 54(1):109, 1997. CO-

DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERa

[Ano97i] Anonymous. Essay review. Annals of Science, 54(1):87–91, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERb

[Ano97j] Anonymous. Essay review. Annals of Science, 54(2):197–206, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERc

[Ano97k] Anonymous. Essay review. Annals of Science, 54(3):301–304, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERd

[Ano97l] Anonymous. Essay review. Annals of Science, 54(3):305–309, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERe

[Ano97m] Anonymous. Essay review. Annals of Science, 54(4):407–421, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1997:ERf

[Ano97n] Anonymous. Essay review. Annals of Science, 54(6):611–622, 1997.

CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:BRa

[Ano98a] Anonymous. Book reviews. Annals of Science, 55(1):105–109, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:BRb

[Ano98b] Anonymous. Book reviews. Annals of Science, 55(2):199–218, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:BRc

[Ano98c] Anonymous. Book reviews. Annals of Science, 55(3):319–328, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:BRd

[Ano98d] Anonymous. Book reviews. Annals of Science, 55(4):421–448, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:ERa

[Ano98e] Anonymous. Essay review. Annals of Science, 55(1):95–103, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1998:ERb

[Ano98f] Anonymous. Essay review. Annals of Science, 55(3):315–318, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1999:BRa

[Ano99a] Anonymous. Book reviews. Annals of Science, 56(1):99–109, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1999:BRb

[Ano99b] Anonymous. Book reviews. Annals of Science, 56(2):211–219, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1999:BRc

[Ano99c] Anonymous. Book reviews. Annals of Science, 56(3):317–329, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1999:BRd

[Ano99d] Anonymous. Book reviews. Annals of Science, 56(4):461–464, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:1999:ERS

[Ano99e] Anonymous. Essay review: Suspension and belief. Annals of Science, 56(2):205–209, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2000:BRa

[Ano00a] Anonymous. Book reviews. Annals of Science, 57(1):101–108, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2000:BRb

[Ano00b] Anonymous. Book reviews. Annals of Science, 57(2):197–217, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2000:BRc

[Ano00c] Anonymous. Book reviews. Annals of Science, 57(3):307–327, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2000:BRd

[Ano00d] Anonymous. Book reviews. Annals of Science, 57(4):455–462, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2001:BRa

[Ano01a] Anonymous. Book reviews. Annals of Science, 58(1):93–109, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2001:BRb

[Ano01b] Anonymous. Book reviews. Annals of Science, 58(2):211–218, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2001:BRc

[Ano01c] Anonymous. Book reviews. Annals of Science, 58(3):323–326, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2001:BRd

[Ano01d] Anonymous. Book reviews. Annals of Science, 58(4):423–452, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2001:ER

[Ano01e] Anonymous. Essay review. Annals of Science, 58(2):203–209, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2002:ASP

[Ano02a] Anonymous. Annals of science prize. Annals of Science, 59(1):109, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2002:BRa

[Ano02b] Anonymous. Book review. Annals of Science, 59(1):97–108, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2002:BRb

[Ano02c] Anonymous. Book reviews. Annals of Science, 59(2):211–219, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2002:BRc

[Ano02d] Anonymous. Book reviews. Annals of Science, 59(3):327–329, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2002:BRd

[Ano02e] Anonymous. Book reviews. Annals of Science, 59(4):423–446, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:ASP

[Ano03a] Anonymous. The annals of science prize. Annals of Science, 60(4): 459, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:BRa

[Ano03b] Anonymous. Book reviews. Annals of Science, 60(1):95–106, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:BRb

[Ano03c] Anonymous. Book reviews. Annals of Science, 60(2):207–218, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:BRc

[Ano03d] Anonymous. Book reviews. Annals of Science, 60(4):437–458, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:ER

[Ano03e] Anonymous. Essay review. Annals of Science, 60(1):85–94, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2003:MVH

[Ano03f] Anonymous. A Manichean view of the history of geology. Annals of Science, 60(4):423–436, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:BRa

[Ano04a] Anonymous. Book reviews. Annals of Science, 61(1):127–139, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:BRb

[Ano04b] Anonymous. Book reviews. Annals of Science, 61(2):227–269, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:BRc

[Ano04c] Anonymous. Book reviews. Annals of Science, 61(3):393–406, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:BRd

[Ano04d] Anonymous. Book reviews. Annals of Science, 61(4):495–520, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:ERR

[Ano04e] Anonymous. Essay review: Romanticism and science. Annals of Science, 61(4):483–487, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2004:ERW

[Ano04f] Anonymous. Essay review: The way of Galton in the world. Annals of Science, 61(4):489–493, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:B

[Ano05a] Anonymous. Biographies. Annals of Science, 62(1):121–139, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:BRa

[Ano05b] Anonymous. Book reviews. Annals of Science, 62(2):265–279, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:BRb

[Ano05c] Anonymous. Book reviews. Annals of Science, 62(3):387–415, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:E

[Ano05d] Anonymous. Erratum. Annals of Science, 62(3):417, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:ERJ

[Ano05e] Anonymous. Essay review: Janis Langins on the Corps Royal du Génie Militaire. Annals of Science, 62(2):255–263, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:ERH

[Ano05f] Anonymous. Essay review: The historicity of sound and hearing. Annals of Science, 62(1):103–105, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:ERW

[Ano05g] Anonymous. Essay review: What does science have to do with music? Annals of Science, 62(1):107–120, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:RLS

[Ano05h] Anonymous. Romantic life and science. Annals of Science, 62(3): 377–385, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:TE

[Ano05i] Anonymous. Technology and engineering. Annals of Science, 62(4): 539–554, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2005:HSC

[Ano05j] Anonymous. To have, to hold, and to share: Compiling scientific knowledge in the information age of enlightenment. Annals of Science, 62(4):533–538, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:AP

[Ano06a] Anonymous. Artful physics. Annals of Science, 63(3):363–370, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:BRa

[Ano06b] Anonymous. Book reviews. Annals of Science, 63(1):119–137, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:BRb

[Ano06c] Anonymous. Book reviews. Annals of Science, 63(2):235–252, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:BRc

[Ano06d] Anonymous. Book reviews. Annals of Science, 63(3):377–393, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:EOL

[Ano06e] Anonymous. Ex oriente lux mundi. Annals of Science, 63(4):483– 492, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:EEE

[Ano06f] Anonymous. Exploration, ecology and Earth and life sciences. Annals of Science, 63(4):503–505, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:GWU

[Ano06g] Anonymous. Gravitational waves under the microscope. Annals of Science, 63(3):371–375, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:MDL

[Ano06h] Anonymous. ‘Models’: Diverse but limited. Annals of Science, 63 (2):227–234, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:NTN

[Ano06i] Anonymous. New trends in Newtonian scholarship: Benefits and pitfalls. Annals of Science, 63(1):111–117, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:RFL

[Ano06j] Anonymous. The remarkable fecundity of Leibniz’s work on infinite series. Annals of Science, 63(2):221–225, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:SY

[Ano06k] Anonymous. Seventy years. Annals of Science, 63(1):1–2, 2006. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2006:GRE

[Ano06l] Anonymous. ‘the geohistorical revolution’: The emergence of geology as an historical science. Annals of Science, 63(4):493–501, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:BRa

[Ano07a] Anonymous. Book reviews. Annals of Science, 64(1):113–146, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:BRb

[Ano07b] Anonymous. Book reviews. Annals of Science, 64(2):281–291, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:BRc

[Ano07c] Anonymous. Book reviews. Annals of Science, 64(3):421–442, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:BRd

[Ano07d] Anonymous. Book reviews. Annals of Science, 64(4):607–620, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:ERB

[Ano07e] Anonymous. Essay review: A Lot About a Little Bit of England. Annals of Science, 64(3):443, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:LAL

[Ano07f] Anonymous. A lot about a little bit of England. Annals of Science, 64(2):251–279, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:MHS

[Ano07g] Anonymous. Making the history of science matter in Ireland. Annals of Science, 64(1):107–112, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:PL

[Ano07h] Anonymous. Priestley and Lavoisier. Annals of Science, 64(4):595– 605, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:SSE

[Ano07i] Anonymous. Seeds sprouting everywhere. Annals of Science, 64(3): 411–420, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:SMH

[Ano07j] Anonymous. State of the medical historical art versus science. Annals of Science, 64(1):101–106, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2007:STA

[Ano07k] Anonymous. (stop) talking about Victorian science. Annals of Science, 64(1):93–100, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:BC

[Ano08a] Anonymous. Blessed are the cheesemakers. Annals of Science, 65(3): 445–451, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:BRa

[Ano08b] Anonymous. Book reviews. Annals of Science, 65(1):157–170, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:BRb

[Ano08c] Anonymous. Book reviews. Annals of Science, 65(2):287–308, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:BRc

[Ano08d] Anonymous. Book reviews. Annals of Science, 65(4):579–597, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:CDA

[Ano08e] Anonymous. Cultures of dissection and anatomies of generation. Annals of Science, 65(3):439–444, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:HCA

[Ano08f] Anonymous. History of chemistry and alchemy. Annals of Science, 65(3):461–464, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:MH

[Ano08g] Anonymous. Medicine and health. Annals of Science, 65(3):455– 457, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:NHL

[Ano08h] Anonymous. The ‘new historiography’ and the limits of alchemy. Annals of Science, 65(1):127–156, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:SB

[Ano08i] Anonymous. On sociological biographies. Annals of Science, 65(4): 573–578, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:PS

[Ano08j] Anonymous. Physical sciences. Annals of Science, 65(3):458–460, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:SAS

[Ano08k] Anonymous. Social aspects of science. Annals of Science, 65(3):453– 455, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:TIA

[Ano08l] Anonymous. Texts, illustrations and artifacts: Intertextuality in the transmission of scientific knowledge. Annals of Science, 65(2):281– 285, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2008:VSS

[Ano08m] Anonymous. Voyaging in strange seas of thought. Annals of Science, 65(2):275–279, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:AI

[Ano09a] Anonymous. Author index. Annals of Science, 66(1S):5–77, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:ASC

[Ano09b] Anonymous. Annals of Science cumulative index volumes 44 to 64 (1987–2007). Annals of Science, 66(1S):3–4, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:C

[Ano09c] Anonymous. Corrigendum. Annals of Science, 66(2):305, 2009. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:E

[Ano09d] Anonymous. Erratum. Annals of Science, 66(1):153, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:RAI

[Ano09e] Anonymous. Reviewed authors index. Annals of Science, 66(1S):79– 140, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:SI

[Ano09f] Anonymous. Subject index. Annals of Science, 66(1S):141–269, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:TAW

[Ano09g] Anonymous. Theory and action in the works of Andreas Libavius and other alchemists. Annals of Science, 66(1):135–139, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:TAP

[Ano09h] Anonymous. A time of air and place. Annals of Science, 66(1):129– 134, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2009:TCT

[Ano09i] Anonymous. A tulip for a cup of tea? Commerce and nature in the Dutch Golden Age. Annals of Science, 66(2):267–276, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2010:CI

[Ano10a] Anonymous. Cover image. Annals of Science, 67(1):e1, 2010. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2010:CPP

[Ano10b] Anonymous. Cover page pdf. Annals of Science, 67(3):e1, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2010:FCPa

[Ano10c] Anonymous. Front cover page. Annals of Science, 67(2):e1, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2010:FCPb

[Ano10d] Anonymous. Front cover page. Annals of Science, 67(4):e1, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2011:FCPa

[Ano11a] Anonymous. Front cover page. Annals of Science, 68(1):e1, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2011:FCPb

[Ano11b] Anonymous. Front cover page. Annals of Science, 68(2):e1, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2011:FCPc

[Ano11c] Anonymous. Front cover page. Annals of Science, 68(3):e1, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2011:NHM

[Ano11d] Anonymous. Navigating the history of mathematics. Annals of Science, 68(3):425–433, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2011:OC

[Ano11e] Anonymous. Online cover. Annals of Science, 68(4):ebi, 2011. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2012:CER

[Ano12a] Anonymous. Corrigendum: Essay review: Einstein’s quest for unity. Annals of Science, 69(4):593, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Del12, Sau14, JL14].

Anonymous:2012:OC

[Ano12b] Anonymous. Online cover. Annals of Science, 69(1):e1, 2012. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2012:OCFa

[Ano12c] Anonymous. Online cover figure. Annals of Science, 69(2):ebi, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2012:OCFb

[Ano12d] Anonymous. Online cover figure. Annals of Science, 69(3):e1, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2012:OCFc

[Ano12e] Anonymous. Online cover figure. Annals of Science, 69(4):e1, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2012:WAS

[Ano12f] Anonymous. Winner of the Annals of Science Prize for 2011. Annals of Science, 69(2):299, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2013:A

[Ano13a] Anonymous. Announcement. Annals of Science, 70(1):147, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2013:E

[Ano13b] Anonymous. Erratum. Annals of Science, 70(4):594, 2013. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [GG13a].

Anonymous:2013:OCa

[Ano13c] Anonymous. Online cover. Annals of Science, 70(2):e1, 2013. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2013:OCb

[Ano13d] Anonymous. Online cover 70:3. Annals of Science, 70(3):e1, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2013:OCc

[Ano13e] Anonymous. Online cover 70:4. Annals of Science, 70(4):e1, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2014:EB

[Ano14a] Anonymous. Editorial board. Annals of Science, 71(4):eb1, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2014:OCb

[Ano14b] Anonymous. Online cover. Annals of Science, 71(2):e1, 2014. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2014:OC

[Ano14c] Anonymous. Online cover. Annals of Science, 71(3):e1, 2014. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2014:OCa

[Ano14d] Anonymous. Online cover 71:1. Annals of Science, 71(1):e1, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2015:DRO

[Ano15] Anonymous. David Roger Oldroyd (20 January 1936–7 November 2014): An appreciation. Annals of Science, 72(1):1, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2016:E

[Ano16a] Anonymous. Erratum. Annals of Science, 73(4):iii, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2016:WTA

[Ano16b] Anonymous. Winner of the Annals of Science Prize for 2015. Annals of Science, 73(3):344–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2018:WAS

[Ano18] Anonymous. Winner of the Annals of Science Prize for 2017. Annals of Science, 75(3):273, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Anonymous:2019:WAS

[Ano19] Anonymous. Winner of the Annals of Science Prize for 2018. Annals of Science, 76(2):239, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Antognazza:2010:P

[Ant10] Maria Rosa Antognazza. Protogaea. Annals of Science, 67(2):281– 283, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Apel:2020:RPW

[Ape20] Thomas Apel. ‘Revolutions, philosophical as well as civil’: French chemistry and American science in Samuel Latham Mitchill’s Medical Repository. Annals of Science, 77(2):189–214, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Appleby:1990:EKE

[App90] John H. Appleby. Erasmus King: Eighteenth-century experimental philosopher. Annals of Science, 47(4):375–392, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Appleby:1996:NPJ

[App96] John H. Appleby. A new perspective on John Rowley, Virtuoso Master of mechanics and hydraulic engineer. Annals of Science, 53(1): 1–27, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Archinard:1990:DUH

[Arc90] Margarida Archinard. The diagram of unequal hours. Annals of Science, 47(2):173–190, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Archinard:2005:CSA

[Arc05] Margarida Archinard. Les Cadrans Solaires Analemmatiques. (French) [the analemmatic sundials]. Annals of Science, 62(3):309– 346, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Archinard:2007:CCS

[Arc07] Margarida Archinard. Une classification des cadrans solaires. Annals of Science, 64(4):471–524, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Arens:2010:VAU

[Are10] Katherine Arens. Vienna in the Age of Uncertainty: Science, liberalism, and private life. Annals of Science, 67(2):285–288, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1972:TFM

[Ari72] Piero Ariotti, Ph.D. From the top to the foot of a mast on a moving ship. Annals of Science, 28(2):191–203, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1973:LKE

[Ari73a] Piero E. Ariotti, Ph.D. A little known early seventeenth century treatise on vision: Benedetto Castelli’s Discorso sopra la Vista (1639, 1669). Translation and critical comments. Annals of Science, 30(1):

1–30, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1973:TAT

[Ari73b] Piero E. Ariotti, Ph.D. Toward absolute time: The undermining and refutation of the Aristotelian conception of time in the sixteenth and seventeenth centuries. Annals of Science, 30(1):31–50, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1974:OAL

[Ari74] Piero E. Ariotti. An overlooked autograph letter of Galileo on the thermometer. Annals of Science, 31(5):457–462, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1979:CHA

[Ari79] Piero E. Ariotti. Christiaan Huygens: Aviation pioneer extraordinary. Annals of Science, 36(6):611–624, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ariotti:1981:BCB

[Ari81] Piero E. Ariotti. Benedetto Castelli’s Discourse on the loadstone (1639–1640): the origin of the notion of elementary magnets similarly aligned. Annals of Science, 38(2):125–140, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1947:DFB

[Arm47] Angus Armitage, M.Sc. The deviation of falling bodies. Annals of Science, 5(4):342–351, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1948:CGB

[Arm48] Angus Armitage, M.Sc., Ph.D. The cosmology of Giordano Bruno. Annals of Science, 6(1):24–31, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1950:BHR

[Arm50] Angus Armitage, M.Sc., Ph.D. ‘Borell’s hypothesis’ and the rise of celestial mechanics. Annals of Science, 6(3):268–282, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1951:MGD

[Arm51] Angus Armitage, M.Sc., Ph.D. Master Georg Dörffel and the rise of cometary astronomy. Annals of Science, 7(4):303–315, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1953:PPA

[Arm53] Angus Armitage, M.Sc., Ph.D. The pilgrimage of Pingré: An astronomer–monk of eighteenth-century France. Annals of Science, 9(1):47–63, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1954:CAP

[Arm54] Angus Armitage, M.Sc., Ph.D. Chappe d’Auteroche: a pathfinder for astronomy. Annals of Science, 10(4):277–293, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armytage:1955:TUJ

[Arm55] W. H. G. Armytage. Technology and Utopianism: J. A. Etzler in England 1840–44. Annals of Science, 11(2):129–136, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1956:AWN

[Arm56a] Angus Armitage, M.Sc., Ph.D. The astronomical work of Nicolas-Louis de Lacaille. Annals of Science, 12(3):163–191, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armytage:1956:EUS

[Arm56b] W. H. G. Armytage, M.A. The early Utopists and science in England. Annals of Science, 12(4):247–254, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armytage:1956:JPE

[Arm56c] W. H. G. Armytage, M.A., Ph.D. Joseph Priestley and Edmund Burke: An unpublished letter. Annals of Science, 12(2):160–161, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armitage:1958:NVL

[Arm58a] Angus Armitage, M.Sc., Ph.D. A naturalist’s vacation. The London letters of J. C. Fabricius. Annals of Science, 14(2):116–131, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Armytage:1958:RWM

[Arm58b] W. H. G. Armytage, M.A. Richard Watson and the Marquess of Rockingham: An unpublished exchange in 1771. Annals of Science, 14(3):155–156, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Arnold:2009:IIS

[Arn09] David Arnold. Indian ink: Script and print in the making of the English East India Company. Annals of Science, 66(3):448–451, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aiton:1981:GKS

[AS81] E. J. Aiton and Eikoh Shimao. Gorai Kinzō’s study of Leibniz and the I ching hexagrams. Annals of Science, 38(1):71–92, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ashcroft:2017:MAS

[Ash17] Richard E. Ashcroft. Multiple autisms: Spectrums of advocacy and genomic science. Annals of Science, 74(3):248–250, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aspden:1969:PE

[Asp69] Harold Aspden. Physics without Einstein. Sabberton Publications, Southampton, UK, 1969. ISBN 0-85056-001-2. xiii + 224 pp. LCCN QC75 .A838p 1969. See critical review [For71a].

Aeit-Touati:2011:FCS

[ATE11] Frédérique Aèit-Touati and Susan Emanuel. Fictions of the cosmos: science and literature in the seventeenth century. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-01122-4 (hardcover). ix + 261 pp. LCCN PN56.C68 A5813 2011. Translated by Susan Emanuel.

Atkinson:1952:WDF

[Atk52] A. D. Atkinson, M.A. William Derham, F.R.S. (1657–1735). Annals of Science, 8(4):368–392, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Attfield:1985:BBD

[Att85] Robin Attfield. Balthasar Bekker and the decline of the Witch-Craze: The old demonology and the new philosophy. Annals of Science, 42 (4):383–395, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Attewell:2009:MIM

[Att09] Guy Attewell. Medieval Islamic medicine. Annals of Science, 66(4): 559–561, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Aubin:2023:PPC

[Aub23] David Aubin. Popularizing precision: cultures of exactness at the Paris Observatory, 1667–1742. Annals of Science, 81(1–2):139–159, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2282783.

Auger:2004:RRU

[Aug04] Jean-François Auger. Le régime de recherche utilitaire du professeurconsultant au cours de la seconde Révolution industrielle. (French) [the utilitarian research program of the professor consultant in the Second Industrial Revolution]. Annals of Science, 61(3):351–374, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Auger:2011:RDC

[Aug11] J.-F. Auger. The rise and decline of colloid science in North America, 1900–1935: The neglected dimension. Annals of Science, 68(3):435– 437, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

AvelardeCarvalho:2020:PIS

[Ave20] Helena Avelar de Carvalho. Preludes to the Inquisition: selfcensorship in medieval astrological discourse. Annals of Science, 77 (1):10–25, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Avicenna:2006:LPN

[AVJ^{+}06] Avicenna, S. Van Riet, Jules L. Janssens, André Allard, and Gérard Verbeke. Liber primus Naturalium: tractatus secundus, De motu et de consimilibus. Académie royale de Belgique, Bruxelles, Belgique, 2006. ISBN 2-8031-0231-5. 87, 148–376 pp. LCCN ????

Badcock:1960:PRS

[Bad60] A. W. Badcock, M.Sc., Ph.D. Physics at the Royal Society, 1660– 1800. I. Change of state. Annals of Science, 16(2):95–115, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Badash:2009:NWT

[Bad09] Lawrence Badash. A nuclear winter’s tale: science and politics in the 1980s. Transformations: studies in the history of science. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01272-3 (hardcover). xiii + 403 pp. LCCN QC794.98 .B325 2009. URL http://www.loc.

gov/catdir/toc/ecip0827/2008039040.html.

Badash:2011:TMG

[Bad11] Lawrence Badash. Trafficking materials and gendered experimental practices: Radium research in early 20th Century Vienna. Annals of Science, 68(4):571–573, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baird:1993:ACB

[Bai93] Davis Baird. Analytical chemistry and the ‘big’ scientific instrumentation revolution. Annals of Science, 50(3):267–290, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baker:1962:UEC

[Bak62] K. M. Baker, B.A. An unpublished essay of Condorcet on technical methods of classification. Annals of Science, 18(2):99–123, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baker:1964:EHT

[Bak64] K. M. Baker, M.A., Ph.D. The early history of the term ‘social science’. Annals of Science, 20(3):211–226, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baker:2021:TSM

[Bak21] Alexi Baker. Thrifty science: making the most of materials in the history of experiment. Annals of Science, 77(3):392–394, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bala:2006:DCB

[Bal06] Arun Bala. The dialogue of civilizations in the birth of modern science. Palgrave Macmillan, New York, NY, USA, 2006.

ISBN 1-4039-7468-3, 0-230-60121-9 (e-book). xii + 230 pp.

LCCN Q125 .B2925 2006. URL http://www.loc.gov/catdir/
enhancements/fy0642/2006043225-b.html;   http://www.loc.
gov/catdir/enhancements/fy0642/2006043225-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0642/2006043225-t.html.

Baldi:2023:HEC

[Bal23] Rossella Baldi. How to ensure a chronometer’s accuracy. Josiah Emery timekeepers and their users. Annals of Science, 81(1–2): 189–207, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2282787.

Banerjee:2014:BRB

[Ban14] Somaditya Banerjee. Book review: Nucleus and Nation: Scientists, International Networks, and Power in India, by Robert S. Anderson. Annals of Science, 71(1):107–110, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barnett:1957:THE

[Bar57] Pamela R. Barnett, M.A. Theodore Haak and the early years of the Royal Society. Annals of Science, 13(4):205–218, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barkla:1973:BRR

[Bar73] H. M. Barkla, M.A.B.Sc. Benjamin Robins and the resistance of air. Annals of Science, 30(1):107–122, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bartholomew:1975:HDD

[Bar75] Michael Bartholomew. Huxley’s defence of Darwin. Annals of Science, 32(6):525–535, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barth:1995:HWA

[Bar95] Michael Barth. Huygens at work: Annotations in his rediscovered personal copy of Hooke’s Micrographia. Annals of Science, 52(6):601– 613, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barton:1998:JBN

[Bar98] Ruth Barton. Just before Nature: The purposes of science and the purposes of popularization in some English popular science journals of the 1860s. Annals of Science, 55(1):1–33, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barnett:2009:WWH

[Bar09] Richard Barnett. From witchcraft to wisdom: a history of obstetrics and gynaecology in the British Isles. Annals of Science, 66(4):561– 563, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barnett:2011:BRJ

[Bar11] Richard Barnett. Book review: John S. Partington, editor, H. G. Wells in Nature. A Reception Reader. ISBN-13 978-3-631-57110-1. Annals of Science, 68(2):288–290, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barnett:2014:BRB

[Bar14] Richard Barnett. Book review: Evolutionary Restraints: The Contentious History of Group Selection, by Mark E. Borrello. Annals of Science, 71(1):106–107, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Basu:1995:CRI

[Bas95] Aparajito Basu. Chemical research in India (1876–1918). Annals of Science, 52(6):591–600, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bashford:2015:BRB

[Bas15] Alison Bashford. Book review: Political Descent: Malthus, Mutualism, and the Politics of Evolution in Victorian England. Annals of Science, 72(3):411–413, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baumgartner:1988:GFC

[Bau88] Frederic J. Baumgartner. Galileo’s French correspondents. Annals of Science, 45(2):169–182, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bayon:1938:WHPb

[Bay38a] H. P. Bayon, M.D., Ph.D. William Harvey, physician and biologist: His precursors, opponents and successors. Part III. Annals of Science, 3(4):435–456, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bayon:1938:WHPa

[Bay38b] H. P. Bayon, M.D., Ph.D. William Harvey, physician and biologist: His precursors, opponents and successors. Parts I & II. Annals of Science, 3(1):59–118, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bayon:1939:WHPa

[Bay39a] H. P. Bayon, M.D., Ph.D. William Harvey, physician and biologist: His precursors, opponents and successors. Part IV. Annals of Science, 4(1):65–106, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bayon:1939:WHPb

[Bay39b] H. P. Bayon, M.D., Ph.D. William Harvey, physician and biologist: His precursors, opponents and successors. Part V. Annals of Science, 4(4):329–389, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bedini:1995:BTO

[BB95] Silvio A. Bedini and Arthur G. Bennett. A treatise on optics by Giovanni Christoforo Bolantio. Annals of Science, 52(2):103–126, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brack-Bernsen:2013:WHC

[BB13] Lis Brack-Bernsen. What and how can we learn from the Babylonian astronomical compendium mul.apin? [book review: Writing science before the Greeks: a naturalistic analysis of the Babylonian astronomical treatise MUL.APIN]. Annals of Science, 70(2):285–289, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bredekamp:2011:G

[BBHN14] Horst Bredekamp, Irene Brückle, Oliver Hahn, and Paul Needham, editors. Galileo’s O. Akademie-Verlag, Berlin, Germany, 2011–2014. ISBN 3-05-005095-0 (set), 3-11-035464-0 (vol. 3), 3-11-035478-0 (vol. 3). ???? pp. LCCN Z156.V4 G35 2011.

Benschop:2000:PPC

[BD00] Ruth Benschop and Douwe Draaisma. In pursuit of precision: The calibration of minds and machines in late nineteenth-century psychology. Annals of Science, 57(1):1–25, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bearman:1976:SSH

[Bea76] David Bearman. Survey of sources for the history of biochemistry and molecular biology: an introductory report. Annals of Science, 33(4):395–397, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beatty:2012:NDL

[Bea12] Heather R. Beatty. Nervous Disease in Late Eighteenth-Century Britain: The Reality of a Fashionable Disorder, volume 6. Pickering and Chatto, London, UK, 2012. ISBN 1-283-29204-1, 1-84893-308-8 (e-book), 1-84893-309-6 (e-book). vii + 241 pp. LCCN

RC530 .B375 2012. URL http://universitypublishingonline.
org/pickeringchatto/ebook.jsf?bid=CBO9781848933095.

Beck:1984:LMS

[Bec84] Daniel A. Beck. Life on the Moon? A short history of the Hansen hypothesis. Annals of Science, 41(5):463–470, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beckers:2001:DIF

[Bec01] Danny J. Beckers. Definitely infinitesimal: Foundations of the calculus in The Netherlands, 1840–1870. Annals of Science, 58(1):1–15, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Becvarova:2009:HFE

[Bec09] Martina Becvárová. How to fix an election honestly! Ivan Petrov Salabashev’s novel voting procedure in Bulgaria, 1879–1880. Annals of Science, 66(3):397–406, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Becker:2011:USW

[Bec11] Barbara J. Becker. Unravelling starlight: William and Margaret Huggins and the rise of the new astronomy. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-00229-X (hardcover). xix + 380

pp. LCCN QB35 .B43 2011. URL http://assets.cambridge.org/
97811070/02296/cover/9781107002296.jpg.

Becker:2013:BRBb

[Bec13a] Barbara J. Becker. Book review: An Observer of Observatories: The Journal of Thomas Bugge’s Tour of Germany, Holland and England in 1777, edited by Kurt Møller Pedersen and Peter De Clercq. Annals of Science, 70(4):570–573, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Becker:2013:BRBa

[Bec13b] Barbara J. Becker. Book review: Observing and Cataloguing Nebulae and Star Clusters, by Wolfgang Steinicke. Annals of Science, 70(4): 569–570, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Becker:2014:BRB

[Bec14] Barbara J. Becker. Book review: From Earth-Bound to Satellite: Telescopes, Skills and Networks. Annals of Science, 71(4):579–581, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beckman:2016:PSJ

[Bec16] Jenny Beckman. The publication strategies of Jöns Jacob Berzelius (1779–1848): negotiating national and linguistic boundaries in chemistry. Annals of Science, 73(2):195–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bedini:1982:OML

[Bed82] Silvio A. Bedini. Obituary: Maria Luisa Righini Bonelli 1917–1981. Annals of Science, 39(4):327–328, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bedini:1993:SLS

[Bed93] Silvio A. Bedini. Of ‘science and liberty’: The scientific instruments of King’s College and eighteenth century Columbia College in New York. Annals of Science, 50(3):201–227, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bedini:1997:TTE

[Bed97] Silvio A. Bedini. The transit in the tower: English astronomical instruments in Colonial America. Annals of Science, 54(2):161–196, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Begley:2021:JHP

[Beg21] Justin Begley. John Hill (1714?–1775) on ‘plant sleep’: experimental physiology and the limits of comparative analysis. Annals of Science, 78(1):41–63, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bellmer:1999:SOG

[Bel99] Elizabeth Henry Bellmer. The statesman and the ophthalmologist: Gladstone and Magnus on the evolution of human colour vision, one small episode of the nineteenth-century Darwinian debate. Annals of Science, 56(1):25–45, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Belhoste:2012:PCA

[Bel12] Bruno Belhoste. A Parisian craftsman among the savants: The joiner André-Jacob Roubo (1739–1791) and his works. Annals of Science, 69(3):395–411, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bentham:1937:SSC

[Ben37] Muriel A. Bentham, M.Sc. Some seventeenth century views concerning the nature of heat and cold. Annals of Science, 2(4):431–450, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bennett:1973:SBT

[Ben73] J. A. Bennett, B.A. A study of Parentalia, with two unpublished letters of Sir Christopher Wren. Annals of Science, 30(2):129–147, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Benham:1974:GAA

[Ben74] W. Benham. The Gauss anagram: An alternative solution. Annals of Science, 31(5):449–455, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bennett:1980:GBA

[Ben80] J. A. Bennett. George Biddell Airy and horology. Annals of Science, 37(3):269–285, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bennett:1991:GSE

[Ben91] J. A. Bennett. Geometry and surveying in early-seventeenth-century England. Annals of Science, 48(4):345–354, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bennett:1995:ORW

[Ben95] J. A. Bennett. Obituary: R. E. W. Maddison (1901–93). Annals of Science, 52(3):306, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bennett:2009:AGS

[Ben09] Jim Bennett. Après Gassendi: son influence et sa réputation, essai, avec l’histoire des collections scientifiques et un catalogue des instruments et appareils concernant les sciences exactes appartenant au Musée Gassandi à Digne-les-Bains. (French) [After Gassendi: his influence and reputation, with the history of the scientific collections and an catalog of instruments and tools on the exact sciences belonging to the Gassandi Museum in Digne-les-Bains]. Annals of Science, 66(4):565–566, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Benbow:2013:NIT

[Ben13a] Peter K. Benbow. Naming infinity: A true story of religious mysticism and mathematical creativity. Annals of Science, 70(3):431–434, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Benson:2013:BRB

[Ben13b] Etienne Benson. Book review: Considering Animals: Contemporary Studies in Human-Animal Relations, edited by Carol Freeman, Elizabeth Leane, and Yvette Watt. Annals of Science, 70(4):588–590, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Benbow:2014:BRB

[Ben14] Peter K. Benbow. Book review: Mistical Wordes and Names Infinite: An Edition and Study of Humfrey Lock’s Treatise on Alchemy, by Peter J. Grund. Annals of Science, 71(2):289–291, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berridge:1979:MMS

[Ber79] Virginia Berridge. Morality and medical science: Concepts of narcotic addiction in Britain, 1820–1926. Annals of Science, 36(1):67–85, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berger:1996:HDT

[Ber96] Michael Berger. Henry David Thoreau’s science in The Dispersion of Seeds. Annals of Science, 53(4):381–397, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berger:1997:CMK

[Ber97] Jutta Berger. Chemische Mechanik und Kinetik: die Bedeutung der mechanischen Wärmetheorie für die Theorie chemischer Reaktionen. (German) [Chemical mechanism and kinetics: the importance of the mechanical theory of heat for the theory of chemical reactions]. Annals of Science, 54(6):567–584, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beretta:2001:LHL

[Ber01] Marco Beretta. Lavoisier and his last printed work: The Mémoires de physique et de chimie (1805). Annals of Science, 58(4):327–356, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beretta:2009:TE

[Ber09] Marco Beretta. Technology and engineering. Annals of Science, 66 (1):150–151, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beretta:2012:RFG

[Ber12] Marco Beretta. The rise and fall of the glassmaker Paul Bosc d’Antic (1753–1784). Annals of Science, 69(3):375–393, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berger:2019:HDT

[Ber19] Michael Berger. Henry David Thoreau: A life. Annals of Science, 76(1):100–103, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Best:1968:DMC

[Bes68] A. E. Best, M.A., Ph.D., B.Litt. The discovery of the mechanism of colour-changes in the chameleon. Annals of Science, 24(2):147– 167, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bewell:2016:NTR

[Bew16] Alan Bewell. Natures in Translation: Romanticism and Colonial Natural History. The Johns Hopkins University Press, Baltimore, MD, USA, 2016. ISBN 1-4214-2096-1 (hardcover), 1-4214-2097-X (electronic). xvii + 393 pp. LCCN PR468.N3 B49 2016.

Beyler:2010:WLB

[Bey10] Richard H. Beyler. William and Lawrence Bragg, father and son: The most extraordinary collaboration in science. Annals of Science, 67(1):137–139, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beyler:2011:WCC

[Bey11] Richard H. Beyler. William Crookes (1832–1919) and the commercialization of science. Annals of Science, 68(2):284–286, 2011. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Beyler:2014:BRB

[Bey14] Richard H. Beyler. Book review: Science Secrets: The Truth about Darwin’s Finches, Einstein’s Wife and, Other Myths, by Alberto A. Martínez. Annals of Science, 71(2):271–273, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bezel:1975:ATC

[Bez75] Nail Bezel. Autobiography and ‘the two cultures’ in the novels of C. P. Snow. Annals of Science, 32(6):555–571, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barker:1994:DVK

[BG94] Peter Barker and Bernard R. Goldstein. Distance and velocity in Kepler’s astronomy. Annals of Science, 51(1):59–73, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Barrow-Green:1999:CST

[BG99] June Barrow-Green. ’A corrective to the spirit of too exclusively pure mathematics’: Robert Smith (1689–1768) and his prizes at Cambridge University. Annals of Science, 56(3):271–316, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Biggs:2021:LPF

[Big21] Norman Biggs. Linear programming from Fibonacci to Farkas. Annals of Science, 78(1):1–21, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bigotti:2022:GDI

[Big22] Fabrizio Bigotti. Gradus Dimetiri: intensity and classification of complexions in 14th-century Italian medicine. Annals of Science, 79(4):419–441, 2022. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2022.2107702.

Bikerman:1965:BHG

[Bik65] J. J. Bikerman. A book Humboldt gave to Schiller. Annals of Science, 21(2):131–134, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bikerman:1971:FML

[Bik71] J. J. Bikerman. Forced marriage and love union: a romantic affair in the history of Russian chemistry. Annals of Science, 27(2):201–204, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bikerman:1975:PLL

[Bik75] J. J. Bikerman. Payment of the learned, and lament of a moneyed, man. History of the commotion created by two letters to Science 25 years ago. Annals of Science, 32(6):573–577, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See E. Bodewig, Science, 112 530 (1950) and 113 330 (1951).

Bilgili:2016:BRS

[Bil16] Alper Bilgili. Book review: Science Among the Ottomans: The Cultural Creation and Exchange of Knowledge, by Miri Shefer-Mossensohn, University of Texas Press, Austin (2015), ISBN 978-1-4773-0359-7. Annals of Science, 73(4):449–450, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bil:2018:BRN

[Bil18] Geoff Bil. Book review: Natures in Translation: Romanticism and Colonial Natural History, by Alan Bewell, Baltimore: Johns Hopkins University Press, 2016, xvii + 393 pp., ISBN 978-1-4214-2096-7. Annals of Science, 75(1):64–65, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Binder:2012:BRM

[Bin12] Christa Binder. Book review: Mathematics and logic. Kirsti Andersen, The Geometry of an Art, The History of the Mathematical Theory of Perspective from Alberti to Monge. Sources and Studies in the History of Mathematics and Physical Sciences, ISBN-10 0-387-25961-9. Annals of Science, 69(2):291–294, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brandon-Jones:1996:CDR

[BJ96] Christine Brandon-Jones. Charles Darwin and the repugnant curators. Annals of Science, 53(5):501–510, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berggren:2000:PGA

[BJ00] J. L. Berggren and Alexander Jones. Ptolemy’s Geography: an annotated translation of the theoretical chapters. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-01042-0, 0-691-09259-1 (paperback). xiii + 192 + 8 pp. LCCN GA881.1 .B47 2000.

Boudia:2013:THR

[BJ13] Soraya Boudia and Nathalie Jas, editors. Toxicants, health and regulation since 1945, volume 9 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-403-3 (hardcover), 1-78144-035-2 (e-book). xiii + 193 pp. LCCN RA1219.3 .T69 2013. URL http://

universitypublishingonline.org/pickeringchatto/ebook.jsf?
bid=CBO9781781440353.

Brock:2021:DKN

[BL21] William H. Brock and David E. Lewis. A different kind of Nierenstein reaction. The Chemical Society’s mistreatment of Maximilian Nierenstein. Annals of Science, 78(2):221–245, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Blackmore:1978:TAM

[Bla78] John T. Blackmore. Three autobiographical manuscripts by Ernst Mach. Annals of Science, 35(4):401–418, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Blay:1988:VOT

[Bla88] Michel Blay. Varignon ou la théorie du mouvement des projectiles ‘comprise en une proposition générale. Annals of Science, 45(6):591– 618, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Blank:2010:MSI

[Bla10] Andreas Blank. Material souls and imagination in Late Aristotelian embryology. Annals of Science, 67(2):187–204, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Blank:2019:ICN

[Bla19] Andreas Blank. Instrumental causes and the natural origin of souls in Antonio Ponce Santacruz’s theory of animal generation. Annals of Science, 76(2):184–209, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Blum:2013:ERT

[Blu13] Paul Richard Blum. Essay review: Theories of life in the Renaissance. Annals of Science, 70(4):539–543, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boato:1994:BRF

[BM94] Giovanni Boato and Natalia Moro. Bancalari’s role in Faraday’s discovery of diamagnetism and the successive progress in the understanding of magnetic properties of matter. Annals of Science, 51(4): 391–412, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Braunack-Mayer:2017:HDS

[BM17] Lydia Braunack-Mayer. How to do science with models: A philosophical primer. Annals of Science, 74(3):250–252, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2008:PRH

[Boa08] Victor Boantza. The phlogistic role of heat in the chemical revolution and the origins of Kirwan’s Ingenious Modifications… Into the Theory of Phlogiston. Annals of Science, 65(3):309–338, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2009:BPU

[Boa09] Victor D. Boantza. The Boyle papers: Understanding the manuscripts of Robert Boyle. Annals of Science, 66(4):570–572, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2012:APF

[Boa12a] Victor D. Boantza. Alchimie et paracelsisme en France à la fin de la Renaissance (1567–1625). (French) [Alchemy and Paracelsism in France at the end of the Renaissance (1567–1625)]. Annals of Science, 69(1):142–145, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2012:BRH

[Boa12b] Victor D. Boantza. Book review: History of chemistry and alchemy: Mathew D. Eddy, The Language of Mineralogy: John Walker, Chemistry and the Edinburgh Medical School, 1750–1800, ISBN-13 978-0-7546-6332-4. Annals of Science, 69(4):579–581, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2016:EJB

[Boa16] Victor D. Boantza. The enlightenment of Joseph Black. Annals of Science, 73(3):328–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boantza:2019:BRL

[Boa19] Victor D. Boantza. Book review: For the love of science: the correspondence of J. H. de Magellan (1722–1790), edited by R. W. Home, I. M. Malaquias and M. F. Thomaz, Bern, Peter Lang, 2017, 2002 pp. (2 volumes), £157 (cloth), ISBN 978-3-0343-1294-3. Annals of Science, 76(3–4):380–382, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bodenheimer:1956:HBT

[Bod56] F. S. Bodenheimer, Ph.D. H. B. Tristram’s collections in natural history, especially of Palestine. Annals of Science, 12(4):278–287, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boersema:2014:BRB

[Boe14] David Boersema. Book review: Michael Polanyi and His Generation: Origins of the Social Construction of Science. Annals of Science, 71 (3):437–439, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bonolis:2014:ISC

[Bon14] Luisa Bonolis. International scientific cooperation during the 1930s. Bruno Rossi and the development of the status of cosmic rays into a branch of physics. Annals of Science, 71(3):355–409, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Borrello:2010:ERC

[Bor10] Mark E. Borrello. Evolutionary restraints: the contentious history of group selection. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-06701-7 (hardcover). xi + 215 pp. LCCN QH376 .B67 2010.

Boschiero:2003:NPC

[Bos03] Luciano Boschiero. Natural philosophical contention inside the accademia del cimento: the properties and effects of heat and cold. Annals of Science, 60(4):329–349, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bouley:2013:BRB

[Bou13] Bradford Bouley. Book review: Histories of Scientific Observation, edited by L. Daston and E. Lunbeck. Annals of Science, 70(4):582– 585, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boulter:2017:BSS

[Bou17] Michael Charles Boulter. Bloomsbury scientists: science and art in the wake of Darwin. UCL Press, London, UK, 2017. ISBN 1-78735-005-3 (paperback). 188 pp. LCCN Q127.G4.

Bowles:1974:PHD

[Bow74] G. Bowles. Physical, human and divine attraction in the life and thought of George Cheyne. Annals of Science, 31(6):473–488, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bowler:1978:HVT

[Bow78] Peter J. Bowler. Hugo de Vries and Thomas Hunt Morgan: The mutation theory and the spirit of Darwinism. Annals of Science, 35 (1):55–73, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bowler:1984:WML

[Bow84] Peter J. Bowler. E. W. MacBride’s Lamarckian eugenics and its implications for the social construction of scientific knowledge. Annals of Science, 41(3):245–260, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See comments [Mac85a, Bow85].

Bowler:1985:RM

[Bow85] Peter J. Bowler. Response to MacKenzie. Annals of Science, 42(4): 420, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Bow84, Mac85a].

Bowler:2009:LS

[Bow09a] Peter J. Bowler. Life sciences. Annals of Science, 66(1):145–147, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bowler:2009:SAP

[Bow09b] Peter J. Bowler. Science for all: the popularization of science in early twentieth-century Britain. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-06863-3 (hardcover). x + 339 pp. LCCN Q225.2.G7 B69 2009.

Bowler:2013:BRP

[Bow13] Peter J. Bowler. Book review: P. M. Harman, The Culture of Nature in Britain, 1680–1890. New Haven: Yale University Press, 2009. xi + 392 pp. 17 plts. £45.00. ISBN-13 978-0-300-15197-8. Annals of Science, 70(1):129–131, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bowler:2016:DSA

[Bow16] Peter J. Bowler. Discovering science from an armchair: Popular science in British magazines of the interwar years. Annals of Science, 73(1):89–107, 2016. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00033790.2015.1061236.

Bowler:2017:BRL

[Bow17] Peter J. Bowler. Book review: The Life Organic: The Theoretical Biology Club and the Roots of Epigenetics. Annals of Science, 74(4): 343–344, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1975:SEV

[Bra75] Margaret Bradley. Scientific education versus military training: The influence of Napoleon Bonaparte on the École Polytechnique. Annals of Science, 32(5):415–449, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1976:FPI

[Bra76] Margaret Bradley. The facilities for practical instruction in science during the early years of the École Polytechnique. Annals of Science, 33(5):425–446, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1979:FBF

[Bra79] Margaret Bradley. The financial basis of French scientific education and scientific institutions in Paris, 1790–1815. Annals of Science, 36(5):451–491, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1981:FRE

[Bra81] Margaret Bradley. Franco–Russian engineering links: The careers of Lamé and Clapeyron, 1820–1830. Annals of Science, 38(3):291– 312, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brackenridge:1982:KEOa

[Bra82a] J. Bruce Brackenridge. Kepler, elliptical orbits, and celestial circularity: a study in the persistence of metaphysical commitment. Part I. Annals of Science, 39(2):117–143, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brackenridge:1982:KEOb

[Bra82b] J. Bruce Brackenridge. Kepler, elliptical orbits, and celestial circularity: a study in the persistence of metaphysical commitment. Part II. Annals of Science, 39(3):265–295, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brackenridge:1988:NMD

[Bra88] J. Bruce Brackenridge. Newton’s mature dynamics: Revolutionary or reactionary? Annals of Science, 45(5):451–476, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brackenridge:1990:NUD

[Bra90] J. Bruce Brackenridge. Newton’s unpublished dynamical principles: a study in simplicity. Annals of Science, 47(1):3–31, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1992:EMS

[Bra92] Margaret Bradley. Engineers as military spies? French engineers come to Britain, 1780–1790. Annals of Science, 49(2):137–161, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:1994:BPI

[Bra94] Margaret Bradley. Bonaparte’s plans to invade England in 1801: The fortunes of Pierre Forfait. Annals of Science, 51(5):453–475, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brassley:1995:ARB

[Bra95] Paul Brassley. Agricultural research in Britain, 1850–1914: Failure, success and development. Annals of Science, 52(5):465–480, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bravo:2002:GEL

[Bra02] M. T. Bravo. The geography of an empire licensed by Providence. Annals of Science, 59(4):413–418, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Branagan:2009:FCG

[Bra09] David Branagan. Four centuries of geological travel: The search for knowledge on foot, bicycle, sledge and camel. Annals of Science, 66(4):572–575, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Braun:2010:TMC

[Bra10] Robyn Braun. Truth machine: The contentious history of DNA fingerprinting. Annals of Science, 67(1):145–147, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bret:1991:DGG

[Bre91] Patrice Bret. Le Dépôt général de la guerre et la formation scientifique des ingénieurs-géographes militaires en France (1789–1830). (French) [The General Ministry of War and the scientific formation of the military engineer-geographers in France (1789–1830)]. Annals of Science, 48(2):113–157, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bret:1993:DPR

[Bre93] Patrice Bret. Le dernier des procédés révolutionnaires: La fabrication et l’expertise de la Poudre Ronde (1795–1830). (French) [The last of the revolutionary processes: Manufacturing and expertise of Round Powder (1795–1830)]. Annals of Science, 50(4):325–347, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bret:2016:LDL

[Bre16] Patrice Bret. The letter, the dictionary and the laboratory: translating chemistry and mineralogy in eighteenth-century France. Annals of Science, 73(2):122–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brianta:2000:ETM

[Bri00] Donata Brianta. Education and training in the mining industry, 1750–1860: European models and the Italian case. Annals of Science, 57(3):267–300, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brianta:2007:EMC

[Bri07] Donata Brianta. Europa mineraria: circolazione delle élites e trasferimento tecnologico, secoli XVIII–XIX, volume 95 of Temi di storia; Istituzioni formative e agenti di sviluppo. F. Angeli, Milano, Italy, 2007. ISBN 88-464-8818-0. 447 pp. LCCN TN55 .B753 2007.

Brinkman:2010:SJD

[Bri10] Paul D. Brinkman. The Second Jurassic Dinosaur Rush: Museums and Paleontology in America at the Turn of the Twentieth Century. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-07472-2. 312 pp. LCCN ????

Brinkman:2013:BRJ

[Bri13] Paul D. Brinkman. Book review: John M. MacKenzie, Museums and Empire: Natural History, Human Cultures and Colonial Identities. Manchester and New York: Manchester University Press, 2009. xv + 286 pp. 15 figs. $89.95. ISBN-13 978-0-7190-8022-7. Annals of Science, 70(1):114–115, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:1936:HSS

[Bro36a] H. Brown. The History of Science Society Report of the St. Louis meeting 1936. Annals of Science, 1(2):233–234, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:1936:UMA

[Bro36b] Harcourt Brown, M.A., Ph.D. The utilitarian motive in the age of Descartes. Annals of Science, 1(2):182–192, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brook:1955:DWC

[Bro55] Michael Brook, M.A., A.L.A. Dr Warwick’s chemistry lectures and the scientific audience in Sheffield (1799–1801). Annals of Science, 11(3):224–237, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:1969:SHPa

[Bro69a] W. H. Brock, M.Sc., Ph.D. Studies in the history of Prout’s hypotheses. Part I. Annals of Science, 25(1):49–80, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:1969:SHPb

[Bro69b] W. H. Brock, M.Sc., Ph.D. Studies in the history of Prout’s hypotheses. Part II. Prout’s lectures of 1814. Annals of Science, 25(2): 127–137, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:1971:DFR

[Bro71] Harcourt Brown, M.A., Ph.D. Doctor François Rabelais: Pantagruel and health. Annals of Science, 27(2):117–134, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brooke:1977:NTP

[Bro77] John Hedley Brooke. Natural theology and the plurality of worlds: Observations on the Brewster–Whewell debate. Annals of Science, 34(3):221–286, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:1978:SPG

[Bro78a] W. H. Brock. The society for the perpetuation of Gmelin: The Cavendish Society, 1846–1872. Annals of Science, 35(6):599–617, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brockliss:1978:MTU

[Bro78b] L. W. B. Brockliss. Medical teaching at the University of Paris, 1600– 1720. Annals of Science, 35(3):221–251, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:1979:GOI

[Bro79a] Joyce Brown. Guild organisation and the instrument-making trade, 1550–1830: the Grocers’ and Clockmakers’ Companies. Annals of Science, 36(1):1–34, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Browne:1979:JBV

[Bro79b] Alice Browne. J. B. van Helmont’s attack on Aristotle. Annals of Science, 36(6):575–591, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brockliss:1981:ADN

[Bro81] L. W. B. Brockliss. Aristotle, Descartes and the new science: Natural philosophy at the University of Paris, 1600–1740. Annals of Science, 38(1):33–69, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brooks:1987:MIC

[Bro87] Randall C. Brooks. Magnetic influence on chronometers, 1798–1834: a case study. Annals of Science, 44(3):245–264, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:1990:CSW

[Bro90] W. H. Brock. The Cavendish Society’s wonderful repertory of chemistry. Annals of Science, 47(1):77–80, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brooks:1992:CDE

[Bro92] John Brooks. The circular dividing engine: Development in England 1739–1843. Annals of Science, 49(2):101–135, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:1993:HBN

[Bro93] W. H. Brock. Humboldt and the British: a note on the character of British science. Annals of Science, 50(4):365–372, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brooks:1995:RCF

[Bro95] Nathan M. Brooks. Russian chemistry in the 1850s: a failed attempt at institutionalization. Annals of Science, 52(6):577–589, 1995. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Broughton:2009:AUS

[Bro09] Peter Broughton. The accuracy and use of sextants and watches in Rupert’s land in the 1790s. Annals of Science, 66(2):209–229, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:2015:RED

[Bro15] C. N. Brown. Ruling engines, diffraction gratings and wavelength measurements before the Rowland era. Annals of Science, 72(1):28– 74, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brock:2016:HCV

[Bro16] W. H. (William Hodson) Brock. The history of chemistry: a very short introduction, volume 456 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-871648-6 (paperback). xvi + 151 pp. LCCN QD11 .B758 2016.

Brown:2018:RED

[Bro18] C. N. Brown. Ruling engines and diffraction gratings before Rowland: the work of Lewis Rutherfurd and William Rogers. Annals of Science, 75(4):330–360, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:2022:RED

[Bro22] C. N. Brown. The ruling engines and diffraction gratings of Henry Augustus Rowland. Annals of Science, 79(1):81–130, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1957:DKTa

[Bru57a] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases. I. Herapath. Annals of Science, 13(3):188–198, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1957:DKTb

[Bru57b] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases II. Waterston. Annals of Science, 13(4):273–282, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1958:DKTa

[Bru58a] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases III. Clausius. Annals of Science, 14(3):185–196, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1958:DKTb

[Bru58b] S. G. Brush, A.B., D.Phil. The development of the kinetic theory of gases IV. Maxwell. Annals of Science, 14(4):243–255, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brush:1979:NCD

[Bru79] Stephen G. Brush. Nineteenth-century debates about the inside of the Earth: Solid, liquid or gas? Annals of Science, 36(3):225–254, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brundtland:2011:ABL

[Bru11] Terje Brundtland. After Boyle and the Leviathan: the second generation of British air pumps. Annals of Science, 68(1):93–124, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bryden:1972:GBA

[Bry72] D. J. Bryden, B.Sc. George Brown, author of the Rotula. Annals of Science, 28(1):1–29, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bryden:1975:SSM

[Bry75] D. J. Bryden. Sir Samuel Morland’s account of the balance barometer, 1678. Annals of Science, 32(4):359–368, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bryden:1990:EOS

[Bry90] D. J. Bryden. The Edinburgh Observatory 1736–1811: a story of failure. Annals of Science, 47(5):445–474, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bryden:1992:EAM

[Bry92] D. J. Bryden. Evidence from advertising for mathematical instrument making in London, 1556–1714. Annals of Science, 49(4):301– 336, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brysse:2010:CAP

[Bry10] Keynyn Brysse. Cryptozoology, archaeology and palaeontology: Histories near the high table. Annals of Science, 67(4):569–575, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brysse:2013:BRB

[Bry13] Keynyn Brysse. Book review: The Second Jurassic Dinosaur Rush: Museums and Paleontology in America at the Turn of the Twentieth Century, by Paul D. Brinkman. Annals of Science, 70(4):552–554, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Baker:1965:OAE

[BS65] K. M. Baker, M.A., Ph.D. and W. A. Smeaton, M.Sc., Ph.D. The origins and authorship of the educational proposals published in 1793 by the Bureau de Consultation des Arts et Métiers and generally ascribed to Lavoisier. Annals of Science, 21(1):33–46, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bryden:1993:SIP

[BS93] D. J. Bryden and D. L. Simms. Spectacles improved to perfection and approved of by the Royal Society. Annals of Science, 50(1): 1–32, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bradley:2007:LEL

[BS07] Robert E. Bradley and Charles Edward Sandifer, editors. Leonhard Euler: Life, Work, and Legacy, volume 5 of Studies in the history and philosophy of mathematics. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-52728-1. viii + 534 pp. LCCN QA29.E8 L465 2007.

Bertomeu-Sanchez:2012:AEV

[BS12] José Ramón Bertomeu-Sánchez. Animal experiments, vital forces and courtrooms: Mateu Orfila, François Magendie and the study of poisons in nineteenth-century France. Annals of Science, 69(1):1– 26, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bertomeu-Sanchez:2015:CMS

[BS15] José Ramón Bertomeu-Sánchez. Chemistry, microscopy and smell: bloodstains and nineteenth-century legal medicine. Annals of Science, 72(4):490–516, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Berger:2021:ONT

[BS21] Susanna Berger and Sara J. Schechner. Observations on Niccolò Tornioli’s The Astronomers. Annals of Science, 78(4):418–462, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bschir:2012:BRP

[Bsc12] Karim Bschir. Book review: Philosophical aspects of science: Alan Chalmers, The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms, ISBN-13 978-90-481-2361-2. Annals of Science, 69(4):575– 577, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bschir:2014:BRBa

[Bsc14a] Karim Bschir. Book review: Cosmopolitics I, by Isabelle Stengers. Annals of Science, 71(1):129–132, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bschir:2014:BRBb

[Bsc14b] Karim Bschir. Book review: Wrestling with Nature. From Omens to Science, by Peter Harrison, Ronald L. Numbers, and Michael H. Shank. Annals of Science, 71(1):144–147, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Brown:1955:WJR

[BT55] John R. Brown and John L. Thornton. William James Russell (1830– 1909) and investigations on London fog. Annals of Science, 11(4): 331–336, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Buchwald:2008:DEJ

[Buc08] Jed Z. Buchwald. Descartes’s experimental journey past the prism and through the invisible world to the rainbow. Annals of Science, 65(1):1–46, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Buckland:2013:NSF

[Buc13] Adelene Buckland. Novel science: fiction and the invention of nineteenth-century geology. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-07968-6 (hardcover), 0-226-92363-0 (e-book). 377 pp. LCCN PR468.G38 B83 2013.

Bunn:1977:QRB

[Bun77] Robert Bunn. Quantitative relations between infinite sets. Annals of Science, 34(2):177–191, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

BurnistonBrown:1960:NLV

[Bur60] G. Burniston Brown, M.Sc., Ph.D. The Newton letters (vols. I and II). Annals of Science, 16(2):117–124, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burstyn:1965:DFE

[Bur65] Harold L. Burstyn, M.S., Ph.D. The deflecting force of the Earth’s rotation from Galileo to Newton. Annals of Science, 21(1):47–80, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:1990:NSD

[Bur90] Charles S. F. Burnett. A new source for Dominicus Gundissalinus’s account of the science of the stars? Annals of Science, 47(4):361– 374, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:1998:KPA

[Bur98] Charles Burnett. King Ptolemy and Alchandreus the Philosopher: The earliest texts on the astrolabe and Arabic astrology at Fleury, Micy and Chartres. Annals of Science, 55(4):329–368, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2000:AKP

[Bur00] Charles Burnett. Addendum to “King Ptolemy and Alchandreus the Philosopher: The Earliest Texts on the Astrolabe and Arabic Astrology at Fleury, Micy and Chartres”. Annals of Science, 57(2): 187, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2009:VAI

[Bur09] Charles Burnett. Volkstümliche Astronomie im islamischen Mittelalter Zur Bestimmung der Gebetszeiten und der Qibla bei, Ibn und al-Farisi. (German) [popular astronomy in the Islamic Middle Ages to determine the prayer times and the Qibla bei, Ibn, and al-Farisi]. Annals of Science, 66(4):584–585, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2010:EMS

[Bur10a] Charles Burnett. Expounding the mathematical seed: a translation of Bhaskara I on the mathematical chapter of the Aryabhatiya. Annals of Science, 67(1):132–134, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2010:SMA

[Bur10b] Charles Burnett. Studies in medieval astronomy and optics. Annals of Science, 67(1):131–132, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burney:2010:AHM

[Bur10c] Ian Burney. Allergy: The history of a modern malady. Annals of Science, 67(2):277–281, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2011:BRA

[Bur11a] Charles Burnett. Book review: Aristoteles Latinus, X 2.1. Meteorologica. Translatio Guillelmi de Morbeka, ISBN-13 978-2-503-53078-9. Annals of Science, 68(4):579–580, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2011:BRM

[Bur11b] Charles Burnett. Book review: Medieval science. Avicenna Latinus, Liber primus naturalium. Tractatus secundus de motu et de consimilibus, ISBN 2-8031-0231-5. Annals of Science, 68(4):577–578, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2012:DIA

[Bur12a] Charles Burnett. The dispensatory of Ibn al-Tilmid. Annals of Science, 69(1):133–135, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2012:EAH

[Bur12b] Charles Burnett. Eastern astrolabes, historic scientific instruments of the Adler Planetarium and Astronomy Museum. Annals of Science, 69(3):446–447, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2012:APE

[Bur12c] Charles Burnett. Les Alchandreana primitifs: Études sur les plus anciens traités astrologiques latins d’origine arabe (Xe siècle). (French) [The primitive Alchandreana: Studies on the oldest Roman astrological treatises of Arab origin (tenth century)]. Annals of Science, 69(1):139–141, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bursten:2012:PDB

[Bur12d] Julia R. Bursten. Pauling’s defence of bent-equivalent bonds: a view of evolving explanatory demands in modern chemistry. Annals of Science, 69(1):69–90, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2014:BRBb

[Bur14a] Charles Burnett. Book review: Le Liber Mahameleth: Édition critique et commentaires, edited by Anne-Marie Vlasschaert. Annals of Science, 71(2):286–287, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2014:BRBa

[Bur14b] Charles Burnett. Book review: Quaestiones super libros. De generatione et corruptione Aristotelis, by John Buridan. Annals of Science, 71(1):110–111, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2014:BRN

[Bur14c] Charles Burnett. Book review: Nicole Oresme, Questiones super geometriam Euclidis, edited by H. L. L. Busard. Annals of Science, 71(1):111–113, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burnett:2015:BRB

[Bur15] Charles Burnett. Book review: Science in Medieval Jewish Cultures. Annals of Science, 72(1):144–147, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Burdett:2017:BRS

[Bur17] Carolyn Burdett. Book review: The Science of Sympathy: Morality, Evolution, and Victorian Civilization. Annals of Science, 74(4):337– 339, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Busard:2010:NOQ

[Bus10] H. L. L. (Hubertus Lambertus Ludovicus) Busard. Nicole Oresme, Questiones super geometriam Euclidis, volume 57 of Boethius; Wissenschaftsgeschichte. Franz Steiner Verlag, Stuttgart, Germany, 2010. ISBN 3-515-09287-0. 199 pp. LCCN QA444 .O7413 2010.

URL http://d-nb.info/1006975152/04.

Butler:1957:RCY

[But57] F. H. C. Butler. Report of the Council for the Year 1956–57. Annals of Science, 13(3):82, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bensaude-Vincent:1983:MRC

[BV83] Bernadette Bensaude-Vincent. Une mythologie révolutionnaire dans la chimie française. Annals of Science, 40(2):189–196, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bensaude-Vincent:1999:APL

[BV99] Bernadette Bensaude-Vincent. Atomism and positivism: a legend about French chemistry. Annals of Science, 56(1):81–94, 1999. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Busard:1973:BAS

[BvK73] H. L. L. Busard and P. S. van Koningsveld. Der Liber de Arcubus Similibus des Ahmed Ibn Jusuf. (German) [The Book of Arcubus Similibus by Ahmed Ibn Jusuf]. Annals of Science, 30(4):381–406, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Byford:2019:RET

[Byf19] Andy Byford. Russian eugenics in transnational and transhistorical perspective: from comparison to translation: Review of With and without Galton: Vasilii Florinskii and the fate of eugenics in Russia, by N. Krementsov, Cambridge, UK, Open Book Publishers, 2018, xxvi + 666 pp., £32.95 (Hardback), ISBN 978-1-78374-512-8. Annals of Science, 76(3–4):355–364, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Byrne:1980:SCC

[Byr80] Patrick H. Byrne. Statistical and causal concepts in Einstein’s early thought. Annals of Science, 37(2):215–228, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Byrne:1981:OEU

[Byr81] Patrick H. Byrne. The origins of Einstein’s use of formal asymmetries. Annals of Science, 38(2):191–206, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cable:1973:ESS

[Cab73] John A. Cable, M.A., M.Ed., Dipl.Deutsch (Marburg), A.L.C.M. Early Scottish science: The vocational provision. Annals of Science, 30(2):179–199, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cahan:1990:DFK

[Cah90] David Cahan. From dust figures to the kinetic theory of gases: August Kundt and the changing nature of experimental physics in the 1860s and 1870s. Annals of Science, 47(2):151–172, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cahan:2010:HGA

[Cah10] David Cahan. Helmholtz in gilded-age America: The International Electrical Congress of 1893 and the relations of science and technology. Annals of Science, 67(1):1–38, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Calinger:1968:FGB

[Cal68] Ronald S. Calinger, M.A. Frederick the Great and the Berlin Academy of Sciences (1740–1766). Annals of Science, 24(3):239– 249, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Calcraft:1980:AHS

[Cal80] L. G. A. Calcraft. Aldous Huxley and the Sheldonian hypothesis. Annals of Science, 37(6):657–671, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cameron:1955:RAV

[Cam55] G. R. Cameron, D.Sc., F.R.C.P., F.R.S. Rudolf Albert v. Koelliker (1817–1905). Annals of Science, 11(2):166–172, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Campbell:1980:DVD

[Cam80] Margaret Campbell. Did de Vries discover the law of segregation independently? Annals of Science, 37(6):639–655, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cambrosio:2009:FNS

[Cam09] A. Cambrosio. La fabrique d’une nouvelle science. La biologie moléculaire à l’âge atomique (1945–1964). (French) [The factory of a new science. Molecular biology in the atomic age]. Annals of Science, 66(3):426–429, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

CamposRibeiro:2023:NAF

[Cam23] Luís Campos Ribeiro. Nautical astrology: a forgotten early modern tradition. Annals of Science, 80(3):199–231, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2179111.

Cantor:1975:CSS

[Can75a] G. N. Cantor. A critique of Shapin’s social interpretation of the Edinburgh phrenology debate. Annals of Science, 32(3):245–256, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cantor:1975:EPD

[Can75b] G. N. Cantor. The Edinburgh phrenology debate: 1803–1828. Annals of Science, 32(3):195–218, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Canales:2009:TSH

[Can09] Jimena Canales. A tenth of a second: a history. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-09318-2 (hardcover). xii + 269 pp. LCCN QB213 .C36 2009.

Cantor:2010:MFM

[Can10] Geoffrey Cantor. Michael Faraday’s mental exercises: An artisan essay circle in Regency London. Annals of Science, 67(2):284–285, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Canales:2015:PPE

[Can15] Jimena Canales. The physicist and the philosopher: Einstein and Bergson and the debate that changed our understanding of time.

Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16534-3 (hardcover), 1-4008-6577-8 (e-book). viii + 479 pp. LCCN BD638 .C326 2015.

Caneva:2019:HCF

[Can19] Kenneth L. Caneva. Helmholtz, the conservation of force and the conservation of vis viva. Annals of Science, 76(1):17–57, 2019. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Capecchi:2012:HVW

[Cap12] Danilo Capecchi. History of virtual work laws: A History of Mechanics Prospectivea. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 88-470-2056-5. xii + 492 pp. LCCN ????

Cardwell:1992:MFH

[Car92] D. S. L. Cardwell. On Michael Faraday, Henry Wilde, and the dynamo. Annals of Science, 49(5):479–487, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carey:1997:CNH

[Car97a] Daniel Carey. Compiling nature’s history: Travellers and travel narratives in the early Royal Society. Annals of Science, 54(3):269–292, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cartwright:1997:RPL

[Car97b] David E. Cartwright. Robert Paul de Lamanon: An unlucky naturalist. Annals of Science, 54(6):585–596, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carneiro:2005:OGS

[Car05] Ana Carneiro. Outside government science, ‘not a single tiny bone to cheer us up!’ The Geological Survey of Portugal (1857–1908), the involvement of common men, and the reaction of civil society to geological research. Annals of Science, 62(2):141–204, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carlson:2010:BTP

[Car10] Robert H. Carlson. Biology is technology: the promise, peril, and new business of engineering life. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-03544-5. 279 pp. LCCN TP248.2 .C37

Carey:2013:LSM

[Car13] Daniel Carey. Locke’s species: Money and philosophy in the 1690s. Annals of Science, 70(3):357–380, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carlyle:2017:BRB

[Car17] Margaret Carlyle. Book review: The Body as Object and Instrument of Knowledge: Embodied Empiricism in Early Modern Science. Annals of Science, 74(4):332–335, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carlyle:2018:BRU

[Car18] Margaret Carlyle. Book review: The Use of Humans in Experiment: Perspectives from the 17^{th} to the 20^{th} Century, edited by E. Dyck and L. Stewart, Leiden, Brill, 2016, xii + 297 pp., EUR 115.00, $149.00, ISBN 978-90-04-28670-2. Annals of Science, 75(2):154–156, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carlyle:2021:MAM

[Car21] Margaret Carlyle. Malleable anatomies. models, makers, and material culture in Eighteenth-Century Italy. Annals of Science, 78(1): 128–131, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carlyle:2022:MFS

[Car22] Margaret Carlyle. Minerva’s French sisters: Women of science in Enlightenment France. Annals of Science, 79(3):413–415, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2066178.

Case:2015:LMU

[Cas15] Stephen Case. ‘land-marks of the universe’: John Herschel against the background of positional astronomy. Annals of Science, 72(4): 417–434, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cavicchi:2006:NCD

[Cav06] Elizabeth Cavicchi. Nineteenth-Century Developments in coiled instruments and experiences with electromagnetic induction. Annals of Science, 63(3):319–361, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cawood:1977:TMD

[Caw77] John Cawood. Terrestrial magnetism and the development of international collaboration in the early nineteenth century. Annals of Science, 34(6):551–587, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Castel-Branco:2022:PML

[CB22] Nuno Castel-Branco. Physico-mathematics and the life sciences: experiencing the mechanism of venous return, 1650s–1680s. Annals of Science, 79(4):442–467, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00033790.2022.2086301.

Cranmer-Byng:1981:CSC

[CBL81] J. L. Cranmer-Byng and Trevor H. Levere. A case study in cultural collision: Scientific apparatus in the Macartney embassy to China, 1793. Annals of Science, 38(5):503–525, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Clow:1947:NEH

[CC47] Archibald Clow, M.A., Ph.D., D.Sc. and Nan L. Clow, M.A., B.Sc. The natural and economic history of kelp. Annals of Science, 5(4): 297–316, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Clow:1956:TFD

[CC56] Archibald Clow, M.A., D.Sc., F.R.S.E. and Nan L. Clow, M.A.B.Sc. The timber famine and the development of technology. Annals of Science, 12(2):85–102, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carden-Coyne:2009:MH

[CC09] Ana Carden-Coyne. Medicine and health. Annals of Science, 66(1): 147–150, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Connor:2011:MRR

[CC11] James Thomas Hamilton Connor and Stephan Michael Curtis, editors. Medicine in the remote and rural north, 1800–2000, volume 3 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-157-3. xiii + 301 pp. LCCN RA395.E853 M43 2011.

Calvert:1957:AYE

[CCC57] H. R. Calvert, John Craggs, and Co. Accounts for the year ending 31st March, 1957. Annals of Science, 13(3):83–84, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cecon:2011:CTC

[Cec11] Kleber Cecon. Chemical translation: The case of Robert Boyle’s experiments on sensible qualities. Annals of Science, 68(2):179–198, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Buchwald:2021:WAS

[CECE21] Jed Z. Buchwald (Co-Editor) and Mordechai Feingold (Co-Editor). Winner of the Annals of Science Best Paper Prize for 2019. Annals of Science, 77(4):555, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cooper:1982:WRG

[CH82] M. L. Cooper and V. M. D. Hall. William Robert Grove and the London Institution, 1841–1845. Annals of Science, 39(3):229–254, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1948:BSP

[Cha48] J. Challinor, M.A. The beginnings of scientific palaeontology in Britain. Annals of Science, 6(1):46–53, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1952:BTE

[Cha52] J. A. Chaldecott, M.Sc., A.Inst.P. Bartolomeo Telioux and the early history of the thermometer. Annals of Science, 8(3):195–201, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1953:ZOD

[Cha53a] J. A. Chaldecott, M.Sc., A.Inst.P. The zograscope or optical diagonal machine. Annals of Science, 9(4):315–322, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1953:EPB

[Cha53b] J. Challinor, M.A. The early progress of British geology. I. From Leland to Woodward, 1538–1728. Annals of Science, 9(2):124–153, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1954:EPBa

[Cha54a] J. Challinor, M.A. The early progress of British geology. II. From Strachey to Michell, 1719–1788. Annals of Science, 10(1):1–19, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1954:EPBb

[Cha54b] J. Challinor, M.A. The early progress of British geology. III. From Hutton to Playfair, 1788–1802. Annals of Science, 10(2):107–148, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1961:SCT

[Cha61] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). I. Annals of Science, 17(3):175–195, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1962:SCT

[Cha62] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). II. Annals of Science, 18(3):147–175, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1963:SCTa

[Cha63a] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). III. Annals of Science, 19(1):49–79, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1963:SCTb

[Cha63b] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). IV. Annals of Science, 19(4):285–297, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1964:SCTa

[Cha64a] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). V. Annals of Science, 20(1):59–80, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1964:SCTb

[Cha64b] John Challinor, M.A., F.G.S. Some correspondence of Thomas Webster, geologist (1773–1844). VI. Annals of Science, 20(2):143–164, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1968:SAP

[Cha68] J. A. Chaldecott, M.Sc., A.Inst.P. Scientific activities in Paris in 1791: Evidence from the diaries of Sir James Hall for 1791, and other contemporary records. Annals of Science, 24(1):21–52, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Challinor:1970:PBG

[Cha70] John Challinor, M.A. The progress of British geology during the early part of the nineteenth century. Annals of Science, 26(3):177– 234, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1971:WMP

[Cha71] J. A. Chaldecott, M.Sc., F.Inst.P. Wollaston, Malacrida, and the platinum thermometer an unsolved problem in the history of pyrometry. Annals of Science, 27(4):409–411, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1972:HLM

[Cha72a] J. A. Chaldecott, M.Sc., F.Inst.P. Hans Loeser’s metallic thermometers of 1746 and 1747. Annals of Science, 28(1):87–100, 1972. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaldecott:1972:PPL

[Cha72b] J. A. Chaldecott, M.Sc., F.Inst.P. The platinum pyrometers of Louis Bernard Guyton de Morveau, F.R.S. (1737–1816). Annals of Science, 28(4):347–368, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chandler:1975:CCN

[Cha75] Philip Chandler. Clairaut’s critique of Newtonian attraction: Some insights into his philosophy of science. Annals of Science, 32(4):369– 378, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chapman:1983:DAS

[Cha83a] Allan Chapman. The design and accuracy of some observatory instruments of the seventeenth century. Annals of Science, 40(5):457– 471, September 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chapman:1983:SAS

[Cha83b] Allan Chapman. A study of the accuracy of scale graduations on a group of European astrolabes. Annals of Science, 40(5):473–488, September 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chapman:1984:TBC

[Cha84] Allan Chapman. Tycho Brahe in China: the Jesuit mission to Peking and the iconography of European instrument-making processes. Annals of Science, 41(5):417–443, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chapin:1988:LLY

[Cha88] Seymour L. Chapin. Lalande and the length of the year; or, how to win a prize and double publish. Annals of Science, 45(2):183– 190, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chayut:1991:JJT

[Cha91] Michael Chayut. J. J. Thomson: The discovery of the electron and the chemists. Annals of Science, 48(6):527–544, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chapman:1995:MJB

[Cha95] Allan Chapman. Out of the meridian: John Bird’s equatorial sector and the new technology of astronomical measurement. Annals of Science, 52(5):431–463, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chalmers:2009:SAP

[Cha09] A. F. (Alan Francis) Chalmers. The Scientist’s Atom and the Philosopher’s Stone: How Science Succeeded and Philosophy Failed to Gain Knowledge of Atoms, volume 279 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2361-5, 90-481-2362-3. ISSN 0068-0346. xii + 287 pp. LCCN QD461 .C375 2009. URL

https://link.springer.com/book/10.1007/978-90-481-2362-9.

Chandelier:2012:BRR

[Cha12a] Joel Chandelier. Book review: Renaissance and Reformation. Hiro Hirai, editor. Cornelius Gemma. Cosmology, Medicine and Natural Philosophy in Renaissance Louvain. ISBN-13 978-88-6227-118-9; ISSN 1127-6045. Annals of Science, 69(3):447–449, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Channell:2012:BRT

[Cha12b] David F. Channell. Book review: Technology and engineering: Aitor Anduaga [Egaña], Wireless and Empire: Geopolitics, Radio Industry and Ionosphere in the British Empire, 1918–1939, ISBN-13 978-0-19-956272-5. Annals of Science, 69(4):591–592, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chaney:2013:BRB

[Cha13] Sarah Chaney. Book review: The Woman Who Walked into the Sea: Huntington’s and the Making of a Genetic Disease. Annals of Science, 70(2):291–293, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Charmantier:2014:BRB

[Cha14] Isabelle Charmantier. Book review: Science, Philosophy and Religion in the Age of the Enlightenment. British and Global Contexts, by John Gascoigne. Annals of Science, 71(1):127–129, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Charenko:2016:EDH

[Cha16] Melissa Charenko. Earth’s deep history: How it was discovered and why it matters. Annals of Science, 73(3):335–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chalmers:2017:OHY

[Cha17] Alan F. Chalmers. One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, volume 51 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer International Publishing AG, Cham, Switzerland, 2017. ISBN 3-319-56528-1 (hardcover), 3-319-56529-X (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). ix + 197 pp. LCCN QA905 .C43 2017. URL

http://link.springer.com/10.1007/978-3-319-56529-3.

Charalampous:2019:OCM

[Cha19] Charis Charalampous. ‘One common matter’ in Descartes’ physics: the Cartesian concepts of matter quantities, weight and gravity. Annals of Science, 76(3–4):324–339, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chassefiere:2021:ABS

[Cha21] E. Chassefière. Aurora borealis systems in the German–Russian world in the first half of the eighteenth century: the cases of Friedrich Christoph Mayer and Leonhard Euler. Annals of Science, 78(2):162– 196, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chen:1998:DEA

[Che98] Xiang Chen. Dispersion, experimental apparatus, and the acceptance of the wave theory of light. Annals of Science, 55(4):401–420, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chitnis:1970:UEN

[Chi70] Anand C. Chitnis, M.A., Ph.D. The University of Edinburgh’s natural history museum and the Huttonian–Wernerian debate. Annals of Science, 26(2):85–94, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chinnici:2008:SDS

[Chi08] Ileana Chinnici. The ‘Società degli Spettroscopisti Italiani’: birth and evolution. Annals of Science, 65(3):393–438, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Christensen:1995:ORP

[Chr95] Dan Ch. Christensen. The Ørsted–Ritter partnership and the birth of Romantic natural philosophy. Annals of Science, 52(2):153–185, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chrisomalis:2010:NNC

[Chr10] Stephen Chrisomalis. Numerical Notation: a Comparative History. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-87818-7 (hardcover). ix + 486 pp. LCCN ????

Christie:2020:AC

[Chr20] John R. R. Christie. Atlantic chemistries, 1600–1820. Annals of Science, 77(2):135–138, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chu:2013:BRB

[Chu13] Pingyi Chu. Book review: Sojourners in a Strange Land: Jesuits and their Scientific Missions in Late Imperial China, by Florence C.

Hsia. Annals of Science, 70(4):562–564, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cifoletti:1986:KQ

[Cif86] Giovanna Cifoletti. Kepler’s De quantitatibus. Annals of Science, 43(3):213–238, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cassebaum:1976:ACC

[CK76] Heinz Cassebaum and George B. Kauffman. The analytical concept of a chemical element in the work of Bergman and Scheele. Annals of Science, 33(5):447–456, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carazza:1989:ABP

[CK89] Bruno Carazza and Helge Kragh. Adolfo Bartoli and the problem of radiant heat. Annals of Science, 46(2):183–194, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Clericuzio:1990:RBC

[Cle90] Antonio Clericuzio. A redefinition of Boyle’s chemistry and corpuscular philosophy. Annals of Science, 47(6):561–589, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Clow:1954:FPD

[Clo54] Archibald Clow, M.A., D.Sc., F.R.I.C., F.R.S.E. Fiscal policy and the development of technology. Annals of Science, 10(4):342–358, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Clow:1955:REW

[Clo55] Archibald Clow, M.A., Ph.D., D.Sc., F.R.S.E. A re-examination of William Walker’s Distinguished Men of Science. Annals of Science, 11(3):183–193, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Canguilhem:2008:KL

[CMM08] Georges Canguilhem, Paola Marrati, and Todd Meyers, editors. Knowledge of life. Fordham University Press, New York, NY, USA, 2008. ISBN 0-8232-2925-4 (hardcover), 0-8232-2926-2 (paperback). xx + 202 pp. LCCN QH341 .C313 2008. Translated by Stefanos Geroulanos and Daniela Ginsburg.

Cobb:2002:MSC

[Cob02] Matthew Cobb. Malpighi, Swammerdam and the colourful silkworm: Replication and visual representation in early modern science. Annals of Science, 59(2):111–147, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cobbold:2018:RAB

[Cob18] Carolyn Ann Cobbold. The rise of alternative bread leavening technologies in the nineteenth century. Annals of Science, 75(1):21–39, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cochrane:1956:FBR

[Coc56] Rexmond C. Cochrane, M.A., Ph.D. Francis Bacon and the rise of the mechanical arts in eighteenth-century England. Annals of Science, 12(2):137–156, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cock:1983:WBR

[Coc83] A. G. Cock. William Bateson’s rejection and eventual acceptance of chromosome theory. Annals of Science, 40(1):19–59, January 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cohen:1936:DHM

[Coh36] Leonora D. Cohen, M.A. Descartes and Henry More on the beastmachine — a translation of their correspondence pertaining to animal automatism. Annals of Science, 1(1):48–61, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cohen:1951:GD

[Coh51] I. Bernard Cohen, Ph.D. Guericke and Dufay. Annals of Science, 7 (2):207–209, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cohen:1987:SSE

[Coh87] H. Floris Cohen. Simon Stevin’s equal division of the octave. Annals of Science, 44(5):471–488, September 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cohen:2010:HMS

[Coh10a] H. Floris Cohen. How modern science came into the world: four civilizations, one 17th-century breakthrough. Amsterdam University Press, Amsterdam, The Netherlands, 2010. ISBN 90-8964-239-0 (hardbound), 90-485-1273-5 (e-book). xl + 784 pp. LCCN Q125

.C537 2010. URL http://www.loc.gov/catdir/toc/fy11pdf03/
2011401935.html.

Cohen:2010:SAP

[Coh10b] Sarah R. Cohen. Searching the animal psyche with Charles Le Brun. Annals of Science, 67(3):353–382, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cohen:2011:SLC

[Coh11] Claudine Cohen. Science, libertinage et clandestinité à l’aube des Lumières: le transformisme de Telliamed. Science, histoire et société travaux et recherches. Presses Universitaires de France, PUF, Paris, France, 2011. ISBN 2-13-057908-6. 433 pp. LCCN Q127.F8. URL

http://www.gbv.de/dms/casalini/60/60718875.pdf.

Cole:1937:LZRa

[Col37a] F. J. Cole, D.Sc., F.R.S. Leeuwenhoek’s zoological researches. Part I. Annals of Science, 2(1):1–46, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cole:1937:LZRb

[Col37b] F. J. Cole, D.Sc., F.R.S. Leeuwenhoek’s zoological researches. Part II. Bibliography and analytical index. Annals of Science, 2(2):185–235, April 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1939:HPB

[Col39] L. J. M. Coleby, M.A., M.Sc., Ph.D. A history of Prussian blue. Annals of Science, 4(2):206–211, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cole:1951:FAB

[Col51] R. J. Cole, B.Sc., F.R.I.C., A.M.I.Chem.E. Friedrich Accum (1769– 1838). A biographical study. Annals of Science, 7(2):128–143, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cole:1952:SAC

[Col52a] R. J. Cole, B.Sc., F.R.I.C., A.M.I.Chem.E. Sir Anthony Carlisle, F.R.S. (1768–1840). Annals of Science, 8(3):255–270, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1952:JFV

[Col52b] L. J. M. Coleby, M.A., M.Sc., Ph.D. John Francis Vigani: First Professor of Chemistry in the University of Cambridge. Annals of Science, 8(1):46–60, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1952:JHF

[Col52c] L. J. M. Coleby, M.A., M.Sc., Ph.D. John Hadley: Fourth Professor of Chemistry in the University of Cambridge. Annals of Science, 8 (4):293–301, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1952:JMP

[Col52d] L. J. M. Coleby, M.A., M.Sc., Ph.D. John Mickleburgh: Professor of Chemistry in the University of Cambridge, 1718–56. Annals of Science, 8(2):165–174, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1953:RWP

[Col53] L. J. M. Coleby, M.A., M.Sc., Ph.D. Richard Watson, Professor of Chemistry in the University of Cambridge, 1764–71. Annals of Science, 9(2):101–123, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cole:1954:BNJ

[Col54a] R. J. Cole, B.Sc., F.R.I.C., A.M.I.Chem.E. Biographical note on John Symmons, F.R.S. Annals of Science, 10(3):272–273, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coleby:1954:IMJ

[Col54b] L. J. M. Coleby, M.A., M.Sc., Ph.D. Isaac Milner and the Jacksonian Chair of Natural Philosophy. Annals of Science, 10(3):234–257, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coley:1964:PCS

[Col64] N. G. Coley, M.Sc. The physico-chemical studies of Amedeo Avogadro. Annals of Science, 20(3):195–210, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coles:1995:AZT

[Col95] B. R. Coles. An absolute zero of temperature: Locke’s enunciation of the concept. Annals of Science, 52(4):411–412, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Comfort:2021:HUM

[Com21] Nathaniel Comfort. Heredity under the microscope: Chromosomes and the study of the human genome. Annals of Science, 78(4):533– 535, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cook:1990:SJC

[Coo90] Harold J. Cook. Sir John Colbatch and Augustan medicine: Experimentalism, character and entrepreneurialism. Annals of Science, 47(5):475–505, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cooter:2009:FIS

[Coo09] Roger Cooter. Figuring it out: Science, gender, and visual culture. Annals of Science, 66(4):580–582, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cooter:2011:CMC

[Coo11] Roger Cooter. Contested medicine: Cancer research and the military. Annals of Science, 68(4):565–567, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cooke:2013:BRA

[Coo13] Roger Cooke. Book review: Amir Alexander, Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Cambridge, MA and London: Harvard University Press, 2010. 307 pp. $28.95. ISBN-13 978-0-674-04661-0. Annals of Science, 70(1):136–138, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Copenhaver:1980:JTS

[Cop80] Brian P. Copenhaver. Jewish theologies of space in the scientific revolution: Henry More, Joseph Raphson, Isaac Newton and their predecessors. Annals of Science, 37(5):489–548, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Copenhaver:1987:SPE

[Cop87] Brian P. Copenhaver. Science and philosophy in early modern Europe: The historiographical significance of the work of Charles B. Schmitt. Annals of Science, 44(5):507–517, September 1987. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cornell:1936:ESR

[Cor36] E. S. Cornell, M.Sc. Early studies in radiant heat. Annals of Science, 1(2):217–225, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cornell:1938:RHSa

[Cor38a] E. S. Cornell, M.Sc. The radiant heat spectrum from Herschel to Melloni. I. The work of Herschel and his contemporaries. Annals of Science, 3(1):119–137, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cornell:1938:RHSb

[Cor38b] E. S. Cornell, M.Sc. The radiant heat spectrum from Herschel to Melloni. II. The work of Melloni and his contemporaries. Annals of Science, 3(4):402–416, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Corsi:1975:ER

[Cor75] Pietro Corsi. Essay review. Annals of Science, 32(6):583–586, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Corneanu:2011:RMB

[Cor11] Sorana Corneanu. Regimens of the mind: Boyle, Locke, and the early modern cultura animi tradition. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-11639-5 (hardcover). ix + 308 pp. LCCN B1131 .C67 2011.

Cosans:2009:OAD

[Cos09] Christopher Ernest Cosans. Owen’s ape and Darwin’s bulldog: beyond Darwinism and creationism. Indiana University Press, Bloomington, 2009. ISBN 0-253-35272-X (hardcover), 0-253-22051-3 (paperback). xxi + 166 pp. LCCN GN281 .C598 2009. URL http://

www.loc.gov/catdir/toc/ecip0827/2008035752.html.

Cotter:1976:GBA

[Cot76] Charles H. Cotter. George Biddell Airy and his mechanical correction of the magnetic compass. Annals of Science, 33(3):263–274, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cotter:1977:EHS

[Cot77] Charles H. Cotter. The early history of ship magnetism: The Airy– Scoresby controversy. Annals of Science, 34(6):589–599, 1977. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cotter:2008:DIA

[Cot08] Donald Cotter. A disciplinary immigrant. Alexander Smith at the University of Chicago, 1894–1911. Annals of Science, 65(2):221– 256, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Coutts:1959:WCW

[Cou59] A. Coutts. William Cruickshank of Woolwich. Annals of Science, 15(2):121–133, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cozzoli:2015:BRB

[Coz15] Daniele Cozzoli. Book review: Toxicants, Health and Regulation since 1945. Annals of Science, 72(1):143–144, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cozzoli:2019:BRC

[Coz19] Daniele Cozzoli. Book review: The circulation of penicillin in Spain: health, wealth and authority, by María Jesús Santesmases, London, Palgrave MacMillan, 2017, i–x; 1–238 pp., $99.99 (hbk), ISBN 978-3-319-69717-8. Annals of Science, 76(3–4):371–372, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carneiro:1997:CAP

[CP97] Ana Carneiro and Natalie Pigeard. Chimistes alsaciens à Paris au 19ème siècle: un réseau, une école?. (French) [Alsace chemists in Paris in the 19th Century: a network, a school?]. Annals of Science, 54(6):533–546, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carneiro:2002:HSP

[CP02] Robert L. Carneiro and Robert G. Perrin. Herbert Spencer’s principles of sociology: a centennial retrospective and appraisal. Annals of Science, 59(3):221–261, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carazza:1996:FMM

[CR96] Bruno Carazza and Nadia Robotti. The first molecular models for an electromagnetic theory of dispersion and some aspects of physics at the end of the nineteenth century. Annals of Science, 53(6):587– 607, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Carazza:2002:EAS

[CR02] Bruno Carazza and Nadia Robotti. Explaining atomic spectra within classical physics: 1897–1913. Annals of Science, 59(3):299–320, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cranefield:1954:CMC

[Cra54] Paul F. Cranefield, Ph.D. Clerk Maxwell’s corrections to the page proofs of A Dynamical Theory of the Electromagnetic Field. Annals of Science, 10(4):359–362, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crawforth:1985:ETC

[Cra85] M. A. Crawforth. Evidence from trade cards for the scientific instrument industry. Annals of Science, 42(5):453–544, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crawforth:1987:IML

[Cra87] M. A. Crawforth. Instrument makers in the London guilds. Annals of Science, 44(4):319–377, July 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crawford:2012:MNW

[Cra12] Matthew James Crawford. Measuring the New World: Enlightenment science and South America. Annals of Science, 69(3):441–443, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crellin:1966:APG

[Cre66] J. K. Crellin, B.Pharm., M.Sc. Airborne particles and the germ theory: 1860–1880. Annals of Science, 22(1):49–60, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Creager:2010:EBS

[Cre10] Angela N. H. Creager. Elizabeth Blackburn and the story of telomeres: Deciphering the ends of DNA. Annals of Science, 67(2):265– 268, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crombie:1948:SRH

[Cro48] A. C. Crombie, B.Sc., Ph.D. Some reflections on the history of science and its conception of nature. Annals of Science, 6(1):54–75, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:1959:UDC

[Cro59] M. P. Crosland, M.Sc., Ph.D. The use of diagrams as chemical ‘equations’ in the lecture notes of William Cullen and Joseph Black. Annals of Science, 15(2):75–90, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:1961:OGL

[Cro61] M. P. Crosland, M.Sc., Ph.D. The origins of Gay-Lussac’s law of combining volumes of gases. Annals of Science, 17(1):1–26, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cross:1996:PCD

[Cro96] Sherrie Cross. Prestige and comfort: The development of social Darwinism in early Meiji Japan, and the role of Edward Sylvester Morse. Annals of Science, 53(4):323–344, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:2003:DBE

[Cro03] Maurice Crosland. Difficult beginnings in experimental science at Oxford: the gothic chemistry laboratory. Annals of Science, 60(4): 399–421, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:2005:ELC

[Cro05] Maurice Crosland. Early laboratories c. 1600–c. 1800 and the location of experimental science. Annals of Science, 62(2):233–253, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:2009:FAS

[Cro09a] Maurice Crosland. The French Academy of Sciences as a patron of the medical sciences in the early Nineteenth Century. Annals of Science, 66(2):247–265, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:2009:MEC

[Cro09b] Maurice Crosland. Materials in eighteenth-century science. A historical ontology. Annals of Science, 66(3):421–422, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crosland:2010:PCS

[Cro10] Maurice Crosland. Pensions for ‘cultivators of science’. Annals of Science, 67(4):527–559, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Crockatt:2016:ETC

[Cro16] Richard Crockatt. Einstein and Twentieth-century Politics: ‘a Salutary Moral Influence’. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2016. ISBN 0-19-878549-6 (hardcover), 0-19-108829-3 (e-book), 0-19-182740-1 (e-book). viii + 236 pp. LCCN QC16.E5 C76 2016.

Carolino:2012:EAH

[CS12] Luís Miguel Carolino and Ana Simões. The eclipse, the astronomer and his audience: Frederico Oom and the total solar eclipse of 28 May 1900 in Portugal. Annals of Science, 69(2):215–238, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cook:1980:ERB

[CSVK80] Harold J. Cook, Nicholas H. Steneck, Arthur J. Vander, and Gordon L. Kane. Early research on the biological effects of microwave radiation: 1940–1960. Annals of Science, 37(3):323–351, 1980. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cecilia:1969:NTH

[CT69] R. Cecilia and H. Tanner Ph.D. Nathaniel Torporley and the Harriot manuscripts. Annals of Science, 25(4):339–349, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cumming:1984:JMM

[Cum84] David A. Cumming. John MacCulloch’s ‘Millstone Survey’ and its consequences. Annals of Science, 41(6):567–591, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Cunningham:2004:DMC

[Cun04] Clifford Cunningham. Discovery of the missing correspondence between Carl Friedrich Gauss and the Rev. Nevil Maskelyne (1802–5). Annals of Science, 61(4):469–481, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Curtis:1987:DE

[Cur87] Ronald Curtis. Darwin as an epistemologist. Annals of Science, 44 (4):379–408, July 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chang:2005:AMW

[CY05] Hasok Chang and Sang Wook Yi. The absolute and its measurement: William Thomson on temperature. Annals of Science, 62(3):281– 308, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Chu:2019:REI

[CZ19] Longfei Chu and Haohao Zhu. Re-examining the impact of European astronomy in seventeenth-century China: a study of Xue Fengzuo’s system of thought and his integration of Chinese and Western knowledge. Annals of Science, 76(3–4):303–323, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dingwall:2011:SMI

[D^{+}11] Helen M. Dingwall et al. Scottish medicine: an illustrated history. Birlinn, Edinburgh, Scotland, 2011. ISBN 1-78027-018-6. xvii + 286

pp. LCCN R495 .D542 2011. URL http://www.loc.gov/catdir/
enhancements/fy1212/2012429543-b.html;   http://www.loc.
gov/catdir/enhancements/fy1212/2012429543-d.html.

deAsua:2021:TMS

[dA21] Miguel de Asúa. Traces on a muddy shore. Science and religion in colonial and early independent Río de la Plata. Annals of Science, 78(2):197–220, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DalPrete:2014:BIN

[Dal14] Ivano Dal Prete. Brokering instruments in Napoleon’s Europe: The Italian journeys of Franz Xaver von Zach (1807–1814). Annals of Science, 71(1):82–101, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Danielson:2010:SLR

[Dan10] Dennis Danielson. Science, literature and rhetoric in early modern England. Annals of Science, 67(2):270–272, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Darden:1985:HVL

[Dar85] Lindley Darden. Hugo de Vries’s lecture plates and the discovery of segregation. Annals of Science, 42(3):233–242, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Darrigol:2012:HOG

[Dar12] Olivier Darrigol. A history of optics from Greek antiquity to the Nineteenth Century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-964437-3 (hardcover). xii + 327 pp. LCCN QC355.3 .D37 2012.

Darden:2013:BOF

[Dar13] Lindley Darden. The Annotated Origin: A facsimile of the first edition of On the Origin of Species, by James T. Costa. Annals of Science, 70(3):417–419, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Daston:2009:ESC

[Das09] L. Daston. The emergence of a scientific culture. Science and the shaping of modernity 1210–1685. Annals of Science, 66(2):284–288, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Daum:2013:DWO

[Dau13] Andreas W. Daum. Ein ‘Diplomat aus den Wäldern des Orinoko’: Alexander von Humboldt als Mittler zwischen Preußen und Frankreich. (German) [A ‘diplomat from the forests of the Orinoco’: Alexander von Humboldt as a mediator between Prussia and France]. Annals of Science, 70(3):428–431, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Daum:2014:BRBb

[Dau14a] Andreas W. Daum. Book review: Alexander von Humboldt–August Böckh: Briefwechsel. Annals of Science, 71(3):439–441, 2014. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Daum:2014:BRBa

[Dau14b] Andreas W. Daum. Book review: Nature’s Interpreter: The Life and Times of Alexander von Humboldt, by Donald McCrory. Annals of Science, 71(1):118–121, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

David:1957:MNM

[Dav57] Florence N. David, D.Sc. Mr Newton, Mr Pepys & Dyse: a historical note. Annals of Science, 13(3):137–147, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davies:1966:ECD

[Dav66] Gordon L. Davies, M.A. The eighteenth-century denudation dilemma and the Huttonian theory of the Earth. Annals of Science, 22(2):129– 138, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davies:1968:TBI

[Dav68] Gordon L. Davies, M.A., Ph.D. The tour of the British Isles made by Louis Agassiz in 1840. Annals of Science, 24(2):131–146, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davies:1978:LDS

[Dav78] Alun C. Davies. The life and death of a scientific instrument: The marine chronometer, 1770–1920. Annals of Science, 35(5):509–525, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davis:1984:WFD

[Dav84] John L. Davis. Weather forecasting and the development of meteorological theory at the Paris Observatory, 1853–1878. Annals of Science, 41(4):359–382, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davies:1990:WAR

[Dav90a] Mansel Davies. W. T. Astbury, Rosie Franklin, and DNA: a memoir. Annals of Science, 47(6):607–618, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davis:1990:AMG

[Dav90b] Audrey B. Davis. American medicine in the gilded age: The first technological era. Annals of Science, 47(2):111–125, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davies:1992:FSS

[Dav92] Mansel Davies. Frederick Soddy: The scientist as prophet. Annals of Science, 49(4):351–367, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davis:1995:RSM

[Dav95] J. L. Davis. The research school of Marie Curie in the Paris faculty, 1907–14. Annals of Science, 52(4):321–355, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Davis:1998:ASR

[Dav98] John L. Davis. Artisans and savants: The role of the Academy of Sciences in the process of electrical innovation in France, 1850–1880. Annals of Science, 55(3):291–314, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBruin:1986:GSE

[dBP86] Joanneke de Bruin and Lewis Pyenson. ‘gentleman-scientist’: Elie van Rijckevorsel and the Dutch overseas effort in exact sciences at the end of the nineteenth century. Annals of Science, 43(5):447– 473, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deClercq:1988:SCE

[dC88] Peter de Clercq. Science at court: the eighteenth-century cabinet of scientific instruments and models of the Dutch stadholders. Annals of Science, 45(2):113–152, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

daCosta:2009:SOI

[dC09a] Palmira Fontes da Costa. Secrecy, ostentation, and the illustration of exotic animals in sixteenth-century Portugal. Annals of Science, 66(1):59–82, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Darwin:2009:ABF

[DC09b] Charles Darwin and James T. Costa. The annotated Origin: a facsimile of the first edition of On the Origin of Species. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 2009. ISBN 0-674-03281-0. xx + ix + 537 pp. LCCN QH365 .O2 2009.

Andrade:1936:RCB

[dCA36] E. N. da C. Andrade, D.Sc., Ph.D., F.R.S. The real character of Bishop Wilkins. Annals of Science, 1(1):4–12, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dagg:2021:CDD

[DD21] Joachim L. Dagg and J. F. Derry. Charles Darwin did not mislead Joseph Hooker in their 1881 correspondence about Leopold von Buch and Karl Ernst von Baer. Annals of Science, 77(3):349–365, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1949:SLJ

[de 49] G. R. de Beer, M.A., D.Sc., F.R.S., F.L.S., F.Z.S. Some letters from Jakob Samuel Wyttenbach to Sir James Edward Smith. Annals of Science, 6(2):105–114, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1950:CV

[de 50a] G. R. de Beer, M.A., D.Sc., F.R.S. Combe–Varin. Annals of Science, 6(3):215–228, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1950:JHH

[de 50b] G. R. de Beer, M.A., D.Sc., F.R.S. Johann Heinrich Hottinger’s description of the ice-mountains of Switzerland, 1703. Annals of Science, 6(4):327–360, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1953:BFK

[de 53a] G. R. de Beer, D.Sc., F.R.S. Bernhard Friedrich Kuhn’s investigations on glaciers. Annals of Science, 9(4):323–341, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1953:HBS

[de 53b] G. R. de Beer, M.A., D.Sc., F.R.S. Haller’s Historia Stirpium. Annals of Science, 9(1):1–46, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1954:JJR

[de 54] Sir Gavin de Beer, D.Sc., Hon.D.-ès-L., F.R.S. Jean-Jacques Rousseau: Botanist. Annals of Science, 10(3):189–223, 1954. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1956:HAM

[de 56] Sir Gavin de Beer, D.Sc., F.R.S., F.S.A. The history of the altimetry of Mont Blanc. Annals of Science, 12(1):3–29, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1957:GWT

[de 57] Sir Gavin de Beer, D.Sc., Sc.D., D.-ès-L., F.R.S., F.S.A. Gregory Watt’s tour on the Continent, 1801. Annals of Science, 13(3):127– 136, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1958:FUL

[de 58] Sir Gavin de Beer, D.Sc., Sc.D., D.-ès-L., F.R.S., F.S.A. Further unpublished letters of Charles Darwin. Annals of Science, 14(2):83– 115, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1960:AM

[de 60] Sir Gavin de Beer, D.Sc., F.S.A., F.R.S. Alexander Moritzi. Annals of Science, 16(4):251–254, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1962:VA

[de 62] Sir Gavin de Beer, D.Sc., F.S.A., F.R.S. The volcanoes of Auvergne. Annals of Science, 18(1):49–61, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deBeer:1964:OMS

[de 64] Sir Gavin de Beer, D.Sc., F.S.A., F.R.S. Other men’s shoulders. Annals of Science, 20(4):303–322, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DeBont:2005:RTE

[De 05] Raf De Bont. Rome and theistic evolutionism: The hidden strategies behind the ‘Dorlodot Affair’, 1920–1926. Annals of Science, 62(4): 457–478, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DeClark:2017:QVQ

[De 17] Sybil Gertrude De Clark. Qualitative vs quantitative conceptions of homogeneity in nineteenth century dimensional analysis. Annals of Science, 74(4):299–325, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Douglas:1950:JPG

[DE50] J. A. Douglas, M.A., D.Sc. and J. M. Edmonds, M.A., B.Sc. John Phillips’s geological maps of the British Isles. Annals of Science, 6 (4):361–375, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1973:JHH

[Dea73] Dennis R. Dean, A.B., A.M., Ph.D. James Hutton and his public, 1785–1802. Annals of Science, 30(1):89–105, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1975:JHR

[Dea75] Dennis R. Dean. James Hutton on religion and geology: the unpublished preface to his Theory of the Earth (1788). Annals of Science, 32(3):187–193, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1979:WG

[Dea79] Dennis R. Dean. The word ‘geology’. Annals of Science, 36(1):35– 43, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1981:AEC

[Dea81] Dennis R. Dean. The age of the Earth controversy: Beginnings to Hutton. Annals of Science, 38(4):435–456, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1983:JPH

[Dea83] Dennis R. Dean. John Playfair and his books. Annals of Science, 40(2):179–187, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1989:BFE

[Dea89a] Dennis R. Dean. Benjamin Franklin and earthquakes. Annals of Science, 46(5):481–495, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1989:NLW

[Dea89b] Dennis R. Dean. New light on William Maclure. Annals of Science, 46(6):549–574, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1991:JMO

[Dea91a] Dennis R. Dean. John Muir and the origin of Yosemite Valley. Annals of Science, 48(5):453–485, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1991:RMF

[Dea91b] Dennis R. Dean. Robert Mallet and the founding of seismology. Annals of Science, 48(1):39–67, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dean:1993:SFE

[Dea93] Dennis R. Dean. The San Francisco earthquake of 1906. Annals of Science, 50(6):501–521, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Debus:1962:EHM

[Deb62] Allen G. Debus, Ph.D. An Elizabethan history of medical chemistry. Annals of Science, 18(1):1–29, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Debus:1967:FAE

[Deb67] Allen G. Debus, B.S., A.M., Ph.D. Fire analysis and the elements in the sixteenth and the seventeenth centuries. Annals of Science, 23(2):127–147, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deiman:1991:MRC

[Dei91] J. C. Deiman. A myth revealed: The case of the ‘Beeldsnyder achromatic objective’. Annals of Science, 48(6):577–581, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DeKosky:1983:WCQ

[DeK83] Robert K. DeKosky. William Crookes and the quest for absolute vacuum in the 1870s. Annals of Science, 40(1):1–18, January 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:1987:EES

[Dek87] E. Dekker. Early explorations of the southern celestial sky. Annals of Science, 44(5):439–470, September 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:1990:LDR

[Dek90] Elly Dekker. The light and the dark: a reassessment of the discovery of the Coalsack Nebula, the Magellanic Clouds and the Southern Cross. Annals of Science, 47(6):529–560, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:1993:ETI

[Dek93] Elly Dekker. Epact tables on instruments: Their definition and use. Annals of Science, 50(4):303–324, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:1995:UMA

[Dek95] Elly Dekker. An unrecorded medieval astrolabe quadrant from c. 1300. Annals of Science, 52(1):1–47, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:1996:CGD

[Dek96] Elly Dekker. The Copernican globe: a delayed conception. Annals of Science, 53(6):541–566, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dekker:2008:HSL

[Dek08] Elly Dekker. ‘with his sharp lok perseth the sonne’: a new quadrant from Canterbury. Annals of Science, 65(2):201–220, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2009:BRB

[Del09a] Robert J. Deltete. Book review: The End of the Certain World: The Life and Science of Max Born, the Nobel Scientist who Ignited the Quantum Revolution. Annals of Science, 66(3):433–436, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Gre05].

Deltete:2009:ML

[Del09b] Robert J. Deltete. Mathematics and logic. Annals of Science, 66(1): 141–143, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2011:ERH

[Del11a] Robert Deltete. Essay review: How Relativity got accepted and how Einstein came to be regarded as its author. Richard Stayley, Einstein’s Generation: The Origins of the Relativity Revolution, ISBN-13 978-0-226-77057-5. Annals of Science, 68(2):261–265, 2011. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2011:BRF

[Del11b] Robert J. Deltete. Book review: Fulvio Melia, Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics, ISBN-13 978-0-226-51951-7. Annals of Science, 68(4):581–583, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Deltete:2012:ERE

[Del12] Robert Deltete. Essay review: Einstein’s quest for unity: Jeroen van Dongen, Einstein’s Unification, ISBN 978-0-521-88346-7. Annals of Science, 69(2):283–288, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See corrigendum [Ano12a].

Deltete:2013:HAA

[Del13] Robert J. Deltete. Heisenberg in the atomic age: Science and the public sphere. Annals of Science, 70(3):419–422, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Despeaux:2008:MSA

[Des08] Sloan Evans Despeaux. Mathematics sent across the channel and the Atlantic: British mathematical contributions to European and American scientific journals, 1835–1900. Annals of Science, 65(1): 73–99, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Determann:2018:LOI

[Det18] Jörg Matthias Determann. The lighthouse and the observatory: Islam, science, and empire in Late Ottoman Egypt. Annals of Science, 75(4):375–377, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DalPrete:2009:IAV

[DG09] Ivano Dal Prete and Dario Generali. On the inventory of Antonio Vallisneri’s correspondence. Annals of Science, 66(3):453–454, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

DiGennaroSplendore:2017:CMM

[Di 17] Barbara Di Gennaro Splendore. Craft, money and mercy: an apothecary’s self-portrait in sixteenth-century Bologna. Annals of Science, 74(2):91–107, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dias:1994:CMS

[Dia94] Penha Maria Cardoso Dias. Clausius and Maxwell: The statistics of molecular collisions (1857–1862). Annals of Science, 51(3):249– 261, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dias:1996:WTH

[Dia96] Penha Maria Cardoso Dias. William Thomson and the heritage of caloric. Annals of Science, 53(5):511–520, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dietz:2012:CCP

[Die12] Bettina Dietz. Contribution and co-production: The collaborative culture of Linnaean botany. Annals of Science, 69(4):551–569, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dietz:2016:ISI

[Die16a] Bettina Dietz. Introduction: Special issue ‘Translating and translations in the history of science’. Annals of Science, 73(2):117–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dietz:2016:LRS

[Die16b] Bettina Dietz. Linnaeus’ restless system: translation as textual engineering in eighteenth-century botany. Annals of Science, 73(2):143– ??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dietz:2021:KTG

[Die21] Bettina Dietz. Knowledge in translation: global patterns of scientific exchange, 1000–1800 CE. Annals of Science, 77(3):388–389, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dijksterhuis:2004:OSB

[Dij04] Fokko Jan Dijksterhuis. Once Snell breaks down: From geometrical to physical optics in the Seventeenth Century. Annals of Science, 61(2):165–185, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dijksterhuis:2013:BRB

[Dij13] Fokko Jan Dijksterhuis. Book review: Materials and Expertise in Early Modern Europe. Between Market and Laboratory, edited by Ursula Klein and Emma C. Spary. Annals of Science, 70(4):559– 562, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dingle:1950:TWA

[Din50] Herbert Dingle, D.Sc., D.I.C., A.R.C.S. Thomas Wright’s astronomical heritage. Annals of Science, 6(4):404–415, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1954:ALL

[DK54a] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein, M.D. Antoine Laurent Lavoisier (1743–1794) and Christopher Columbus (1446?– 1506). Annals of Science, 10(1):63–68, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1954:BSI

[DK54b] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein M.D. A bibliographical study of the introduction of Lavoisier’s Traité élémentaire de chimie into Great Britain and America. Annals of Science, 10(4): 321–338, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1954:LBB

[DK54c] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein, M.D. A letter from Berthollet to Blagden relating to the experiments for a largescale synthesis of water carried out by Lavoisier and Meusnier in 1785. Annals of Science, 10(1):58–62, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1955:BFAa

[DK55a] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein M.D. Benjamin Franklin (1706–1790) and Antoine Laurent Lavoisier (1743– 1794) Part I. Franklin and the new chemistry. Annals of Science, 11(2):103–128, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1955:BFAb

[DK55b] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein M.D. Benjamin Franklin (1706–1790) and Antoine Laurent Lavoisier (1743–1794). Part II. Joint investigations. Annals of Science, 11(4):271–302, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1957:BFA

[DK57] Denis I. Duveen, F.R.I.C. and Herbert S. Klickstein, M.D. Benjamin Franklin (1706–1790) and Antoine Laurent Lavoisier (1743–1794). Part III. Documentation. Annals of Science, 13(1):30–46, 1957. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dunbar:2011:UC

[DK11] Deanne Dunbar and Howard I. Kushner. Universalism in context. Annals of Science, 68(1):125–132, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Daston:2011:HSO

[DL11] Lorraine Daston and Elizabeth Lunbeck, editors. Histories of scientific observation. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-13677-9 (hardcover), 0-226-13678-7 (paperback). 460 + 1 pp. LCCN Q174.8 .H57 2011.

Dobbin:1936:CCM

[Dob36] Leonard Dobbin, Ph.D. A Cullen chemical manuscript of 1753. Annals of Science, 1(2):138–156, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dobbin:1945:HDP

[Dob45] Leonard Dobbin, Ph.D. The history of the discovery of phosgene. Annals of Science, 5(3):270–287, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dolby:1977:TTN

[Dol77] R. G. A. Dolby. The transmission of two new scientific disciplines from Europe to North America in the late nineteenth century. Annals of Science, 34(3):287–310, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Domingues:2009:BRM

[Dom09] J. Caramalho Domingues. Book review: Mathematics and logic. Robert E. Bradley and C. Edward Sandifer, editors, Leonhard Euler: Life, Work and Legacy. Studies in the History and Philosophy of Mathematics, 5. ISBN-13 978-0-444-52728-8. Annals of Science, 66 (3):446–448, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Donald:1936:HCNa

[Don36a] M. B. Donald, M.Sc., F.I.C., M.I.Chem.E. History of the Chile nitrate industry. I. Annals of Science, 1(1):29–47, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Donald:1936:HCNb

[Don36b] M. B. Donald, M.Sc., F.I.C., M.I.Chem.E. History of the Chile nitrate industry. II. Annals of Science, 1(2):193–216, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Donald:1950:BKC

[Don50] M. B. Donald, M.Sc., F.R.I.C., M.I.Chem.E., A.R.C.S. Burchard Kranich (c. 1515–1578), miner and queen’s physician, Cornish mining stamps, antimony and, Frobisher’s gold. Annals of Science, 6(3):308– 322, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See note [Don51].

Donald:1951:FNB

[Don51] M. B. Donald, M.Sc., F.R.I.C., M.I.Chem.E., A.R.C.S. A further note on Burchard Kranich. Annals of Science, 7(1):107–108, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Don50].

Donovan:1979:HSU

[Don79] Arthur Donovan. The history of science in undergraduate education. Annals of Science, 36(2):177–181, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Donnelly:1988:CEE

[Don88] J. F. Donnelly. Chemical engineering in England, 1880–1922. Annals of Science, 45(6):555–590, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Doody:2010:PER

[Doo10] Aude Doody. Pliny’s Encyclopedia: the reception of the natural history. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-49103-7 (hardcover). viii + 194 pp. LCCN Q143.P64

D66 2010. URL http://assets.cambridge.org/97805214/91037/
cover/9780521491037.jpg.

Dorling:1976:GMC

[Dor76] Alison R. Dorling. The Graves Mathematical Collection in University College London. Annals of Science, 33(3):307–309, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Drubba:1954:FES

[DR54] H. Drubba, Dip.-Phys. and H. H. Rust, Dr.Phil. On the first echosounding experiment. Annals of Science, 10(1):28–32, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Drake:1981:ADG

[Dra81] Stillman Drake. Alleged departures from Galileo’s Law of Descent. Annals of Science, 38(3):339–342, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Drake:1982:AGE

[Dra82] Stillman Drake. Analysis of Galileo’s experimental data. Annals of Science, 39(4):389–397, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Drake:1986:GPA

[Dra86] Stillman Drake. Galileo and the projection argument. Annals of Science, 43(1):77–79, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Drake:1988:TAD

[Dra88] Stillman Drake. The tower argument in the Dialogue. Annals of Science, 45(3):295–302, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dritsas:2012:GR

[Dri12] Lawrence Dritsas. Geography and revolution. Annals of Science, 69(2):289–291, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dittmann:1997:HST

[DS97] Frank Dittmann and Jérôme Ségal. Hermann Schmidt (1894–1968) et la théorie générale de la régulation: Une cybernétique allemande en 1940?. (French) [Hermann Schmidt (1894–1968) and the general theory of regulation: a German cybernetics in 1940?]. Annals of Science, 54(6):547–565, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dyck:2016:UHE

[DS16] Erika Dyck and Larry Stewart, editors. The uses of humans in experiment: perspectives from the 17th to the 20th century, volume 95 of Clio medica: perspectives in medical humanities. Brill Rodopi, Leiden, The Netherlands, 2016. ISBN 90-04-28670-5 (hardcover), 90-04-28671-3 (e-book). ISSN 0045-7183 (print), 1875-6689 (electronic). xii + 297 pp. LCCN R853.H8 U85 2016.

Dubbey:1963:IDN

[Dub63] J. M. Dubbey, M.Sc. The introduction of the differential notation to Great Britain. Annals of Science, 19(1):37–48, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dubbey:1966:CCE

[Dub66] J. M. Dubbey, M.Sc., A.K.C. Cauchy’s contribution to the establishment of the calculus. Annals of Science, 22(1):61–67, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duck:1988:NGC

[Duc88] Michael J. Duck. Newton and Goethe on colour: Physical and physiological considerations. Annals of Science, 45(5):507–519, 1988. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duck:1997:QFG

[Duc97] Michael Duck. A question of faith: Goethe’s belief in the immutability of light. Annals of Science, 54(4):397–406, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ducheyne:2009:SHH

[Duc09] Steffen Ducheyne. The secret history of Hermes: Hermeticism from ancient times to modern times. Annals of Science, 66(2):294–295, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ducheyne:2013:BRB

[Duc13] Steffen Ducheyne. Book review: Mathématiques et connaissance du monde réel avant Galilée. Series: Histoire des savoirs, by Sabine Rommevaux. Annals of Science, 70(4):574–575, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ducheyne:2014:BRB

[Duc14] Steffen Ducheyne. Book review: La via delle acque (1500–1700). Appropriazione delle arti e trasformazione delle matematiche, by Cesare Maffioli. Annals of Science, 71(1):104–105, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ducheyne:2017:DSN

[Duc17] Steffen Ducheyne. Different shades of Newton: Herman Boerhaave on Newton mathematicus, philosophus, and optico-chemicus. Annals of Science, 74(2):108–125, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ducheyne:2018:TRM

[Duc18] S. Ducheyne. Thomas Reid on mathematics and natural philosophy. Annals of Science, 75(4):369–371, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duncan:1962:STAa

[Dun62a] A. M. Duncan, M.A., M.Sc., Dip.Ed. Some theoretical aspects of eighteenth-century tables of affinity. I. Annals of Science, 18(3):177– 194, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duncan:1962:STAb

[Dun62b] A. M. Duncan, M.A., M.Sc., Dip.Ed. Some theoretical aspects of eighteenth-century tables of affinity. II. Annals of Science, 18(4): 217–232, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duncan:1967:WKB

[Dun67] A. M. Duncan, M.A., M.Sc. William Keir’s De Attractione Chemica (1778) and the concepts of chemical saturation, attraction and repulsion. Annals of Science, 23(2):149–173, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dunn:1994:TPA

[Dun94] Richard Dunn. The true place of astrology among the mathematical arts of late Tudor England. Annals of Science, 51(2):151–163, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dunn:2013:BRB

[Dun13] Rachel Dunn. Book review: Popularizing Science and Technology in the European Periphery, 1800–2000, edited by F. Papanelopoulou, A. Nieto-Galan, and E. Perdiguero; Science for All: The Popularisation of Science in Early Twentieth Century Britain, by P. J. Bowler. Annals of Science, 70(4):590–593, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duner:2021:IIT

[Dun21] David Dunér. From influence to inhabitation: the transformation of astrobiology in the early modern period. Annals of Science, 77(3): 389–392, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dunn:2023:BSE

[Dun23] Richard Dunn. On being sufficiently exact: assessing navigational instruments in the eighteenth century. Annals of Science, 81(1–2): 208–234, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2282779.

Dupre:2011:MHI

[Dup11] Sven Dupré. The mindful hand: Inquiry and invention from the Late Renaissance to early industrialization. Annals of Science, 68(2):291– 294, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duris:2010:LRS

[Dur10] P. Duris. L’introuvable révolution scientifique. Francesco Redi et la génération spontanée. (French) [the unreachable scientific revolution. Francesco Redi and spontaneous generation]. Annals of Science, 67 (4):431–455, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1952:URW

[Duv52] Denis Duveen, F.R.I.C. An unpublished report on the waterproofing of shoe leather by Lavoisier and Hassenfratz. Annals of Science, 8 (2):162–164, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Duveen:1954:AFS

[Duv54] Denis I. Duveen, F.R.I.C. Augustin François Silvestre and the Société Philomathique. Annals of Science, 10(4):339–341, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

deVries:2021:EFI

[dV21] Lyke de Vries. The experimental fire: Inventing English alchemy, 1300–1700. Annals of Science, 78(3):395–397, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heathcote:1936:ENC

[dVH36] N. H. de Vaudrey Heathcote, B.Sc. Early nautical charts. Annals of Science, 1(1):13–28, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heathcote:1950:GSG

[dVH50] N. H. de V. Heathcote, B.Sc., Ph.D. Guericke’s sulphur globe. Annals of Science, 6(3):293–305, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heathcote:1965:JCI

[dVH65] Niels H. de V. Heathcote, B.Sc., Ph.D. N. J. Callan: inventor of the induction coil. Annals of Science, 21(3):145–167, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heathcote:1967:EME

[dVH67] Niels H. de V. Heathcote, B.Sc., Ph.D. The early meaning of electricity: Some Pseudodoxia Epidemica. I. Annals of Science, 23(4): 261–275, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heathcote:1957:R

[dVHDB57] N. H. de V. Heathcote, F. N. David, and G. Burniston Brown. Reviews. Annals of Science, 13(3):199–203, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dyment:1937:SEC

[Dym37] S. A. Dyment, M.Sc. Some eighteenth century ideas concerning aqueous vapour and evaporation. Annals of Science, 2(4):465–473, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Dym:2011:FFL

[Dym11] Warren Alexander Dym. Freiberg and the frontier: Louis Janin, German engineering, and ‘civilisation’ in the American west. Annals of Science, 68(3):295–323, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Earles:1960:EIV

[Ear60] M. P. Earles, B.Pharm., M.Sc., Ph.D. The experimental investigation of viper venom by Felice Fontana (1730–1805). Annals of Science, 16(4):255–268, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Earles:1961:ETM

[Ear61] M. P. Earles, B.Pharm., M.Sc., Ph.D., F.P.S. Early theories of the mode of action of drugs and poisons. Annals of Science, 17(2):97– 110, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Earles:1963:EDP

[Ear63] M. P. Earles, B.Pharm., M.Sc., Ph.D., F.P.S. Experiments with drugs and poisons in the seventeenth and eighteenth centuries. Annals of Science, 19(4):241–254, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eastwood:1986:ALL

[Eas86] Bruce Eastwood. Alhazen, Leonardo, and late-medieval speculation on the inversion of images in the eye. Annals of Science, 43(5):413– 446, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ebenstein:1939:ERD

[Ebe39] William Ebenstein, Ll.D., Ph.D. The early reception of the doctrine of evolution in the United States. Annals of Science, 4(3):306–318, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eccarius:1976:ALC

[Ecc76] W. Eccarius. August Leopold Crelle als Herausgeber wissenschaftlicher Fachzeitschriften. (German) [August Leopold Crelle as a publisher of scientific journals]. Annals of Science, 33(3):229– 261, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eccarius:1980:GZJ

[Ecc80] W. Eccarius. Der Gegensatz zwischen Julius Plücker und Jakob Steiner im Lichte ihrer Beziehungen zu August Leopold Crelle Hintergründe eines wissenschaftlichen Meinungsstreites. (German) [The contrast between Julius Plücker and Jakob Steiner in the light of their relations with August Leopold Crelle: backgrounds of a scientific opinion dispute]. Annals of Science, 37(2):189–213, 1980. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eckert:2010:HH

[Eck10] Michael Eckert. Heinrich Hertz. Hamburger Köpfe. Ellert and Richter, Hamburg, Germany, 2010. ISBN 3-8319-0371-9 (cloth). 168 pp. LCCN QC16.H4 E29 2010.

Eddy:2008:LMJ

[Edd08] Matthew Eddy, editor. The language of mineralogy: John Walker, chemistry and the Edinburgh Medical School, 1750–1800. Science, technology, and culture, 1700–1945. Ashgate, Aldershot, Hants, England and Farnham, England, 2008. ISBN 0-7546-6332-9 (hardcover). xx + 309 pp. LCCN QH31.W16 E33 2008. URL http://www.loc.

gov/catdir/toc/ecip0819/2008023651.html.

Eddy:2013:BRM

[Edd13] John H. Eddy. Book review: Marc J. Ratcliff, The Quest for the Invisible: Microscopy in the Enlightenment. Farnham, England and Burlington, Vermont: Ashgate Pub. Limited, 2009. xii + 315 pp. 6 charts. 14 tables. 34 reprints. $124.95. ISBN-13 978-0-7546-6150-4. Annals of Science, 70(1):107–109, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edmonds:1975:FGL

[Edm75] J. M. Edmonds. The first geological lecture course at the University of London, 1831. Annals of Science, 32(3):257–275, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:1972:OLL

[Edw72a] Nicholas Edwards, Ph.D. One of the last letters of Adam Sedgwick, geologist (1785–1873). Annals of Science, 28(2):109–112, 1972. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:1972:SCT

[Edw72b] Nicholas Edwards, Ph.D. Some correspondence of Thomas Webster (circa 1772–1844), concerning the Royal Institution. Annals of Science, 28(1):43–60, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:2004:GBO

[Edw04] Martin Edwards. Good, bad or offal? The evaluation of raw pancreas therapy and the rhetoric of control in the therapeutic trial, 1925. Annals of Science, 61(1):79–98, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:2009:NEP

[Edw09a] Elizabeth Edwards. Nature exposed: Photography as eyewitness in Victorian science. Annals of Science, 66(3):436–439, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:2009:PTT

[Edw09b] Martin Edwards. Penicillin: Triumph and tragedy. Annals of Science, 66(2):295–297, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Edwards:2010:WMC

[Edw10] Paul N. Edwards. The world in a machine: computer models, data networks, and global atmospheric politics. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01392-4 (hardcover). xxvii + 518 pp. LCCN QC995 .E296 2010.

Eyles:1938:DIF

[EE38] V. A. Eyles, B.Sc. and Joan M. Eyles, B.Sc., F.G.S. On the different issues of the first geological map of England and Wales. Annals of Science, 3(2):190–212, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eyles:1951:SGC

[EE51] V. A. Eyles, B.Sc., F.G.S., F.R.S.E. and Joan M. Eyles, B.Sc., F.G.S. Some geological correspondence of James Hutton. Annals of Science, 7(4):316–339, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Effros:2021:GAP

[Eff21] Bonnie Effros. Germany’s ancient pasts: archaeology and historical interpretation since 1700. Annals of Science, 77(3):394–397, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Egerton:1968:WSA

[Ege68] M. C. Egerton, M.Sc. William Strutt and the application of convection to the heating of buildings. Annals of Science, 24(1):73–87, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Egmond:2016:EDI

[Egm16] Florike Egmond. Eye for Detail: Images of Plants and Animals in Art and Science, 1500-1630. Reaktion Books, London, UK, 2016. ISBN 1-78023-640-9. 288 pp. LCCN QH46.5 .E46 2017.

Egmond:2018:BRP

[Egm18] Florike Egmond. Book review: The Paper Zoo: 500 Years of Animals in Art, by Charlotte Sleigh, Chicago: University of Chicago Press, 256 pp., ISBN 978-0-226-44712-4. Annals of Science, 75(2):150–151, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Egorova:2019:HWS

[Ego19] Yulia Egorova. History within: the science, culture, and politics of bones, organisms, and molecules. Annals of Science, 76(1):109– 111, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ehrlich:1995:MAS

[Ehr95] Mark E. Ehrlich. Mechanism and activity in the scientific revolution: The case of Robert Hooke. Annals of Science, 52(2):127–151, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eire:2018:LSM

[Eir18] Carlos M. N. Eire. Levitation: The science, myth, and magic of suspension. Annals of Science, 75(4):368–369, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eisenstaedt:2005:BRP

[Eis05] Jean Eisenstaedt. Light and Relativity, a previously unknown eighteenth-century manuscript by Robert Blair (1748–1828). Annals of Science, 62(3):347–376, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ekholm:2010:FHR

[Ekh10] Karin J. Ekholm. Fabricius’s and Harvey’s representations of animal generation. Annals of Science, 67(3):329–352, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Elazar:2011:HFC

[Ela11] Michael Elazar. Honoré Fabri and the Concept of Impetus: a Bridge Between Conceptual Frameworks, volume 288 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 94-007-1604-4 (hardcover), 94-007-1605-2 (e-book). ISSN 0068-0346. xxix + 259 pp. LCCN Q175 .E43 2011 v.288; QA802 .E43

2011. URL http://www.loc.gov/catdir/enhancements/fy1202/
2011928368-d.html;   http://www.loc.gov/catdir/enhancements/
fy1202/2011928368-t.html;   https://link.springer.com/book/
10.1007/978-94-007-1605-6.

Elazar:2013:DSJ

[Ela13] Michael Elazar. A dispute over superposition: John Wallis, Honoré Fabri, and Giovanni Alfonso Borelli. Annals of Science, 70(2):175– 195, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Elliott:1983:CBO

[Ell83] Graham F. Elliott. The contribution of British oil interests in the middle east to palaeontology. Annals of Science, 40(3):273–279, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Elliott:2000:BPS

[Ell00] Paul Elliott. The birth of public science in the English provinces: Natural philosophy in Derby, c. 1690–1760. Annals of Science, 57 (1):61–100, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Emerson:1988:SRS

[Eme88] Roger L. Emerson. Sir Robert Sibbald, Kt, the Royal Society of Scotland and the origins of the Scottish enlightenment. Annals of Science, 45(1):41–72, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Emerson:2002:SIA

[Eme02] Roger L. Emerson. The scientific interests of Archibald Campbell, 1st Earl of Ilay and 3rd Duke of Argyll (1682–1761). Annals of Science, 59(1):21–56, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Emerson:2013:EDL

[Eme13] Roger L. Emerson. An Enlightened Duke: the life of Archibald Campbell (1692–1761), Earl of Ilay, 3rd Duke of Argyll. Perspectives: Scottish studies of the long eighteenth century. Humming Earth, Kilkerran, UK, 2013. ISBN 1-84622-040-8 (hardcover), 1-84622-039-4 (paperback). xvii + 553 pp. LCCN DA814.A84 E44 2013.

Endersby:2016:CDL

[End16] Jim Endersby. Charles Darwin’s looking glass: The theory of evolution and the life of its author in contemporary British fiction and nonfiction. Annals of Science, 73(3):337–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Endersby:2017:BRH

[End17] Jim Endersby. Book review: Heredity Explored: Between Public Domain and Experimental Science, 1850–1930. Annals of Science, 74 (2):176–178, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Endersby:2018:MMD

[End18] Jim Endersby. Of making many Darwins. Annals of Science, 75(4): 361–367, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Erlichson:1992:IIC

[Erl92] Herman Erlichson. The instantaneous impulse construction as a formula for central force motion on an arbitrary plane curve with respect to an arbitrary force centre in the plane of that curve. Annals of Science, 49(4):369–375, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Erlichson:1994:GPP

[Erl94] Herman Erlichson. Galileo’s pendulums and planes. Annals of Science, 51(3):263–272, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Evans:1978:CBB

[Eva78] Gillian R. Evans. A commentary on Boethius’s Arithmetica of the twelfth or thirteenth century. Annals of Science, 35(2):131–141, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Evans:2017:DWS

[Eva17] Samantha Evans, editor. Darwin and women: a selection of letters. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-15886-9 (hardcover), 1-316-67003-1 (e-book), 1-316-61154-X. xxviii + 270 pp. LCCN QH31.D2 D3625 2017.

Everest:2015:BRB

[Eve15] James Everest. Book review: The Secrets of Alchemy. Annals of Science, 72(2):272–273, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eyles:1937:JMF

[Eyl37] V. A. Eyles, B.Sc. John Macculloch, F.R.S., and his geological map: An account of the first geological survey of Scotland. Annals of Science, 2(1):114–129, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eyles:1963:ECS

[Eyl63] V. A. Eyles, D.Sc., F.G.S., F.R.S.E. The evolution of a chemist Sir James Hall, Bt., F.R.S., P.R.S.E., of Dunglass, Haddingtonshire, (1761–1832), and his relations with Joseph Black, Antoine Lavoisier, and other scientists of the period. Annals of Science, 19(3):153– 182, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Eyles:1967:WSS

[Eyl67] Joan M. Eyles, B.Sc., F.G.S. William Smith: The sale of his geological collection to the British Museum. Annals of Science, 23(3): 177–212, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fancher:2009:MTG

[Fan09] Raymond E. Fancher. Monkey trials and gorilla sermons: Evolution and Christianity from Darwin to intelligent design. Annals of Science, 66(3):429–431, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farrar:1969:RLC

[Far69] W. V. Farrar, B.Sc., Ph.D. Richard Laming and the coal-gas industry, with his views on the structure of matter. Annals of Science, 25(3):243–253, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farrar:1972:SPS

[Far72] W. V. Farrar, B.Sc., Ph.D. The Society for the Promotion of Scientific Industry, 1872–1876. Annals of Science, 29(1):81–86, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farrall:1979:HEB

[Far79] Lyndsay A. Farrall. The history of eugenics: a bibliographical review. Annals of Science, 36(2):111–123, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farr:1983:ROO

[Far83] A. D. Farr. Religious opposition to obstetric anaesthesia: a myth? Annals of Science, 40(2):159–177, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farr:1984:R

[Far84] A. D. Farr. Reply. Annals of Science, 41(2):180, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farber:2018:HDV

[Far18] Paul Lawrence Farber. Henry Dresser and Victorian ornithology: Birds, books and business. Annals of Science, 75(3):265–266, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farhat:2019:CDR

[Far19] Georges Farhat. Can diplopia reshape our views of perspective? Annals of Science, 76(2):221–227, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fauque:2005:OLC

[Fau05] Danielle Fauque. Organisation des laboratoires de chimie à Paris sous le ministère Duruy (1863–1869): Cas des laboratoires de Fremy et de Wurtz. (French) [Organization of the chemistry laboratories in Paris under the Dury ministry (1863–1869): case of the laboratories of Fremy and Wurtz]. Annals of Science, 62(4):501–531, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Boistel:2022:MRI

[FB22] Fernando B. Figueiredo and Guy Boistel. Monteiro da Rocha and the international debate in the 1760s on astronomical methods to find the longitude at sea: his proposals and criticisms to Lacaille’s lunar-distance method. Annals of Science, 79(2):215–258, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2059567.

Fulton:1936:BSG

[FC36] John F. Fulton, M.D. and Harvey Cushing, M.D. A bibliographical study of the Galvani and the Aldini writings on animal electricity. Annals of Science, 1(3):239–268, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feingold:2003:NSJ

[Fei03] Mordechai Feingold, editor. The new science and Jesuit science: seventeenth century perspectives, volume 6 of Archimedes. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-0848-1. ix + 270 pp.

LCCN BL245 .N48 2003. URL http://www.loc.gov/catdir/
enhancements/fy0813/2002032017-d.html;   http://www.loc.
gov/catdir/enhancements/fy0813/2002032017-t.html.

Feingold:2019:RYI

[Fei19] Mordechai Feingold. The religion of the young Isaac Newton. Annals of Science, 76(2):210–218, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fellmann:2003:C

[Fel03] E. A. Fellmann. Corrigendum. Annals of Science, 60(4):461, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ferraro:2002:CFM

[Fer02] Giovanni Ferraro. Convergence and formal manipulation of series from the origins of calculus to about 1730. Annals of Science, 59(2): 179–199, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1971:SREa

[Feu71a] Lewis S. Feuer, B.Sc., Ph.D. The social roots of Einstein’s Theory of Relativity. Part I. Annals of Science, 27(3):277–298, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1971:SREb

[Feu71b] Lewis S. Feuer, B.Sc., Ph.D. The social roots of Einstein’s Theory of Relativity. Part II. Annals of Science, 27(4):313–344, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1975:DMC

[Feu75] Lewis S. Feuer. Is the ‘Darwin–Marx correspondence’ authentic? Annals of Science, 32(1):1–12, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feuer:1976:DMC

[Feu76] Lewis S. Feuer. The ‘Darwin–Marx correspondence’: a correction and revision. Annals of Science, 33(4):383–394, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farrar:1968:TAM

[FF68] W. V. Farrar, B.Sc., Ph.D. and Kathleen R. Farrar, B.Sc., Ph.D. Thomas Allan, mineralogist: An autobiographical fragment. Annals of Science, 24(2):115–120, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Feingold:1983:ECB

[FG83] Mordechai Feingold and Penelope M. Gouk. An early critique of Bacon’s Sylva Sylvarum: Edmund Chilmead’s treatise on sound. Annals of Science, 40(2):139–157, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Farinelli:1974:AFA

[FGR74] Dott. Maurizio Farinelli, A. L. B. Gale, B.Sc., and A. J. B. Robertson, M.A., Ph.D., D.Sc. Ambrogio Fusinieri and the adsorption theory of heterogeneous catalysis. Annals of Science, 31(1):19–20, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Figala:1972:TBE

[Fig72] Karin Figala, Dr.rer.nat. Tycho Brahe’s elixier. Annals of Science, 28(2):139–176, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Finocchiaro:2002:DG

[Fin02] Maurice A. Finocchiaro. Drake on Galileo. Annals of Science, 59(1): 83–88, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Finlay:2009:FVC

[Fin09] Mark R. Finlay. Fermentation: Vital or chemical process? Annals of Science, 66(3):419–421, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Finlay:2012:PSS

[Fin12] Mark Finlay. Power struggles: Scientific authority and the creation of practical electricity before Edison. Annals of Science, 69(1):149– 152, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Findlen:2022:PRT

[Fin22] Paula Findlen. On pestilence: a Renaissance treatise on plague. Annals of Science, 79(4):513–515, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2022.2114547.

Fisher:1936:MWB

[Fis36] R. A. Fisher, M.A., Sc.D., F.R.S. Has Mendel’s work been rediscovered? Annals of Science, 1(2):115–137, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fitzgerald:2017:BRE

[Fit17] Deborah Fitzgerald. Book review: Evolution Made to Order: Plant Breeding and Technological Innovation in Twentieth-Century America. Annals of Science, 74(4):341–342, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fyfe:2007:SMN

[FL07] Aileen Fyfe and Bernard V. Lightman, editors. Science in the marketplace: nineteenth-century sites and experiences. University of Chicago Press, Chicago, IL, USA and London, UK, 2007. ISBN 0-226-27650-3 (hardcover). x + 410 pp. LCCN

Q175.52.G7 S355 2007. URL http://www.loc.gov/catdir/
enhancements/fy0732/2007000102-d.html;   http://www.loc.
gov/catdir/enhancements/fy0827/2007000102-b.html;   http:
//www.loc.gov/catdir/toc/ecip077/2007000102.html.

Fleming:1974:NGD

[Fle74] Robin S. Fleming. Newton, gases, and Daltonian chemistry: The foundations of combination in definite proportion. Annals of Science, 31(6):561–574, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Flesher:1997:ROH

[Fle97] Mary Mosher Flesher. Repetitive order and the human walking apparatus: Prussian military science versus the Webers’ locomotion research. Annals of Science, 54(5):463–487, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Flinn:1959:TAT

[Fli59] Michael W. Flinn, M.A. Timber and the advance of technology: a reconsideration. Annals of Science, 15(2):109–120, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Flipse:2008:ASR

[Fli08] Abraham C. Flipse. Against the science–religion conflict: the genesis of a Calvinist science faculty in The Netherlands in the early Twentieth Century. Annals of Science, 65(3):363–391, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Florescu:1960:ODS

[Flo60] Radu R. Florescu, M.A., B.Litt., Ph.D. The origin and development of science in Rumania. Annals of Science, 16(1):43–58, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Flohr:1996:PFW

[Flo96] Carsten Flohr. The plague fighter: Wu Lien-teh and the beginning of the Chinese public health system. Annals of Science, 53(4):361– 380, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Freeman:2011:CAC

[FLW11] Carol Freeman, Elizabeth Leane, and Yvette Watt, editors. Considering animals: contemporary studies in human-animal relations. Ashgate Publishers, Farnham, Surrey, England, 2011. ISBN 1-4094-0013-1 (hardcover), 0-7546-9863-7 (e-book). xvi + 236 pp. LCCN QL85 .C667 2011.

Folkerts:2011:JFB

[Fol11] Menso Folkerts. Jacopo da Firenze’s Tractatus Algorismi and early Italian abbacus culture. Annals of Science, 68(2):282–284, 2011. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Folse:2014:BRB

[Fol14a] Henry Folse. Book review: Thought Experiments in Methodological and Historical Contexts, edited by K. Ierodiakonou and S. Roux. Annals of Science, 71(2):277–280, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Folse:2014:BSG

[Fol14b] Henry Folse. Higher Speculations: Grand Theories and Failed Revolutions in Physics and Cosmology, by Helge Kragh. Annals of Science, 71(1):123–127, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Foote:1952:MMS

[Foo52] George A. Foote, M.A., Ph.D. Mechanism, materialism, and science in England, 1800–1850. Annals of Science, 8(2):152–161, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Foote:1969:HSG

[Foo69] Edward T. Foote, S.J., A.M., Ph.D. Harvey: Spontaneous generation and the egg. Annals of Science, 25(2):139–163, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1961:HSR

[For61] Eric Gray Forbes, Ph.D., F.R.A.S. A history of the solar red shift problem. Annals of Science, 17(3):129–164, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1966:ODM

[For66a] Eric Gray Forbes, M.Sc., Ph.D., F.R.A.S. The origin and development of the marine chronometer. Annals of Science, 22(1):1–25, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1966:TML

[For66b] Eric Gray Forbes, M.Sc., Ph.D., F.R.A.S. Tobias Mayer’s lunar tables. Annals of Science, 22(2):105–116, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1970:TMM

[For70a] Eric G. Forbes, M.Sc., Ph.D. Tobias Mayer’s method of measuring the areas of irregular polygons. Annals of Science, 26(4):319–329, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1970:TMT

[For70b] Eric G. Forbes, M.Sc., Ph.D. Tobias Mayer’s theory of colour-mixing and its application to artistic reproductions. Annals of Science, 26 (2):95–114, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1971:BRB

[For71a] Eric G. Forbes. Book review: Physics without Einstein, by Harold Aspden. Annals of Science, 27(3):302–303, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1971:CBC

[For71b] Eric G. Forbes, B.Sc., Ph.D. The correspondence between Carl Friedrich Gauss and the Rev. Nevil Maskelyne (1802–5). Annals of Science, 27(3):213–237, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1971:TMN

[For71c] Eric G. Forbes, M.Sc., Ph.D. Tobias Mayer’s new astrolabe (1759): Its principles and construction. Annals of Science, 27(2):109–116, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1972:GCL

[For72a] Eric G. Forbes, B.Sc., Ph.D. Georg Christoph Lichtenberg and the Opera Inedita of Tobias Mayer. Annals of Science, 28(1):31–42, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forbes:1972:TMM

[For72b] Eric G. Forbes, M.Sc., Ph.D. Tobias Mayer’s method for calculating the circumstances of a solar eclipse. Annals of Science, 28(2):177– 189, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Forman:1998:MBM

[For98] Paul Forman. Molecular beam measurements of nuclear moments before magnetic resonance. Part I: I. I. Rabi and deflecting magnets to 1938. Annals of Science, 55(2):111–160, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fors:2015:LMC

[For15] Hjalmar Fors. The limits of matter: chemistry, mining, and Enlightenment. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2015. ISBN 0-226-19499-X (hardcover), 0-226-19504-X (e-book). 241 pp. LCCN BD646 .F67 2015.

Foster:1937:OHY

[Fos37] Charles Foster, M.A., B.Sc. One hundred years of science teaching in Great Britain. Annals of Science, 2(3):335–344, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fournier:1989:HDS

[Fou89] Marian Fournier. Huygens’ designs for a simple microscope. Annals of Science, 46(6):575–596, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fox:1996:TEP

[Fox96] Robert Fox. Thomas Edison’s Parisian campaign: Incandescent lighting and the hidden face of technology transfer. Annals of Science, 53(2):157–193, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fox:2001:SO

[Fox01] Robert Fox. Science at Oxford. Annals of Science, 58(3):319–322, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1936:SSH

[Fra36] K. J. Franklin, D.M., M.R.C.P. A short sketch of the history of the Oxford medical school. Annals of Science, 1(4):431–446, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1938:SHI

[Fra38] K. J. Franklin, D.M., M.R.C.P. A short history of the international congresses of physiologists. Annals of Science, 3(3):241–335, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1939:IEH

[Fra39a] K. J. Franklin, D.M., M.R.C.P. An introduction to the earlier history of phlebitis. Annals of Science, 4(1):47–60, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1939:STC

[Fra39b] K. J. Franklin, D.M., M.R.C.P. Some textual changes in successive editions of Richard Lower’s Tractatus de corde item de motu & colore sanguinis et chyli in eum transitu. Annals of Science, 4(3):283–294, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1941:SGK

[Fra41] K. J. Franklin, D.M., F.R.C.P. A survey of the growth of knowledge about certain parts of the foetal cardio-vascular apparatus, and about the foetal circulation, in Man and some other mammals. Part I: Galen to Harvey. Annals of Science, 5(1):57–89, August 15, 1941. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1945:JMH

[Fra45] K. J. Franklin, D.M., F.R.C.P. Jean Méry (1645–1722) and his ideas on the foetal blood flow. Annals of Science, 5(3):203–228, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franklin:1947:ALG

[Fra47] K. J. Franklin. An autograph letter by Gilbert White. Annals of Science, 5(4):370–372, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Franceschetti:2009:ANE

[Fra09] Donald R. Franceschetti. The artificial and the natural: An evolving polarity. Annals of Science, 66(4):582–584, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Frampton:2010:MMG

[Fra10] Sally Frampton. The making of Mr. Gray’s anatomy: Bodies, books, fortune and fame. Annals of Science, 67(1):142–144, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Frampton:2013:BRC

[Fra13] Sally Frampton. Book review: Christoph Gradmann, Laboratory Disease: Robert Koch’s Medical Bacteriology. Translated from the German by Elborg Forster. Baltimore, MD: Johns Hopkins University Press, 2009. 318 pp. $35.00 / £18.00. ISBN-13 978-0-801-89313-1. Annals of Science, 70(1):133–136, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Frampton:2014:BRB

[Fra14a] Sally Frampton. Book review: Possessing the Dead: The Artful Science of Anatomy, by Helen MacDonald. Annals of Science, 71(2): 283–286, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fraser:2014:BRB

[Fra14b] Craig Fraser. Book review: Turbulent Times in Mathematics: The Life of J. C. Fields and the History of the Fields Medal. Annals of Science, 71(4):590–592, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Freudenthal:2011:SMJ

[Fre11] Gad Freudenthal, editor. Science in medieval Jewish cultures. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-00145-5. xii + 547 pp. LCCN Q128 .S35 2011. URL http://www.loc.gov/

catdir/enhancements/fy1209/2010037107-b.html;   http://www.
loc.gov/catdir/enhancements/fy1209/2010037107-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1209/2010037107-t.html.

Frumer:2016:BWR

[Fru16] Yulia Frumer. Before words: reading Western astronomical texts in early nineteenth-century Japan. Annals of Science, 73(2):170– ??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Flood:2011:MVB

[FRW11] Raymond Flood, Adrian C. (Adrian Clifford) Rice, and Robin Wilson, editors. Mathematics in Victorian Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960139-9. ix + 466 pp. LCCN QA26 .M37

2011. URL http://www.loc.gov/catdir/enhancements/fy1211/
2011938072-b.html;   http://www.loc.gov/catdir/enhancements/
fy1211/2011938072-d.html;   http://www.loc.gov/catdir/enhancements/
fy1211/2011938072-t.html.

Feingold:2021:PCC

[FS21] Mordechai Feingold and Andrej Svorenčík. A preliminary census of copies of the first edition of Newton’s Principia (1687). Annals of Science, 77(3):253–348, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL https://www.livescience.

com/200-more-copies-newton-principia.html.

Fullmer:1961:SC

[Ful61] J. Z. Fullmer, M.S.Chem., Ph.D. On spontaneous combustion. Annals of Science, 17(2):65–80, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fullmer:1964:HDG

[Ful64] J. Z. Fullmer, M.S.Chem., Ph.D. Humphry Davy and the gunpowder manufactory. Annals of Science, 20(3):165–194, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Fussell:1958:CNL

[Fus58] G. E. Fussell. Crop nutrition in the late Stuart age (1660–1714). Annals of Science, 14(3):173–184, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Field:1985:GBP

[FW85] J. V. Field and M. T. Wright. Gears from the Byzantines: a portable sundial with calendrical gearing. Annals of Science, 42(2):87–138, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Finn:2009:CUS

[FY09] Bernard S. Finn and Daqing Yang, editors. Communications under the seas: the evolving cable network and its implications. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2009. ISBN 0-262-01286-3 (hardcover). 303 pp. LCCN TK5103.15 .C66 2009.

Gabriel:1978:IDV

[Gab78] Gottfried Gabriel. Implizite Definitionen — Eine Verwechselungsgeschichte. (German) [Implicit definitions — confusing story]. Annals of Science, 35(4):419–423, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Galvez:1988:RFA

[Gál88] Antonio Gálvez. The role of the French Academy of Sciences in the clarification of the issue of spontaneous generation in the midnineteenth century. Annals of Science, 45(4):345–365, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gambaro:2021:GHM

[Gam21] Ivana Gambaro. Geo-heliocentric models and the Society of Jesus: from Clavius’s resistance to Dechales’s Mathesis Regia. Annals of Science, 78(3):265–294, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Garber:1976:TM

[Gar76] Elizabeth Garber. Thermodynamics and meteorology (1850–1900). Annals of Science, 33(1):51–65, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Garwood:2004:GCC

[Gar04] Christine Garwood. Green crusaders or captives of industry? The British Alkali Inspectorate and the ethics of environmental decision making, 1864–95. Annals of Science, 61(1):99–117, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Garber:2010:BPS

[Gar10] Elizabeth Garber. The beginnings of piezoelectricity: a study in mundane physics. Annals of Science, 67(2):261–265, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gass:2007:SII

[Gas07] Gillian Gass. Spheres of influence: Illustration, notation, and John Dalton’s conceptual toolbox, 1803–1835. Annals of Science, 64(3): 349–382, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gascoigne:2010:SPR

[Gas10] John Gascoigne. Science, philosophy and religion in the age of the Enlightenment: British and global contexts, volume 946 of Variorum collected studies series. Ashgate/Variorum, Farnham, UK, 2010. ISBN 1-4094-0058-1 (hardcover). [various] pp. LCCN B802 .G37 2010.

Gatley:2011:CTC

[Gat11] Katrina Gatley. Cancer in the Twentieth Century. Annals of Science, 68(1):144–146, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gattei:2018:BRR

[Gat18] Stefano Gattei. Book review: Reading Galileo: Scribal Technologies and the Two New Sciences, by Renée Raphael, Baltimore, Johns Hopkins University Press, 2017, xii + 266 pp., $54.95, ISBN 978-1-4214-2177-3. Annals of Science, 75(2):151–153, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gauvin:2012:INW

[Gau12a] Jean-François Gauvin. The instrument that never was: Inventing, manufacturing, and branding Réaumur’s thermometer during the Enlightenment. Annals of Science, 69(4):515–549, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gauvin:2012:TTP

[Gau12b] Jean-François Gauvin. Toward a theory of practice (but not yet there). Annals of Science, 69(3):435–440, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gaukroger:2014:BRB

[Gau14] Stephen Gaukroger. Book review: How Modern Science Came into the World. Four Civilizations, One 17th Century Breakthrough, by H. Floris Cohen. Annals of Science, 71(1):102–104, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gay:2000:APC

[Gay00] Hannah Gay. Association and practice: The City and guilds of London Institute for the Advancement of Technical Education. Annals of Science, 57(4):369–398, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gay:2010:CED

[Gay10] Hannah Gay. Chemist, entomologist, Darwinian, and man of affairs: Raphael Meldola and the making of a scientific career. Annals of Science, 67(1):79–119, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gay:2013:QPH

[Gay13a] Hannah Gay. A questionable project: Herbert McLeod and the making of the Fourth series of the Royal Society Catalogue of Scientific Papers, 1901–25. Annals of Science, 70(2):149–174, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gay:2013:SCH

[Gay13b] Hannah Gay. The Silwood Circle: a history of ecology and the making of scientific careers in Late Twentieth-Century Britain. Imperial College Press, London, UK, 2013. ISBN 1-84816-989-2 (hardcover), 1-78326-292-3 (paperback). ix + 430 + 16 pp. LCCN QH541.264.G7 G28 2013.

Gal:2019:BKN

[GE19] Ofer Gal and Cindy Hodoba Eric. Between Kepler and Newton: Hooke’s ‘principles of congruity and incongruity ’ and the naturalization of mathematics. Annals of Science, 76(3–4):241–266, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gee:1993:EDM

[Gee93] Brian Gee. The early development of the magneto-electric machine. Annals of Science, 50(2):101–133, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Geller:2021:MAA

[Gel21] Mark J. Geller. Medicine in ancient Assur, a microhistorical study of the Neo-Assyrian healer Kis.ir-Assur. Ancient magic and divination 18. Annals of Science, 78(3):393–395, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Genin:2007:DSD

[Gen07] Christophe Genin. Le dosage spectrométrique du mercure dans l’air vers 1930. (French) [The spectrometric dosage of mercury in the air around 1930]. Annals of Science, 64(4):579–593, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gendron:2022:MJM

[Gen22] François Gendron. The M de Jussieu’s ‘mirror of the Incas’: an Ecuadorian archaeological artefact in the mineralogical collection of René-Just Haüy (1743–1822). Annals of Science, 79(2):259–273, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2028900.

George:1938:GAS

[Geo38a] Albert J. George. The genesis of the académie des sciences. Annals of Science, 3(4):372–401, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

George:1938:SCA

[Geo38b] Albert Joseph George. A seventeenth-century amateur of science: Jean Chapelain. Annals of Science, 3(2):217–236, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

George:1952:SMD

[Geo52] Philip George, M.A., Ph.D. The scientific movement and the development of chemistry in England, as seen in the papers published in the Philosophical Transactions from 1664/5 until 1750. Annals of Science, 8(4):302–322, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Germanese:2023:ISA

[Ger23] Donatella Germanese. The ingredients of a successful atomic exhibition in Cold War Italy. Annals of Science, 80(1):10–37, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2164794.

Gessner:2023:STO

[Ges23] Samuel Gessner. ‘Si te omnimoda delectat precisio’: early astronomical instruments with scales and the multiple meanings of precision in the sixteenth century. Annals of Science, 81(1–2):30–59, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2285845.

Grattan-Guinness:1974:RAS

[GG74] I. Grattan-Guinness. The Russell Archives: Some new light on Russell’s logicism. Annals of Science, 31(5):387–406, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1975:SNP

[GG75a] I. Grattan-Guinness. Some new policies for Annals of Science. Annals of Science, 32(1):i–ii, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1975:WLR

[GG75b] I. Grattan-Guinness. Wiener on the logics of Russell and Schröder. An account of his doctoral thesis, and of his discussion of it with Russell. Annals of Science, 32(2):103–132, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1976:EPC

[GG76] I. Grattan-Guinness. An extension of the period covered by Annals of Science. Annals of Science, 33(1):1, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1977:HSJ

[GG77] I. Grattan-Guinness. History of science journals: ‘to be useful, and to the living’? Annals of Science, 34(2):193–202, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gavroglu:1984:SMH

[GG84a] Kostas Gavroglu and Yorgos Goudaroulis. Some methodological and historical considerations in low temperature physics: The case of superconductivity 1911–57. Annals of Science, 41(2):135–149, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1984:WWA

[GG84b] I. Grattan-Guinness. Work for the workers: Advances in engineering mechanics and instruction in France, 1800–1830. Annals of Science, 41(1):1–33, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gavroglou:1986:SMH

[GG86] Kostas Gavroglou and Yorgos Goudaroulis. Some methodological and historical considerations in low temperature physics. II: The case of superfluidity. Annals of Science, 43(2):137–146, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gavroglu:1988:UMQ

[GG88] Kostas Gavroglu and Yorgos Goudaroulis. Understanding macroscopic quantum phenomena: The history of superfluidity 1941–1955. Annals of Science, 45(4):367–385, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Galdabini:1991:MFE

[GG91a] Silvana Galdabini and Giuseppe Giuliani. Magnetic field effects and dualistic theory of metallic conduction in Italy (1911–1926): cultural heritage, creativity, epistemological beliefs, and national scientific community. Annals of Science, 48(1):21–37, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1991:EAA

[GG91b] I. Grattan-Guinness. Eric Aiton: an appreciation. Annals of Science, 48(3):305–308, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1992:CNN

[GG92] I. Grattan-Guinness. Counting the notes: Numerology in the works of Mozart, especially Die Zauberflöte. Annals of Science, 49(3):201– 232, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1993:CIG

[GG93] I. Grattan-Guinness. Cottage industry or ghetto? the British Society for the History of Mathematics, 1971–1992. Annals of Science, 50 (5):483–490, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1994:SNF

[GG94] I. Grattan-Guinness. Some numerological features of Beethoven’s output. Annals of Science, 51(2):103–135, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1997:BPB

[GG97] I. Grattan-Guinness. Benjamin Peirce’s Linear Associative Algebra (1870): New light on its preparation and ‘publication’. In fond memory of Max H. Fisch (1900–95). Annals of Science, 54(6):597–606, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2002:TCA

[GG02] I. Grattan-Guinness. Truths and contradictions about Karl Popper. Annals of Science, 59(1):89–96, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2010:HTJ

[GG10] I. Grattan-Guinness. How to take over a journal without trying: Annals of Science 1974. Annals of Science, 67(2):239–242, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Graney:2011:TDS

[GG11a] Christopher M. Graney and Timothy P. Grayson. On the telescopic disks of stars: a review and analysis of stellar observations from the early seventeenth through the middle nineteenth centuries. Annals of Science, 68(3):351–373, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2011:JMK

[GG11b] I. Grattan-Guinness. 6000 Jahre Mathematik. Eine kulturgeschichtliche Zeitreise. (German) [6000 years of mathematics. A cultural and historical journey through time]. Annals of Science, 68(2):281–282, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2012:HHB

[GG12] Ivor Grattan-Guinness. Of the human heart. A biography of Benjamin Peirce. Annals of Science, 69(1):141–142, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2013:BRBb

[GG13a] I. Grattan-Guinness. Book review: Numerical Notation: a Comparative History. Annals of Science, 70(2):294–295, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See erratum [Ano13b].

Grattan-Guinness:2013:BRBa

[GG13b] I. Grattan-Guinness. Book review: The Development of Modern Logic. Annals of Science, 70(2):294, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:2014:BRL

[GG14] I. Grattan-Guinness. Iris Runge: A Life at the Crossroads of Mathematics, Science, and Industry, by Renate Tobies. Annals of Science, 71(2):294–295, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Granek:2008:SAF

[GH08] Galina Granek and Giora Hon. Searching for asses, finding a kingdom: The story of the invention of the scanning tunnelling microscope (STM). Annals of Science, 65(1):101–125, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Giannini:2023:EEA

[Gia23] Giulia Giannini. Establishing an experimental agenda at the Accademia del Cimento: Carlo Rinaldini’s book lists. Annals of Science, 80(2):112–142, 2023. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2023.2168060.

Gibbs:1938:NN

[Gib38] F. W. Gibbs, M.Sc. On ‘nitre’ and ‘natron’. Annals of Science, 3(2): 213–216, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1939:HMS

[Gib39] F. W. Gibbs, M.Sc. The history of the manufacture of soap. Annals of Science, 4(2):169–190, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1948:FTC

[Gib48] F. W. Gibbs, M.Sc. The furnaces and thermometers of Cornelis Drebbel. Annals of Science, 6(1):32–43, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1950:RTI

[Gib50] F. W. Gibbs, M.Sc., Ph.D. The rise of the tinplate industry. I. The tinplate workers. Annals of Science, 6(4):390–403, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:HSJ

[Gib51a] F. W. Gibbs, M.Sc., Ph.D. Historical survey of the japanning trade. 1. Eastern and Western lacquer. Annals of Science, 7(4):401–416, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:PSR

[Gib51b] F. W. Gibbs, M.Sc., Ph.D. Peter Shaw and the revival of chemistry. Annals of Science, 7(3):211–237, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:RTIa

[Gib51c] F. W. Gibbs, M.Sc., Ph.D. The rise of the tinplate industry. II. Early tinplate manufacture to 1700. Annals of Science, 7(1):25–42, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:RTIb

[Gib51d] F. W. Gibbs, M.Sc., Ph.D. The rise of the tinplate industry. III. John Hanbury (1664–1734). Annals of Science, 7(1):43–61, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:RTIc

[Gib51e] F. W. Gibbs, M.Sc., Ph.D. The rise of the tinplate industry. IV. An eighteenth century tinplate mill. Annals of Science, 7(2):113– 127, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1951:RDS

[Gib51f] F. W. Gibbs, M.Sc., Ph.D. Robert Dossie (1717–1777) and the society of arts. Annals of Science, 7(2):149–172, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1952:NWL

[Gib52a] F. W. Gibbs, M.Sc., Ph.D. A notebook of William Lewis and Alexander Chisholm. Annals of Science, 8(3):202–220, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1952:PCI

[Gib52b] F. W. Gibbs, M.Sc., Ph.D. Prelude to chemistry in industry. Annals of Science, 8(3):271–281, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1952:WLM

[Gib52c] F. W. Gibbs, M.Sc., Ph.D. William Lewis, M.B., F.R.S. (1708– 1781). Annals of Science, 8(2):122–151, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1953:HSJa

[Gib53a] F. W. Gibbs, M.Sc., Ph.D., F.R.I.C. Historical survey of the japanning trade. II. Early British japanning. Annals of Science, 9(1): 88–95, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1953:HSJb

[Gib53b] F. W. Gibbs, M.Sc., Ph.D., F.R.I.C. Historical survey of the japanning trade. III. Pontypool and Usk. Annals of Science, 9(3):197–213, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1953:HSJc

[Gib53c] F. W. Gibbs, M.Sc., Ph.D., F.R.I.C. Historical survey of the japanning trade. IV. The Midlands. Annals of Science, 9(3):214–232, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1953:RDF

[Gib53d] F. W. Gibbs, M.Sc., Ph.D., F.R.I.C. Robert Dossie (1717–77). A further bibliographical note. Annals of Science, 9(2):191–193, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1955:RTI

[Gib55] F. W. Gibbs, D.Sc., Ph.D., F.R.I.C. The rise of the tinplate industry. V. Cockshutt on tinplate manufacture. Annals of Science, 11(2):145– 153, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gibbs:1957:BB

[Gib57] F. W. Gibbs, Ph.D., D.Sc., F.R.I.C. Boerhaave and the botanists. Annals of Science, 13(1):47–61, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gielas:2016:BRM

[Gie16] Anna Gielas. Book review: Melinda Baldwin, Making “Nature”. The History of a Scientific Journal. Chicago, IL: Chicago University Press, 2015. 309 pp + plates. $45.00; £31.50, ISBN 978-0-226-26145-4. Annals of Science, 73(3):338–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gielas:2019:SJA

[Gie19] Anna Gielas. The scientific journal. Authorship and the politics of knowledge in the nineteenth century. Annals of Science, 76(2):236– 238, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gill:1976:VEP

[Gil76] Sydney Gill. A voltaic enigma and a possible solution to it. Annals of Science, 33(4):351–370, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gillispie:2002:SRB

[Gil02] Charles C. Gillispie. Some recent ’big pictures’ in the history of science. Annals of Science, 59(4):409–412, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gittins:1966:MAB

[Git66] L. Gittins, B.Sc., (Econ.), Ph.D. The manufacture of alkali in Britain, 1779–1789. Annals of Science, 22(3):175–190, 1966. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guillem-Llobat:2011:SIP

[GL11] Ximo Guillem-Llobat. The sugar industry, political authorities, and scientific institutions in the regulation of saccharin: Valencia (1888– 1939). Annals of Science, 68(3):401–424, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gluch:2023:PPQ

[Glu23a] Sibylle Gluch. Promises of precision: questioning precision in ‘precision’ instruments. Annals of Science, 81(1–2):1–9, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2282777.

Gluch:2023:TTC

[Glu23b] Sibylle Gluch. Time troubles: clocks and practices of precision in early eighteenth-century observatories. Annals of Science, 81(1–2): 160–188, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2282773.

Giuliani:1994:IPC

[GM94] Giuseppe Giuliani and Paolantonio Marazzini. The Italian physics community and the crisis of classical physics: New radiations, quanta and Relativity (1896–1925). Annals of Science, 51(4):355–390, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Goddu:2009:MCM

[God09] A. Goddu. Mechanics and cosmology in the medieval and early modern period. Annals of Science, 66(2):281–284, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Goeing:2018:SAP

[Goe18] Anja-Silvia Goeing. Science in the archives: Pasts, presents, futures. Annals of Science, 75(3):268–270, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Golinski:2019:BRL

[Gol19] Jan Golinski. Book review: The life and legend of James Watt: collaboration, natural philosophy, and the improvement of the steam engine, by David Philip Miller, Pittsburgh, PA, University of Pittsburgh Press, 2019, xix + 420 pp., 24 b&w illus., $50 (hardcover), ISBN 978-0-8229-4558-1. Annals of Science, 76(3–4):382–384, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Goodison:1967:DQP

[Goo67] Nicholas Goodison, M.A. Daniel Quare and the portable barometer. Annals of Science, 23(4):287–293, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gooding:1980:MVM

[Goo80] David Gooding. Metaphysics versus measurement: The conversion and conservation of force in Faraday’s physics. Annals of Science, 37(1):1–29, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gooday:2008:DET

[Goo08] Graeme Gooday. Domesticating Electricity: Technology, Uncertainty and Gender, 1880–1914. Pickering and Chatto, London, UK, 2008. ISBN 1-85196-975-6 (hardcover). viii + 292 pp. LCCN ????

Goodrum:2010:AAR

[Goo10] Matthew R. Goodrum. Adam’s ancestors: Race, religion, and the politics of human origins. Annals of Science, 67(2):296–298, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Goodrum:2011:GGI

[Goo11] Matthew R. Goodrum. God or gorilla: Images of evolution in the Jazz Age. Annals of Science, 68(4):575–577, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gordon:1987:SCM

[Gor87] Robert B. Gordon. Sixteenth-century metalworking technology used in the manufacture of two German astrolabes. Annals of Science, 44 (1):71–84, January 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gordin:2016:DMG

[Gor16] Michael D. Gordin. The Dostoevsky Machine in Georgetown: scientific translation in the Cold War. Annals of Science, 73(2):208–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gordin:2017:SKS

[Gor17] Michael D. Gordin. Sharing knowledge, shaping Europe: US technological collaboration and nonproliferation. Annals of Science, 74(3): 245–247, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gossick:1976:HKS

[Gos76] B. R. Gossick. Heaviside and Kelvin: a study in contrasts. Annals of Science, 33(3):275–287, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gossick:1977:WHM

[Gos77] B. R. Gossick. Where is Heaviside’s manuscript for volume 4 of his Electromagnetic Theory. Annals of Science, 34(6):601–606, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gosline:2015:BRB

[Gos15] Sheldon Lee Gosline. Book review: A Man in a Hurry: The Extraordinary Life and Times of Edward Payson Weston. Annals of Science, 72(1):133–134, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gouk:1980:RAM

[Gou80] Penelope Gouk. The role of acoustics and music theory in the scientific work of Robert Hooke. Annals of Science, 37(5):573–605, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gouk:1983:UAS

[Gou83] Penelope M. Gouk. The union of arts and sciences in the eighteenth century: Lorenz Spengler (1720–1807), artistic turner and natural scientist. Annals of Science, 40(5):411–436, September 1983. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Goulding:2010:DHR

[Gou10] Robert Goulding. Defending Hypatia: Ramus, Savile, and the Renaissance rediscovery of mathematical history, volume 25 of Archimedes, 1385-0180. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 90-481-3542-7, 90-481-3541-9. xx + 201 pp. LCCN QA24 .G68 2010. URL http:

//public.eblib.com/EBLPublic/PublicView.do?ptiID=646196;
 http://site.ebrary.com/id/10395443;   http://swb.eblib.
com/patron/FullRecord.aspx?p=646196.

Goulding:2022:HOD

[Gou22] Robert Goulding. The harvest of optics: Descartes, Mydorge, and their paths to a theory of refraction. Annals of Science, 79(2): 164–214, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2022.2026479.

Goulding:2023:OPI

[Gou23] Robert Goulding. The optical papers of Isaac Newton volume II: The Opticks and related papers ca. 1688–1717. Annals of Science, 80(3):299–302, 2023. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2023.2178673.

Govier:2021:ASR

[Gov21] Mark Adrian Govier. Allegiance and supremacy: Religion and the Royal Society’s 3rd charter of 1669. Annals of Science, 78(4):463– 483, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gower:1982:APF

[Gow82] Barry Gower. Astronomy and probability: Forbes versus Michell on the distribution of the stars. Annals of Science, 39(2):145–160, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grabo:1939:SRM

[Gra39] Carl Grabo. Science and the romantic movement. Annals of Science, 4(2):191–205, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Graham:1970:FMS

[Gra70] George Graham, M.D., F.R.C.P. The formation of the medical and surgical professorial units in the London teaching hospitals. Annals of Science, 26(1):1–22, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1971:TBG

[Gra71] I. Grattan-Guinness, M.A., M.Sc., Ph.D. Towards a biography of Georg Cantor. Annals of Science, 27(4):345–391, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1972:MUW

[Gra72a] I. Grattan-Guinness, M.A., M.Sc., Ph.D. A mathematical union: William Henry and Grace Chisholm Young. Annals of Science, 29 (2):105–185, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grattan-Guinness:1972:UMT

[Gra72b] I. Grattan-Guinness, M.A., M.Sc., Ph.D. University mathematics at the turn of the century unpublished recollections of W. H. Young. Annals of Science, 28(4):369–384, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grapi:2001:MBC

[Gra01] Pere Grapí. The marginalization of Berthollet’s chemical affinities in the French textbook tradition at the beginning of the Nineteenth Century. Annals of Science, 58(2):111–135, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Green:1957:TCB

[Gre57] J. H. S. Green, B.Sc., Ph.D., A.R.I.C. Thomas Clark (1801–1867). A biographical study. Annals of Science, 13(3):164–179, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gregory:1977:AFH

[Gre77] Anita Gregory. Anatomy of a fraud: Harry Price and the medium Rudi Schneider. Annals of Science, 34(5):449–549, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Greenberg:1981:AFF

[Gre81] John L. Greenberg. Alexis Fontaine’s Fluxio-differential method and the origins of the calculus of several variables. Annals of Science, 38(3):251–290, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Greenberg:1982:AFI

[Gre82] John L. Greenberg. Alexis Fontaine’s integration of ordinary differential equations and the origins of the calculus of several variables. Annals of Science, 39(1):1–36, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Greenberg:1984:DLE

[Gre84] John L. Greenberg. Degrees of longitude and the Earth’s shape: The diffusion of a scientific idea in Paris in the 1730s. Annals of Science, 41(2):151–158, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Greenberg:1987:ERM

[Gre87] J. L. Greenberg. Essay review: The measurement of the Earth: Antonio Lafuente and Antonio J. Delgado, La Geometrizacion de la Tierra: Observaciones y Resultados de la Expedicion Geodesica Hispano–Francesca al Virreinato del Peru (1735–1744). Annals of Science, 44(3):289–295, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Greenspan:2005:ECW

[Gre05] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Greener:2022:OTC

[Gre22] James Greener. Offering themselves by chance: Newcomen’s starting materials. Annals of Science, 79(3):320–363, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2091157.

Gross:1972:APS

[Gro72a] Walter E. Gross, Ph.D. The American Philosophical Society and the growth of meteorology in the United States: 1835–1850. Annals of Science, 29(4):321–338, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gross:1972:TIC

[Gro72b] Walter E. Gross, Ph.D. Theory and invention: The case of Charles Tablot Porter and his steam engine governor. Annals of Science, 29(3):257–269, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gross:1974:APS

[Gro74a] Walter E. Gross. The American Philosophical Society and the rise of astronomy in the United States in the middle of the nineteenth century. Annals of Science, 31(5):407–427, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gross:1974:GAP

[Gro74b] Walter E. Gross. Geophysics in the American philosophical society 1835–1850. Annals of Science, 31(5):429–447, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gross:1974:RMO

[Gro74c] Walter E. Gross. Relativity of motion: From Occam to Galileo. Annals of Science, 31(6):529–545, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gross:2014:BRB

[Gro14] Ari Gross. Book review: Feynman, by J. Ottaviani and L. Myrick. Annals of Science, 71(2):292–294, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grund:2011:MWN

[Gru11] Peter Grund. Misticall wordes and names infinite: an edition and study of Humfrey Lock’s Treatise on Alchemy, volume 367 of Medieval and Renaissance texts and studies. ACMRS (Arizona Center for Medieval and Renaissance Studies), Tempe, Ariz., 2011. ISBN 0-86698-415-1. xii + 350 pp. LCCN QD25 .L63 2011.

Galilei:2010:S

[GS10] Galileo Galilei and Christoph Scheiner. On sunspots. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-70715-6 (hardcover), 0-226-70716-4 (paperback). xi + 418 pp. LCCN QB525 .G22 2010. Translated by Eileen Adair Reeves and Albert Van Helden.

Guerrini:2009:SCI

[Gue09] Anita Guerrini. Sustainability or collapse? An integrated history and future of people on Earth. Annals of Science, 66(2):292–294, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guerrini:2010:KAK

[Gue10] Anita Guerrini. The king’s animals and the king’s books: the illustrations for the Paris Academy’s Histoire des animaux. Annals of Science, 67(3):383–404, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guerrini:2013:BRB

[Gue13] Anita Guerrini. Book review: Knowing Global Environments. New Historical Perspectives on the Field Sciences, edited by Jeremy Vetter. Annals of Science, 70(4):550–551, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guerrini:2021:TCW

[Gue21] Anita Guerrini. The transmutations of chymistry. Wilhelm Homberg and the Académie Royale des Sciences. Annals of Science, 78(3):398– 400, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guerrini:2023:BRM

[Gue23] Anita Guerrini. Book review: Making physicians. Tradition, teaching, and trials at Leiden University, 1575–1639, vol. 1, by Evan Ragland, Leiden and Boston, Brill, 2022, Clio Medica, vol. 106, xiv + 457 pp., 13 plts., $143.00 (hardback), ISBN: 978-90-04-46511-4. Annals of Science, 80(4):418–420, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2209568.

Guicciardini:1995:JBJ

[Gui95] Niccol‘o Guicciardini. Johann Bernoulli, John Keill and the inverse problem of central forces. Annals of Science, 52(6):537–575, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guicciardini:2011:JWV

[Gui11a] Niccolò Guicciardini. John Wallis. Una vita per un progetto. (Italian) [John Wallis. A life for a project]. Annals of Science, 68(4):561– 563, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guinness:2011:MES

[Gui11b] I. Grattan Guinness. Mathematics emerging. A sourcebook 1540– 1900. Annals of Science, 68(1):133–134, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Guicciardini:2015:ENG

[Gui15] Niccolò Guicciardini. Editing Newton in Geneva and Rome: The annotated edition of the Principia by Calandrini, Le Seur and Jacquier. Annals of Science, 72(3):337–380, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Gansten:2011:STY

[GW11] Martin Gansten and Ola Wikander. Sahl and the Tajika Yogas: Indian transformations of Arabic astrology. Annals of Science, 68(4): 531–546, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Grigelis:2011:SSE

[GWN^{+}11] Algimantas Grigelis, Zbigniew Wójcik, Wojciech Narebski, Leonora Zivile Gelumbauskaite, and Jan Kozák. Stanisław Staszic: An early surveyor of the geology of Central and Eastern Europe. Annals of Science, 68(2):199–228, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill-Andrews:2015:BRB

[HA15] Oliver Hill-Andrews. Book review: Uncommon Contexts: Encounters Between Science and Literature. Annals of Science, 72(3):413–415, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill-Andrews:2018:BRB

[HA18] Oliver Hill-Andrews. Book review: Bloomsbury Scientists: Science and Art in the Wake of Darwin, by Michael Charles Boulter, London: UCL Press, 188 pp., ISBN 978-1-78735-005-2. Annals of Science, 75 (2):159–160, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill-Andrews:2019:EVS

[HA19] Oliver Hill-Andrews. Eugenics: a very short introduction. Annals of Science, 76(1):107–109, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Haaparanta:2009:DML

[Haa09] Leila Haaparanta, editor. The Development of Modern Logic. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-513731-0 (hardcover). x + 994 pp. LCCN BC15 .D48 2009.

Haas:2019:MMS

[Haa19] Robert Haas. Maligned for mathematics: Sir Thomas Urquhart and his Trissotetras. Annals of Science, 76(2):113–156, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hackmann:1979:RBC

[Hac79a] W. D. Hackmann. The relationship between concept and instrument design in eighteenth-century experimental science. Annals of Science, 36(3):205–224, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hackmann:1979:UAR

[Hac79b] W. D. Hackmann. Underwater acoustics and the Royal Navy, 1893– 1930. Annals of Science, 36(3):255–278, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hahm:2015:BRB

[Hah15] David E. Hahm. Book review: What Did the Romans Know? An Inquiry into Science and Worldmaking. Annals of Science, 72(1): 134–137, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hakfoort:1982:NBH

[Hak82] C. Hakfoort. Nicolas Béguelin and his search for a crucial experiment on the nature of light (1772). Annals of Science, 39(3):297–310, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hakfoort:1992:SDW

[Hak92] C. Hakfoort. Science deified: Wilhelm Osstwald’s energeticist worldview and the history of scientism. Annals of Science, 49(6):525– 544, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hall:1955:FOE

[Hal55] A. R. Hall, M.A., Ph.D. Further optical experiments of Isaac Newton. Annals of Science, 11(1):27–43, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hall:1957:NCC

[Hal57] A. Rupert Hall, M.A., Ph.D. Newton on the calculation of central forces. Annals of Science, 13(1):62–71, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hall:1978:SP

[Hal78] Bert S. Hall. The scholastic pendulum. Annals of Science, 35(5):441– 462, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hall:2013:RBC

[Hal13] Lesley A. Hall. Reconstructing the body: Classicism, modernism, and the First World War. Annals of Science, 70(3):425–428, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hale:2014:PDM

[Hal14a] Piers J. Hale. Political descent: Malthus, mutualism, and the politics of evolution in Victorian England. University of Chicago Press, Chicago, IL, USA and London, UK, 2014. ISBN 0-226-10849-X (hardcover), 0-226-10852-X (e-book). 442 pp. LCCN JC336 .H254 2014.

Hall:2014:BRB

[Hal14b] Lesley A. Hall. Book review: Managing the Body: Beauty, Health, and Fitness in Britain, 1880–1939, by Ina Zweiniger-Bargielowska.

Annals of Science, 71(1):113–116, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hale:2018:BRG

[Hal18] Piers J. Hale. Book review: Global Spencerism: The Communication and Appropriation of a British Evolutionist, edited by B. Lightman, Leiden: Brill, 2016. xiv + 409 pp., EUR 115, $149, ISBN 978-90-04-26399-4. Annals of Science, 75(2):156–159, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Halm:2020:EPP

[Hal20] Christopher Halm. ‘Enquiries on Plaister of Paris’: a material history of early agrochemical knowledge in the United States of America, 1785–1812. Annals of Science, 77(2):169–188, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hamlin:1982:JGJ

[Ham82] Christopher Hamlin. James Geikie, James Croll, and the eventful ice age. Annals of Science, 39(6):565–583, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hamm:2001:HWH

[Ham01] E. P. Hamm. Of ’histories written by the hand of nature itself’. Annals of Science, 58(3):311–317, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hamerla:2003:EWM

[Ham03] R. Richard Hamerla. Edward Williams Morley and the atomic weight of oxygen: the death of Prout’s hypothesis revisited. Annals of Science, 60(4):351–372, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hamilton:2012:HOT

[Ham12] David Hamilton. A history of organ transplantation: ancient legends to modern practice. University of Pittsburgh Press, Pittsburgh, PA, USA, 2012. ISBN 0-8229-4413-8 (hardcover). xx + 556 pp. LCCN RD120.6 .H36 2012. With a foreword by Clyde F. Barker and Thomas E. Starzl.

Hansen:2021:PLE

[Han21] Bert Hansen. Pasteur’s lifelong engagement with the fine arts: uncovering a scientist’s passion and personality. Annals of Science, 78 (3):334–386, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hartog:1941:NVP

[Har41] Sir Philip J. Hartog, K.B.E., C.I.E., Ll.D. The newer views of Priestley and Lavoisier. Annals of Science, 5(1):1–56, August 15, 1941. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harrison:1957:BHM

[Har57] John Anthony Harrison. Blind Henry Moyes, ‘an excellent lecturer in philosophy’. Annals of Science, 13(2):109–125, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harrison:1958:LB

[Har58] Thomas P. Harrison, B.A., Ph.D. Longolius on birds. Annals of Science, 14(4):257–268, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hardin:1966:SWR

[Har66] Clyde L. Hardin, A.B., A.M., Ph.D. The scientific work of the reverend John Michell. Annals of Science, 22(1):27–47, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harley:1971:OCC

[Har71] R. D. Harley, B.A., Ph.D. Oil colour containers: Development work by artists and colourmen in the nineteenth century. Annals of Science, 27(1):1–12, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harder:1972:CCE

[Har72] Allen Harder, M.A., Ph.D. The Copernican character of Einstein’s cosmology. Annals of Science, 29(4):339–347, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harder:1974:MSR

[Har74] Allen J. Harder. E. A. Milne, scientific revolutions and the growth of knowledge. Annals of Science, 31(4):351–363, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harry:1984:HMW

[Har84] Owen G. Harry. The Hon. Mrs Ward and ‘A windfall for the microscope’, of 1856 and 1864. Annals of Science, 41(5):471–482, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harwood:1985:GES

[Har85] Jonathan Harwood. Geneticists and the evolutionary synthesis in interwar Germany. Annals of Science, 42(3):279–301, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harman:1996:CPA

[Har96] P. M. Harman. The collected papers of Albert Einstein, vol 5, the Swiss years: Correspondence, 1902–1914 — M. J. Klein, A. J. Kox, R. Schulmann. Annals of Science, 53(1):93–94, January 1996. CODEN ANNSA8. ISSN 0003-3790.

Harper:2006:MSP

[Har06] Kristine C. Harper. Meteorology’s struggle for professional recognition in the USA (1900–1950). Annals of Science, 63(2):179–199, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harman:2009:CNB

[Har09a] P. M. (Peter Michael) Harman. The culture of nature in Britain, 1680–1860. Yale University Press, New Haven, CT, USA, 2009. ISBN 0-300-15197-7 (hardcover). xi + 393 pp. LCCN BH301.N3 H36 2009.

Harvey:2009:CLH

[Har09b] Olivia Harvey. Culturing life: How cells became technologies. Annals of Science, 66(4):585–588, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harris:2012:MHE

[Har12] Nick Harris. A man in a hurry: the extraordinary life and times of Edward Payson Weston, the world’s greatest walker. deCoubertin, London, UK, 2012. ISBN 0-9564313-7-2. 276 pp. LCCN GV1072.W47 H37 2012.

Hawes:1968:NEA

[Haw68] Joan L. Hawes, M.Sc. Newton and the ‘electrical attraction unexcited’. Annals of Science, 24(2):121–130, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hawes:1971:NTE

[Haw71] Joan L. Hawes, M.Sc., Ph.D. Newton’s two electricities. Annals of Science, 27(1):95–103, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hays:1964:SBS

[Hay64] J. N. Hays, B.A. Science and Brougham’s society. Annals of Science, 20(3):227–241, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hays:1981:RFD

[Hay81] J. N. Hays. The rise and fall of Dionysius Lardner. Annals of Science, 38(5):527–542, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hayward:2018:BRR

[Hay18] Rhodri Hayward. Book review: Ruling Minds: Psychology in the British Empire, by Erik Linstrum, Cambridge, MA and London, Harvard University Press, 2016. viii + 309 pp., $41.00, £32.95, EUR 37.00, ISBN 978-0-674-08866-5. Annals of Science, 75(2):161–163, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harris:1974:GGS

[HB74] J. Harris and W. H. Brock. From Giessen to Gower street: Towards a biography of Alexander William Williamson (1824–1904). Annals of Science, 31(2):95–130, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Huggins:2014:SCW

[HB14] Sir William Huggins and Barbara J. Becker, editors. Selected correspondence of William Huggins. Pickering and Chatto, London, UK, 2014. ISBN 1-84893-415-7 (hardback: set). xlix + 641 (vol. 1), vii + 691 (vol. 2) pp. LCCN QB460.72.H84 A4 2014. Two volumes.

Hird:2008:DPJ

[HBP08] Jonathan Hird, Jonathan Betts, and Derek Pratt. The diamond pallets of John Harrison’s fourth longitude timekeeper — H4. Annals of Science, 65(2):171–200, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hedesan:2022:PTW

[Hed22] Georgiana D. Hedesan. The poison trials: wonder drugs, experiment, and the battle for authority in renaissance science. Annals of Science, 79(3):408–410, 2022. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2022.2063388.

Heimann:1967:MCS

[Hei67a] P. M. Heimann, B.A.B.Sc. Moseley and celtium: The search for a missing element. Annals of Science, 23(4):249–260, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heimann:1967:RNN

[Hei67b] P. M. Heimann, B.A.B.Sc. Rutherford, Nagaoka, and the nuclear atom. Annals of Science, 23(4):299–303, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heilbron:1976:ML

[Hei76] J. L. Heilbron. Moseley letters. Annals of Science, 33(3):309, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heilbron:2014:BRB

[Hei14] J. L. Heilbron. Book review: Between Raphael and Galileo. Mutio Oddi and the Mathematical Culture of Late Renaissance Italy, by A. Marr. Annals of Science, 71(1):140–142, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Heilbron:2019:BRP

[Hei19] J. L. Heilbron. Book review: Photons. The history and mental models of light quanta, Klaus Hentschel, Cham, Switzerland, Springer International, 2018, xiii + 231 pp. EUR 69,99 (hardback), ISBN: 978-3-319-95251-2. Annals of Science, 76(3–4):387–389, 2019. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henderson:1978:PGB

[Hen78] J. Frank Henderson. The position of the glycosidic bond in purine nucleosides: The conservative influence of a convention of chemical nomenclature. Annals of Science, 35(3):299–323, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henry:1979:FPC

[Hen79] John Henry. Francesco Patrizi da Cherso’s concept of space and its later influence. Annals of Science, 36(6):549–573, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hendry:1980:DAW

[Hen80] John Hendry. The development of attitudes to the wave-particle duality of light and quantum theory, 1900–1920. Annals of Science, 37(1):59–79, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hendry:1982:MHP

[Hen82] John Hendry. Mayer, Herschel and Prévost on the solar motion. Annals of Science, 39(1):61–75, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:1986:KES

[Hen86] Klaus Hentschel. Die Korrespondenz Einstein–Schlick: Zum Verhältnis der Physik zur Philosophie. (German) [The Einstein– Schlick correspondence: On the relationship of physics to philosophy]. Annals of Science, 43(5):475–488, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hendry:1987:SOC

[Hen87] John Hendry. The scientific origins of controlled fusion technology. Annals of Science, 44(2):143–168, March 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:1988:KDM

[Hen88] Klaus Hentschel. Die Korrespondenz Duhem–Mach: Zur ‘Modellbeladenheit’ von Wissenschaftsgeschichte. (German) [The Duhem– Mach correspondence: The ‘model laden’ history of science]. Annals of Science, 45(1):73–91, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hensel:1989:EAB

[Hen89] Susann Hensel. Zu einigen Aspekten der Berufung von Mathematikern an die Technischen Hochschulen Deutschlands im letzten Drittel des 19. Jahrhunderts. (German) [On some aspects of the vocation of mathematicians at the technical universities in Germany in the last third of the 19th century]. Annals of Science, 46(4):387–416, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:1992:GBP

[Hen92] Klaus Hentschel. Grebe/Bachems photometrische Analyse der Linienprofile und die Gravitations-Rotverschiebung: 1919 bis 1922. (German) [Grebe/Bachem’s photometric analysis of the line profiles and the gravitational redshift: 1919 to 1922]. Annals of Science, 49 (1):21–46, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:1996:MGR

[Hen96] Klaus Hentschel. Measurements of gravitational redshift between 1959 and 1971. Annals of Science, 53(3):269–295, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2000:ERH

[Hen00] K. Hentschel. Essay review: Heisenberg, German culture, and other such horrifying things. Paul Lawrence Rose, Heisenberg and the Nazi Atomic Bomb Project, 1939–1945, ISBN 0-520-21077-8. Annals of Science, 57(3):301–306, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2002:SP

[Hen02] Klaus Hentschel. Spectroscopic portraiture. Annals of Science, 59(1): 57–82, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2007:GMH

[Hen07] Klaus Hentschel. Gauss, Meyerstein and Hanoverian metrology. Annals of Science, 64(1):41–75, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2009:AIS

[Hen09] Klaus Hentschel. Aufrecht im Sturm der Zeit. Der Physiker James Franck 1882–1964. (German) [Upright in stormy times. The physicist James Franck 1882–1964]. Annals of Science, 66(4):578–580, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2014:BRB

[Hen14] Klaus Hentschel. Book review: A History of Optics from Greek Antiquity to the Nineteenth Century. Annals of Science, 71(4):586–588, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henry:2017:BRC

[Hen17] John Henry. Book review: Copernicus: A Very Short Introduction. Annals of Science, 74(2):169–171, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henderson:2018:RMG

[Hen18] Gabriel Henderson. Resurrecting Maunder’s ghost: John ‘Jack’ Eddy, the Maunder Minimum, and the rise of a dilettante astrophysicist. Annals of Science, 75(3):234–254, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henry:2021:NAS

[Hen21a] John Henry. Newton the alchemist: science, enigma, and the quest for nature’s ‘secret fire’. Annals of Science, 77(4):549–552, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2021:WLP

[Hen21b] Klaus Hentschel. Weighing light and pondering historiographies. Annals of Science, 77(3):383–387, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Henry:2023:NAA

[Hen23] John Henry. Newton’s ‘De Aere et Aethere’ and the introduction of interparticulate forces into his physics. Annals of Science, 80(3): 232–267, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2192721.

Herring:2017:BRP

[Her17] Emily Herring. Book review: The Physicist and the Philosopher: Einstein, Bergson and the Debate that Changed our Understanding of Time. Annals of Science, 74(1):86–87, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Herring:2018:GDB

[Her18] Emily Herring. ‘Great is Darwin and Bergson his poet’: Julian Huxley’s other evolutionary synthesis. Annals of Science, 75(1):40–54, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hesse:1952:BPL

[Hes52] Mary B. Hesse, M.Sc., Ph.D. Boole’s philosophy of logic. Annals of Science, 8(1):61–81, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hestmark:2017:TNE

[Hes17] Geir Hestmark. Tracings of the north of Europe: Robert Chambers in search of the Ice Age. Annals of Science, 74(4):262–281, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hesketh:2021:MJT

[Hes21] Ian Hesketh. The making of John Tyndall’s Darwinian revolution. Annals of Science, 77(4):524–548, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hetherington:1972:FMS

[Het72] Norriss S. Hetherington, Ph.D. The first measurements of stellar parallax. Annals of Science, 28(4):319–325, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hetherington:1976:CAV

[Het76] Norriss S. Hetherington. Cleveland Abbe and a view of science in mid-nineteenth-century America. Annals of Science, 33(1):31–49, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hetherington:1983:MNC

[Het83] Norriss S. Hetherington. Mid-nineteenth-century American astronomy: Science in a developing nation. Annals of Science, 40(1):61–80, January 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Higgins:1953:CS

[Hig53] Kathleen Higgins, M.A.B.Sc. The classification of sundials. Annals of Science, 9(4):342–358, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Higgitt:2017:BRC

[Hig17a] Rebekah Higgitt. Book review: A Companion to the History of Science, by, Bernard Lightman, editor, Chichester, West Sussex: Wiley Blackwell, 2016. 624 pp. $150.00, £120.00, ISBN 978-1-118-62077-9. Annals of Science, 74(2):165–166, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Higgitt:2017:FTE

[Hig17b] Rebekah Higgitt. Framing the transit: expeditionary culture and identities in Lieutenant E. J. W. Noble’s caricatures of the 1874 transit of Venus expedition to Honolulu. Annals of Science, 74(3):214– 239, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill:1984:PAG

[Hil84] David K. Hill. The projection argument in Galileo and Copernicus: Rhetorical strategy in the defence of the new system. Annals of Science, 41(2):109–133, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill:1985:ABM

[Hil85] Donald R. Hill. Al-Biruni’s mechanical calendar. Annals of Science, 42(2):139–163, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill:1986:CBM

[Hil86] C. R. Hill. The cabinet of Bonnier de la Mosson (1702–1744). Annals of Science, 43(2):147–174, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hill:1991:SCB

[Hil91] R. D. Hill. Some comments on Forming new physics communities: Australia and Japan, 1914–1950, by R. W. Home and M. Watanabe. Annals of Science, 48(6):583–587, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hilbert:1999:HIN

[Hil99] Martin Hilbert. Herschel’s investigation of the nature of radiant heat: The limitations of experiment. Annals of Science, 56(4):357–378, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hirai:2007:ICC

[Hir07] Hiro Hirai. Interprétation chymique de la création et origine corpusculaire de la vie chez Athanasius Kircher. (French) [chemical interpretation of the corpuscular creation and origin of life according to Athanasius Kircher]. Annals of Science, 64(2):217–234, 2007. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hirai:2008:CGC

[Hir08] Hiro Hirai, editor. Cornelius Gemma: cosmology, medicine, and natural philosophy in renaissance Louvain, volume 24 of Bruniana and campanelliana. Supplementi, 1127-6045; Studi; 10. Serra, Pisa, Italy, 2008. ISBN 88-6227-118-2. ISSN 1127-6045. 153 pp. LCCN B785.G384 C67 2008.

Hjermitslev:2010:DCD

[Hje10] Hans Henrik Hjermitslev. Danes commemorating Darwin: apes and evolution at the 1909 anniversary. Annals of Science, 67(4):485– 525, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hayakawa:1991:SPP

[HL91] Satio Hayakawa and Morris F. Low. Science policy and politics in post-war Japan: the establishment of the KEK high energy physics laboratory. Annals of Science, 48(3):207–229, 1991. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Harrison:2011:WNO

[HNS11] Peter Harrison, Ronald L. Numbers, and Michael H. Shank, editors. Wrestling with nature: from omens to science. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-31781-1 (hardcover), 0-226-31783-8 (paperback). x + 416 pp. LCCN Q125 .W86 2011.

Hoare:1955:EDR

[Hoa55] Cecil A. Hoare, D.Sc., F.R.S. Erasmus Darwin in Russia. Annals of Science, 11(3):255–256, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoch:1983:RCE

[Hoc83] Paul K. Hoch. The reception of central European refugee physicists of the 1930s: U.S.S.R., U.K., U.S.A. Annals of Science, 40(3):217–246, May 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoff:1936:GPG

[Hof36] Hebbel E. Hoff, Ph.D. Galvani and the pre-Galvanian electrophysiologists. Annals of Science, 1(2):157–172, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hollinger:1976:CPS

[Hol76] David A. Hollinger. Comments on papers by sharlin and Wall. Annals of Science, 33(5):476–480, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Holmberg:1996:MAV

[Hol96] Gustav Holmberg. Mechanizing the astronomer’s vision: On the role of photography in Swedish astronomy, c. 1880–1914. Annals of Science, 53(6):609–616, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Holmberg:2007:TTT

[Hol07] Gustav Holmberg. Taming tempests through telegraphy and media appearances: Science communication and the construction of a Swedish storm-warning system before the Great War. Annals of Science, 64(1):77–91, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Holmes:2017:BRC

[Hol17a] John Holmes. Book review: Charles Darwin and the Church of Wordsworth. Annals of Science, 74(2):174–176, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Holmes:2017:CTC

[Hol17b] Matthew Holmes. Changing techniques in crop plant classification: molecularization at the National Institute of Agricultural Botany during the 1980s. Annals of Science, 74(2):149–164, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Home:1979:NBN

[Hom79] R. W. Home. Nollet and Boerhaave: a note on eighteenth-century ideas about electricity and fire. Annals of Science, 36(2):171–175, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Home:1988:LEA

[Hom88] R. W. Home. Leonhard Euler’s ‘anti-Newtonian’ theory of light. Annals of Science, 45(5):521–533, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Honigmann:1984:ESH

[Hon84] Peter Honigmann. Entstehung und Schicksal von Humboldts Magnetischen ‘Verein’ (1829–1834) im Zusammenhang mit seiner Rußlandreise. (German) [Origin and fate of Humboldt’s magnetic ‘association’ (1829–1834) in connection with his visit to Russia]. Annals of Science, 41(1):57–86, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hon:1987:KAP

[Hon87] Giora Hon. On Kepler’s awareness of the problem of experimental error. Annals of Science, 44(6):545–591, November 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hong:1995:EAO

[Hon95] Sungook Hong. Efficiency and authority in the ‘open versus closed’ transformer controversy. Annals of Science, 52(1):49–76, 1995. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hong:1996:SCE

[Hon96] Sungook Hong. Styles and credit in early radio engineering: Fleming and Marconi on the first transatlantic wireless telegraphy. Annals of Science, 53(5):431–465, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1957:CMA

[Hop57a] I. B. Hopley, M.Sc., Ph.D. Clerk Maxwell’s apparatus for the measurement of surface tension. Annals of Science, 13(3):180–187, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1957:MWE

[Hop57b] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance I. The determination of the absolute unit of resistance. Annals of Science, 13(4):265–272, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1958:MWE

[Hop58] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance. II. Proposals for the re-determination of the B.A. unit of 1863. Annals of Science, 14(3):197–210, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1959:MDN

[Hop59a] I. B. Hopley, M.Sc., Ph.D. Maxwell’s determination of the number of electrostatic units in one electromagnetic unit of electricity. Annals of Science, 15(2):91–108, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hopley:1959:MWE

[Hop59b] I. B. Hopley, M.Sc., Ph.D. Maxwell’s work on electrical resistance III. Improvement on Mance’s method for the measurement of battery resistance. Annals of Science, 15(1):51–55, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoppen:1981:MEM

[Hop81] Alison Hoppen. Military engineers in Malta, 1530–1798. Annals of Science, 38(4):413–433, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hornix:1987:ERT

[Hor87] W. J. Hornix. Essay review: The tales of Hofmann: William Hodson Brock, editor, Justus von Liebig und August Wilhelm Hofmann in ihren Briefen (1841–1873). Annals of Science, 44(5):519–524, September 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hormogon:1991:HST

[Hor91] Mariano Hormogón. History of science and technology today in Spain. Annals of Science, 48(2):187–190, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Horstmann:2000:HSR

[Hor00] Frank Horstmann. Hobbes und das Sinusgesetz der Refraktion. (German) [Hobbes and the sine law of refraction]. Annals of Science, 57 (4):415–440, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Howarth:2002:FGF

[How02] Richard J. Howarth. Fitting geomagnetic fields before the invention of least squares: I. Henry Bond’s predictions (1636, 1668) of the change in magnetic declination in London. Annals of Science, 59(4): 391–408, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Howarth:2003:FGF

[How03] Richard J. Howarth. Fitting geomagnetic fields before the invention of least squares: II. William Whiston’s isoclinic maps of Southern England (1719 and 1721). Annals of Science, 60(1):63–84, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Howe:2019:BRA

[How19] Joshua Howe. Book review: The ascent of John Tyndall: Victorian scientist, mountaineer, and public intellectual, by Roland Jackson, Oxford, Oxford University Press, 2018. xx + 556 pp., 26 plts, £25.00, ISBN 0-19-878895-9. Annals of Science, 76(3–4):385–387, 2019. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoyrup:2023:NHG

[Høy23] Jens Høyrup. A new history of Greek mathematics. Annals of Science, 80(3):293–296, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00033790.2023.2195864.

Hunter:2003:LPR

[HP03] Michael Hunter and Lawrence M. Principe. The lost papers of Robert Boyle. Annals of Science, 60(3):269–311, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Habib:1989:ISR

[HR89] S. Irfan Habib and Dhruv Raina. The introduction of scientific rationality into India: a study of Master Ramchandra — Urdu journalist, mathematician and educationalist. Annals of Science, 46(6):597–610, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hepler-Smith:2022:EVH

[HS22] Evan Hepler-Smith. The elements: a visual history of their discovery. Annals of Science, 79(3):406–408, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2022.2087903.

Hsia:2009:SSL

[Hsi09] Florence C. Hsia. Sojourners in a Strange Land: Jesuits and Their Scientific Missions in Late Imperial China. Science and Its Conceptual Foundations S. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-35561-6. 291 pp. LCCN

BV2290 .H75 2009. URL http://public.eblib.com/EBLPublic/
PublicView.do?ptiID=680714;   http://www.myilibrary.com?
id=307838;   http://www.myilibrary.com?id=307838\%26ref=
toc.

Harris:2023:QPB

[HT23] Arthur Harris and Liba Taub. Quantification and precision: a brief look at some ancient accounts. Annals of Science, 81(1–2):10–29, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2282778.

Huang:2022:APE

[Hua22] Yijie Huang. Anatomizing the pulse: Edmund King’s analogy, observation and conception of the tubular body. Annals of Science, 79(3):292–319, 2022. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2022.2093971.

Hudson:2003:ERB

[Hud03] Giles Hudson. Essay review of The Ambassadors’ Secret: Holbein and the World of the Renaissance by John North. Annals of Science, 60(2):201–205, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:1951:EJT

[Hug51a] Edward Hughes, M.A. The early journal of Thomas Wright of Durham. Annals of Science, 7(1):1–24, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:1951:SEE

[Hug51b] Arthur Hughes, M.A., Ph.D. Science in English encyclopædias, 1704– 1875. I. Annals of Science, 7(4):340–370, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:1952:SEE

[Hug52] Arthur Hughes, M.A., Ph.D. Science in English encyclopædias, 1704– 1875. II. Theories of the elementary composition of matter. Annals of Science, 8(4):323–367, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:1953:SEE

[Hug53] Arthur Hughes, M.A., Ph.D.Sc.D. Science in English encyclopædias, 1704–1875. III. Meteorology. Annals of Science, 9(3):233–264, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:1955:SEE

[Hug55] Arthur Hughes, Sc.D. Science in English encyclopædias, 1704–1875. IV. Theories of the Earth. Annals of Science, 11(1):74–92, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hughes:2011:GBB

[Hug11] Sally Smith Hughes. Genentech: the beginnings of biotech. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-35918-2 (hardcover). xv + 213 pp. LCCN HD9999.B444 H85 2011.

Hui:2011:IMI

[Hui11] A. E. Hui. Instruments of music, instruments of science: Hermann von Helmholtz’s musical practices, his classicism, and his Beethoven sonata. Annals of Science, 68(2):149–177, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hull:1985:BCS

[Hul85] David L. Hull. Bias and commitment in science: Phenetics and cladistics. Annals of Science, 42(3):319–338, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hull:2002:FTR

[Hul02] Andrew J. Hull. Food for thought?: The relations between the Royal Society Food Committees and Government, 1915–19. Annals of Science, 59(3):263–298, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Humphreys:1937:DCM

[Hum37] A. W. Humphreys, B.A.B.Sc. The development of the conception and measurement of electric current. Annals of Science, 2(2):164–178, April 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hunter:1999:UCA

[Hun99] Patti Wilger Hunter. An unofficial community: American mathematical statisticians before 1935. Annals of Science, 56(1):47–68, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hunter:2014:BRB

[Hun14] Patti Wilger Hunter. Book review: Mathematics in Victorian Britain. Annals of Science, 71(3):445–448, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hunter:2023:MSR

[Hun23] Michael Hunter. Magic, science and religion in Early Modern Europe. Annals of Science, 80(1):77–78, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2022.2127898.

Hutchison:1973:UER

[Hut73] Keith Hutchison, M.Sc., Dipl.Hist.Phil.Sc. Der ursprung der entropiefunktion bei Rankine und Clausius. (German) [The origin of the entropy by Rankine and Clausius]. Annals of Science, 30(3):341– 364, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hutton:1977:SRC

[Hut77] Sarah Hutton. Some Renaissance critiques of Aristotle’s theory of time. Annals of Science, 34(4):345–363, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Home:1987:PAJ

[HW87] R. W. Home and Masao Watanabe. Physics in Australia and Japan to 1914: a comparison. Annals of Science, 44(3):215–235, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Home:1990:FNP

[HW90] R. W. Home and Masao Watanabe. Forming new physics communities: Australia and Japan, 1914–1950. Annals of Science, 47(4): 317–345, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hoch:1987:SOH

[HY87] P. K. Hoch and E. J. Yoxen. Schrödinger at Oxford: a hypothetical national cultural synthesis which failed. Annals of Science, 44 (6):593–616, November 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hyde:2019:CCC

[Hyd19] Elizabeth Hyde. Cultivating commerce: cultures of botany in Britain and France, 1760–1815. Annals of Science, 76(1):98–100, 2019. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hon:2007:GLS

[HZ07] Giora Hon and Yaakov Zik. Geometry of light and shadow: Francesco Maurolyco (1494–1575) and the pinhole camera. Annals of Science, 64(4):549–578, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Iliffe:2012:E

[IL12] Rob Iliffe and Trevor Levere. Editorial. Annals of Science, 69(4): 451, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Illy:1989:EEV

[Ill89] József Illy. Einstein und der Eötvös-Versuch: Ein Brief Albert Einsteins an Willy Wien. (German) [Einstein and the Eötvös experiment: a letter between Albert Einstein and Willy Wien]. Annals of Science, 46(4):417–422, ???? 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00033798900200311.

Indorato:1989:PRC

[IM89] Luigi Indorato and Guido Masotto. Poincaré’s role in the Crémieu– Pender controversy over electric convection. Annals of Science, 46 (2):117–163, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Iliffe:2015:WAS

[IM15] Rob Iliffe and David Miller. Winner of the Annals of Science Prize for 2014. Annals of Science, 72(3):275, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Iliffe:2017:WAS

[IM17] Rob Iliffe and David Philip Miller. Winner of the Annals of Science Prize for 2016 . Annals of Science, 74(3):253, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1972:CSG

[Ink72a] Ian Inkster, B.A. Charles Sylvester and the great railroad debate. Annals of Science, 28(2):113–120, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1972:NIS

[Ink72b] Ian Inkster, B.A. A note on itinerant science lecturers, 1790–1850. Annals of Science, 28(3):235–236, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1975:SMI

[Ink75] Ian Inkster. Science and the Mechanics’ Institutes, 1820–1850: The case of Sheffield. Annals of Science, 32(5):451–474, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1977:SSM

[Ink77] Ian Inkster. Science and society in the metropolis: a preliminary examination of the social and institutional context of the Askesian Society of London, 1796–1807. Annals of Science, 34(1):1–32, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1978:RGA

[Ink78] Ian Inkster. Robert Goodacre’s astronomy lectures (1823–1825), and the structure of scientific culture in Philadelphia. Annals of Science, 35(4):353–363, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1988:PBT

[Ink88] Ian Inkster. Prometheus bound: Technology and industrialization in Japan, China and India prior to 1914 — a political economy approach. Annals of Science, 45(4):399–426, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1989:ATA

[Ink89] Ian Inkster. Appropriate technology, alternative technology and the Chinese model: Terminology and analysis. Annals of Science, 46(3): 263–276, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Inkster:1991:STE

[Ink91] Ian Inkster. Science, technology and economic development — Japanese historical experience in context. Annals of Science, 48(6): 545–563, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ierodiakonou:2011:TEM

[IR11] Katerina Ierodiakonou and Sophie Roux, editors. Thought experiments in methodological and historical contexts, volume 19 (or 15??) of History of science and medicine library, 1872-0684; Medieval and early modern science. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20176-9 (hardcover). LCCN BD265 .T46 2011.

Ishibashi:2014:PAT

[Ish14] Yuto Ishibashi. In pursuit of accurate timekeeping: Liverpool and Victorian electrical horology. Annals of Science, 71(4):474–496, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jacquot:1952:SCCa

[Jac52a] Jean Jacquot. Sir Charles Cavendish and his learned friends: a contribution to the history of scientific relations between England and the Continent in the earlier part of the 17th century. Annals of Science, 8(1):13–27, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jacquot:1952:SCCb

[Jac52b] Jean Jacquot. Sir Charles Cavendish and his learned friends: a contribution to the history of scientific relations between England and the Continent in the earlier part of the 17th century. Annals of Science, 8(2):175–191, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jackson:1999:GMT

[Jac99] Patrick N. Wyse Jackson. Geological museums and their collections: Rich sources for historians of geology. Annals of Science, 56(4):417– 431, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jacyna:2009:HSI

[Jac09] L. S. Jacyna. Herbert Spencer and the invention of modern life. Annals of Science, 66(3):431–433, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jacyna:2012:MMH

[Jac12] Stephen Jacyna. Marking the mind: a history of memory. Annals of Science, 69(1):138–139, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jacyna:2014:BRB

[Jac14] L. S. Jacyna. Book review: The Sciences of the Soul: the Early Modern Origins of Psychology. Annals of Science, 71(4):585–586, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jackson:2015:CIC

[Jac15a] Catherine M. Jackson. Chemical identity crisis: Glass and glassblowing in the identification of organic compounds: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):187–205, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jackson:2015:JTE

[Jac15b] Roland Jackson. John Tyndall and the early history of diamagnetism. Annals of Science, 72(4):435–489, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jaecks:2010:IEC

[Jae10] Duane H. Jaecks. An investigation of the eighteenth-century achromatic telescope. Annals of Science, 67(2):149–186, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jalobeanu:2023:FBP

[Jal23] Dana Jalobeanu. Francis Bacon and the practices of measurement. Annals of Science, 81(1–2):79–99, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2282772.

James:1978:SSS

[Jam78] John James. Significance in sacred sites: The churches around Positano. Annals of Science, 35(2):103–130, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

James:2012:BRN

[Jam12] Frank A. J. L. James. Book review: The nineteenth century: Aileen Fyfe and Bernard Lightman, editors, Science in the Marketplace: Nineteenth-Century Sites and Experiences/Victorian Popularizers of Science: Designing Nature for New Audiences, ISBN-13 978-0-226-27650-2 and Bernard Lightman, Victorian Popularizers of Science: Designing Nature for New Audiences, ISBN-13 978-0-226-48119-7. Annals of Science, 69(4):584–586, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jarrell:1998:VBH

[Jar98] Richard A. Jarrell. Visionary or bureaucrat? T. H. Huxley, the science and art department and science teaching for the working class. Annals of Science, 55(3):219–240, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jardine:2018:BRH

[Jar18a] Boris Jardine. Book review: How Scientific Instruments Have Changed Hands, by A. D. Morrison-Low, Sara Schechner, and Paolo Brenni, Leiden: Brill, 2017, xxxii + 239 pp., ISBN 978-90-04-32492-3. Annals of Science, 75(1):60–61, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jardine:2018:ISH

[Jar18b] Boris Jardine. Instruments of statecraft: Humphrey Cole, Elizabethan economic policy and the rise of practical mathematics. Annals of Science, 75(4):304–329, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jardine:2023:SLP

[Jar23] Boris Jardine. The social life of precision instruments: artisans’ trials in early-modern England, 1550–1700. Annals of Science, 81(1–2): 100–123, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2282784.

Jayawardene:1978:WSM

[Jay78] S. A. Jayawardene. Western scientific manuscripts before 1600: a checklist of published catalogues: To the memory of Carl B. Boyer (1906–1976). Annals of Science, 35(2):143–172, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jenkins:2016:PER

[Jen16] Bill Jenkins. The platypus in Edinburgh: Robert Jameson, Robert Knox and the place of the Ornithorhynchus in nature, 1821–24. Annals of Science, 73(4):425–441, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jayawardene:1979:BNH

[JL79] S. A. Jayawardene and Jennifer Lawes. Biographical notices of historians of science: a checklist. Annals of Science, 36(4):315–394, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Janssen:2014:CCE

[JL14] Michel Janssen and Christopher Lehner, editors. The Cambridge Companion to Einstein. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). xvi + 562 pp. LCCN QC16.E5 C36 2014.

Johnston:1991:MPI

[Joh91] Stephen Johnston. Mathematical practitioners and instruments in Elizabethan England. Annals of Science, 48(4):319–344, 1991. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Johnston:2009:HSB

[Joh09] Sean Johnston. History of science: a beginner’s guide. Oneworld beginners guides. Oneworld, Oxford, UK, 2009. ISBN 1-85168-681-9 (paperback). viii + 218 pp. LCCN Q126 .J64 2009.

Johnson:2010:MTD

[Joh10] Ryan Johnson. The making of a tropical disease: a short history of malaria. Annals of Science, 67(1):139–142, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Johnson:2014:BRB

[Joh14] Jeffrey Allan Johnson. Book review: Image and Reality: Kekulé, Kopp, and the Scientific Imagination, by Alan J. Rocke. Annals of Science, 71(1):121–123, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Johnson:2015:BRB

[Joh15a] Jeffrey Allan Johnson. Book review: Progressive Enlightenment: The Origins of the Gaslight Industry, 1780–1820. Annals of Science, 72 (2):270–272, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Johnson:2015:DCL

[Joh15b] Jeffrey Allan Johnson. Dilemmas of 19th-century liberalism among German academic chemists: Shaping a national science policy from Hofmann to Fischer, 1865–1919: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):224–241, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Johnston:2018:VIA

[Joh18] Sean F. Johnston. Vaunting the independent amateur: Scientific American and the representation of lay scientists. Annals of Science, 75(2):97–119, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1937:VNH

[Jon37] W. P. Jones, Ph.D. The vogue of natural history in England, 1750– 1770. Annals of Science, 2(3):345–352, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1950:SRB

[Jon50] H. Whitmore Jones, M.A., Ph.D. Some reflections on the beginnings of experimental science. Annals of Science, 6(3):283–292, 1950. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1974:SCG

[Jon74] H. W. Jones. A seventeenth-century geometrical debate. Annals of Science, 31(4):307–333, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1982:JHF

[Jon82] Jean Jones. James Hutton and the Forth and Clyde canal. Annals of Science, 39(3):255–263, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1983:JHE

[Jon83] Jean Jones. James Hutton: Exploration and oceanography. Annals of Science, 40(1):81–94, January 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1984:GCJ

[Jon84] Jean Jones. The geological collection of James Hutton. Annals of Science, 41(3):223–244, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1985:JHA

[Jon85] Jean Jones. James Hutton’s agricultural research and his life as a farmer. Annals of Science, 42(6):573–601, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1986:MSC

[Jon86] Greta Jones. The mushroom-shaped cloud: British scientists’ opposition to nuclear weapons policy, 1945–57. Annals of Science, 43(1): 1–26, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1995:WEB

[Jon95] Greta Jones. Women and eugenics in Britain: The case of Mary Scharlieb, Elizabeth Sloan Chesser, and Stella Browne. Annals of Science, 52(5):481–502, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1999:ERP

[Jon99] Jean Jones. Essay review: Priestley enlightened. Annals of Science, 56(1):95–98, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:2006:PAP

[Jon06] Alexander Jones. Ptolemy’s ancient planetary observations. Annals of Science, 63(3):255–290, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:2013:BRB

[Jon13] Paul R. Jones. Book review: The Early Britannica: The Growth of an Outstanding Encyclopedia, Voltaire Foundation, University of Oxford, 2009. xiii + 307 pp., £65.00 / EUR 76 / $104. ISBN-13 978-0-7294-0981-0. Annals of Science, 70(1):119–122, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:2019:PMI

[Jon19] Claire L. Jones. The patent medicines industry in Georgian England. Annals of Science, 76(2):230–233, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jordan:1982:ASI

[Jor82] D. W. Jordan. The adoption of self-induction by telephony, 1886– 1889. Annals of Science, 39(5):433–461, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jorink:2012:ASE

[JR12] Eric Jorink and B. A. M. (Bart A. M.) Ramakers, editors. Art and science in the early modern Netherlands. WBooks, Zwolle, The Netherlands, 2012. ISBN 90-400-7808-4 (hardcover). 367 pp. LCCN N72.S3 A63 2011.

Jones:1994:CBJ

[JTR94] Jean Jones, Hugh S. Torrens, and Eric Robinson. The correspondence between James Hutton (1726–1797) and James Watt (1736–1819) with two letters from Hutton to George Clerk-Maxwell (1715–1784): Part I. Annals of Science, 51(6):637–653, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jones:1995:CBJ

[JTR95] Jean Jones, Hugh S. Torrens, and Eric Robinson. The correspondence between James Hutton (1726–1797) and James Watt (1736–1819) with two letters from Hutton to George Clerk-Maxwell (1715–1784). Part II. Annals of Science, 52(4):357–382, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Jurkowitz:2010:HEE

[Jur10] Edward Jurkowitz. Helmholtz’s early empiricism and the Erhaltung der Kraft [Conservation of Energy]. Annals of Science, 67(1):39– 78, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Juste:2011:NTK

[Jus11] D. Juste. Non-transferable knowledge: Arabic and Hebrew onomancy into Latin. Annals of Science, 68(4):517–529, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Juste:2012:NTK

[Jus12] D. Juste. Non-transferable knowledge: Arabic and Hebrew onomancy into Latin. Annals of Science, 69(2):301, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kaalund:2017:FDJ

[Kaa17] Nanna Katrine Lüders Kaalund. A frosty disagreement: John Tyndall, James David Forbes, and the early formation of the X-Club. Annals of Science, 74(4):282–298, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kahn:2001:EAA

[Kah01] Didier Kahn. Entre atomisme, alchimie et théologie: La réception des thèses d’Antoine de Villon et Étienne de Clave contre Aristote, Paracelse et les cabalistes’ (24–25 août 1624). (French) [Between atomism, alchemy, and theology. The reception of the theses of Antoine de Villon and Étienne de Clave against Aristotle, Paracelsus, and the cabalists (24–25 August 1624)]. Annals of Science, 58(3): 241–286, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kahl:2010:IBB

[Kah10] Oliver Kahl. Ibn Baklarish’s Book of Simples: Medical remedies between three faiths in Twelfth-Century Spain. Annals of Science, 67(2):291–294, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kaiser:2010:DSS

[Kai10] David Kaiser. Disrupting science: Social movements, American scientists, and the politics of the military, 1945–1975. Annals of Science, 67(1):135–137, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kangro:1970:UBG

[Kan70] Hans Kangro, Dr.rer.nat. Ultrarotstrahlung bis zur Grenze elektrisch erzeugter Wellen, Das Lebenswerk von Heinrich Rubens. (German) [Infrared radiation at the border, electrically generated waves, the life’s work of Heinrich Rubens]. Annals of Science, 26(3):235–259, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kangro:1971:UBG

[Kan71] Hans Kangro. Ultrarotstrahlung bis zur Grenze elektrisch erzeugter Wellen, das Lebenswerk von Heinrich Rubens. (German) [Infrared radiation at the border, electrically generated waves, the life’s work of Heinrich Rubens]. Annals of Science, 27(2):165–200, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Karnfelt:2014:KLM

[Kär14] Johan Kärnfelt. Knut Lundmark, meteors and an early Swedish crowdsourcing experiment. Annals of Science, 71(4):449–473, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Katzir:2006:TDV

[Kat06] Shaul Katzir. Thermodynamic deduction versus quantum revolution: The failure of Richardson’s theory of the photoelectric effect. Annals of Science, 63(4):447–469, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Katzir:2016:PFS

[Kat16] Shaul Katzir. Pursuing frequency standards and control: the invention of quartz clock technologies. Annals of Science, 73(1):1–39, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1008044.

Katzir:2023:UCL

[Kat23] Shaul Katzir. The use of the conservation of living force before Helmholtz. Annals of Science, 80(4):337–356, 2023. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2205429.

Kauffman:1975:CWB

[Kau75] George B. Kauffman. Christian Wilhelm Blomstrand (1826–1897) Swedish chemist and mineralogist. Annals of Science, 32(1):13–37, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kauffman:1979:HCC

[Kau79] George B. Kauffman. History in the chemistry curriculum: pros and cons. Annals of Science, 36(4):395–402, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kedar:2016:NIN

[Ked16] Yael Kedar. The nomological image of nature: explaining the tide in the thirteenth century. Annals of Science, 73(1):68–88, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1034781.

Keevil:1952:SCS

[Kee52] J. J. Keevil, D.S.O., M.D., F.S.A. Sir Charles Scarburgh. Annals of Science, 8(2):113–121, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Keith:1984:SEA

[Kei84] S. T. Keith. Scientists as entrepreneurs: Arthur Tyndall and the rise of Bristol physics. Annals of Science, 41(4):335–357, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kelly:2011:WMB

[Kel11] Catherine Kelly. War and the militarization of British Army medicine, 1793–1830, volume 5 of Studies for the Society for the Social History of Medicine. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-183-2, 1-84893-184-0 (e-book). xi + 220 pp. LCCN RC971 .K395 2011.

Kemsley:1968:RIR

[Kem68] Douglas S. Kemsley, M.A. (Oxon.), Ph.D., M.Sc. Religious influences in the rise of modern science: a review and criticism, particularly of the ‘Protestant–Puritan ethic’ theory. Annals of Science, 24(3):199– 226, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kemeny:2017:WMW

[Kem17] Max Kemeny. What motion is: William Neile and the laws of motion. Annals of Science, 74(3):179–191, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kennaway:1952:SRA

[Ken52] Professor Sir Ernest Kennaway, F.R.S. Some recollections of Albrecht Kossel, Professor of Physiology in Heidelberg, 1901–1924. Annals of Science, 8(4):393–397, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Keyser:1990:BSC

[Key90] Barbara Whitney Keyser. Between science and craft: The case of Berthollet and dyeing. Annals of Science, 47(3):213–260, 1990. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kent:1960:SNC

[KH60] Andrew Kent, M.A., Ph.D. and Owen Hannaway, B.Sc. Some new considerations on Beguin and Libavius. Annals of Science, 16(4):241– 250, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kidwell:1981:PSE

[Kid81] Peggy Aldrich Kidwell. Prelude to solar energy: Pouillet, Herschel, Forbes and the solar constant. Annals of Science, 38(4):457–476, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kidd:2013:ERS

[Kid13] Ian James Kidd. Essay review: Science and the making of modernity. D. M. Knight, The Making of Modern Science: Science, Technology, Medicine and Modernity: 1789–1914. Cambridge: Polity, 2009. xii + 370 pp. 30 plts. $25.75 / £24.99. ISBN-13 978-0-7456-3676-4. S. F. Johnson, History of Science: A Beginner’s Guide: Oxford:

Oneworld, 2009. viii + 218 pp. $14.95 / £9.99. ISBN-13 978-1-85168-681-0. Annals of Science, 70(1):101–104, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kidwell:2014:BRB

[Kid14] Peggy Aldrich Kidwell. Book review: The History of the History of Mathematics: Case Studies for the Seventeenth, Eighteenth and Nineteenth Centuries. Annals of Science, 71(4):595–597, 2014. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kim:1995:ETP

[Kim95] Dong-Won Kim. The emergence of theoretical physics in Japan: Japanese physics community between the two World Wars. Annals of Science, 52(4):383–402, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kim:1999:TCH

[Kim99] Yung Sik Kim. Towards a ’comparative history of the foundations of science’: Language and logic in traditional China. Annals of Science, 56(4):451–460, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kim:2006:PSH

[Kim06] Mi Gyung Kim. ‘public’ science: Hydrogen balloons and Lavoisier’s decomposition of water. Annals of Science, 63(3):291–318, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kimler:2012:CSC

[Kim12] William Kimler. Case studies, controversy and the ‘Fieldworker’s Regress’. Annals of Science, 69(1):127–132, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

King:2004:BRB

[Kin04] David A. King. Book reviews: Towards a History from Antiquity to the Renaissance of Sundials and Other Instruments for Reckoning Time by the Sun and Stars, by Hester Higton, Sundials — An Illustrated History of Portable Dials, London: Philip Wilson, 2001, with contributions from Silke Ackermann, Richard Dunn, Kiyoshi T. Akada and Anthony Turner, Sundials at Greenwich — A Catalogue of the Sundials, Horary Quadrants and Nocturnals in the National Maritime Museum, Greenwich Oxford: Oxford University Press, and Greenwich: National Maritime Museum, 2002. Annals of Science, 61 (3):375–388, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kinder:2019:RMP

[Kin19a] John M. Kinder. Rethinking modern prostheses in Anglo-American commodity cultures, 1820–1939. Annals of Science, 76(1):105–107, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kingsland:2019:BRI

[Kin19b] Sharon Kingsland. Book review: Inside science: Stories from the field in human and animal science, by R. E. Kohler, Chicago and London, University of Chicago Press, 2019, 245 pp., $35.00 (hardback), ISBN 978-0-226-61798-5. Annals of Science, 76(3–4):367–369, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kipnis:1987:LGD

[Kip87] Naum Kipnis. Luigi Galvani and the debate on animal electricity, 1791–1800. Annals of Science, 44(2):107–142, March 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kitteringham:1982:SPS

[Kit82] Guy Kitteringham. Science in provincial society: The case of Liverpool in the early nineteenth century. Annals of Science, 39(4):329– 348, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kargon:2002:KUS

[KK02] Robert H. Kargon and Scott G. Knowles. Knowledge for use: Science, higher learning, and America’s new industrial heartland, 1880–1915. Annals of Science, 59(1):1–20, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2007:CDL

[KL07] Helge S. Kragh and Dominique Lambert. The context of discovery: Lemaître and the origin of the primeval-atom Universe. Annals of Science, 64(4):445–470, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kafker:2009:EBG

[KL09] Frank A. Kafker and Jeff Loveland, editors. The early Britannica (1768–1803): the growth of an outstanding encyclopedia, volume 2009:10 of SVEC. Voltaire Foundation, Oxford, UK, 2009. ISBN 0-7294-0981-3. xiii + 349 pp. LCCN AE1 .E26 2009.

Klein:2012:ASE

[Kle12a] Ursula Klein. Artisanal-scientific experts in eighteenth-century France and Germany. Annals of Science, 69(3):303–306, 2012. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Klein:2012:PMO

[Kle12b] Ursula Klein. The Prussian mining official Alexander von Humboldt. Annals of Science, 69(1):27–68, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Klein:2012:SOP

[Kle12c] Ursula Klein. Savant officials in the Prussian mining administration. Annals of Science, 69(3):349–374, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Klein:2021:SIG

[Kle21] Ursula Klein. Science, industry, and the German Bildungsbürgertum. Annals of Science, 77(3):366–376, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kline:2015:CMW

[Kli15] Ronald R. Kline. The cybernetics moment: or why we call our age the information age. New studies in American intellectual and cultural history. The Johns Hopkins University Press, Baltimore, MD, USA, 2015. ISBN 1-4214-1671-9 (hardcover), 1-4214-1672-7 (electronic). xi + 336 pp. LCCN Q360 .K56 2015. URL http://www.h-net.org/

reviews/showrev.php?id=46341.

Klosterman:1985:RSCa

[Klo85a] Leo J. Klosterman. A research school of chemistry in the nineteenth century: Jean Baptiste Dumas and his research students. Part I. Annals of Science, 42(1):1–40, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Klosterman:1985:RSCb

[Klo85b] Leo J. Klosterman. A research school of chemistry in the nineteenth century: Jean Baptiste Dumas and his research students. Part II. Annals of Science, 42(1):41–80, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kellert:2006:SP

[KLW06] Stephen H. Kellert, Helen E. Longino, and C. Kenneth Waters, editors. Scientific pluralism, volume 19 of Minnesota studies in the philosophy of science. University of Minnesota Press, Minneapolis, MN, 2006. ISBN 0-8166-4763-1 (hardcover). xxix + 248 pp.

LCCN Q175 .S42367 2006. URL http://www.loc.gov/catdir/
enhancements/fy0666/2006018614-d.html;   http://www.loc.
gov/catdir/toc/ecip0615/2006018614.html.

Kragh:2013:SGP

[KN13] Helge Kragh and Kristian Hvidtfelt Nielsen. Spreading the gospel: a popular book on the Bohr atom in its historical context. Annals of Science, 70(2):257–283, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knight:1975:THS

[Kni75] D. M. Knight. Teaching the history of science in some British universities and polytechnics. Annals of Science, 32(2):169–173, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knight:1976:ACB

[Kni76] David Knight. Agriculture and chemistry in Britain around 1800. Annals of Science, 33(2):187–196, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knight:1977:NSN

[Kni77] D. M. Knight. A note on sumptuous natural histories. Annals of Science, 34(3):311–314, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knight:2009:MMS

[Kni09] David M. Knight. The making of modern science: science, technology, medicine and modernity: 1789–1914. Polity, Cambridge, UK, 2009. ISBN 0-7456-3675-6 (hardcover), 0-7456-3676-4 (paperback). xii + 370 pp. LCCN Q125 .K566 2009.

Knight:2013:BRB

[Kni13] David M. Knight. Book review: Geography and Science in Britain, 1831–1939: A Study of the British Association for the Advancement of Science, by Charles W. Withers. Annals of Science, 70(4):548– 550, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knight:2014:VSS

[Kni14] David M. Knight. Voyaging in Strange Seas: the Great Revolution in Science. Yale University Press, New Haven, CT, USA, 2014. ISBN 0-300-17379-2. 10 + 329 + 32 pp. LCCN Q125 .K578 2014.

Knight:2017:BRH

[Kni17] David Knight. Book review: The History of Chemistry: A Very Short Introduction. Annals of Science, 74(1):83–84, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

KnowlesMiddleton:1964:CM

[Kno64] W. E. Knowles Middleton, D.Sc., F.R.S.C. Chemistry and meteorology, 1700–1825. Annals of Science, 20(2):125–141, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Knobloch:1989:LEM

[Kno89] Eberhard Knobloch. Leonhard Eulers Mathematische Notizbücher. (German) [Leonhard Euler’s mathematical notebooks]. Annals of Science, 46(3):277–302, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kassler:1983:RHT

[KO83] J. C. Kassler and D. R. Oldroyd. Robert Hooke’s Trinity College ‘musick scripts’, his music theory and the role of music in his cosmology. Annals of Science, 40(6):559–595, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Koberer:2017:BRN

[Köb17] Wolfgang Köberer. Book review: Navigation. A Very Short Introduction. Annals of Science, 74(4):331–332, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kojevnikov:2002:LCP

[Koj02] Alexei Kojevnikov. The last century of physics. Annals of Science, 59(4):419–422, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kollerstrom:1999:PHC

[Kol99] Nicholas Kollerstrom. The path of Halley’s Comet, and Newton’s late apprehension of the Law of Gravity. Annals of Science, 56(4):331– 356, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Konecny:2012:HEB

[Kon12] Peter Konecný. The hybrid expert in the ‘Bergstaat’: Anton von Ruprecht as a Professor of Chemistry and Mining and as a mining official, 1779–1814. Annals of Science, 69(3):335–347, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kooijmans:2010:DDA

[Koo10] L. Kooijmans. Death Defied: the Anatomy Lessons of Frederik Ruysch. Brill, Leiden, The Netherlands, 2010. ISBN 1-282-95290-0. LCCN QM16.R89 K6613 2010eb.

Kottler:1978:CDB

[Kot78] Malcolm J. Kottler. Charles Darwin’s biological species concept and theory of geographic speciation: the transmutation notebooks. Annals of Science, 35(3):275–297, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kottler:1979:HVR

[Kot79] Malcolm J. Kottler. Hugo de Vries and the rediscovery of Mendel’s laws. Annals of Science, 36(5):517–538, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kox:1990:HLC

[Kox90] A. J. Kox. H. A. Lorentz’s contributions to kinetic gas theory. Annals of Science, 47(6):591–606, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:1982:JTC

[Kra82] Helge Kragh. Julius Thomsen and 19th-century speculations on the complexity of atoms. Annals of Science, 39(1):37–60, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:1995:CIE

[Kra95] Helge Kragh. From curiosity to industry: The early history of cryolite soda manufacture. Annals of Science, 52(3):285–301, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2000:CUH

[Kra00] Helge Kragh. The chemistry of the Universe: Historical roots of modern cosmochemistry. Annals of Science, 57(4):353–368, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2009:CHC

[Kra09a] Helge Kragh. Contemporary history of cosmology and the controversy over the multiverse. Annals of Science, 66(4):529–551, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2009:SER

[Kra09b] Helge Kragh. The solar element: a reconsideration of Helium’s early history. Annals of Science, 66(2):157–182, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2011:FPF

[Kra11a] Helge Kragh. Fermilab: Physics, the frontier & megascience. Annals of Science, 68(3):437–440, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2011:HSG

[Kra11b] Helge Kragh. Higher speculations: grand theories and failed revolutions in physics and cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-959988-2. 411

pp. LCCN QB981 .K735 2011. URL http://www.loc.gov/catdir/
enhancements/fy1111/2011377091-b.html;   http://www.loc.
gov/catdir/enhancements/fy1111/2011377091-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1111/2011377091-t.html.

Kragh:2013:BRM

[Kra13] Helge Kragh. Book review: Michael Eckert, Heinrich Hertz. Hamburg: Ellert & Richter Verlag, 2010. 168 pp. 45 figures. EUR 14.00. ISBN-13 978-3-8319-0371-9. Annals of Science, 70(1):141–143, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2017:LSS

[Kra17] Helge Kragh. ‘Let the stars shine in peace!’ Niels Bohr and stellar energy, 1929–1934. Annals of Science, 74(2):126–148, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kragh:2021:CQM

[Kra21] Helge Kragh. Constructing quantum mechanics, volume 1: The Scaffold 1900–1923. Annals of Science, 77(4):552–554, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kretzmann:1938:GTU

[Kre38] Edwin M. J. Kretzmann, M.A., Ph.D. German technological Utopias of the pre-war period. Annals of Science, 3(4):417–430, 1938. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Krementsov:2011:BPM

[Kre11] Nikolai Krementsov. From ‘beastly philosophy’ to medical genetics: Eugenics in Russia and the Soviet Union. Annals of Science, 68(1): 61–92, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kremer:2013:ECS

[Kre13] Richard L. Kremer. European collections of scientific instruments, 1550–1750. Annals of Science, 70(3):423–425, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kremer:2023:SPL

[Kre23] Richard L. Kremer. Searching for precision: Lorenz Eichstadt’s Tabulae harmonicae coelestium motuum (Stetin 1644) and astronomical prediction after Kepler. Annals of Science, 81(1–2):60–78, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2284340.

Krige:2019:BRG

[Kri19] John Krige. Book review: Greening the alliance: the diplomacy of NATO’s science and environmental initiatives, by S. Turchetti, Chicago and London, University of Chicago Press, 2019, 256 pp., 19 plts, $37.50 (paperback), ISBN 978-0-226-59579-5. Annals of Science, 76(3–4):391–393, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Krige:2022:CWS

[Kri22] John Krige. Cold War social science: transnational entanglements. Annals of Science, 79(3):416–417, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2022.2087904.

Klein:2010:MEE

[KS10] Ursula Klein and E. C. (Emma C.) Spary. Materials and Expertise in Early Modern Europe: Between Market and Laboratory. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-43968-2. 408 pp. LCCN ????

Ku:2006:BAT

[Ku06] Ja Hyon Ku. British acoustics and its transformation from the 1860s to the 1910s. Annals of Science, 63(4):395–423, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ku:2009:UFI

[Ku09] Ja Hyon Ku. Uses and forms of instruments: Resonator and tuning fork in Rayleigh’s acoustical experiments. Annals of Science, 66(3): 371–395, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ku:2013:AMM

[Ku13] Ja Hyon Ku. Alfred M. Mayer and acoustics in Nineteenth-Century America. Annals of Science, 70(2):229–256, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kunitzsch:2000:NAB

[Kun00] Paul Kunitzsch. A note on Ascelinus’ Table of Astrolabe Stars. Annals of Science, 57(2):181–185, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kurzer:1997:SPC

[Kur97] Frederick Kurzer. Samuel Parkes: Chemist, author, reformer — a biography. Annals of Science, 54(5):431–462, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kurzer:1999:LWE

[Kur99] Frederick Kurzer. The life and work of Edward Charles Howard FRS. Annals of Science, 56(2):113–141, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kurzer:2000:HSI

[Kur00] Frederick Kurzer. A history of the Surrey Institution. Annals of Science, 57(2):109–141, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kurzer:2001:CCL

[Kur01] Frederick Kurzer. Chemistry and chemists at the London Institution 1807–1912. Annals of Science, 58(2):163–201, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kurzer:2003:WHP

[Kur03] Frederick Kurzer. William Hasledine Pepys FRS: a life in scientific research, learned societies and technical enterprise. Annals of Science, 60(2):137–183, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kusukawa:2010:SGP

[Kus10] S. Kusukawa. The sources of Gessner’s pictures for the Historia animalium. Annals of Science, 67(3):303–328, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kusukawa:2018:BRE

[Kus18] Sachiko Kusukawa. Book review: Eye for Detail: Images of Plants and Animals in Art and Science 1500–1630, by Florike Egmond, London: Reaktion Books, 288 pp., ISBN 978-1-78023-640-7. Annals of Science, 75(2):149–150, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Kvicalova:2022:POH

[Kvi22] Anna Kvicalova. Purkyne’s Opistophone: the hearing ‘deaf’, auditory attention and organic subjectivity in Prague psychophysical experiments, ca 1850s. Annals of Science, 79(1):60–80, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

LaBerge:2011:MEH

[La 11] Ann F. La Berge. The microscope and the eye: a history of reflections, 1740–1870. Annals of Science, 68(1):139–141, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Allan:2022:SCC

[LA22] C. S. Leedham and V. L. Allan. Scientific computing in the Cavendish Laboratory and the pioneering women computors. Annals of Science, 79(4):497–512, 2022. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2022.2106382.

Lachapelle:2011:ISS

[Lac11] Sofie Lachapelle. Investigating the supernatural: from spiritism and occultism to psychical research and metapsychics in France, 1853– 1931. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0013-8 (hardcover). 198 pp. LCCN BF1434.F8 L33 2011.

Laita:1977:IBS

[Lai77] Luis M. Laita. The influence of Boole’s search for a universal method in analysis on the creation of his logic. Annals of Science, 34(2):163– 176, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Laita:1979:IBL

[Lai79] Luis M. Laita. Influences on Boole’s logic: The controversy between William Hamilton and Augustus De Morgan. Annals of Science, 36 (1):45–65, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Laidler:2004:ERB

[Lai04] Keith J. Laidler. Essay review: From Classical to Modern Chemistry: The Instrumental Revolution, Edited by Peter J. T. Morris, editor, Royal Society of Chemistry in association with the Science Museum. London: xvi + 347 pp. ISBN 0-85404-479-5. Annals of Science, 61 (2):219–225, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lalli:2012:RME

[Lal12] Roberto Lalli. The reception of Miller’s ether-drift experiments in the USA: The history of a controversy in Relativity revolution. Annals of Science, 69(2):153–214, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lamprey:1997:ETG

[Lam97] John P. Lamprey. An examination of two groups of Georg Hartmann sixteenth-century astrolabes and the tables used in their manufacture. Annals of Science, 54(2):111–142, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Langins:1983:HPB

[Lan83] Janis Langins. Hydrogen production for ballooning during the French Revolution: An early example of chemical process development. Annals of Science, 40(6):531–558, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Langer:2013:BRS

[Lan13] Bernard Langer. Book review: Simon Werrett, Fireworks: Pyrotechnic Arts and Sciences in European History. Chicago and London: University of Chicago Press, 2010. Maps / viii + 359 pp. $45.00 (cloth), ISBN-13 978-0-226-89377-8. Annals of Science, 70(1):127– 129, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Larwood:1961:SEI

[Lar61] H. J. C. Larwood, M.Sc., M.Ed. Science and education in India before the Mutiny. Annals of Science, 17(2):81–96, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Larson:1991:SAS

[Lar91] Edward J. Larson. Science in the American South through the eyes of four natural historians, 1750–1850. Annals of Science, 48(3):231– 240, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lattis:2014:BRB

[Lat14] James Lattis. Book review: On Sunspots, by Galileo Galilei and Christoph Scheiner. Annals of Science, 71(2):288–289, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lattis:2015:BRH

[Lat15] James Lattis. Book review: Horst Bredekamp, general editor, Galileo’s O. Volume 1: Irene Brückle and Oliver Hahn, editors, Galileo’s Sidereus Nuncius: A Comparison of the Proof Copy (New York) with other Paradigmatic Copies. Volume 2: Paul Needham, Galileo Makes a Book: The First Edition of Sidereus Nuncius. Venice 1610. Berlin: Akademie Verlag, 2011, ISBN 978-3-05-005095-9. Annals of Science, 72(1):126–128, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Laudan:1966:CMP

[Lau66] Laurens Laudan, M.A., Ph.D. The clock metaphor and probabilism: The impact of Descartes on English methodological thought, 1650– 65. Annals of Science, 22(2):73–104, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Laube:2021:BRA

[Lau21] Stefan Laube. Book review: Anna Zieglerin and the Lion’s blood: alchemy and end times in reformation Germany, by Tara Nummedal, Philadelphia, University of Pennsylvania Press, 2019, xi + 288 pp., 16 fig., £37.00 (Hardback); $39.95, ISBN 978-0-81225-089-3. Annals of Science, 78(2):250–252, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawson:2008:BMS

[Law08] Gordon S. Lawson. Benjamin Moore, science, and medical planning in early Twentieth-Century Britain. Annals of Science, 65(4):487– 517, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawrence:2009:DUM

[Law09a] Christopher Lawrence. Degeneration under the microscope at the fin de siècle. Annals of Science, 66(4):455–471, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawrence:2009:GB

[Law09b] Christopher Lawrence. The Grand Banks. Annals of Science, 66(3): 407–417, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawrence:2013:BRJ

[Law13] Christopher Lawrence. Book review: Jimena Canales, A Tenth of a Second. A History. Chicago: The University of Chicago Press, 2010. 288 pp. Illus. $35.00. ISBN-13 978-0-226-09318-5. Annals of Science, 70(1):122–123, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawrence:2015:TWH

[Law15] Christopher Lawrence. Today, will T. H. Huxley dig the garden or paper the parlour? Annals of Science, 72(2):258–261, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lawrence:2021:SMI

[Law21] Christopher Lawrence. The sense of movement. An intellectual history. Annals of Science, 78(4):531–533, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

LeGrand:1972:LOT

[Le 72] H. E. Le Grand, A.B., Ph.D. Lavoisier’s oxygen theory of acidity. Annals of Science, 29(1):1–18, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

LeGrand:1974:ICM

[Le 74] H. E. Le Grand. Ideas on the composition of muriatic acid and their relevance to the oxygen theory of acidity. Annals of Science, 31(3):

213–225, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lea:2018:CTH

[Lea18] Andrew S. Lea. The Chinese typewriter: A history. Annals of Science, 75(3):260–262, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leegwater:2013:BRB

[Lee13] Arie Leegwater. Book review: Collected Papers on the Philosophy of Chemistry, by Eric R. Scerri; Selected Papers on the Periodic Table, by Eric R. Scerri. Annals of Science, 70(4):567–568, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lee:2023:CRC

[Lee23] Jia Hui Lee. Colonial rodent control in Tanganyika and the application of ecological frameworks. Annals of Science, 80(2):83–111, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2181399.

Leggett:2017:BRS

[Leg17] Don Leggett. Book review: Scientists’ expertise as performance: between state and society, 1860–1960. Annals of Science, 74(1):84– 86, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lehman:2012:PJM

[Leh12a] Christine Lehman. Pierre-Joseph Macquer: an eighteenth-century artisanal-scientific expert. Annals of Science, 69(3):307–333, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lehoux:2012:WDR

[Leh12b] Daryn Lehoux. What did the Romans know?: an inquiry into science and worldmaking. University of Chicago Press, Chicago, IL, USA and London, UK, 2012. ISBN 0-226-47114-4 (hardcover). xi + 275 pp. LCCN Q124.95 .L44 2012.

Leicester:1973:ETM

[Lei73] Henry M. Leicester, Ph.D. The electrical theories of M. V. Lomonosov. Annals of Science, 30(3):299–310, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leicester:1976:GIM

[Lei76] Henry M. Leicester. The geochemical ideas of Mikhail Lomonosov. Annals of Science, 33(4):341–350, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leitao:2020:ISW

[Lei20] Henrique Leitão. Inquisition and science: where do we stand now? Annals of Science, 77(1):127–133, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lepeltier:2006:EMI

[Lep06] Thomas Lepeltier. Edward Milne’s influence on modern cosmology. Annals of Science, 63(4):471–481, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leung:2008:TCA

[Leu08] Danny C. K. Leung. Thomas Clifford Allbutt and comparative pathology. Annals of Science, 65(4):547–571, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:1978:CPV

[Lev78] Trevor Levere. S. T. Coleridge: a poet’s view of science. Annals of Science, 35(1):33–44, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:1986:MIC

[Lev86] Trevor H. Levere. Magnetic instruments in the Canadian Arctic expeditions of Franklin, Lefroy, and Nares. Annals of Science, 43(1): 57–76, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:1994:CAE

[Lev94] Trevor H. Levere. Chronometers on the Arctic expeditions of John Ross and William Edward Parry: With notes on a letter from Messrs. William Parkinson & William James Frodsham. Annals of Science, 51(2):165–175, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:1999:MI

[Lev99] Trevor Levere. More introductions. Annals of Science, 56(2):111, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:2009:JUT

[Lev09a] T. H. Levere. Journals under threat: a joint response from history of science, technology and medicine editors. Annals of Science, 66(1): 1–3, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:2009:EIC

[Lev09b] Trevor Levere. Editor’s introduction to the cumulative index volumes 44–64. Annals of Science, 66(1S):1, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levere:2014:NET

[Lev14a] Trevor H. Levere. A new editorial team. Annals of Science, 71(1): 1, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levitin:2014:EHR

[Lev14b] Dmitri Levitin. The experimentalist as humanist: Robert Boyle on the history of philosophy. Annals of Science, 71(2):149–182, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levitin:2021:ING

[Lev21] Dmitri Levitin. Isaac Newton’s ‘De gravitatione et aequipondio fluidorum’: its purpose in historical context. Annals of Science, 78(2): 133–161, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lewis:1976:FRS

[Lew76] Christopher J. T. Lewis. The fortunes of Richard Swineshead in the time of Galileo. Annals of Science, 33(6):561–584, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lewis:1977:HGB

[Lew77] Albert C. Lewis. H. Grassmann’s 1844 Ausdehnungslehre and Schleiermacher’s Dialektik. Annals of Science, 34(2):103–162, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leu:2008:CGP

[LGKW08] Urs B. Leu, Conrad Gessner, Raffael Keller, and Sandra Weidmann. Conrad Gessner’s private library, volume 5 of History of science and medicine library. Brill, Leiden, The Netherlands, 2008. ISBN 90-04-16723-4. xii + 310 + 20 pp. LCCN Z997.G42.

Li:1994:ICT

[Li94] Yadong Li. The introduction of chemical theories into nineteenthcentury China. Annals of Science, 51(5):517–530, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Li:2012:PPM

[Li12] Shang-Jen Li. People’s Peking Man: Popular science and human identity in Twentieth-Century China. Annals of Science, 69(3):443– 446, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lightman:2010:VPS

[Lig10] Bernard V. Lightman. Victorian popularizers of science: designing nature for new audiences. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-48119-0 (paperback). xvi + 545 pp. LCCN Q127.G4 L54 2010.

Lightman:2016:GSC

[Lig16] Bernard V. Lightman, editor. Global Spencerism: the communication and appropriation of a British evolutionist, volume 1 of Cultural dynamics of science. Brill, Leiden, The Netherlands, 2016. ISBN 90-04-26399-3 (hardcover), 90-04-26400-0 (e-book). ISSN 2351-9932. xiii + 309 pp. LCCN B1657 .G56 2016.

Lilley:1948:NJ

[Lil48] S. Lilley, M.Sc., Ph.D. ‘Nicholson’s Journal’ (1797–1813). Annals of Science, 6(1):78–101, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lindeboom:1963:PLI

[Lin63] G. A. Lindeboom. Pitcairne’s Leyden interlude described from the documents. Annals of Science, 19(4):273–284, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Linthorst:2010:CIS

[Lin10] Johan Alfredo Linthorst. Chemistry: The impure science. Annals of Science, 67(4):579–581, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Linstrum:2016:RMP

[Lin16] Erik Linstrum. Ruling minds: psychology in the British Empire. Harvard University Press, Cambridge, MA, USA, 2016. ISBN 0-674-08866-2. 309 pp. LCCN BF108.G7 L56 2016.

Lindee:2022:BRS

[Lin22] M. Susan Lindee. Book review: Science on a mission: how military funding shaped what we do and don’t know about the ocean, by Naomi Oreskes, Chicago and London, University of Chicago Press, 2021, 738 pp., $40.00 (hardback), ISBN 978–0-226-73238-I. Annals of Science, 79(1):133–135, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Laboulle:1937:SQS

[LL37] M. J. J. Laboulle, B.A. and H. Levy, M.A., D.Sc. Some questions of scientific method in the ‘Siècle des Lumières’. Annals of Science, 2 (2):153–163, April 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lucas:2006:BSP

[LMBM06] A. M. Lucas, Sara Maroske, and Andrew Brown-May. Bringing science to the public: Ferdinand von Mueller and botanical education in Victorian Victoria. Annals of Science, 63(1):25–57, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lomb:2011:TVP

[Lom11] Nick Lomb. Transit of Venus: 1631 to the present. Experiment, New York, NY, USA, 2011. ISBN 1-61519-055-4. 228 pp. LCCN QB509 .L66 2012.

Long:2019:VTN

[Lon19] Kathleen Long. The voyage of thought: navigating knowledge across the sixteenth-century world. Annals of Science, 76(1):96–98, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lopez:1960:STG

[Lop60] Claude A. Lopez, M.A. Saltpetre, tin and gunpowder: Addenda to the correspondence of Lavoisier and Franklin. Annals of Science, 16 (2):83–94, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Love:1974:HBE

[Lov74] Rosaleen Love. Herman Boerhaave and the element-instrument concept of Fire. Annals of Science, 31(6):547–559, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Love:1979:AEL

[Lov79] Rosaleen Love. ‘Alice in Eugenics-Land’: Feminism and eugenics in the scientific careers of Alice Lee and Ethel Elderton. Annals of Science, 36(2):145–158, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Loveland:2011:GFC

[Lov11] Jeff Loveland. Good faith collaboration: The culture of Wikipedia. Annals of Science, 68(4):555–558, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Loveland:2012:OC

[Lov12] Jeff Loveland. Oeuvres complètes. Annals of Science, 69(1):145– 149, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Loveland:2013:BRA

[Lov13] Jeff Loveland. Book review: Aude Doody, Pliny’s Encyclopedia: The Reception of the Natural History. Cambridge: Cambridge University Press, 2010. viii + 194 pp. $95.00 / £55.00. ISBN-13 978-0-521-49103-7. Annals of Science, 70(1):143–146, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lowe:1975:WHM

[Low75] Victor Lowe. A. N. Whitehead on his mathematical goals: a letter of 1912. Annals of Science, 32(2):85–101, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Low:1990:JSW

[Low90] Morris Fraser Low. Japan’s secret war? ‘Instant’ scientific manpower and Japan’s World War II. Atomic bomb project. Annals of Science, 47(4):347–360, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Low:1996:MRB

[Low96] Morris F. Low. Medical representations of the body in Japan: Gender, class, and discourse in the eighteenth century. Annals of Science, 53(4):345–359, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Low:2003:BRB

[Low03] M. F. Low. Book review: Current Perspectives in the History of Science in East Asia, edited by Y. S. Kim and F. Bray. Annals of Science, 60(3):327–328, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Levene:2011:CGM

[LPST11] Alysa Levene, Martin Powell, John Stewart, and Becky Taylor. Cradle to grave: municipal medicine in interwar England and Wales, volume 4 of Studies in the history of medicine. Peter Lang, Bern, Switzerland, 2011. ISBN 3-03910-904-9. xii + 258 pp. LCCN RA487 .C73 2011.

Leone:2000:SSL

[LR00] Matteo Leone and Nadia Robotti. Stellar, solar and laboratory spectra: The history of Lockyer’s proto-elements. Annals of Science, 57(3):241–266, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Linder:1968:SBB

[LS68] Bertel Linder, M.A., Ll.B. and W. A. Smeaton, D.Sc. Schwediauer, Bentham and Beddoes: Translators of Bergman and Scheele. Annals of Science, 24(4):259–273, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lodwig:1974:ICL

[LS74] T. H. Lodwig, M.Sc. and W. A. Smeaton, D.Sc. The ice calorimeter of Lavoisier and Laplace and some of its critics. Annals of Science, 31(1):1–18, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lamy:2015:LCD

[LS15a] Jérôme Lamy and Frédéric Soulu. L’émergence contrariée du chronographe imprimant dans les observatoires français (fin 19^{e}–début 20^{e} s.). Annals of Science, 72(1):75–98, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Loveland:2015:PEB

[LS15b] Jeff Loveland and Stéphane Schmitt. Poinsinet’s edition of the Naturalis historia (1771–1782) and the revival of Pliny in the sciences of the Enlightenment. Annals of Science, 72(1):2–27, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Leber:2023:MHC

[LS23] Christoffer Leber and Claus Spenninger. The many histories of the conflict thesis: the science vs. religion narrative in nineteenth-century Germany. Annals of Science, 80(4):390–417, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2187086.

Lucier:2008:SSC

[Luc08] Paul Lucier. Scientists and swindlers: consulting on coal and oil in America, 1820–1890. Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-9003-9 (hardcover). xiii + 426 pp. LCCN Q127.U6 L83 2008.

Luckhurst:2019:BRH

[Luc19] Roger Luckhurst. Book review: A history of the future: prophets of progress from H. G. Wells to Isaac Asimov, by Peter J. Bowler, Cambridge, Cambridge University Press, 2017, x + 287 pp., 9 colour plts, 18 black and white plts, £19.99, ISBN 978-1-316-60262-1. Annals of Science, 76(3–4):389–391, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lundgreen:1990:EEE

[Lun90] Peter Lundgreen. Engineering education in Europe and the U.S.A., 1750–1930: The rise to dominance of school culture and the engineering professions. Annals of Science, 47(1):33–75, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lusk:2014:BRB

[Lus14] Greg Lusk. Book review: A Vast Machine: Computer Models, Climate Data, and the Politics of Global Warming, by P. N. Edwards. Annals of Science, 71(2):295–298, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lusinchi:2018:GFP

[Lus18] Dominic Lusinchi. ‘The Great Fiasco’ of the 1948 presidential election polls: status recognition and norms conflict in social science. Annals of Science, 75(2):120–144, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lynn:2013:BRB

[Lyn13] Michael R. Lynn. Book review: Science, libertinage et clandestinité à l’aube des Lumières: Le transformisme de Telliamed, by Claudine Cohen. Annals of Science, 70(4):575–577, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lynn:2014:BRB

[Lyn14] Michael R. Lynn. Book review: Investigating the Supernatural: From Spiritism and Occultism to Psychical Research and Metapsychics in France, 1853–1931, by Sofie Lachapelle. Annals of Science, 71(1): 142–144, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Lynn:2015:BRB

[Lyn15] Michael R. Lynn. Book review: Female Sexuality and Cultural Degradation in Enlightenment France: Medicine and Literature. Annals of Science, 72(1):129–130, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ma:2021:MML

[Ma21] Xi Ma. Mineral and mineralogy in late Qing China: translations and conceptualizations, 1860s–1910s. Annals of Science, 78(1):64– 91, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLean:1972:HRS

[Mac72a] J. MacLean, D.Sc. On harmonic ratios in spectra. Annals of Science, 28(2):121–137, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLean:1972:STG

[Mac72b] J. MacLean, D.Sc. Science and theology at Groningen University (1698–1702). Annals of Science, 29(2):187–201, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLean:1973:NSJ

[Mac73] J. MacLean, D.Sc. Natural science in Japan. I. Before 1830. Annals of Science, 30(3):257–298, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLachlan:1976:GEP

[Mac76] James MacLachlan. Galileo’s experiments with pendulums: Real and imaginary. Annals of Science, 33(2):173–185, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacKenzie:1979:KPP

[Mac79] D. MacKenzie. Karl Pearson and the professional middle class. Annals of Science, 36(2):125–143, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLachlan:1982:NRH

[Mac82] James MacLachlan. Note on R. H. Naylor’s error in analysing experimental data. Annals of Science, 39(4):381–384, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacKenzie:1985:PIS

[Mac85a] Donald MacKenzie. The political ‘implications’ of scientific theories: a comment on Bowler. Annals of Science, 42(4):417–419, 1985. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Bow84] and response [Bow85].

MacRoberts:1985:WMP

[Mac85b] Michael H. MacRoberts. Was Mendel’s paper on Pisum neglected or unknown? Annals of Science, 42(3):339–345, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLeod:1989:ASA

[Mac89] Roy M. MacLeod. The ‘arsenal’ in the strand: Australian chemists and the British munitions effort 1916–1919. Annals of Science, 46(1): 45–67, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacLeod:1993:CGW

[Mac93] Roy MacLeod. The chemists go to war: The mobilization of civilian chemists and the British war effort, 1914–1918. Annals of Science, 50(5):455–481, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MacKenzie:2009:MEN

[Mac09] John M. (John MacDonald) MacKenzie. Museums and empire: natural history, human cultures and colonial identities. Studies in imperialism. Manchester University Press, Manchester, UK, 2009. ISBN 0-7190-8022-3 (hardcover). xv + 286 pp. LCCN ????

MacDonald:2010:PDA

[Mac10] Helen (Helen Patricia) MacDonald. Possessing the dead: the artful science of anatomy. Melbourne University Press, Carlton, Victoria, Australia, 2010. ISBN 0-522-85735-3 (paperback). vii + 289 pp. LCCN GT3190 .M33 2010.

Macdonald:2018:OEY

[Mac18] Lee T. Macdonald. The origins and early years of the Magnetic and Meteorological Department at Greenwich Observatory, 1834–1848. Annals of Science, 75(3):201–233, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Macdonald:2019:MSP

[Mac19] Lee T. Macdonald. Making stars physical: the astronomy of Sir John Herschel. Annals of Science, 76(2):227–230, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Macdonald:2020:FUP

[Mac20a] J. Marc Macdonald. Failed utopias and practical chemistry: the Priestleys, the Du Ponts, and the transmission of transatlantic science, 1770–1820. Annals of Science, 77(2):215–252, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Macuglia:2020:NIC

[Mac20b] Daniele Macuglia. Newtonianism and information control in Rome at the wake of the eighteenth century. Annals of Science, 77(1):108– 126, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maclean:2023:RMS

[Mac23] Ian Maclean. Renaissance medicine: a short history of European medicine in the sixteenth century. Annals of Science, 80(2):195–197, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2130979.

Maddison:1955:NSM

[Mad55] R. E. W. Maddison, B.Sc., Ph.D. Notes on some members of the Hanckwitz family in England. Annals of Science, 11(1):64–73, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1957:SFA

[Mad57] R. E. W. Maddison, B.Sc., Ph.D. A summary of former accounts of the life and work of Robert Boyle. Annals of Science, 13(2):90– 108, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1959:PHR

[Mad59] R. E. W. Maddison, B.Sc., Ph.D. The portraiture of the honourable Robert Boyle, F.R.S. Annals of Science, 15(3–4):141–214, 1959. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1961:EPW

[Mad61a] R. E. W. Maddison, B.Sc., Ph.D. The earliest published writing of Robert Boyle. Annals of Science, 17(3):165–173, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1961:PFB

[Mad61b] R. E. W. Maddison, B.Sc., Ph.D. The plagiary of Francis Boyle. Annals of Science, 17(2):111–120, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1962:FER

[Mad62] R. E. W. Maddison, B.Sc., Ph.D. The first edition of Robert Boyle’s medicinal experiments. Annals of Science, 18(1):43–47, 1962. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1964:BH

[Mad64] R. E. W. Maddison, B.Sc., Ph.D. Boyle’s Hell. Annals of Science, 20(2):101–110, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1988:LRB

[Mad88a] R. E. W. Maddison. The life of Robert Boyle: Addenda. Annals of Science, 45(2):193–195, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1988:PRB

[Mad88b] R. E. W. Maddison. ‘the portraiture of Robert Boyle’: Addenda. Annals of Science, 45(2):196–198, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maddison:1988:ULJ

[Mad88c] R. E. W. Maddison. An unpublished letter of Justus Liebig. Annals of Science, 45(2):191–192, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maehle:2005:QDD

[Mae05] Andreas-Holger Maehle. The quantification and differentiation of the drug receptor theory, c. 1910–1960. Annals of Science, 62(4):479– 500, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maffioli:2010:DAA

[Maf10] Cesare Maffioli. La via delle acque, 1500–1700: appropriazione delle arti e trasformazione delle matematiche. (Italian) [The path of the water, 1500–1700: appropriation of arts and transformation of mathematics], volume 4 of Hyperchen. L. S. Olschki, Firenze, Italia, 2010. ISBN 88-222-6008-2. xxii + 394 + 1 pp. LCCN TC145 .M3155 2010.

43.00 EUR. URL http://www.loc.gov/catdir/toc/casalini09/
10482326.pdf.

Maffioli:2013:FEC

[Maf13] Cesare S. Maffioli. A fruitful exchange/conflict: Engineers and mathematicians in early modern Italy. Annals of Science, 70(2):197–228, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Magnello:1998:KPM

[Mag98] M. Eileen Magnello. Karl Pearson’s mathematization of inheritance: From ancestral heredity to Mendelian genetics (1895–1909). Annals of Science, 55(1):35–94, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mallatt:1982:DBF

[Mal82] Jon M. Mallatt. Dr Beringer’s fossils: a study in the evolution of scientific world view. Annals of Science, 39(4):371–380, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Malley:1994:TCL

[Mal94] Marjorie Malley. Thermodynamics and cold light. Annals of Science, 51(3):203–224, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Malet:2003:KT

[Mal03] Antoni Malet. Kepler and the telescope. Annals of Science, 60(2): 107–136, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Malet:2009:JMA

[Mal09] Antoni Malet. José María Albareda (1902–1966) and the formation of the Spanish Consejo Superior de Investigaciones Científicas. Annals of Science, 66(3):307–332, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MaltaRomeiras:2020:ICS

[Mal20a] Francisco Malta Romeiras. The Inquisition and the censorship of science in early modern Europe: Introduction. Annals of Science, 77 (1):1–9, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

MaltaRomeiras:2020:PIP

[Mal20b] Francisco Malta Romeiras. Putting the Indices into practice: censoring science in early modern Portugal. Annals of Science, 77(1): 71–95, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Manning:1993:JEI

[Man93] Robert J. Manning. John Elliot and the inhabited sun. Annals of Science, 50(4):349–364, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martin:1976:GMH

[Mar76] R. N. D. Martin. The genesis of a mediaeval historian: Pierre Duhem and the origins of statics. Annals of Science, 33(2):119–129, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinsen:2001:DEC

[Mar01] Hanna E. H. Martinsen. Dyeing in the Eighteenth Century. Annals of Science, 58(4):399–402, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Marcum:2006:SVS

[Mar06] James A. Marcum. ‘soup’ vs. ‘sparks’: Alexander Forbes and the synaptic transmission controversy. Annals of Science, 63(2):139–156, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinsen:2009:SIP

[Mar09] Hanna E. H. Martinsen. Scientific institutions and practice in France and Britain, c. 1700–c. 1870. Annals of Science, 66(2):290–292, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Marr:2011:BRG

[Mar11a] Alexander Marr. Between Raphael and Galileo: Mutio Oddi and the mathematical culture of late Renaissance Italy. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-50628-2 (hardcover). xiii + 359 pp. LCCN QA27.I8 M35 2011.

Martinez:2011:ERM

[Mar11b] Alberto A. Martínez. Essay review: The myriad pieces of Einstein’s remains: John Stachel, Robert Schulmann, Diana Kormos Buchwald, et al, editors, The Collected Papers of Albert Einstein. Annals of Science, 68(2):267–280, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinez:2011:SST

[Mar11c] Alberto Antonio Martínez. Science secrets: the truth about Darwin’s finches, Einstein’s wife, and other myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2011. ISBN 0-8229-4407-3 (hardcover). xviii + 324 pp. LCCN Q173 .M316 2011.

Margocsy:2013:FMM

[Mar13] Dániel Margócsy. The fuzzy metrics of money: The finances of travel and the reception of curiosities in Early Modern Europe. Annals of Science, 70(3):381–404, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Margocsy:2015:BRB

[Mar15] Dániel Margócsy. Book review: Androids in the Enlightenment: Mechanics, Artisans, and Cultures of the Self. Annals of Science, 72(3): 407–409, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martin:2016:OBW

[Mar16a] Alison E. Martin. Outward bound: women translators and scientific travel writing, 1780–1800. Annals of Science, 73(2):157–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinez:2016:GBH

[Mar16b] Alberto A. Martinez. Giordano Bruno and the heresy of many worlds. Annals of Science, 73(4):345–374, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martinez:2019:MIG

[Mar19] Alberto A. Martinez. Melchior Inchofer, Giordano Bruno, and the soul of the world. Annals of Science, 76(3–4):267–302, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martin:2021:BRF

[Mar21a] Craig Martin. Book review: Forbidden knowledge: medicine, science, and censorship in early modern Italy, by Hannah Marcus, Chicago, University of Chicago Press, 2020, xi + 356 pp., 36 fig., $45.00 (Hardback), ISBN 978-0-226-73658-7. Annals of Science, 78(2):261–264, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Martin:2021:FBJ

[Mar21b] Craig Martin. Francis Bacon, José de Acosta, and traditions of natural histories of winds. Annals of Science, 77(4):445–468, 2021. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mason:1953:SRPa

[Mas53a] S. F. Mason, Ph.D. The scientific revolution and the Protestant Reformation. I. Calvin and Servetus in relation to the new astronomy and the theory of the circulation of the blood. Annals of Science, 9 (1):64–87, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mason:1953:SRPb

[Mas53b] S. F. Mason, Ph.D. The scientific revolution and the Protestant Reformation. II. Lutheranism in relation to iatrochemistry and the German nature-philosophy. Annals of Science, 9(2):154–175, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Massa:1973:GBT

[Mas73] Daniel Massa, M.A. Giordano Bruno and the top-sail experiment. Annals of Science, 30(2):201–211, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Matheson:1954:TPM

[Mat54] Colin Matheson, M.A., B.Sc., F.M.A. Thomas Pennant and the Morris brothers. Annals of Science, 10(3):258–271, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Matthews:1976:DSA

[Mat76] M. H. Matthews. The development of the synthetic alkali industry in Great Britain by 1823. Annals of Science, 33(4):371–382, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Matta:2011:HGB

[Mat11] Christina Matta. H. G. Bronn, Ernst Haeckel, and the origins of German Darwinism. Annals of Science, 68(4):568–570, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mauskopf:1969:TBD

[Mau69] Seymour H. Mauskopf. Thomson before Dalton: Thomas Thomson’s considerations of the issue of combining weight proportions prior to his acceptance of Dalton’s chemical atomic theory. Annals of Science, 25(3):229–242, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maulitz:1990:CFD

[Mau90] Russell C. Maulitz. In the clinic: Framing disease at the Paris hospital. Annals of Science, 47(2):127–137, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mauskopf:2015:WBD

[Mau15] Seymour H. Mauskopf. William Barlow and the determination of atomic arrangement in crystals: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):206–223, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mauger:2018:WTE

[Mau18] Matthew Mauger. Writing tea’s empire. Annals of Science, 75(3): 255–259, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Maydom:2021:BRD

[May21] Katrina Maydom. Book review: Drugs on the Page: Pharmacopoeias and Healing Knowledge in the Early Modern Atlantic World, edited by M. J. Crawford and J. M. Gabriel, Pittsburgh: Pittsburgh University Press, 2019, ix + 374 pp., $50.00, ISBN 0-822-94562-2. Annals of Science, 78(2):259–261, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mazzotti:2010:BRE

[Maz10] Massimo Mazzotti. Book review: Earth sciences. Donata Brianta, Europa mineraria. Circolazione delle élites e trasferimento tecnologico (secoli XVIII–XIX), ISBN-13 978-88-464-8818-3. Annals of Science, 67(2):268–270, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mazumdar:2012:HPC

[Maz12] Pauline M. H. Mazumdar. Heredity produced: At the crossroads of biology, politics and culture 1500–1870. Annals of Science, 69(1): 136–137, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McAlpin:2012:FSC

[McA12] Mary McAlpin. Female sexuality and cultural degradation in Enlightenment France: medicine and literature. Ashgate, Farnham, Surrey, UK, 2012. ISBN 1-4094-2241-0 (hardcover), 1-4094-2242-9 (e-book). xii + 195 pp. LCCN PQ265 .M33 2012.

McBryde:1991:PDE

[McB91] W. A. E. McBryde. Petroleum deodorized: Early Canadian history of the ‘doctor sweetening’ process. Annals of Science, 48(2):103– 111, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1936:SEB

[McC36a] Grant McColley, Ph.D. The second edition of The discovery of a world in the moone. Annals of Science, 1(3):330–334, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1936:SCD

[McC36b] Grant McColley, Ph.D. The seventeenth-century doctrine of a plurality of worlds. Annals of Science, 1(4):385–430, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1937:FST

[McC37] Grant McColley, Ph.D. A facsimile of Salusbury’s translation of Didacus à Stunica’s Commentary upon Job. Annals of Science, 2(2):

179–182, April 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1938:RWC

[McC38] Grant McColley, Ph.D. The Ross–Wilkins controversy. Annals of Science, 3(2):153–189, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1939:DBJ

[McC39a] Grant McColley, Ph.D. The debt of Bishop John Wilkins to the Apologia pro Galileo of Tommaso Campanella. Annals of Science, 4 (2):150–168, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1939:JHB

[McC39b] Grant McColley, Ph.D. J. H. and the Astronomia Crystallina. Annals of Science, 4(3):319–321, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McColley:1939:NHB

[McC39c] Grant McColley, Ph.D. Nicholas Hill and the Philosophia Epicurea. Annals of Science, 4(4):390–405, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McCosh:1975:BVV

[McC75] F. W. J. McCosh. Boussingault versus Ville: The social, political and scientific aspects of their disputes. Annals of Science, 32(5): 475–490, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mcclellan:1979:SPT

[Mcc79] James E. Mcclellan, III. The scientific press in transition: Rozier’s journal and the scientific societies in the 1770s. Annals of Science, 36(5):425–449, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McCann:1983:PLP

[McC83] Alison McCann. A private laboratory at Petworth House, Sussex, in the late Eighteenth Century. Annals of Science, 40(6):635–655, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McConnell:1986:SLW

[McC86] Anita McConnell. The scientific life of William Scoresby Jnr, with a catalogue of his instruments and apparatus in the Whitby Museum. Annals of Science, 43(3):257–286, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McConnell:1989:AAS

[McC89] Anita McConnell. Aluminium and its alloys for scientific instruments, 1855–1900. Annals of Science, 46(6):611–620, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McConnell:1991:CSJ

[McC91] Anita McConnell. La Condamine’s scientific journey down the River Amazon, 1743–1744. Annals of Science, 48(1):1–19, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McConnell:1994:BPA

[McC94] Anita McConnell. Bankruptcy proceedings against William Harris, optician, of Cornhill, 1830. Annals of Science, 51(3):273–279, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McConnell:2005:OMB

[McC05] Anita McConnell. Origins of the marine barometer. Annals of Science, 62(1):83–101, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McCrory:2010:NIL

[McC10] Donald McCrory. Nature’s interpreter: the life and times of Alexander von Humboldt. Lutterworth Press, Cambridge, UK, 2010. ISBN 0-7188-9231-3. 242 pp. LCCN Q143.H9 M33 2010.

McCrae:2013:BRB

[McC13] Morrice McCrae. Book review: Medicine in the Remote and Rural North, 1800–2000, edited by J. T. H. Connor and Stephen Curtis. Annals of Science, 70(4):580–582, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McCrae:2014:BRB

[McC14] Morrice McCrae. Book review: Cradle to Grave, by A. Levene, M. Powell, J. Stewart, and B. Taylor. Annals of Science, 71(2):280– 283, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McCaskey:2021:HTM

[McC21] John P. McCaskey. History of ‘temperature’: maturation of a measurement concept. Annals of Science, 77(4):399–444, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McDonald:1972:PCA

[McD72] John F. McDonald, M.A. Properties and causes: An approach to the problem of hypothesis in the scientific methodology of Sir Isaac Newton. Annals of Science, 28(3):217–233, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McEvoy:2010:CUO

[McE10a] John G. McEvoy. Chemist: Understanding the origins of the steam age. Annals of Science, 67(4):581–583, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McEvoy:2010:HCR

[McE10b] John G. McEvoy. The historiography of the chemical revolution: patterns of interpretation in the history of science. Pickering and Chatto, London, UK, 2010. ISBN 1-84893-030-5. xiii + 328 pp. LCCN ????

McGuire:1978:EAN

[McG78] J. E. McGuire. Existence, actuality and necessity: Newton on space and time. Annals of Science, 35(5):463–508, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1936:BTC

[McK36a] Douglas McKie, Ph.D., B.Sc. Béraut’s theory of calcination (1747). Annals of Science, 1(3):269–293, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1936:TCM

[McK36b] Douglas McKie, Ph.D., B.Sc. On Thos. Cochrane’s ms. notes of Black’s chemical lectures, 1767–8. Annals of Science, 1(1):101–110, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1957:OAF

[McK57] Douglas McKie, D.Sc., Ph.D., F.R.I.C., F.R.S.E. The “Observations” of the Abbé François Rozier (1734–93). I. Annals of Science, 13(2):73–89, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1959:SMC

[McK59] Douglas McKie, D.Sc., Ph.D., F.R.S.E. On some ms. copies of Black’s chemical lectures. II. Annals of Science, 15(2):65–73, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1960:SMC

[McK60] Douglas McKie, D.Sc., Ph.D., F.R.S.E. On some ms. copies of Black’s chemical lectures. III. Annals of Science, 16(1):1–9, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1962:SMC

[McK62] Douglas McKie, D.Sc., Ph.D., F.R.S.E. On some ms. copies of Black’s chemical lectures. IV. Annals of Science, 18(2):87–97, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1965:SMC

[McK65] Douglas McKie, D.Sc., Ph.D., F.S.A., F.R.S.E. On some ms. copies of Black’s chemical lectures. V. Annals of Science, 21(4):209–255, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1967:SMC

[McK67] Douglas McKie, D.Sc., Ph.D., F.S.A., F.R.S.E. On some MS. copies of Black’s chemical lectures. VI. Annals of Science, 23(1):1–33, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McOuat:2000:NHF

[McO00] Gordon McOuat. Networks, hybrids and forms of life. Annals of Science, 57(2):189–195, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mcvaugh:2021:BRG

[Mcv21] Michael Mcvaugh. Book review: Galen: A Thinking Doctor in Imperial Rome, by Vivian Nutton, London and New York, Routledge, 2020, xiii + 209 pp., 12 plts., $39.95 (Paperback), ISBN 978-0-367-35723-8. Annals of Science, 78(2):246–248, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1958:WCB

[MdVH58] Douglas McKie, D.Sc., Ph.D., F.R.I.C., F.R.S.E. and Niels H. de V. Heathcote, B.Sc., Ph.D. William Cleghorn’s De Igne (1779).

Annals of Science, 14(1):1–82, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meadows:1966:DAV

[Mea66] A. J. Meadows, M.A., D.Phil. The discovery of an atmosphere on Venus. Annals of Science, 22(2):117–127, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meadows:1970:ODE

[Mea70] A. J. Meadows, D.Phil. Observational defects in eighteenth-century British telescopes. Annals of Science, 26(4):305–317, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meadows:1985:OEG

[Mea85] Jack Meadows. Obituary: Eric Gray Forbes 1933–1984. Annals of Science, 42(6):547–548, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meadows:2009:SSP

[Mea09] A. J. Meadows. Science for sale: The perils, rewards, and delusions of campus capitalism. Annals of Science, 66(3):441–444, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meiklejohn:1954:BD

[Mei54] M. F. M. Meiklejohn, M.A. The birds of Dante. Annals of Science, 10(1):33–43, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meinel:1979:THM

[Mei79] Christoph Meinel. Teaching the history of medicine, science and technology in the Federal Republic of Germany and in West Berlin. Annals of Science, 36(3):279–289, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meijer:1985:HVN

[Mei85] Onno G. Meijer. Hugo de Vries no Mendelian? Annals of Science, 42(3):189–232, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meiring:2021:TRM

[Mei21] Henry-James Meiring. Thomas Robert Malthus, naturalist of the mind. Annals of Science, 77(4):495–523, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Melmore:1953:NPO

[Mel53] Sidney Melmore, B.Sc., F.R.A.S. Nathaniel Pigott’s observatory 1781–1793. Annals of Science, 9(3):281–286, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meli:1988:LEB

[Mel88] Domenico Bertoloni Meli. Leibniz’s excerpts from the Principia mathematica. Annals of Science, 45(5):477–505, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Melia:2009:CEC

[Mel09] Fulvio Melia. Cracking the Einstein code: Relativity and the birth of black hole physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-51951-1. xi + 137 pp. LCCN QC173.6 .M434 2009. With an afterword by Roy Kerr.

Meli:2010:RIS

[Mel10] Domenico Bertoloni Meli. The representation of insects in the seventeenth century: a comparative approach. Annals of Science, 67(3): 405–429, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mellyn:2021:HIE

[Mel21] Elizabeth W. Mellyn. Healers, innovators, entrepreneurs: women in early modern healthcare: Forgotten Healers: women and the pursuit of health in late Renaissance Italy, by Sharon Strocchia, Cambridge, MA, Harvard University Press, 2020, ix + 330 pp., $49.95, £39.95, EUR 45.00, ISBN 978-0-674-24174-9. Annals of Science, 78(2):252– 259, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mertens:1998:LRW

[Mer98] Joost Mertens. From the lecture room to the workshop: John Frederic Daniell, the constant battery and electrometallurgy around 1840. Annals of Science, 55(3):241–261, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meskens:1996:MEL

[Mes96] Ad Meskens. Mathematics education in late sixteenth-century Antwerp. Annals of Science, 53(2):137–155, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meskens:1997:MCC

[Mes97] Ad Meskens. Michiel Coignet’s contribution to the development of the sector. Annals of Science, 54(2):143–160, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Metzger:1956:TS

[Met56] Charles R. Metzger, Ph.D. Thoreau on science. Annals of Science, 12(3):206–211, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Meli:2010:RAE

[MG10] Domenico Bertoloni Meli and Anita Guerrini. The representation of animals in the early modern period. Annals of Science, 67(3):299– 301, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Marsden:2013:UCE

[MHO13] Ben Marsden, Hazel Hutchison, and Ralph O’Connor, editors. Uncommon contexts: encounters between science and literature, 1800– 1914, volume 23 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2013. ISBN 1-84893-362-2 (hardcover). xiii + 239 pp. LCCN PR468.S34 U53 2013.

Micault:2013:CLW

[Mic13] Natalie Pigeard Micault. The Curie’s lab and its women (1906–1934). Le laboratoire Curie et ses femmes (1906–1934). Annals of Science, 70(1):71–100, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millington:1942:HYH

[Mil42] E. C. Millington, M.Sc. History of the Young–Helmholtz theory of colour vision. Annals of Science, 5(2):167–176, 1942. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millington:1945:TCS

[Mil45] E. C. Millington, M.Sc., Ph.D. Theories of cohesion in the seventeenth century. Annals of Science, 5(3):253–269, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millington:1947:SCC

[Mil47] E. C. Millington, M.Sc., Ph.D. Studies in capillarity and cohesion in the eighteenth century. Annals of Science, 5(4):352–369, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millburn:1974:WSE

[Mil74] John R. Millburn. William Stukeley and the early history of the Orrery. Annals of Science, 31(6):511–528, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:1981:AES

[Mil81a] Arthur I. Miller. Albert Einstein’s Special Theory of Relativity: emergence (1905) and early interpretation, 1905–1911. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-04680-6, 0-201-04679-2 (paperback). xxviii + 466 pp. LCCN QC173.52 .M54.

Miller:1981:UIC

[Mil81b] Arthur I. Miller. Unipolar induction: a case study of the interaction between science and technology. Annals of Science, 38(2):155–189, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millburn:1983:LEC

[Mil83] John R. Millburn. The London evening courses of Benjamin Martin and James Ferguson, eighteenth-century lecturers on experimental philosophy. Annals of Science, 40(5):437–455, September 1983. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millburn:1985:JFL

[Mil85] John R. Millburn. James Ferguson’s lecture tour of the English Midlands in 1771. Annals of Science, 42(4):397–415, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1986:PHS

[Mil86] Allan A. Mills. Portable heliostats (solar illuminators). Annals of Science, 43(4):369–406, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Millburn:1988:OOI

[Mil88a] John R. Millburn. The office of ordnance and the instrument-making trade in the mid-eighteenth century. Annals of Science, 45(3):221– 293, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1988:MCB

[Mil88b] A. A. Mills. The mercury clock of the Libros del Saber. Annals of Science, 45(4):329–344, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1991:CB

[Mil91a] Allan A. Mills. Canada balsam. Annals of Science, 48(2):173–185, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1991:OPG

[Mil91b] Eric L. Mills. The oceanography of the Pacific: George F. McEwen, H. U. Sverdrup and the origin of physical oceanography on the west coast of North America. Annals of Science, 48(3):241–266, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1993:SHS

[Mil93] Allan A. Mills. Seasonal-hour sundials on vertical and horizontal planes, with an explanation of the scratch dial. Annals of Science, 50(1):83–93, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1996:ASS

[Mil96] Allan A. Mills. Altitude sundials for seasonal and equal hours. Annals of Science, 53(1):75–84, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:2001:ENH

[Mil01] Eric L. Mills. Enlightened natural history or the beginnings of oceanic science? Annals of Science, 58(4):403–408, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2002:DDB

[Mil02] David Philip Miller. ’distributing discovery’ between Watt and Cavendish: a reassessment of the nineteenth-century ’water controversy’. Annals of Science, 59(2):149–178, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:2003:EVB

[Mil03] Allan A. Mills. Early voltaic batteries: an evaluation in modern units and application to the work of Davy and Faraday. Annals of Science, 60(4):373–398, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:2004:EHI

[Mil04a] Allan A. Mills. The early history of insulated copper wire. Annals of Science, 61(4):453–467, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:2004:LHP

[Mil04b] Allan A. Mills. The lodestone: History, physics, and formation. Annals of Science, 61(3):273–319, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2008:SCS

[Mil08] David Philip Miller. Seeing the chemical steam through the historical fog: Watt’s steam engine as chemistry. Annals of Science, 65(1):47– 72, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2011:PED

[Mil11a] David Philip Miller. The political economy of discovery stories: The case of Dr Irving Langmuir and General Electric. Annals of Science, 68(1):27–60, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2011:SMK

[Mil11b] David Philip Miller. The singular and the making of knowledge at the Royal Society of London in the Eighteenth Century. Annals of Science, 68(4):563–565, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Milligan:2011:PB

[Mil11c] Barry Milligan. The poet as botanist. Annals of Science, 68(2):290– 291, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2012:BRJ

[Mil12] David Philip Miller. Book review: John G. McEvoy, The Historiography of the Chemical Revolution: Patterns of Interpretation in the History of Science, ISBN-13 978-1-84893-030-8. Annals of Science, 69(4):581–584, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:2013:BRB

[Mil13] Eric Mills. Book reviews: Bernard Finn and Daqing Yang, editors, Communications under the Seas: The Evolving Cable Network and its Implications. Dibner Institute Studies in the History of Science and Technology. Cambridge, MA and London: MIT Press, 2009. vii + 303 pp. $40.00. ISBN-13 978-0-262-01286-7. Ted Rowe, Connecting the Continents: Heart’s Content and the Atlantic Cable. St. John’s, NL: Creative Publishers, 2009. vii + 158 pp. $19.95 (Canadian) pbk. ISBN-13 978-1-897174-42-5. Annals of Science, 70(1):138–141, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2015:IOG

[MIL15] David Philip Miller, Rob Iliffe, and Trevor Levere. Ivor Owen Grattan-Guinness (1941–2014). Annals of Science, 72(3):276–278, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Miller:2016:BRL

[Mil16] David Philip Miller. Book review: The Limits of Matter. Chemistry, Mining & Enlightenment. Annals of Science, 73(1):110–112, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1064997.

Miller:2017:SSS

[Mil17] David Philip Miller. The story of ‘Scientist: The Story of a Word’. Annals of Science, 74(4):255–261, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Minor:2021:ACV

[Min21] Adriana Minor. Atoms in the campus: Van de Graaff accelerators and the making of two major Latin American universities in 1950s Brazil and Mexico. Annals of Science, 78(4):504–530, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mitchell:1978:MGS

[Mit78] Ian Mitchell. Marxism and German scientific materialism. Annals of Science, 35(4):379–400, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mittelstrass:1979:PCM

[Mit79] J. Mittelstrass. The Philosopher’s conception of Mathesis Universalis from Descartes to Leibniz. Annals of Science, 36(6):593–610, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mitchell:2012:MFE

[Mit12] Daniel Jon Mitchell. Measurement in French experimental physics from Regnault to Lippmann. Rhetoric and theoretical practice. Annals of Science, 69(4):453–482, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mills:1989:TLC

[MJ89] A. A. Mills and M. L. Jones. Three lenses by Constantine Huygens in the possession of the Royal Society of London. Annals of Science, 46(2):173–182, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

McKie:1960:SLJ

[MK60] Douglas McKie, D.Sc., Ph.D., F.R.S.E. and David Kennedy, M.Sc. On some letters of Joseph Black and others. Annals of Science, 16 (3):129–170, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morrison-Low:2012:EBS

[ML^{+}12] A. D. Morrison-Low et al., editors. From Earth-bound to satellite: telescopes, skills, and networks, volume 23 / 2 of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-21150-0 (hardcover). xxix + 265 pp. LCCN QB88 .F86 2012.

Morrison-Low:2017:HSI

[MLSB17] A. D. Morrison-Low, Sara Schechner, and Paolo Brenni, editors. How Scientific Instruments Have Changed Hands, volume 56(5) of History of science and medicine library; Scientific instruments and collections. Brill, Leiden, The Netherlands, 2017. ISBN 90-04-32492-5 (hardcover), 90-04-32493-3 (e-book). xxxii + 239 pp. LCCN Q185 .H75 2017.

Martin-Nielsen:2013:DMR

[MN13] Janet Martin-Nielsen. ‘the deepest and most rewarding hole ever drilled’: Ice cores and the Cold War in Greenland. Annals of Science, 70(1):47–70, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Molella:1984:ESJ

[Mol84] Arthur P. Molella. At the edge of science: Joseph Henry, ‘visionary theorizers’, and the Smithsonian Institution. Annals of Science, 41 (5):445–461, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Molland:1996:AAA

[Mol96] George Molland. Addressing ancient authority: Thomas Bradwardine and Prisca Sapientia. Annals of Science, 53(3):213–233, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Monaldi:2005:LOS

[Mon05] Daniela Monaldi. Life of μ: The observation of the spontaneous decay of mesotrons and its consequences, 1938–1947. Annals of Science, 62 (4):419–455, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Moore:1987:ERB

[Moo87] J. R. Moore. Essay review: Born-again social Darwinism: Michael Ruse, Taking Darwin Seriously: A Naturalistic Approach to Philosophy. Annals of Science, 44(4):409–417, July 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morgan:1936:SSL

[Mor36] J. R. Morgan, Ph.D., M.Sc., F.I.C. The search for a safety-lamp in mines. Annals of Science, 1(3):302–329, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morgan:1968:GHE

[Mor68] Paul Morgan, M.A., F.S.A. George Hartgill: An Elizabethan parson– astronomer and his library. Annals of Science, 24(4):295–311, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morgan:1986:NPA

[Mor86] Neil Morgan. A note on the proposed amalgamation of the Lister Institute of Preventive Medicine and the Medical Research Committee: Philanthropy and state support of medical research, 1914. Annals of Science, 43(3):287–289, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morrell:1988:EYG

[Mor88] J. B. Morrell. The early Yorkshire Geological and Polytechnic Society: a reconsideration. Annals of Science, 45(2):153–167, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morgan:1990:SBR

[Mor90] Neil Morgan. The strategy of biological research programmes: Reassessing the ‘dark age’ of biochemistry, 1910–1930. Annals of Science, 47(2):139–150, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morus:2001:HHS

[Mor01] Iwan Rhys Morus. Henry’s house: The Smithsonian Institution and the making of American public science. Annals of Science, 58(4):409– 416, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morris:2002:CMC

[Mor02] Peter John Turnbull Morris, editor. From classical to modern chemistry: the instrumental revolution. Royal Society of Chemistry in association with the Science Museum, Cambridge, 2002. ISBN 0-85404-

479-5. LCCN QD53 .F76 2002. URL http://www.loc.gov/catdir/
enhancements/fy1109/2003446259-d.html;   http://www.loc.
gov/catdir/enhancements/fy1109/2003446259-t.html.

Morange:2008:LE

[Mor08] Michel Morange. Life explained. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-13732-X. xv + 205 pp.

LCCN QH331 .M6313 2008. URL http://www.loc.gov/catdir/
toc/fy0904/2008017943.html.

Morus:2009:PJH

[Mor09] Iwan Rhys Morus. The papers of Joseph Henry. Annals of Science, 66(4):568–570, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Morris:2010:SNP

[Mor10] Peter John Turnbull Morris, editor. Science for the nation: perspectives on the history of the Science Museum. Palgrave Macmillan, Basingstoke, UK, 2010. ISBN 0-230-23009-1 (hardcover). xxi + 350 + 16 pp. LCCN Q105.G7 S36 2010.

Mosley:2014:BRB

[Mos14a] Adam Mosley. Book review: Sacred Words and Worlds: Geography, Religion, and Scholarship, 1550–1700. Annals of Science, 71(4):577– 579, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Moss:2014:BRB

[Mos14b] Michael Moss. Book review: War and the Militarization of British Army Medicine, 1793–1830, by Catherine Kelly. Annals of Science, 71(1):116–118, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mozzochi:2004:ERQ

[Moz04] Charles J. Mozzochi. Essay review: Quest and conquest: Proof of Fermat’s last theorem. Annals of Science, 61(1):119–126, 2004. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Moore-Sheeley:2022:BRI

[MS22] Kirsten Moore-Sheeley. Book review: Immunization: How Vaccines Became Controversial, 2nd ed., by Stuart Blume, London, Reaktion Books, Ltd., 2021, 280 pp., £9.99 (Paperback), ISBN 978-1-78914-504-5. Annals of Science, 79(1):131–133, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Suarez-Diaz:2023:PGM

[MSD23] Gisela Mateos and Edna Suárez-Díaz. The photographers’ gaze: the Mobile Radioisotope Exhibition in Latin America (1960–1965). Annals of Science, 80(1):62–76, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2168061.

Marmoy:1972:APB

[MT72] C. F. A. Marmoy, F.L.A. and John L. Thornton, F.L.A. The anatomical and physiological bibliography of George E. Day (1815–1872). Annals of Science, 28(3):285–291, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mucchielli:1998:OPU

[Muc98] Laurent Mucchielli. Aux origines de la psychologie universitaire en France (1870–1900): enjeux intellectuels, contexte politique, réseaux et stratégies d’alliance autour de la Revue philosophique de Théodule Ribot. Annals of Science, 55(3):263–289, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mulligan:1973:ALS

[Mul73] Lotte Mulligan, Ph.D. Anglicanism, latitudinarianism and science in seventeenth century England. Annals of Science, 30(2):213–219, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Multhauf:1976:HMA

[Mul76] Robert P. Multhauf. A history of magnesia alba. Annals of Science, 33(2):197–200, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Mulligan:1992:RHM

[Mul92] Lotte Mulligan. Robert Hooke’s Memoranda: Memory and natural history. Annals of Science, 49(1):47–61, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Moreno:2002:RDS

[MVK02] Roberto Moreno, Koenraad Van Cleempoel, and David King. A recently discovered sixteenth-century Spanish astrolabe. Annals of Science, 59(4):331–362, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Muller-Wille:2013:SHL

[MW13] S. Müller-Wille. Systems and how Linnaeus looked at them in retrospect. Annals of Science, 70(3):305–317, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Muller-Wille:2007:LSM

[MWO07] S. Müller-Wille and V. Orel. From Linnaean species to Mendelian factors: Elements of hybridism, 1751–1870. Annals of Science, 64(2): 171–215, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nakajima:1984:TKM

[Nak84] Hideto Nakajima. Two kinds of modification theory of light: Some new observations on the Newton–Hooke controversy of 1672 concerning the nature of light. Annals of Science, 41(3):261–278, 1984. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nappi:2009:IIL

[Nap09] Carla Nappi. Inventing the indigenous: Local knowledge and natural history in Early Modern Europe. Annals of Science, 66(4):566–568, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nauenberg:2014:BLG

[Nau14] Michael Nauenberg. Barrow, Leibniz and the geometrical proof of the Fundamental Theorem of the Calculus. Annals of Science, 71(3): 335–354, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nauenberg:2019:VNA

[Nau19] Michael Nauenberg. Visiting Newton’s atelier before the Principia, 1679–1684. Annals of Science, 76(1):1–16, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Naylor:1976:GSP

[Nay76] R. H. Naylor. Galileo: the search for the parabolic trajectory. Annals of Science, 33(2):153–172, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Naylor:1977:GTM

[Nay77] R. H. Naylor. Galileo’s theory of motion: Processes of conceptual change in the period 1604–1610. Annals of Science, 34(4):365–392, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Naylor:1980:REG

[Nay80] R. H. Naylor. The role of experiment in Galileo’s early work on the law of fall. Annals of Science, 37(4):363–378, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Naylor:1982:GLF

[Nay82] R. H. Naylor. Galileo’s law of fall: Absolute truth or approximation. Annals of Science, 39(4):384–389, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Naylor:1983:GEE

[Nay83] R. H. Naylor. Galileo’s early experiments on projectile trajectories. Annals of Science, 40(4):391–395, July 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1950:CRJ

[Nea50] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. I. The journal and its editors. Annals of Science, 6(4):416–421, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1951:CRJa

[Nea51a] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. II. The phlogiston theory. Annals of Science, 7(1):101–106, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1951:CRJb

[Nea51b] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. III. Pierre Bayen. Annals of Science, 7(2):144–148, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1951:CRJc

[Nea51c] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. IV. C. L. Berthollet. V. [Unknown]. Annals of Science, 7(3):284– 299, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1951:CRJd

[Nea51d] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. VI and VII. Annals of Science, 7(4):393–400, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neave:1952:CRJ

[Nea52] E. W. J. Neave, M.C., M.Sc., Ph.D. Chemistry in Rozier’s journal. VIII and IX. Annals of Science, 8(1):28–45, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Newbury:1938:HCS

[New38] N. F. Newbury, M.A., M.Sc., A.I.C. The history of the common salt industry on Merseyside. Annals of Science, 3(1):138–148, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Newth:1952:LLI

[New52] D. R. Newth, B.Sc., Ph.D. Lamarck in 1800: a lecture on the invertebrate animals and a note on fossils taken from the Système des Animaux sans Vertèbres by J. B. Lamarck. Annals of Science, 8(3): 229–254, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Newman:1996:ASR

[New96] William R. Newman. The alchemical sources of Robert Boyle’s corpuscular philosophy. Annals of Science, 53(6):567–585, 1996. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nuttall:1973:MJL

[NF73] R. H. Nuttall, Ph.D. and A. Frank. Makers of jewel lenses in Scotland in the early nineteenth century. Annals of Science, 30(4):407–416, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nieto-Galan:1997:CPC

[NG97] Agustí Nieto-Galan. Calico printing and chemical knowledge in Lancashire in the early nineteenth century: the life and ‘colours’ of John Mercer. Annals of Science, 54(1):1–28, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nieto-Galan:2009:LFS

[NG09] Agustí Nieto-Galan. ‘… not fundamental in a state of full civilization’: The Sociedad Astronómica de Barcelona (1910–1921) and its popularization programme. Annals of Science, 66(4):497–528, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Needham:1975:PTM

[NGDP75] Joseph Needham, Lu Gwei-Djen, and Manfred Porkert. Problems of translation and modernisation of ancient Chinese technical terms. Annals of Science, 32(5):491–502, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicolson:1939:EAN

[Nic39] Marjorie Nicolson, Ph.D., Litt.D. English almanacs and the ‘new astronomy’. Annals of Science, 4(1):1–33, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nickelsen:2006:CCE

[Nic06] Kärin Nickelsen. The challenge of colour: Eighteenth-century botanists and the hand-colouring of illustrations. Annals of Science, 63(1):3–23, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicholson:2011:GRE

[Nic11a] Daniel J. Nicholson. Genesis redux: Essays in the history and philosophy of artificial life. Annals of Science, 68(1):136–139, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicolson:2011:AHB

[Nic11b] Malcolm Nicolson. Alexander von Humboldt and the botanical exploration of the Americas. Annals of Science, 68(3):440–441, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicholson:2013:BRB

[Nic13] Daniel J. Nicholson. Book review: Life Explained, by Michel Morange. Annals of Science, 70(4):554–557, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicholson:2014:BRB

[Nic14] Daniel J. Nicholson. Book review: Knowledge of Life. Annals of Science, 71(3):434–437, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nielsen:2010:WNB

[Nie10] Kristian Hvidtfelt Nielsen. We have never been scientists. Annals of Science, 67(4):561–567, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nitsch:1996:KGS

[Nit96] Edgar Nitsch. Keuper 1820–34: Geburt eines stratigraphischen Begriffes. (German) [Keuper 1820–34: Birth of a stratigraphic term]. Annals of Science, 53(5):489–500, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicolson:1937:SBS

[NM37a] Marjorie Nicolson, Litt.D., Ph.D. and Nora M. Mohler, Ph.D. The scientific background of Swift’s Voyage to Laputa. Annals of Science, 2(3):299–334, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nicolson:1937:SFI

[NM37b] Marjorie Nicolson, Litt.D., Ph.D. and Nora M. Mohler, Ph.D. Swift’s ‘flying island’ in the Voyage to Laputa. Annals of Science, 2(4):405– 430, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Newbury:2006:ACL

[NNS^{+}06] B. D. Newbury, M. R. Notis, B. Stephenson, G. S. Cargill III, and G. B. Stephenson. The astrolabe craftsmen of Lahore and early brass metallurgy. Annals of Science, 63(2):201–213, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nofre:2014:MTE

[Nof14] David Nofre. Managing the technological edge: The UNESCO International Computation Centre and the limits to the transfer of computer technology, 1946–61. Annals of Science, 71(3):410–431, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

North:1942:PCR

[Nor42] F. J. North, D.Sc., F.G.S. Paviland cave, the ‘Red Lady’, the Deluge, and William Buckland. Annals of Science, 5(2):91–128, 1942. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Norton:1975:MPG

[Nor75] B. Norton. Metaphysics and population genetics: Karl Pearson and the background to Fisher’s multi-factorial theory of inheritance. Annals of Science, 32(6):537–553, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

North:1982:O

[Nor82] J. D. North. Obituary. Annals of Science, 39(2):115–116, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nordlund:2001:GWN

[Nor01] Christer Nordlund. ’on going up in the world’: Nation, region and the land elevation debate in Sweden. Annals of Science, 58(1):17– 50, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nordlund:2012:BRI

[Nor12] Christer Nordlund. Book review: Institutions: Education, libraries, museums: Håkon With Andersen, Britta Brenna, Magne Njåstad, and Astrid Wale, Æemula lauri: The Royal Norwegian Society of Sciences and Letters, 1760–2010, ISBN-13 978-0-88135-383-9. Annals of Science, 69(4):589–590, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Norris:2016:RDA

[Nor16] Jacob Norris. Reading Darwin in Arabic, 1860–1950. Annals of Science, 73(3):333–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nothaft:2017:ACR

[Not17] C. Philipp E. Nothaft. Astronomy and calendar reform at the curia of Pope Clement VI: a new source. Annals of Science, 74(1):1–24, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nothaft:2018:HBT

[Not18] C. Philipp E. Nothaft. Henry Bate’s Tabule Machlinenses:the earliest astronomical tables by a Latin author. Annals of Science, 75(4):275– 303, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nothaft:2021:GMT

[Not21] C. Philipp E. Nothaft. Guillaume des Moustiers’ treatise on the armillary instrument (1264) and the practice of astronomical observation in medieval Europe. Annals of Science, 78(4):401–417, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neugebauer:1952:HAN

[NS52] Professor O. Neugebauer, Ph.D., Ll.D. and Olaf Schmidt, Ph.D. Hindu astronomy at Newminster in 1428. Annals of Science, 8(3): 221–228, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Neville:1981:MBC

[NS81] Roy G. Neville and W. A. Smeaton. Macquer’s Dictionnaire de Chymie: a bibliographical study. Annals of Science, 38(6):613–662, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nutton:2014:BRB

[Nut14] Vivian Nutton. Book review: Regimens of the Mind. Boyle, Locke and the Early Modern Cultura Animi Tradition. Annals of Science, 71(4):581–583, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nutton:2021:BRC

[Nut21] Vivian Nutton. Book review: Cutting words: polemical dimensions of Galen’s anatomical experiments. Studies in Ancient Medicine 55, by Luis Alejandro Salas, Leiden and Boston, Brill, 2020, x + 328 pp., EUR 119.00 (Hardback); $143.00, ISBN 978-90-04-43918-4. Annals of Science, 78(2):248–250, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nye:1981:BAA

[Nye81] Mary Jo Nye. Berthelot’s anti-atomism: a ‘matter of taste’? Annals of Science, 38(5):585–590, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nye:1989:CEP

[Nye89] Mary Jo Nye. Chemical explanation and physical dynamics: Two research schools at the First Solvay Chemistry Conferences, 1922– 1928. Annals of Science, 46(5):461–480, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nye:2011:MPH

[Nye11] Mary Jo Nye. Michael Polanyi and his generation: origins of the social construction of science. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-61063-2 (hardcover). xxi + 405 pp. LCCN B945.P584 N94 2011.

Nye:2018:MSB

[Nye18] Robert A. Nye. Masculinity and science in Britain, 1830–1918. Annals of Science, 75(3):266–268, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Nyhart:2009:MNR

[Nyh09] Lynn K. Nyhart. Modern nature: the rise of the biological perspective in Germany. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-61089-6 (hardcover). xiv + 423 pp.

LCCN QH540.83.G3 N94 2009. URL http://www.loc.gov/catdir/
toc/ecip0822/2008026875.html.

Nyhart:2016:ERI

[Nyh16] Lynn K. Nyhart. Essay review: Interpreting visual cultures of science. Annals of Science, 73(4):442–446, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

OBrien:1965:SHIa

[O’B65a] John J. O’Brien, M.A.B.Sc. Samuel Hartlib’s influence on Robert Boyle’s scientific development Part I. The Stalbridge period. Annals of Science, 21(1):1–14, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

OBrien:1965:SHIb

[O’B65b] John J. O’Brien, M.A.B.Sc. Samuel Hartlib’s influence on Robert Boyle’s scientific development. Part II. Boyle in Oxford. Annals of Science, 21(4):257–276, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ockenden:1937:GBL

[Ock37] L. C. Ockenden, B.Sc. The great batteries of the London institution. Annals of Science, 2(2):183–184, April 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Orozco-Echeverri:2021:MCN

[OEMB21] Sergio H. Orozco-Echeverri and Sebastián Molina-Betancur. A mestizo cosmographer in the New Kingdom of Granada: astronomy and chronology in Sánchez de Cozar Guanientá’s Tratado (c. 1696). Annals of Science, 78(3):295–333, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olby:1968:ERM

[OG68] Robert Olby, M.A., D.Phil. and Peter Gautrey. Eleven references to Mendel before 1900. Annals of Science, 24(1):7–20, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ogilvie:1975:CHC

[Ogi75] Marilyn Bailey Ogilvie. Caroline Herschel’s contributions to astronomy. Annals of Science, 32(2):149–161, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olby:1985:EI

[OH85] R. C. Olby and M. J. S. Hodge. Editors’ introduction. Annals of Science, 42(3):187–188, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1997:GJC

[OH97] David Oldroyd and Beryl Hamilton. Geikie and Judd, and controversies about the igneous rocks of the Scottish Hebrides: Theory, practice, and power in the geological community. Annals of Science, 54(3):221–268, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ohnesorge:2023:PPP

[Ohn23] Miguel Ohnesorge. The promises and pitfalls of precision: random and systematic error in physical geodesy, c. 1800–1910. Annals of Science, 81(1–2):258–284, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00033790.2023.2284335.

Oldroyd:1996:BFW

[OJY96] David Oldroyd and Yang Jing-Yi. On being the first Western geologist in China: The work of Raphael Pumpelly (1837–1923). Annals of Science, 53(2):107–136, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:2003:BWG

[OJY03] David Oldroyd and Yang Jing-Yi. Bailey Willis (1857–1949): Geological theorizing and Chinese geology. Annals of Science, 60(1): 1–37, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olby:1989:SBE

[Olb89] Robert Olby. Scientists and bureaucrats in the establishment of the John Innes horticultural institution under William Bateson. Annals of Science, 46(5):497–510, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olby:1991:SIS

[Olb91] Robert Olby. Social imperialism and state support for agricultural research in Edwardian Britain. Annals of Science, 48(6):509–526, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1972:EDC

[Old72a] D. R. Oldroyd, M.A., M.Sc. Edward Daniel Clarke, 1769–1822, and his rôle in the history of the blow-pipe. Annals of Science, 29(3):213– 235, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1972:NCC

[Old72b] D. R. Oldroyd, M.A., M.Sc. Nineteenth century controversies concerning the mesozoic/tertiary boundary in New Zealand. Annals of Science, 29(1):39–57, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1972:TLK

[Old72c] D. R. Oldroyd, M.A., M.Sc. Two little known copies of Black’s lecture notes. Annals of Science, 29(1):35–37, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1974:SPM

[Old74] D. R. Oldroyd. Some phlogistic mineralogical schemes, illustrative of the evolution of the concept of ‘earth’ in the 17th and 18th centuries. Annals of Science, 31(4):269–305, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1980:SAG

[Old80] D. R. Oldroyd. Sir Archibald Geikie (1835–1924), geologist, romantic aesthete, and historian of geology: The problem of Whig historiography of science. Annals of Science, 37(4):441–462, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1991:ACB

[Old91] D. R. Oldroyd. The Archaean controversy in Britain: Part I. The Rocks of St David’s. Annals of Science, 48(5):407–452, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1992:ACB

[Old92] D. R. Oldroyd. The Archaean controversy in Britain: Part II. The Malverns and Shropshire. Annals of Science, 49(5):401–460, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1993:ACB

[Old93] D. R. Oldroyd. The Archaean controversy in Britain: Part III. The rocks of Anglesey and Caernarvonshire. Annals of Science, 50(6): 523–584, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1994:ACB

[Old94] D. R. Oldroyd. The Archaean controversy in Britain: Part IV. Some general theoretical and social issues. Annals of Science, 51(6):571– 592, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1997:SYB

[Old97] David Oldroyd. Some youthful beliefs of Sir Archibald Geikie, PRS, and the first publication of his On the study of the sciences. Annals of Science, 54(1):69–86, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1999:EIA

[Old99a] David Oldroyd. Early ideas about glaciation in the English Lake District: The problem of making sense of glaciation in a glaciated region. Annals of Science, 56(2):175–203, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1999:NWS

[Old99b] David Oldroyd. Non-written sources in the study of the history of geology: Pros and cons, in the light of the views of Collingwood and Foucault. Annals of Science, 56(4):395–415, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:2009:SB

[Old09] David Oldroyd. A Skeleton’s ‘biography’. Annals of Science, 66(4): 553–558, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:2010:GRE

[Old10] David Oldroyd. Geohistory revisited and expanded. Annals of Science, 67(2):249–259, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olesko:2023:HCE

[Ole23] Kathryn M. Olesko. Helmholtz and the conservation of energy: contexts of creation and reception. Annals of Science, 80(1):78–81, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2150307.

Oliver:1969:WBC

[Oli69] J. Oliver, B.A., Ph.D. William Borlase’s contribution to eighteenthcentury meteorology and climatology. Annals of Science, 25(4):275– 317, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olson:1969:NLC

[Ols69] Richard G. Olson, A.M., Ph.D. A note on Leslie’s cube in the study of radiant heat. Annals of Science, 25(3):203–208, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olson:1970:CRS

[Ols70] Richard G. Olson, A.M., Ph.D. Count Rumford, Sir John Leslie, and the study of the nature and propagation of heat at the beginning of the nineteenth century. Annals of Science, 26(4):273–304, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Olson:1971:GCD

[Ols71] Richard G. Olson, A.M., Ph.D. The Gould controversy at Dudley Observatory: Public and professional values in conflict. Annals of Science, 27(3):265–276, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldroyd:1995:NAR

[OM95] D. R. Oldroyd and G. McKenna. A note on Andrew Ramsay’s unpublished report on the St David’s area, recently discovered. Annals of Science, 52(2):193–196, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oldfield:1951:HOM

[OO51] R. C. Oldfield, M.A. and Lady Kathleen Oldfield, M.A. Hartley’s Observations on Man. Annals of Science, 7(4):371–381, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Orange:1975:ITB

[Ora75] A. D. Orange. The idols of the theatre: The British Association and its early critics. Annals of Science, 32(3):277–294, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Orr:2014:FDA

[Orr14] Mary Orr. Fish with a different angle: The Fresh-Water Fishes of Great Britain, by Mrs Sarah Bowdich (1791–1856). Annals of Science, 71(2):206–240, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oedingen:1997:FCD

[OS97] Christina Oedingen and Joseph W. Staerk. First cure for diphtheria by antitoxin as early as 1891. Annals of Science, 54(6):607–610, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Osman:1958:AVI

[Osm58] W. A. Osman, M.Sc., A.R.C.S. Alessandro Volta and the inflammable-air eudiometer. Annals of Science, 14(4):215–242, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ospovat:1980:IRG

[Osp80] Alexander M. Ospovat. The importance of regional geology in the geological theories of Abraham Gottlob Werner: a contrary opinion. Annals of Science, 37(4):433–440, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Oster:1992:SCR

[Ost92] Malcolm Oster. The scholar and the craftsman revisited: Robert Boyle as aristocrat and artisan. Annals of Science, 49(3):255–276, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Otsubo:2005:BTW

[Ots05] Sumiko Otsubo. Between two worlds: Yamanouchi Shigeo and eugenics in early twentieth-century Japan. Annals of Science, 62(2): 205–231, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ottaviani:2011:F

[Ott11] Jim Ottaviani. Feynman. First Second, New York, NY, USA, 2011. ISBN 1-59643-259-4. ???? pp. LCCN QC16.F49 O88 2011. Illustrated by Leland Myrick.

Owen:1955:DE

[Owe55] G. E. Owen, Ph.B., Ph.D. The discovery of the electron. Annals of Science, 11(2):173–182, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ozden:2014:PMW

[Ozd14] Canay Ozden. The Pontifex Minimus: William Willcocks and engineering British colonialism. Annals of Science, 71(2):183–205, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Page:2003:LEB

[Pag03] Frederick G. Page. Lime in the early bleaching industry of Britain 1633–1828: Its prohibition and repeal. Annals of Science, 60(2):185– 200, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Palter:1983:ERH

[Pal83] Robert M. Palter. Essay review: History, philosophy, and physics in Einstein’s special relativity paper: Arthur Miller, Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–1911). Annals of Science, 40(6):657–662, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). See [Mil81a].

Palmieri:2012:BRE

[Pal12] Paolo Palmieri. Book review: Editions and selections: William R. Shea and Tiziana Bascelli, editors, Galileo’s Sidereus Nuncius or Sidereal Message, ISBN-13 978-0-88135-375-4. Annals of Science, 69(4):586–587, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Palladino:2013:BBH

[Pal13] Paolo Palladino. Biopolitics between history and philosophy. Annals of Science, 70(3):405–415, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Palomo:2021:NIO

[Pal21] Miguel Palomo. New insight into the origins of the calculus war. Annals of Science, 78(1):22–40, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pal:2023:LRI

[Pal23] Carol Pal. Lady Ranelagh: the incomparable life of Robert Boyle’s sister. Annals of Science, 80(3):296–299, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2196981.

Pantalony:2005:RKW

[Pan05] David Pantalony. Rudolph Koenig’s workshop of sound: Instruments, theories, and the debate over combination tones. Annals of Science, 62(1):57–82, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pantalony:2009:ASR

[Pan09a] David Pantalony, editor. Altered sensations: Rudolph Koenig’s acoustical workshop in nineteenth-century Paris, volume 24 of Archimedes, 1385-0180. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 90-481-2815-3 (print), 90-481-2816-1 (e-book). xxxvi + 372 pp.

LCCN ML3805 .P26 2009. URL http://www.loc.gov/catdir/
enhancements/fy1006/2009928017-d.html;   http://www.loc.
gov/catdir/enhancements/fy1006/2009928017-t.html.

Pantalony:2009:WMH

[Pan09b] David Pantalony. The Whipple Museum of the History of Science: Instruments and interpretations, to celebrate the 60th anniversary of R. S. Whipple’s gift to the University of Cambridge. Annals of Science, 66(2):277–279, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Papanelopoulou:2009:IAR

[Pap09] Faidra Papanelopoulou. The International Association of Refrigeration through the correspondence of Heike Kamerlingh Onnes and Charles-Édouard Guillaume, 1908–1914. Annals of Science, 66(3): 345–370, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1936:JBV

[Par36] J. R. Partington, M.B.E., D.Sc. Joan Baptista van Helmont. Annals of Science, 1(4):359–384, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1939:OAT

[Par39] J. R. Partington, M.B.E., D.Sc. The origins of the atomic theory. Annals of Science, 4(3):245–282, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1942:EHS

[Par42] J. R. Partington, M.B.E., D.Sc. The early history of strontium. Annals of Science, 5(2):157–166, 1942. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1945:SMS

[Par45] J. R. Partington, M.B.E., D.Sc. Some milestones in the study of hydrocarbon flames. Annals of Science, 5(3):229–252, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1949:TT

[Par49] J. R. Partington, M.B.E., D.Sc. Thomas Thomson, 1773–1852. Annals of Science, 6(2):115–126, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1951:EHS

[Par51a] J. R. Partington, M.B.E., D.Sc. The early history of strontium. Part II. Annals of Science, 7(1):95–100, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1951:JBR

[Par51b] J. R. Partington, M.B.E., D.Sc. Jeremias Benjamin Richter and the law of reciprocal proportions. I. Annals of Science, 7(2):173– 198, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1953:JBR

[Par53a] J. R. Partington, M.B.E., D.Sc. Jeremias Benjamin Richter and the Law of Reciprocal Proportions. II. Annals of Science, 9(4):289– 314, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1953:LMC

[Par53b] J. R. Partington, M.B.E., D.Sc. Lavoisier’s memoir on the composition of nitric acid. Annals of Science, 9(1):96–98, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1955:LN

[Par55] J. R. Partington, M.B.E., D.Sc. Lignum nephriticum. Annals of Science, 11(1):1–26, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Parshall:1984:EHM

[Par84] Karen Hunger Parshall. Eliakim Hastings Moore and the founding of a mathematical community in America, 1892–1902. Annals of Science, 41(4):313–333, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Park:2003:HQC

[Par03] Buhm Soon Park. The ‘hyperbola of quantum chemistry’: the changing practice and identity of a scientific discipline in the early years of electronic digital computers, 1945–65. Annals of Science, 60(3):219–247, 2003. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.ingentaconnect.com/
content/tandf/tasc/2003/00000060/00000003/art00001.

Parsons:2013:BRS

[Par13] Thad Parsons. Book review: Sam Alberti, Nature and Culture: Objects, Disciplines and the Manchester Museum. Manchester and New York: Manchester University Press, 2009. xiii + 239 pp. 28 figures. $89.95 / £60.00. ISBN-13 978-0-7190-8114-9. Annals of Science, 70 (1):111–114, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Park:2016:CFL

[Par16] Hyung Wook Park. Constructing failure: Leonard Hayflick, biomedicine, and the problems with tissue culture. Annals of Science, 73(3):303–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Passer:1951:ESE

[Pas51] Harold C. Passer, M.A., Ph.D. Electrical science and the early development of the electrical manufacturing industry in the United States. Annals of Science, 7(4):382–392, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1936:RBE

[Pat36a] T. S. Patterson, D.Sc., Ph.D. Richard Boyle, Esq., M.P., F.R.S., and — incidentally — some of his relatives. Annals of Science, 1(1): 62–65, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1936:VHI

[Pat36b] T. S. Patterson, D.Sc., Ph.D. Van Helmont’s ice and water experiments. Annals of Science, 1(4):462–467, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1937:JBH

[Pat37] T. S. Patterson, D.Sc., Ph.D. Jean Beguin and his Tyrocinium Chymicum [Rudiments of chemistry]. Annals of Science, 2(3):243– 298, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1938:DOR

[Pat38] T. S. Patterson, D.Sc., Ph.D. The discovery of the optical rotatory power of tartaric acid. Annals of Science, 3(4):431–434, 1938. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1939:NBV

[Pat39] T. S. Patterson, D.Sc., Ph.D. Note on Blaise de Vigenère, John Ferguson, and benzoic acid. Annals of Science, 4(1):61–64, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patton:2017:BRS

[Pat17] Maryam M. Patton. Book review: Success and Suppression: Arabic Sciences and Philosophy in the Renaissance. Annals of Science, 74 (2):167–169, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Paul:1978:AWW

[Pau78] E. Robert Paul. Alexander W. Williamson on the atomic theory: a study of nineteenth-century British atomism. Annals of Science, 35 (1):17–31, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pavuk:2017:BEG

[Pav17] Alexander Pavuk. Biologist Edwin Grant Conklin and the idea of the religious direction of human evolution in the early 1920s. Annals of Science, 74(1):64–82, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Payne:2018:PRS

[Pay18] Christiana Payne. The pre-Raphaelites and science. Annals of Science, 75(4):373–375, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1945:HOC

[PB45] T. S. Patterson, D.Sc., Ph.D., Ll.D. and Charles Buchanan, B.Sc., Ph.D. Historical and other considerations regarding the crystal form of sodium-ammonium d — and l-tartrate, potassium d — and ltartrate, potassium-ammonium d — and l-tartrate, and potassium racemate — I. Annals of Science, 5(3):288–295, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1947:HOC

[PB47] T. S. Patterson, D.Sc., Ph.D., Ll.D. and Charles Buchanan, B.Sc., Ph.D. Historical and other considerations regarding the crystal form of sodium-ammonium d — and l-tartrate, potassium d — and ltartrate, potassium-ammonium d — and l-tartrate, and potassium racemate. II. Annals of Science, 5(4):317–324, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1948:HOC

[PB48] T. S. Patterson, D.Sc., Ph.D., Ll.D. and Charles Buchanan, B.Sc., Ph.D. Historical and other considerations regarding the crystal form of sodium-ammonium d — and l-tartrate, potassium d — and l-tartrate, potassium-ammonium d — and l-tartrate, and potassium racemate. Addendum on the isomorphism of potassium- and potassium-ammonium tartrate. Annals of Science, 6(1):76–77, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pedersen:2010:OOJ

[PdC10] Kurt Møller Pedersen and P. R. (Peter Richard) de Clercq, editors. An observer of observatories: the journal of Thomas Bugge’s tour of Germany, Holland and England in 1777. Aarhusuniversitetsforlag, Århus, Danmark, 2010. ISBN 87-7934-311-2. xxiii + 220 pp. LCCN QB36.B94 A313 2010.

Pedersen:1975:AWH

[Ped75] Kurt Møller Pedersen. Argumente für und wider das heliozentrische Weltbild. (German) [Arguments for and against the heliocentric world view]. Annals of Science, 32(2):163–167, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pelseneer:1936:NSU

[Pel36] J. Pelseneer. Notes on some unpublished letters from Faraday to Quetelet. Annals of Science, 1(4):447–452, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pelis:1997:BCN

[Pel97] Kim Pelis. Blood clots: the nineteenth-century debate over the substance and means of transfusion in Britain. Annals of Science, 54(4): 331–360, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Peoux:2010:ARR

[Péo10] Gérald Péoux. Atmospheric refraction and the Ramus circle: Aspects of a late sixteenth-century dispute. Annals of Science, 67(4):457– 484, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pepper:1976:HMT

[Pep76] Jon V. Pepper. Harriot’s manuscript on the theory of impacts. Annals of Science, 33(2):131–151, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Perrin:1983:JBA

[Per83] Carleton E. Perrin. Joseph Black and the absolute levity of phlogiston. Annals of Science, 40(2):109–137, March 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Perrin:1986:TSI

[Per86] Carleton E. Perrin. Of theory shifts and industrial innovations: The relations of J. A. C. Chaptal and A. L. Lavoisier. Annals of Science, 43(6):511–542, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Petsche:2014:CSS

[Pet14] Hans-Joachim Petsche. The ‘chemistry of space’: The sources of Hermann Grassmann’s scientific achievements. Annals of Science, 71(4):522–576, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Piazza:2021:GOP

[Pia21] Vincent Roy-Di Piazza. ‘Ghosts from other planets’: plurality of worlds, afterlife and satire in Emanuel Swedenborg’s De Telluribus in mundo nostro solari (1758). Annals of Science, 77(4):469–494, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pisano:2013:BRB

[Pis13] Raffaele Pisano. Book review: The English Galileo: Thomas Harriot’s Work on Motion as an Example of Preclassical Mechanics. Volume 1: The interpretation. Volume 2: Sources, by Matthias Schemmel. Annals of Science, 70(4):585–588, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pisano:2014:BRBb

[Pis14a] Raffaele Pisano. Book review: History of Virtual Work Laws. A History of Mechanics Prospective. Annals of Science, 71(3):441–442, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pisano:2014:BRBa

[Pis14b] Raffaele Pisano. Book review: Honoré Fabri and the Concept of Impetus: A Bridge between Conceptual Frameworks, by Michael Elazar. Annals of Science, 71(2):276–277, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pappas:1987:PBP

[PK87] Vasilis Pappas and Ioannis Karas. The printed book of physics: The dissemination of scientific thought in Greece 1750–1821 before the Greek revolution. Annals of Science, 44(3):237–244, May 1987. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Patterson:1948:RVB

[PLC48] T. S. Patterson, D.Sc., Ph.D., Ll.D., J. D. Loudon, D.Sc., Ph.D., and Adeline O. M. Cook, M.A. Robertus Vallensis’ De veritate et antiquitate artis chemicæ [The authentic and ancient art of chemistry]. Annals of Science, 6(1):1–23, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Plotkin:1978:ECP

[Plo78] Howard Plotkin. Edward C. Pickering, the Henry Draper Memorial, and the beginnings of astrophysics in America. Annals of Science, 35(4):365–377, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Plotkin:1980:HTF

[Plo80] Howard Plotkin. Henry Tappan, Franz Brünnow, and the founding of the Ann Arbor School of Astronomers, 1852–1863. Annals of Science, 37(3):287–302, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Plomp:1999:LTI

[Plo99] Reinier Plomp. A longitude timekeeper by Isaac Thuret with the balance spring invented by Christiaan Huygens. Annals of Science, 56(4):379–394, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Plofker:2011:YIT

[Plo11] Kim Plofker. ‘Yavana’ and ‘Indian’: Transmission and foreign identity in the exact sciences. Annals of Science, 68(4):467–476, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1937:HSP

[PM37] J. R. Partington, M.B.E., D.Sc. and Douglas McKie, D.Sc., Ph.D. Historical studies on the phlogiston theory. I. The levity of phlogiston. Annals of Science, 2(4):361–404, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1938:HSPa

[PM38a] J. R. Partington, M.B.E., D.Sc. and Douglas McKie, D.Sc., Ph.D. Historical studies on the phlogiston theory. II. The negative weight of phlogiston. Annals of Science, 3(1):1–58, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1938:HSPb

[PM38b] J. R. Partington, M.B.E., D.Sc. and Douglas McKie, D.Sc., Ph.D. Historical studies on the phlogiston theory. III. Light and heat in combustion. Annals of Science, 3(4):337–371, 1938. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1939:HSP

[PM39] J. R. Partington, M.B.E., D.Sc. and Douglas McKie, D.Sc., Ph.D. Historical studies on the phlogiston theory. IV. Last phases of the theory. Annals of Science, 4(2):113–149, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:1950:SJE

[PM50] J. R. Partington, M.B.E., D.Sc. and Douglas McKie, D.Sc., Ph.D. Sir John Eliot, Bart. (1736–86), and John Elliot (1747–87). Annals of Science, 6(3):262–267, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Papanelopoulou:2009:PST

[PNGP09] Faidra Papanelopoulou, Agustí Nieto-Galan, and Enrique Perdiguero, editors. Popularizing science and technology in the European periphery, 1800–2000. Science, technology, and culture, 1700-1945. Ashgate Publishers, Burlington, VT, 2009. ISBN 0-7546-6269-1 (hardcover). xviii + 284 pp. LCCN Q225.2.E85 P67 2009. URL http://www.

loc.gov/catdir/toc/ecip0819/2008023647.html.

Pomata:2011:WEA

[Pom11] Gianna Pomata. A word of the empirics: The ancient concept of observation and its recovery in early modern medicine. Annals of Science, 68(1):1–25, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Popplow:2012:KME

[Pop12] Marcus Popplow. Knowledge management to exploit agrarian resources as part of late-eighteenth-century cultures of innovation: Friedrich Casimir Medicus and Franz von Paula Schrank. Annals of Science, 69(3):413–433, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Porter:1978:GHT

[Por78] Roy Porter. George Hoggart Toulmin’s theory of man and the Earth in the light of the development of British geology. Annals of Science, 35(4):339–352, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Porter:1981:PMA

[Por81] Theodore M. Porter. The promotion of mining and the advancement of science: the chemical revolution of mineralogy. Annals of Science, 38(5):543–570, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Porter:2001:MFP

[Por01] Theodore M. Porter. Modern facts and postmodern interpretations. Annals of Science, 58(4):417–422, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pormann:2010:AAC

[Por10] Peter E. Pormann. Arabic astronomy and the Copernican ‘Revolution’. Annals of Science, 67(2):243–248, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pormann:2011:FAP

[Por11] Peter E. Pormann. The formation of the Arabic pharmacology between tradition and innovation. Annals of Science, 68(4):493–515, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Porter:2016:BRK

[Por16] Theodore M. Porter. Book review: Kuhn’s Structure of Scientific Revolutions at Fifty: Reflections on a Science Classic, Robert J. (Robert John) Richards and Lorraine Daston, University of Chicago Press, Chicago (2016), ISBN 978-0-226-31703-8. Annals of Science, 73(4):450–452, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Potter:1953:ALC

[Pot53] Owen Potter, M.Sc. Auguste Laurent’s contributions to chemistry. Annals of Science, 9(3):271–280, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Potters:2022:CPR

[Pot22] Jan Potters. Conceptualizing paradigms: on reading Kuhn’s history of the quantum. Annals of Science, 79(3):386–405, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2063945.

Pourciau:1991:NPI

[Pou91] Bruce H. Pourciau. On Newton’s proof that inverse-square orbits must be conics. Annals of Science, 48(2):159–172, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pourciau:2001:NTG

[Pou01] Bruce Pourciau. A new translation of and guide to Newton’s Principia. Annals of Science, 58(1):85–91, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Porter:1977:RBG

[PP77] Roy Porter and Kate Poulton. Research in British geology 1660– 1800: a survey and thematic bibliography. Annals of Science, 34(1): 33–42, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Prasad:2010:MAM

[Pra10] Amit Prasad. Magnetic appeal: MRI and the myth of transparency. Annals of Science, 67(4):577–578, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

PresasiPuig:2008:LTS

[Pre08] Albert Presas i Puig. On a lecture trip to Spain: the scientific relations between Germany and Spain during the Entente Boycott (1919– 1926). Annals of Science, 65(4):529–546, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Preston:2016:BRM

[Pre16] Christopher D. Preston. Book review: Methodus Plantarum Nova: John Ray (1682). Annals of Science, 73(1):108–110, 2016. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1064601.

Price:1952:EOI

[Pri52] Derek J. Price, Ph.D. The early observatory instruments of Trinity College, Cambridge. Annals of Science, 8(1):1–12, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Price:1954:CAS

[Pri54] Derek J. Price, Ph.D. A collection of armillary spheres and other antique scientific instruments. Annals of Science, 10(2):172–187, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Prince:1955:SEW

[Pri55] A. Prince, B.Met., A.I.M. The significance of the early work of Fredrik Rudberg on alloy constitution. Annals of Science, 11(1): 58–63, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Price:2010:PRM

[Pri10] Katy Price. Pythagoras’ revenge: a mathematical mystery. Annals of Science, 67(4):583–585, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Principe:2013:SA

[Pri13] Lawrence Principe. The secrets of alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-68295-1 (hardcover), 0-226-92378-9 (e-book). v + 281 + 8 pp. LCCN QD13 .P75 2013.

Pronskikh:2018:BRW

[Pro18] Vitaly Pronskikh. Book review: What Makes a Good Experiment? Reasons and Roles in Science, A. Franklin, Pittsburgh, PA: University of Pittsburgh Press, 2016. 384 pp. $55.00, ISBN 978-0-8229-4441-6. Annals of Science, 75(1):62–64, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Prull:2009:LS

[Prü09] Cay-Rüdiger Prüll. Life sciences. Annals of Science, 66(1):143–145, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pulte:1994:CGJ

[Pul94] Helmut Pulte. C. G. J. Jacobis Vermächtnis einer ‘konventionalen’ analytischen Mechanik: Vorgeschichte, Nachschriften und Inhalt seiner letzten Mechanik-Vorlesung. (German) [C. G. J. Jacobi’s legacy of a conventional analytical mechanics: history, transcripts and content of his last mechanics lecture]. Annals of Science, 51(5): 497–516, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Pumfrey:1987:MMR

[Pum87] Stephen Pumfrey. Mechanizing magnetism in restoration England — the decline of magnetic philosophy. Annals of Science, 44(1):1–21, January 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Partington:2008:HGW

[PW08] John S. Partington and H. G. (Herbert George) Wells. H. G. Wells in Nature, 1893–1946: a reception reader, volume 3 of Alph: Arbeiten zur literarischen Phantastik. Peter Lang, Frankfurt am Main, Germany, 2008. ISBN 3-631-57110-0. ISSN 1864-323X. x + 514 pp. LCCN PR5777 .H265 2008.

Quan:1968:GPC

[Qua68] Stanislaus Quan, M.A. Galileo and the problem of concentric circles: a refutation, and the solution. Annals of Science, 24(4):313–338, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quan:1970:GPI

[Qua70] Stanislaus Quan, M.A. Galileo and the problem of infinity: a refutation, and the solution. Annals of Science, 26(2):115–151, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quan:1972:GPI

[Qua72] Stanislaus Quan, M.A. Galileo and the problem of infinity. Part II. The dialectical arguments, and the solution. Annals of Science, 28(3):237–284, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quan:1974:GTP

[Qua74] Stanislaus Quan. Galileo and the Theorem of Pythagoras. Annals of Science, 31(3):227–261, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quedec:1990:ILN

[Qué90] Pierre Quédec. Itinéraire de Louis Néel des aimants aux grenats ferrimagnétiques. (French) [The route of Louis Néel to magnets of ferrimagnetic garnets]. Annals of Science, 47(3):261–275, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quick:2009:NLD

[Qui09] Tom Quick. On the nature of limbs: a discourse. Annals of Science, 66(2):302–304, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quick:2012:BRB

[Qui12] Tom Quick. Book review: Biographies: Nicolaas A. Rupke, Richard Owen: Biology without Darwin, ISBN-13 978-0-226-73177-3. Annals of Science, 69(4):587–588, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Quick:2013:BRC

[Qui13] Tom Quick. Book review: Christopher E. Cosans, Owen’s Ape and Darwin’s Bulldog: Beyond Darwinism and Creationism. Bloomington, Indiana University Press, 2009. xxi + 162 pp. 29 plts. $21.95. ISBN-13 978-0-253-22051-6. Annals of Science, 70(1):105–107, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rabinovitch:1977:OME

[Rab77] Nachum L. Rabinovitch. The one and the many: Early stochastic reasoning in philosophy. Annals of Science, 34(4):331–344, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rader:2018:SMT

[Rad18] Karen Rader. Science museums in transition: cultures of display in nineteenth-century Britain and America. Annals of Science, 75(3): 270–272, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rampelt:2010:RCS

[Ram10] Jason M. Rampelt. Religion as a cause in scientific research. Annals of Science, 67(1):121–130, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ramberg:2015:CRI

[Ram15] Peter J. Ramberg. Chemical research and instruction in Zürich, 1833–1872: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):170–186, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rancke-Madsen:1963:PHR

[Ran63] E. Rancke-Madsen, Dr.Phil. A painting of the Hon. Robert Boyle in Danish possession. Annals of Science, 19(2):147–148, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Raphael:2012:PGD

[Rap12] Renée J. Raphael. Printing Galileo’s Discorsi: a collaborative affair. Annals of Science, 69(4):483–513, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Raphael:2015:GBT

[Rap15a] Renee Jennifer Raphael. Galileo’s Discorsi as a tool for the analytical art. Annals of Science, 72(1):99–123, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Raposo:2015:TWE

[Rap15b] Pedro M. P. Raposo. Time, weather and empires: The Campos Rodrigues Observatory in Lourenço Marques, Mozambique (1905– 1930). Annals of Science, 72(3):279–305, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Raphael:2017:RGS

[Rap17] Renée Jennifer Raphael. Reading Galileo: scribal technologies and the two new sciences. The Johns Hopkins University Press, Baltimore, MD, USA, 2017. ISBN 1-4214-2177-1 (hardcover), 1-4214-2178-X (electronic). ix + 265 pp. LCCN QB36.G2 R2155 2017.

Rasmussen:2017:LGS

[Ras17] Nicolas Rasmussen. Life’s greatest secret: The race to crack the genetic code. Annals of Science, 74(3):247–248, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ratcliff:2009:QIM

[Rat09] Marc Ratcliff. The quest for the invisible: microscopy in the Enlightenment. Ashgate, Aldershot, Hants, England and Farnham, England, 2009. ISBN 0-7546-6150-4. xii + 315 pp.

LCCN QH204 .R38 2009. URL http://www.loc.gov/catdir/toc/
ecip081/2007039630.html.

Rawson:1974:PDM

[Raw74] Don C. Rawson. The process of discovery: Mendeleev and the periodic law. Annals of Science, 31(3):181–204, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ray:1682:MPN

[Ray82] John Ray. Methodus plantarum nova, brevitatis and perspicuitatis causa synoptice in tabulis exhibita . . .. (Latin) []. Impensis Henrici Faithorne and Joannis Kersey, Londini, 1682. 22 + 166 + 34 pp. LCCN QK41 .R21.

Raynaud:2015:FDS

[Ray15] Dominique Raynaud. Un fragment du De speculis comburentibus de Regiomontanus copié par Toscanelli et inséré dans les carnets de Leonardo (Codex Atlanticus, 611rb/915ra). Annals of Science, 72(3): 306–336, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Raymer:2018:BRD

[Ray18] Emilie J. Raymer. Book review: Darwin and Women: A Selection of Letters, edited by Samantha Evans, New York: Cambridge University Press, 2017, xxviii + 270 pp, ISBN 978-1-107-15886-3. Annals of Science, 75(1):66–67, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robotti:2001:MPC

[RB01] Nadia Robotti and Massimiliano Badino. Max Planck and the ’constants of nature’. Annals of Science, 58(2):137–162, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ruiz-Castell:2008:SIE

[RC08] Pedro Ruiz-Castell. Scientific instruments for education in early Twentieth-Century Spain. Annals of Science, 65(4):519–527, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Richards:2016:KSS

[RD16] Robert J. (Robert John) Richards and Lorraine Daston, editors. Kuhn’s Structure of Scientific Revolutions at fifty: reflections on a science classic. University of Chicago Press, Chicago, IL, USA and London, UK, 2016. ISBN 0-226-31703-X (hardcover), 0-226-31720-X (paperback), 0-226-31717-X (e-book). 202 pp. LCCN Q175 .K84 2016.

Reeds:1976:RHB

[Ree76] Karen Meier Reeds. Renaissance humanism and botany. Annals of Science, 33(6):519–542, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rees:1980:ASF

[Ree80] Graham Rees. Atomism and ‘subtlety’ in Francis Bacon’s philosophy. Annals of Science, 37(5):549–571, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rees:1981:UMF

[Ree81] Graham Rees. An unpublished manuscript by Francis Bacon: Sylva Sylvarum drafts and other working notes. Annals of Science, 38(4): 377–412, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reeves:2008:GGT

[Ree08] Eileen Adair Reeves. Galileo’s glassworks: the telescope and the mirror. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02667-5. 231 pp. LCCN QB85.8 .R44 2008. URL http://www.

loc.gov/catdir/toc/ecip0710/2007005089.html.

Reeves:2014:BRBb

[Ree14a] Carole Reeves. Book review: Nervous Disease in Late Eighteenth Century Britain: The Reality of a Fashionable Disorder / Desperate Housewives, Neuroses and the Domestic Environment, 1945–1970. Annals of Science, 71(4):592–595, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reeves:2014:BRBa

[Ree14b] Carole Reeves. Book review: Scottish Medicine: An Illustrated History. Annals of Science, 71(3):443–445, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Regal:2009:EDR

[Reg09] Brian Regal. Entering dubious realms: Grover Krantz, science, and Sasquatch. Annals of Science, 66(1):83–102, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Regier:2019:BRP

[Reg19] Jonathan Regier. Book review: The pursuit of harmony: Kepler on cosmos, confession, and community, by Aviva Rothman, Chicago, University of Chicago Press, 2017, vii + 355 pp., $55.00 (hbk), ISBN 978-0-226-49697-9. Annals of Science, 76(3–4):375–377, 2019. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reingold:1981:RMU

[Rei81] Nathan Reingold. Refugee mathematicians in the United States of America, 1933–1941: reception and reaction. Annals of Science, 38 (3):313–338, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reid:1989:NLO

[Rei89] John S. Reid. A new look at old linear measures. Annals of Science, 46(3):221–248, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reinhardt:2004:CPM

[Rei04] Carsten Reinhardt. Chemistry in a physical mode: Molecular spectroscopy and the emergence of NMR. Annals of Science, 61(1):1–32, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reich:2014:BRB

[Rei14] Ulrich Reich. Book review: From Cardano’s Great Art to Lagrange’s Reflections: Filling a Gap in the History of Algebra, by Jacqueline Stedall. Annals of Science, 71(2):273–275, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rentetzi:2011:PRS

[Ren11] Maria Rentetzi. Packaging radium, selling science: Boxes, bottles and other mundane things in the world of science. Annals of Science, 68 (3):375–399, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rentetzi:2018:BRA

[Ren18] Maria Rentetzi. Book review: Analogies that shape the recent history of radiation: Strange Glow: The Story of Radiation, by Timothy Jorgensen, Princeton University Press, 2016, 512 pp., $35.00, £27.95, ISBN 978-0-691-16503-5. Radium and the Secret of Life, by Luis Campos, University of Chicago Press, 2015, 352 pp., $35.00, ISBN 978-0-226-41874-2. Annals of Science, 75(1):55–59, 2018. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Germanese:2023:SDD

[RG23] Maria Rentetzi and Donatella Germanese. Science diplomacy on display: mobile atomic exhibitions in the Cold War: Introduction to special issue. Annals of Science, 80(1):1–9, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2166114.

Roll-Hansen:1985:NPL

[RH85] Nils Roll-Hansen. A new perspective on Lysenko? Annals of Science, 42(3):261–278, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Riehm:2011:TTM

[RH11] Elaine McKinnon Riehm and Frances Hoffman. Turbulent times in mathematics: the life of J. C. Fields and the history of the Fields Medal. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-6914-0. xii + 257 + 16 pp. LCCN QA29.F54 R53 2011.

Ribeiro:2020:BHD

[Rib20] Luís Campos Ribeiro. The bounded heavens: defining the limits of astrological practice in the Iberian indices. Annals of Science, 77(1): 50–70, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Richards:1986:DLB

[Ric86] Stewart Richards. Drawing the life-blood of physiology: Vivisection and the physiologists’ dilemma, 1870–1900. Annals of Science, 43(1):

27–56, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ritterbush:1962:JLS

[Rit62] Philip C. Ritterbush, D.Phil. John Lindsay and the sensitive plant. Annals of Science, 18(4):233–254, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ritvo:2011:PAB

[Rit11] Harriet Ritvo. Picturing animals in Britain 1750–1850. Annals of Science, 68(2):286–288, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ramberg:2015:IAO

[RN15] Peter J. Ramberg and Mary Jo Nye. Introduction: Atomism and organic chemistry in context: Essays in honour of Alan J. Rocke. Annals of Science, 72(2):149–152, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reinhardt:1982:KTA

[RO82] O. Reinhardt and D. R. Oldroyd. Kant’s thoughts on the ageing of the Earth. Annals of Science, 39(4):349–369, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reinhardt:1983:KTE

[RO83] O. Reinhardt and D. R. Oldroyd. Kant’s theory of earthquakes and volcanic action. Annals of Science, 40(3):247–272, May 1983. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Reinhardt:1984:AHK

[RO84] O. Reinhardt and D. R. Oldroyd. By analogy with the heavens: Kant’s theory of the Earth. Annals of Science, 41(3):203–221, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1953:DPS

[Rob53a] Eric Robinson, M.A. The Derby Philosophical Society. Annals of Science, 9(4):359–367, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1953:MBP

[Rob53b] Eric Robinson, M.A. Matthew Boulton, patron of the arts. Annals of Science, 9(4):368–376, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1954:EDB

[Rob54a] Eric Robinson, M.A. Erasmus Darwin’s Botanic Garden and contemporary opinion. Annals of Science, 10(4):314–320, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1954:TLP

[Rob54b] Eric Robinson, M.A. Thomas Love Peacock: Critic of scientific progress. Annals of Science, 10(1):69–77, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1954:TCI

[Rob54c] Eric Robinson, M.A. Training captains of industry: The education of Matthew Robinson Boulton [1770–1842] and the younger James Watt [1769–1848]. Annals of Science, 10(4):301–313, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1955:RRF

[Rob55a] Eric Robinson, M.A. R. E. Raspe, Franklin’s ‘Club of thirteen’, and the Lunar Society. Annals of Science, 11(2):142–144, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1955:TBM

[Rob55b] Eric Robinson, M.A. Thomas Beddoes, M.D., and the reform of science teaching in Oxford. Annals of Science, 11(2):137–141, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1956:LSI

[Rob56] Eric Robinson, M.A. The Lunar Society and the improvement of scientific instruments: I. Annals of Science, 12(4):296–304, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1957:LSI

[Rob57] Eric Robinson, M.A. The Lunar Society and the improvement of scientific instruments: II. Annals of Science, 13(1):1–8, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1962:PCE

[Rob62] Eric Robinson, M.A. The profession of civil engineer in the eighteenth century: a portrait of Thomas Yeoman, F.R.S., 1704(?)–1781. Annals of Science, 18(4):195–215, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robinson:1963:BDT

[Rob63] Eric Robinson, M.A. Benjamin Donn (1729–1798), teacher of mathematics and navigation. Annals of Science, 19(1):27–36, 1963. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robison:1974:GMJ

[Rob74] Wade L. Robison. Galileo on the moons of Jupiter. Annals of Science, 31(2):165–169, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robotti:1995:JJT

[Rob95] Nadia Robotti. J. J. Thomson at the Cavendish Laboratory: The history of an electric charge measurement. Annals of Science, 52(3): 265–284, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robson:2010:BRM

[Rob10] Eleanor Robson. Book review: Mathematics and logic. Christine Proust, Tablettes mathématiques de Nippur Varia Anatolica, XVIII, ISBN-13 978-2-906-05392-9. Annals of Science, 67(2):288–290, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Roche:1981:RAE

[Roc81] John J. Roche. The radius astronomicus in England. Annals of Science, 38(1):1–32, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rocke:1985:HEE

[Roc85] Alan J. Rocke. Hypothesis and experiment in the early development of Kekulé’s benzene theory. Annals of Science, 42(4):355–381, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rocke:2010:IRK

[Roc10] Alan J. Rocke. Image and reality: Kekulé, Kopp, and the scientific imagination. Synthesis. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-72332-

1. xxvi + 375 pp. LCCN QD248 .R63 2010. URL http:/

/catdir.loc.gov/catdir/enhancements/fy1012/2009025749-b.
html;   http://catdir.loc.gov/catdir/enhancements/fy1012/
2009025749-d.html;   http://catdir.loc.gov/catdir/enhancements/
fy1012/2009025749-t.html.

Roero:1990:LTV

[Roe90] Clara Silvia Roero. Leibniz and the Temple of Viviani: Leibniz’s prompt reply to the challenge and the repercussions in the field of mathematics. Annals of Science, 47(5):423–443, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rogers:1972:DME

[Rog72] G. A. J. Rogers, Ph.D. Descartes and the method of English science. Annals of Science, 29(3):237–255, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Romer:1978:LAT

[Rom78] Alfred Romer. Liberal arts technology. Annals of Science, 35(4):425– 426, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rommevaux:2006:EBM

[Rom06] Sabine Rommevaux. Un exemple de Question mathématique au Moyen Âge. (French) [An example of mathematical Question of the Middle Ages]. Annals of Science, 63(4):425–445, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ross:1962:SSW

[Ros62] Sydney Ross, B.Sc., Ph.D. Scientist: The story of a word. Annals of Science, 18(2):65–85, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rose:1998:HNA

[Ros98] Paul Lawrence Rose. Heisenberg and the Nazi atomic bomb project: a study in German culture. University of California Press, Berkeley, CA, USA, 1998. ISBN 0-520-21077-8, 0-585-32190-6 (e-book). xx + 352 pp. LCCN QC16.H35 R67 1998. URL http://preterhuman.

net/texts/religion.occult.new_age/occult.conspiracy.and.
related/Rose,%20Paul%20Lawrence%20-%20Heisenberg%20and%20the%
20Nazi%20Atomic%20Bomb%20Project.pdf.

Rossiianov:2021:PLW

[Ros21] Kirill Rossiianov. The problem of Lysenkoism: why we cannot explain it away? Annals of Science, 78(1):117–125, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rothery:1968:AC

[Rot68] E. J. Rothery, B.Sc. Æneas Coffey (1780–1852). Annals of Science, 24(1):53–71, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Roux:2022:MVL

[Rou22] Sophie Roux. Mechanism. A visual, lexical and conceptual history. Annals of Science, 79(3):411–413, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2022.2078506.

Rowbottom:1950:EPW

[Row50a] Margaret E. Rowbottom, M.Sc. The earliest published writing of Robert Boyle. Annals of Science, 6(4):376–389, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rowe:1950:LWG

[Row50b] J. S. Rowe, B.Sc. The life and work of George Fownes, F.R.S. (1815– 49). Annals of Science, 6(4):422–435, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rowe:2009:CCH

[Row09] Ted Rowe. Connecting the continents: Heart’s Content and the Atlantic cable. Creative Publishers, St. John’s, NL, Canada, 2009. ISBN 1-897174-42-X (paperback). vii + 158 pp. LCCN

TK5611 .R69 2009. URL http://www.loc.gov/catdir/toc/
fy1001/2009424702.html.

Rowe:2018:BRE

[Row18] David E. Rowe. Book review: Einstein and Twentieth-Century Politics: ‘A Salutary Moral Influence’, R. Crockatt, Oxford: Oxford University Press, 2016. 256 pp. $40.00, £25.00, ISBN 978-0-19-878549-1. Annals of Science, 75(1):69–71, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Robotti:1998:ZDM

[RP98] Nadia Robotti and Francesca Pastorino. Zeeman’s discovery and the mass of the electron. Annals of Science, 55(2):161–183, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Roderick:1971:PEC

[RS71] Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. and Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. Private enterprise and chemical training in nineteenth century Liverpool. Annals of Science, 27(1):85–93, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Roderick:1974:SSE

[RS74] Gordon W. Roderick and Michael D. Stephens. Science and secondary education in nineteenth century Liverpool. Annals of Science, 31(2):131–163, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1989:CAC

[RS89] John L. Russell and Sj. Catholic astronomers and the Copernican system after the condemnation of Galileo. Annals of Science, 46(4): 365–386, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ramberg:2001:YJH

[RS01] Peter J. Ramberg and Geert J. Somsen. The young J. H. van ’t Hoff: The background to the publication of his 1874 pamphlet on the tetrahedral carbon atom, together with a new English translation. Annals of Science, 58(1):51–74, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Roberts:2009:BCA

[RS09] Gerrylynn K. Roberts and Anna E. Simmons. British chemists abroad, 1887–1971: the dynamics of chemists’ careers. Annals of Science, 66(1):103–128, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rodrigues:2017:HPR

[RS17] Ana Duarte Rodrigues and Ana Simões. Horticulture in Portugal 1850–1900: The role of science and public utility in shaping knowledge. Annals of Science, 74(3):192–213, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rudd:2007:CAE

[Rud07] M. Eugene Rudd. Chromatic aberration of eyepieces in early telescopes. Annals of Science, 64(1):1–18, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rupke:2009:ROB

[Rup09] Nicolaas A. Rupke. Richard Owen: biology without Darwin. University of Chicago Press, Chicago, IL, USA and London, UK, revised edition, 2009. ISBN 0-226-73177-4. xxii + 344 pp. LCCN QH31.O94 R86 2009.

Russell-Wood:1949:BNW

[Rus49] J. Russell-Wood, M.Sc., Ph.D. A biographical note on William Brownrigg, M.D., F.R.S. (1711–1800). Annals of Science, 6(2):186– 196, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell-Wood:1950:SWW

[Rus50] J. Russell-Wood, M.Sc., Ph.D. The scientific work of William Brownrigg, M.D., F.R.S. (1711–1800). I. Annals of Science, 6(4):436–447, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell-Wood:1951:SWWa

[Rus51a] J. Russell-Wood, M.Sc., Ph.D. The scientific work of William Brownrigg, M.D., F.R.S. (1711–1800). II. Annals of Science, 7(1):77–94, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell-Wood:1951:SWWb

[Rus51b] J. Russell-Wood, M.Sc., Ph.D. The scientific work of William Brownrigg, M.D., F.R.S. (1711–1800). III. Annals of Science, 7(2):199–206, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1959:ETSa

[Rus59a] Colin A. Russell, M.Sc., A.R.I.C. The electrochemical theory of Sir Humphry Davy. Part I: The voltaic pile and electrolysis. Annals of Science, 15(1):1–13, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1959:ETSb

[Rus59b] Colin A. Russell, M.Sc., A.R.I.C. The electrochemical theory of Sir Humphry Davy. Part II: Electrical interpretations of chemistry. Annals of Science, 15(1):15–25, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1963:ETBa

[Rus63a] Colin A. Russell, M.Sc., Ph.D., F.R.I.C. The electrochemical theory of Berzelius. Part I: Origins of the theory. Annals of Science, 19(2):

117–126, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1963:ETBb

[Rus63b] Colin A. Russell, M.Sc., Ph.D., F.R.I.C. The electrochemical theory of Berzelius. Part II: An electrochemical view of matter. Annals of Science, 19(2):127–145, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1963:ETS

[Rus63c] Colin A. Russell, M.Sc., Ph.D., F.R.I.C. The electrochemical theory of Sir Humphry Davy. Part III: The evidence of the Royal Institution manuscripts. Annals of Science, 19(4):255–271, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1978:EFC

[Rus78] Colin A. Russell. Edward Frankland and the Cheapside chemists of Lancaster: an early Victorian pharmaceutical apprenticeship. Annals of Science, 35(3):253–273, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ruse:1980:CDG

[Rus80] Michael Ruse. Charles Darwin and group selection. Annals of Science, 37(6):615–630, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Russell:1995:WWC

[Rus95] John L. Russell. What was the crime of Galileo? Annals of Science, 52(4):403–410, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ruse:2004:BRB

[Rus04] Michael Ruse. Book reviews: I May Be Old Fashioned but…, Evelyn Fox Keller, Making Sense of Life: Explaining Biological Development with Models, Metaphors, and Machines Cambridge, MA: Harvard University Press, 2002. Annals of Science, 61(3):389–392, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rushton:2021:CHC

[Rus21] Alan R. Rushton. Counting human chromosomes before 1960: preconceptions, perceptions and predilections. Annals of Science, 78(1): 92–116, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Rushton:2022:CGC

[Rus22] Alan R. Rushton. Cambridge geneticists and the chromosome theory of inheritance: William Bateson, Leonard Doncaster and Reginald Punnett 1879–1940. Annals of Science, 79(4):468–496, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2113141.

Rutenberg:1939:EHP

[Rut39] D. Rutenberg, B.Sc. The early history of the potentiometer system of electrical measurement. Annals of Science, 4(2):212–243, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ryan:1991:SIR

[Rya91] W. F. Ryan. Scientific instruments in Russia from the middle ages to Peter the Great. Annals of Science, 48(4):367–384, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ryan:2023:NCR

[Rya23] Siobhan Moira Ryan. Norwegian climatology, the Republic of Letters and the Nordic Enlightenment. Annals of Science, 80(4):303–336, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2023.2209095.

SanchezDaza:2020:API

[S´ 20] Mariana Sánchez Daza. Alchemical and Paracelsian ideas in the Arte de los Metales. Annals of Science, 77(2):139–154, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sakuri:2012:BRL

[Sak12] Ayako Sakuri. Book review: Life sciences: Lynn K. Nyhart, Modern Nature. The Rise of the Biological Perspective in Germany, ISBN-13 978-0-226-61089-4. Annals of Science, 69(4):577–579, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Salman:1992:WHA

[Sal92] Phillips Salman. William Harvey and art misplaced. Annals of Science, 49(1):3–19, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Salviati:2009:TOT

[Sal09] Niccolò Guicciardini Corsi Salviati. Thinking with objects: The transformation of mechanics in the Seventeenth Century. Annals of Science, 66(2):279–281, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sanders:1984:CBT

[San84] P. M. Sanders. Charles de Bovelles’s treatise on the regular polyhedra (Paris, 1511). Annals of Science, 41(6):513–566, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sanabria:2000:ERP

[San00] Sergio L. Sanabria. Essay review: Perils of certitude in the structural analysis of historic masonry buildings. Annals of Science, 57(4):447– 453, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sandilands:2017:ERM

[San17] Catriona Sandilands. Essay review: Making kin, making trouble: Donna Haraway’s critical ongoingness. Annals of Science, 74(4):326– 330, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sarno:1969:SCW

[Sar69] Ronald A. Sarno, S.J., A.B., M.A. A sixteenth-century war of ideas: Science against the Church. Annals of Science, 25(3):209–227, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sargent:2002:NWR

[Sar02] Rose-Mary Sargent. A new way to read Boyle’s works. Annals of Science, 59(3):321–326, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sauermann:1976:EGI

[Sau76] Lotte Sauermann. Eberhard Gieseler, Ingenieur und Physiker. (German) [Eberhard Gieseler, engineer and physicist]. Annals of Science, 33(1):81–103, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sauer:2014:EUF

[Sau14] Tilman Sauer. Einstein’s unified field theory program. In Janssen and Lehner [JL14], chapter 9, pages 281–305. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). LCCN QC16.E5 C36 2014.

URL http://philsci-archive.pitt.edu/3293/1/uft.pdf.

Savelli:2015:BRB

[Sav15] Mat Savelli. Book review: Competing with the Soviets: Science, Technology, and the State in Cold War America. Annals of Science, 72(1):141–142, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shea:2009:GBN

[SB09] William R. Shea and Tiziana Bascelli. Galileo’s Sidereus Nuncius, or, a sidereal message. Science History Publications, Sagamore Beach, MA, USA, 2009. ISBN 0-88135-375-2. viii + 124 pp. LCCN QB41 .G173 2009.

Schuster:2013:DSM

[SB13] John A. Schuster and Judit Brody. Descartes and sunspots: Matters of fact and systematizing strategies in the Principia Philosophiae. Annals of Science, 70(1):1–45, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharlin:1975:SCT

[SBB^{+}75] Harold Issadore Sharlin, Stephen G. Brush, Harold L. Burstyn, Sandra Herbert, Michael S. Mahoney, and Nathan Sivin. A study and critique of the teaching of the history of science and technology. Interim report by the Committee on Undergraduate Education of the History of Science Society (U.S.A.). Annals of Science, 32(1):55–70, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Streijger:2010:JBQ

[SBT10] Michiel Streijger, Paul J. J. M. Bakker, and J. M. M. H. Thijssen, editors. John Buridan: Quaestiones super libros De generatione et corruptione Aristotelis: a critical edition with an introduction, volume 17 (or 14??) of History of science and medicine library; Medieval and early modern science. Brill, Leiden, The Netherlands, 2010. ISBN 90-04-18504-6 (hardcover). vii + 269 pp. LCCN Q151.A63 B872 2010.

Simpson:2004:MET

[SC04] A. D. C. Simpson and R. D. Connor. The mass of the English troy pound in the Eighteenth Century. Annals of Science, 61(3):321– 349, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Strickland:2015:LOH

[SC15] Lloyd Strickland and Michael Church. Leibniz’s observations on hydrology: An unpublished letter on the Great Lombardy Flood of 1705. Annals of Science, 72(4):517–532, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scerri:1994:PNH

[Sce94] Eric R. Scerri. Prediction of the nature of hafnium from chemistry, Bohr’s theory and quantum theory. Annals of Science, 51(2):137– 150, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scerri:2008:CPP

[Sce08] Eric R. Scerri, editor. Collected papers on philosophy of chemistry. Imperial College Press, London, UK, 2008. ISBN 1-84816-137-9. x +

235 pp. LCCN QD6 .S34 2008. URL http://www.loc.gov/catdir/
toc/fy0903/2008278137.html.

Scerri:2009:SPP

[Sce09] Eric R. Scerri, editor. Selected papers on the periodic table. Imperial College Press, London, UK, 2009. ISBN 1-84816-425-4. ix + 143 pp. LCCN QD467 .S347 2009.

Schlueter:1936:FAH

[Sch36] Robert E. Schlueter, M.D. The first Aphorism of Hippocrates as explained by Paracelsus. Annals of Science, 1(4):453–461, 1936. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scheer:1939:DCT

[Sch39] Bradley T. Scheer. The development of the concept of tissue respiration. Annals of Science, 4(3):295–305, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schofield:1954:JWL

[Sch54] Robert E. Schofield, M.S. James Watt’s letter to Joseph Priestley, 26 April 1783. Annals of Science, 10(4):294–300, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schofield:1956:MLS

[Sch56] Robert E. Schofield, M.S., Ph.D. Membership of the Lunar Society of Birmingham. Annals of Science, 12(2):118–136, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schofield:1957:SBJ

[Sch57] Robert E. Schofield, M.S., Ph.D. The scientific background of Joseph Priestley. Annals of Science, 13(3):148–163, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schallenberg:1975:EAS

[Sch75] Richard H. Schallenberg. Evolution, adaptation and survival: the very slow death of the American charcoal iron industry. Annals of Science, 32(4):341–358, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schmitt:1976:JCA

[Sch76] Charles B. Schmitt. John Case on art and nature. Annals of Science, 33(6):543–559, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schwartz:1980:TUL

[Sch80] Joel S. Schwartz. Three unpublished letters to Charles Darwin: the solution to a ‘geometrico–geological’ problem. Annals of Science, 37(6):631–637, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schettino:1989:NII

[Sch89] Edvige Schettino. A new instrument for infrared radiation measurements: the thermopile of Macedonio Melloni. Annals of Science, 46(5):511–517, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schillinger:1990:DSS

[Sch90] Klaus Schillinger. The development of Saxon scientific instrumentmaking skills from the Sixteenth Century to the Thirty Years War. Annals of Science, 47(3):277–289, 1990. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schemmel:2008:EGTa

[Sch08a] Matthias Schemmel. The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 1: Interpretation, volume 268(1) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346. xx + 762 (two volumes) pp. LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.springer.

com/book/10.1007/978-1-4020-5499-0.

Schemmel:2008:EGTb

[Sch08b] Matthias Schemmel. The English Galileo: Thomas Harriot’s work on motion as an example of preclassical mechanics. Volume 2: Sources, volume 268(2) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-5498-X (set: cased), 1-4020-5499-8 (set: e-book). ISSN 0068-0346. xx + 762 (two volumes) pp. LCCN QC133 .S36 2008; Q174 .B67 v.268. URL https://link.

springer.com/book/10.1007/978-1-4020-5499-0.

Schweizer:2008:SEN

[Sch08c] Claudia Schweizer. Stratigraphy in the early nineteenth century: a transdisciplinary approach, with special reference to Central Europe. Annals of Science, 65(2):257–274, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schiebinger:2009:CDW

[Sch09] L. Schiebinger. The correspondence of Dr. William Hunter, 1740– 1783. Annals of Science, 66(3):424–426, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schmidt:2011:BBK

[Sch11] W. H. Schmidt. Von der Bernoullischen Brachistochrone zum Kalibrator-Konzept. (German) [On the Bernoulli brachistochrone for calibrator concepts]. Annals of Science, 68(1):134–136, 2011. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schmit:2013:ODP

[Sch13a] Christophe Schmit. Sur l’origine du ‘Principe Général’ de Jean Le Rond D’Alembert. Annals of Science, 70(4):493–530, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schupbach:2013:BRB

[Sch13b] William Schupbach. Book review: Death Defied. The Anatomy Lessons of Frederik Ruysch, by L. Kooijmans. Annals of Science, 70(4):577–580, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schuster:2013:DAP

[Sch13c] John Andrew Schuster. Descartes-agonistes: physico-mathematics, method and corpuscular-mechanism 1618–33, volume 27 of Studies in history and philosophy of science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 94-007-4745-4 (hardcover). xix + 631 pp. LCCN B1875 .S387 2013.

Schaffer:2014:SL

[Sch14] Simon Schaffer. Swedenborg’s lunars. Annals of Science, 71(1):2– 26, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schneider:2018:FMI

[Sch18a] Jeremy Robin Schneider. The first mite: insect genealogy in Hooke’s Micrographia. Annals of Science, 75(3):165–200, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schuster:2018:BRO

[Sch18b] John Schuster. Book review: One Hundred Years of Pressure: Hydrostatics from Stevin to Newton, by Alan F. Chalmers, Dordrecht, Springer, 2017, ix + 197 pp., $99.99, £72.00, ISBN 978-3-319-56528-6 (Archimedes: New Studies in the History and Philosophy of Science and Technology, vol. 51, ISSN 1385-0180). Annals of Science, 75(2): 145–148, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schuster:2019:YRD

[Sch19] John A. Schuster. The young René Descartes — lawyer, military engineer, courtier, diplomat … and, we might add, ambitious ‘savant’. Annals of Science, 76(1):87–95, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schickore:2021:PSC

[Sch21] Jutta Schickore. The place and significance of comparative trials in German agricultural writings around 1800. Annals of Science, 78(4): 484–503, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scott:1936:JWH

[Sco36] J. F. Scott, B.A., Ph.D., M.Sc. John Wallis as a historian of mathematics. Annals of Science, 1(3):335–357, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scott:1981:RKJ

[Sco81] E. L. Scott. Richard Kirwan, J. H. de Magellan, and the early history of specific heat. Annals of Science, 38(2):141–153, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Siegfried:1968:CNA

[SD68] Robert Siegfried, Ph.D. and Betty Jo Dobbs, M.A. Composition, a neglected aspect of the chemical revolution. Annals of Science, 24(4):275–293, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:2013:KSE

[SD13] Justin E. H. Smith and James Delbourgo. In kind: Species of exchange in early modern science. Annals of Science, 70(3):299–304, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Searle:1979:EPB

[Sea79] G. R. Searle. Eugenics and politics in Britain in the 1930s. Annals of Science, 36(2):159–169, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Seed:2018:NEE

[See18] Patricia Seed. Navigational enterprises in Europe and its empires, 1730–1850. Annals of Science, 75(4):371–373, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Segre:1983:TCB

[Seg83] Michael Segre. Torricell’s correspondence on ballistics. Annals of Science, 40(5):489–499, September 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Seidel:2013:BRL

[Sei13] Robert W. Seidel. Book review: Lawrence Badash, A Nuclear Winter’s Tale. Science and Politics in the 1980s. Cambridge, Massachusetts: MIT Press, 2009. 424 pp. $40.00 / £29.95. ISBN-13 978-0-262-01272-0. Annals of Science, 70(1):124–126, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Serrano:2012:BNE

[Ser12] Elena Serrano. The Spectacle de la Nature in eighteenth-century Spain: From French households to Spanish workshops. Annals of Science, 69(2):257–282, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stott:1989:TPS

[SH89] Carole Stott and David W. Hughes. Two Piazzi Smyth comet paintings. Annals of Science, 46(2):165–172, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharlin:1972:BSC

[Sha72] Harold I. Sharlin, B.S., M.A., Ph.D. On being scientific: a critique of evolutionary geology and biology in the nineteenth century. Annals of Science, 29(3):271–285, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharp:1973:WCE

[Sha73] Lindsay Sharp, B.A. Walter Charleton’s early life 1620–1659, and relationship to natural philosophy in mid-seventeenth century England. Annals of Science, 30(3):311–340, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shapin:1975:PKS

[Sha75a] Steven Shapin. Phrenological knowledge and the social structure of early nineteenth-century Edinburgh. Annals of Science, 32(3):219– 243, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharlin:1975:WTD

[Sha75b] Harold Issadore Sharlin. William Thomson’s dynamical theory: An insight into a scientist’s thinking. Annals of Science, 32(2):133–147, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharlin:1976:HSS

[Sha76] Harold Issadore Sharlin. Herbert Spencer and scientism. Annals of Science, 33(5):457–465, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shalev:2012:SWW

[Sha12] Zur Shalev. Sacred words and worlds: geography, religion, and scholarship, 1550–1700, volume 21 / 2 of History of Science and Medicine Library; Scientific and learned cultures and their institutions. Brill, Leiden, The Netherlands, 2012. ISBN 90-04-20935-2. ISSN 1872-0684. xxi + 319 pp. LCCN BL65.G4 S53 2012.

Shackelford:2019:BRB

[Sha19a] Jole Shackelford. Book review: Burned alive: Giordano Bruno, Galileo and the Inquisition, by Alberto A. Martínez, London, Reaktion Books, 2018, 348 pp., 21 illustrations, £25.00 (hbk), ISBN 976-1-78023-8968. Annals of Science, 76(3–4):372–375, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shaheen:2019:ERW

[Sha19b] Jonathan Shaheen. Essay rview: The well-ordered universe: the philosophy of Margaret Cavendish, by D. Boyle, New York, Oxford University Press, 2018, x + 273 pp., $74.00, £53.00, ISBN 978-0-19-023480-5. Annals of Science, 76(3–4):340–346, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sheynin:1998:STB

[She98] Oscar Sheynin. Stochastic thinking in the Bible and the Talmud. Annals of Science, 55(2):185–198, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shea:2009:JHE

[She09] William R. Shea. The Jewel House: Elizabethan London and the scientific revolution. Annals of Science, 66(3):439–441, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shea:2011:ADC

[She11] William R. Shea. The Accademia del Cimento and its European context. Annals of Science, 68(4):573–575, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shimao:1981:DJ

[Shi81] Eikoh Shimao. Darwinism in Japan, 1877–1927. Annals of Science, 38(1):93–102, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shimao:1989:SAJ

[Shi89] Eikoh Shimao. Some aspects of Japanese science, 1868–1945. Annals of Science, 46(1):69–91, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Senter:2013:SYO

[SHM13] Phil Senter, LaRhonda C. Hill, and Brandon J. Moton. Solution to a 440-year-old zoological mystery: The case of Aldrovandi’s dragon.

Annals of Science, 70(4):531–537, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shteir:2017:OCH

[Sht17] Ann Shteir. Orchid: A cultural history. Annals of Science, 74(3):243– 245, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shugg:1968:HAE

[Shu68] Wallace Shugg, Ph.D. Humanitarian attitudes in the early animal experiments of the Royal Society. Annals of Science, 24(3):227– 238, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sidoli:2007:WWC

[Sid07] Nathan Sidoli. What we can learn from a diagram: The case of Aristarchus’s On The Sizes and Distances of the Sun and Moon. Annals of Science, 64(4):525–547, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Siebert:2009:PNI

[Sie09] Harald Siebert. Peirescs Nebel im Sternbild Orion — eine neue Textgrundlage für die Geschichte von M42. (German) [Peiresc nebula in the Constellation Orion — a new textual basis for the story of M42]. Annals of Science, 66(2):231–246, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Silk:2018:MFW

[Sil18] Natasha Silk. Medicine in First World War Europe: soldiers, medics, pacifists. Annals of Science, 75(4):377–379, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Simpkins:1966:EEE

[Sim66] Diana M. Simpkins, B.Sc., A.L.A. Early editions of Euclid in England. Annals of Science, 22(4):225–249, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Simpson:1992:GPE

[Sim92] A. D. C. Simpson. Grain packing in early standard capacity measures: Evidence from the Scottish dry capacity standards. Annals of Science, 49(4):337–350, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sima:1993:PST

[Sím93] Zdislav Síma. Prague sextants of Tycho Brahe. Annals of Science, 50(5):445–453, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Simms:2004:NCS

[Sim04] D. L. Simms. Newton’s contribution to the science of heat. Annals of Science, 61(1):33–77, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Simon:2013:BRB

[Sim13] Jonathan Simon. Book review: Aux origines de la chimie organique: Méthodes et pratiques des pharmaciens et des chimistes (1785–1835), by Sacha Tomic. Annals of Science, 70(4):564–566, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Singer:1949:SJP

[Sin49] Dorothea Waley Singer. Sir John Pringle and his circle. Part I. Life. Annals of Science, 6(2):127–180, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Singer:1950:SJPa

[Sin50a] Dorothea Waley Singer. Sir John Pringle and his circle. Part II. Public health. Annals of Science, 6(3):229–247, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Singer:1950:SJPb

[Sin50b] Dorothea Waley Singer. Sir John Pringle and his circle. Part III. Copley discourses. Annals of Science, 6(3):248–261, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stasiewicz-Jasiukowa:1978:THS

[SJ78] Irena Stasiewicz-Jasiukowa. The teaching of the history of science and technology in Polish higher education. Annals of Science, 35(1): 75–80, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sarrett:2011:CPD

[SK11] Jennifer C. Sarrett and Howard I. Kushner. A case for preserving the diversity of madness. Annals of Science, 68(4):547–554, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sharlin:1977:CPH

[SL77] Harold Issadore Sharlin and Robert A. Leacock. A course in physics and history: matching an unlikely pair. Annals of Science, 34(1): 57–62, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Slaton:2003:ERB

[Sla03] A. E. Slaton. Essay review: Technology and the Dream: Reflections on the Black Experience at MIT, 1941–1999 by C. G. Williams. Annals of Science, 60(3):321–325, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Slater:2019:BRS

[Sla19] John Slater. Book review: The Spanish disquiet: the biblical natural philosophy of Benito Arias Montano, by M. M. Portuondo, Chicago and London, The University of Chicago Press, 2019, xv + 426 pp., 13 plts, $65.00 (hardback), ISBN 978-0-226-59226-8. Annals of Science, 76(3–4):377–380, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sleep:1969:SWH

[Sle69] Mark C. W. Sleep, M.Sc., A.R.S.M. Sir William Hamilton (1730– 1803): His work and influence in geology. Annals of Science, 25(4): 319–338, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sleigh:2017:PZY

[Sle17] Charlotte Sleigh. The Paper Zoo: 500 Years of Art and Science. University of Chicago Press, Chicago, IL, USA and London, UK, 2017. ISBN 0-226-44712-X (hardcover), 0-226-44726-X (e-book). 256 pp. LCCN ????

Sloan:2003:WPD

[Slo03] Phillip R. Sloan. Whewell’s philosophy of discovery and the archetype of the vertebrate skeleton: The role of German philosophy of science in Richard Owen’s biology. Annals of Science, 60(1):39–61, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Slowik:2011:MAE

[Slo11] Edward S. Slowik. Mechanics from Aristotle to Einstein. Annals of Science, 68(1):142–144, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Savours:1982:HRO

[SM82] Ann Savours and Anita McConnell. The history of the Rossbank Observatory, Tasmania. Annals of Science, 39(6):527–564, 1982. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Shefer-Mossensohn:2015:SAO

[SM15] Miri Shefer-Mossensohn. Science among the Ottomans: the cultural creation and exchange of knowledge. University of Texas Press, Austin, TX, USA, 2015. ISBN 1-4773-0359-6 (hardcover), 1-4773-0360-X (library e-book), 1-4773-0361-8 (nonlibrary e-book). xiv + 246 pp. LCCN Q127.T9 S54 2015.

Suay-Matallana:2016:BCM

[SM16] Ignacio Suay-Matallana. Between chemistry, medicine and leisure: Antonio Casares and the study of mineral waters and Spanish spas in the nineteenth century. Annals of Science, 73(3):289–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1954:CPJ

[Sme54a] W. A. Smeaton, M.Sc., A.R.C.S. The contributions of P.-J. Macquer, T. O. Bergman and L. B. Guyton de Morveau to the reform of chemical nomenclature. Annals of Science, 10(2):87–106, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1954:EHL

[Sme54b] W. A. Smeaton, M.Sc., A.R.C.S. The early history of laboratory instruction in chemistry at the ecole polytechnique, Paris, and elsewhere. Annals of Science, 10(3):224–233, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1955:EYLa

[Sme55a] W. A. Smeaton, M.Sc., A.R.C.S. The early years of the Lycée and the Lycée des Arts. A chapter in the lives of A. L. Lavoisier and A. F. de Fourcroy. Annals of Science, 11(3):257–267, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1955:EYLb

[Sme55b] W. A. Smeaton, M.Sc., A.R.C.S. The early years of the Lycée and the Lycée des Arts. A chapter in the lives of A. L. Lavoisier and A. F. de Fourcroy. Annals of Science, 11(4):309–319, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1955:JFP

[Sme55c] W. A. Smeaton, M.Sc., A.R.C.S. Jean-François Pilâtre de Rozier. The first aeronaut. Annals of Science, 11(4):349–355, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1956:LMSb

[Sme56a] W. A. Smeaton, M.Sc., A.R.C.S. Lavoisier’s membership of the Société royale d’Agriculture and the Comité d’Agriculture. Annals of Science, 12(4):267–277, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1956:LMSa

[Sme56b] W. A. Smeaton, M.Sc., A.R.C.S. Lavoisier’s membership of the Société royale de médecine. Annals of Science, 12(3):228–244, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1956:TUP

[Sme56c] W. A. Smeaton, M.Sc., A.R.C.S. Two unrecorded publications of the régie des poudres et salpêtres probably written by Lavoisier. Annals of Science, 12(2):157–159, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1957:LCJ

[Sme57a] W. A. Smeaton, M.Sc., Ph.D., A.R.C.S. L’avant-coureur: The journal in which some of Lavoisier’s earliest research was reported. Annals of Science, 13(4):219–234, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1957:LMA

[Sme57b] W. A. Smeaton, M.Sc., Ph.D., A.R.C.S. Lavoisier’s membership of the assembly of representatives of the Commune of Paris, 1789–1790. Annals of Science, 13(4):235–248, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1967:SUE

[Sme67] W. A. Smeaton, D.Sc. Some unrecorded editions of Fourcroy’s Philosophie chimique. Annals of Science, 23(4):295–298, 1967. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1978:CWH

[Sme78] W. A. Smeaton. The chemical work of Horace Bénédict de Saussure (1740–1799), with the text of a letter written to him by madame Lavoisier. Annals of Science, 35(1):1–16, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smeaton:1987:SLB

[Sme87] W. A. Smeaton. Some large burning lenses and their use by eighteenth-century French and British chemists. Annals of Science, 44(3):265–276, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:1976:MPE

[Smi76] Crosbie Smith. ‘mechanical philosophy’ and the emergence of physics in Britain: 1800–1850. Annals of Science, 33(1):3–29, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:1977:HPV

[Smi77] Cyril Stanley Smith. A highly personal view of science and its history. Annals of Science, 34(1):49–56, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:1980:EUW

[Smi80] Crosbie Smith. Engineering the universe: William Thomson and Fleeming Jenkin on the nature of matter. Annals of Science, 37(4): 387–412, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:1987:POS

[Smi87] Diana C. F. Smith. The progress of the Orcades survey, with biographical notes on Murdoch Mackenzie senior (1712–1797). Annals of Science, 44(3):277–288, May 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:2013:SGL

[Smi13] Justin E. H. Smith. ‘A series of generations’: Leibniz on race. Annals of Science, 70(3):319–335, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:2015:BRB

[Smi15] Jonathan Smith. Book review: Novel Science: Fiction and the Invention of Nineteenth-Century Geology. Annals of Science, 72(2): 268–270, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:2018:BRD

[Smi18] Jonathan Smith. Book review: Darwin and the Making of Sexual Selection, by E. Richards, Chicago and London, University of Chicago Press, 2017, xxxiv + 672 pp., 48 halftones, $47.50, £35.50, ISBN 978-0-226-43690-6. Annals of Science, 75(1):67–69, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Smith:2022:MMV

[Smi22] Pamela H. Smith. Minerva meets Vulcan: Scientific and technological literature — 1450–1750. Annals of Science, 79(4):515–518, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2113430.

Smidrkalova:2023:CCA

[Smi23a] Michaela Smidrkalová. Celebrating the Czechoslovak atom: from ‘Atoms for Peace’ to Expo 58. Annals of Science, 80(1):38–61, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2164616.

Smith:2023:FSR

[Smi23b] Nathan Edward Charles Smith. Fertile substrate: the rise, fall, and succession of popular microscopy in Great Britain. Annals of Science, 80(3):268–292, 2023. CODEN ANNSA8. ISSN 0003-3790 (print),

1464-505X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00033790.2023.2181398.

Snelders:1975:RND

[Sne75] H. A. M. Snelders. The reception in The Netherlands of the discoveries of electromagnetism and electrodynamics (1820–1822). Annals of Science, 32(1):39–54, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Snelders:1976:MME

[Sne76] H. A. M. Snelders. A. M. Mayer’s experiments with floating magnets and their use in the atomic theories of matter. Annals of Science, 33 (1):67–80, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Snelders:1981:JFW

[Sne81] H. A. M. Snelders. James F. W. Johnston’s influence on agricultural chemistry in The Netherlands. Annals of Science, 38(5):571–584, 1981. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Snelders:1984:EEU

[Sne84] H. A. M. Snelders. The electromagnetic experiments of the Utrecht physicist Gerrit Moll (1785–1838). Annals of Science, 41(1):35–55, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Snobelen:2001:MHN

[Sno01] Stephen D. Snobelen. Mathematicians, historians and Newton’s Principia. Annals of Science, 58(1):75–84, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sokol:1974:THS

[Sok74] Barnett J. Sokol. Thomas Harriot — Sir Walter Ralegh’s tutor — on population. Annals of Science, 31(3):205–212, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sokol:1994:PAT

[Sok94] B. J. Sokol. The problem of assessing Thomas Harriot’s A briefe and true report of his discoveries in North America. Annals of Science, 51(1):1–16, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sokol:1996:PAA

[Sok96] B. J. Sokol. Poet in the atomic age: Robert Frost’s That Millikan Mote expanded. Annals of Science, 53(4):399–411, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sonntag:1974:LFB

[Son74] Otto Sonntag. Liebig on Francis Bacon and the utility of science. Annals of Science, 31(5):373–386, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sonntag:1975:AHA

[Son75] Otto Sonntag. Albrecht von Haller on academies and the advancement of science: the case of Göttingen. Annals of Science, 32(4): 379–391, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sheets-Pyenson:1985:PSP

[SP85] Susan Sheets-Pyenson. Popular science periodicals in Paris and London: The emergence of a low scientific culture, 1820–1875. Annals of Science, 42(6):549–572, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sheets-Pyenson:1992:HRJ

[SP92] Susan Sheets-Pyenson. Horse race: John William Dawson, Charles Lyell, and the competition over the Edinburgh natural history chair in 1854–1855. Annals of Science, 49(5):461–477, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Solovey:2011:PSS

[SP11] Mark Solovey and Jefferson D. Pooley. The price of success: Sociologist Harry Alpert, the NSF’s first social science policy architect. Annals of Science, 68(2):229–260, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sparrow:1955:CRS

[Spa55] W. J. Sparrow, M.A.B.Sc., Ph.D. Count Rumford as a spy. Annals of Science, 11(4):320–330, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spelda:2016:VFT

[Spe16] Daniel Spelda. Veritas filia temporis: The origins of the idea of scientific progress. Annals of Science, 73(4):375–391, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spiers:1968:SHD

[Spi68] C. H. Spiers, M.A., Ph.D., B.Sc., D.I.C., F.S.A. Sir Humphry Davy and the leather industry. Annals of Science, 24(2):99–113, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spiers:1969:WTB

[Spi69] C. H. Spiers, M.A., Ph.D., B.Sc., F.S.A. William Thomas Brande, leather expert. Annals of Science, 25(3):179–201, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spriggs:1942:JHH

[Spr42] Edmund Anthony Spriggs, B.A. John Hunter and his approach to pathology. Annals of Science, 5(2):177–184, 1942. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spring:1977:NTG

[Spr77] Robin J. Spring. A note on Thomas Graham, surgeon, author of botanical lectures delivered at the Royal Polytechnic Institution, London. Annals of Science, 34(1):43–47, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Spurgin:1987:GLG

[Spu87] C. B. Spurgin. Gay-Lussac’s gas-expansivity experiments and the traditional mis-teaching of ‘Charles’s Law’. Annals of Science, 44 (5):489–505, September 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1971:ETE

[SR71a] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. Education and training for English engineers in the late Nineteenth Century and early Twentieth Century. Annals of Science, 27(2):143–163, June 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1971:STN

[SR71b] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. Science training for the Nineteenth Century English amateur: The Penzance Natural History and Antiquarian Society. Annals of Science, 27(2):135–141, June 1971. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:LVS

[SR72a] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.C.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. The later Victorians and scientific and technical education. Annals of Science, 28(4):385–400, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:SEN

[SR72b] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., M.A., Ph.D., A.Inst.P. Supplementary education in a nineteenth-century British mining area. Annals of Science, 29(1): 59–79, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:BAS

[SR72c] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. British artisan scientific and technical education in the early nineteenth century. Annals of Science, 29(1):87–98, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:ML

[SR72d] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. The Muspratts of Liverpool. Annals of Science, 29(3):287–311, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:NCV

[SR72e] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. Nineteenth century ventures in Liverpool’s scientific education. Annals of Science, 28(1):61–86, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1972:SWC

[SR72f] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. Science, the working classes and mechanics’ institutes. Annals of Science, 29(4):349–360, 1972. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Schove:1973:WSD

[SR73a] D. J. Schove, Ph.D., F.R.Met.S. and David Reynolds, B.A. Weather in Scotland, 1659–1660: The diary of Andrew Hay. Annals of Science, 30(2):165–177, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1973:CAE

[SR73b] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., M.A., Ph.D., M.Inst.P. Changing attitudes to education in England & Wales 1833–1902: The governmental reports, with particular reference to science & technical studies. Annals of Science, 30(2):149–164, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1973:AEA

[SR73c] Michael D. Stephens, M.A., M.Ed., Ph.D., F.R.G.S. and Gordon W. Roderick, B.Sc., Ph.D., A.Inst.P. American and English attitudes to scientific education during the nineteenth-century. Annals of Science, 30(4):435–456, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1974:NCE

[SR74a] Michael D. Stephens and Gordon W. Roderick. Nineteenth century educational finance: The literary and philosophical societies. Annals of Science, 31(4):335–349, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1974:SSI

[SR74b] Michael D. Stephens and Gordon W. Roderick. Science, self improvement and the first industrial revolution. Annals of Science, 31(5):463–470, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scheler:1958:ALR

[SS58] Lucien Scheler and W. A. Smeaton, M.Sc., Ph.D., A.R.C.S. An account of Lavoisier’s reconciliation with the Church a short time before his death. Annals of Science, 14(2):148–153, 1958. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sierksma:1999:GLI

[SS99] Gerard Sierksma and Wybe Sierksma. The great leap to the infinitely small. Johann Bernoulli: Mathematician and philosopher. Annals of Science, 56(4):433–449, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sera-Shriar:2013:OHD

[SS13] Efram Sera-Shriar. Observing human difference: James Hunt, Thomas Huxley and competing disciplinary strategies in the 1860s. Annals of Science, 70(4):461–491, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sera-Shriar:2015:CMF

[SS15] Efram Sera-Shriar. The correspondence of Michael Faraday. Annals of Science, 72(3):401–406, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Scott-Smith:2023:KFG

[SS23] Giles Scott-Smith. Book review: Knowledge flows in a global age: a transnational approach, d. John Krige, Chicago and London, University of Chicago Press, 2022, 368 pp., $135 (Hardback), ISBN: 978-0-226-81994-5; $45 (Paperback), IBSN: 978-0-226-82038-5. Annals of Science, 80(4):423–425, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00033790.2023.2195418.

Street:1951:DA

[ST51] H. E. Street, Ph.D. and G. E. Trease, B.Pharm., F.R.I.C., F.L.S. The discovery of asparagine. Annals of Science, 7(1):70–76, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stafford:1988:RMS

[Sta88a] Robert A. Stafford. Roderick Murchison and the structure of Africa: a geological prediction and its consequences for British expansion. Annals of Science, 45(1):1–40, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stamhuis:1988:MRL

[Sta88b] Ida H. Stamhuis. The mathematician Rehuel Lobatto advocates life insurances in The Netherlands in the period 1830–1860. Annals of Science, 45(6):619–641, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Staley:2013:BRB

[Sta13] Richard Staley. Book review: Domesticating Electricity: Technology, Uncertainty and Gender, 1880–1914. Annals of Science, 70(2):295– 297, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stanley:2017:SSI

[Sta17] Matthew Stanley. Strange science: Investigating the limits of knowledge in the Victorian Age. Annals of Science, 74(3):240–241, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stan:2019:SCN

[Sta19] Marius Stan. Standing colossus: Newton and the French: Essay review of J. B. Shank, Before Voltaire: the French Origins of “Newtonian” Mechanics, 1680–715. University of Chicago Press, 2018. Cloth, x + 444 pp., ill. ISBN 978-0-226-50929-7. $55.00. Annals of Science, 76(3–4):347–354, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stanley:2021:BRE

[Sta21] Matthew Stanley. Book review: Einstein in Bohemia, by M. Gordin, Princeton, NJ, Princeton University Press, 2020, xi + 343 pp., $29.95. Hardback. ISBN 978-0-691-17737-3. Annals of Science, 78 (1):131–132, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sangiacomo:2022:MEE

[STDH22] Andrea Sangiacomo, Raluca Tanasescu, Silvia Donker, and Hugo Hogenbirk. Mapping the evolution of early modern natural philosophy: corpus collection and authority acknowledgement. Annals of Science, 79(1):1–39, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stearns:1936:CCE

[Ste36a] Raymond Phineas Stearns, A.M., Ph.D. The course of Capt. Edmond Halley in the year 1700. Annals of Science, 1(3):294–301, 1936. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stearns:1936:PSB

[Ste36b] Raymond Phineas Stearns, A.M., Ph.D. The production of sugar in Barbados c. 1667. Annals of Science, 1(2):173–181, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stecher:1961:DIL

[Ste61] Robert M. Stecher, M.D. The Darwin–Innes letters the correspondence of an evolutionist with his vicar, 1848–1884. Annals of Science, 17(4):201–258, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stecher:1969:DBLa

[Ste69a] Robert M. Stecher, M.D. The Darwin–Bates letters correspondence between two nineteenth-century travellers and naturalists. Part I. Annals of Science, 25(1):1–47, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stecher:1969:DBLb

[Ste69b] Robert M. Stecher, M.D. The Darwin–Bates letters correspondence between two nineteenth-century travellers and naturalists. Part II. The Darwin–Bates correspondence. Annals of Science, 25(2):95–125, 1969. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stenzel:1976:BVB

[Ste76] Rudolf Stenzel. Begründung für die Verschmelzung der Reichsanstalt für Mass und Gewicht mit der Physikalisch-Technischen Reichsanstalt in Berlin im Jahre 1923. (German) [rationale for the merger of the Imperial Institute of Weights and Measures of the Physical-Technical Institute in Berlin in 1923]. Annals of Science, 33(3):289–306, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stephens:1980:JLD

[Ste80] Lester D. Stephens. Joseph LeConte and the development of the physiology and psychology of vision in the United States. Annals of Science, 37(3):303–321, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stedall:2000:ATL

[Ste00a] Jacqueline Anne Stedall. Ariadne’s thread: The life and times of Oughtred’s Clavis [mathematicae]. Annals of Science, 57(1):27–60, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stewart:2000:SFD

[Ste00b] John Stewart. ’science fights death’: David Stark Murray, science, and socialism in Interwar Britain. Annals of Science, 57(2):143– 161, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Steele:2007:CMB

[Ste07] J. M. Steele. Celestial measurement in Babylonian astronomy. Annals of Science, 64(3):293–325, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Steele:2009:ECI

[Ste09a] John M. Steele. Explaining the Cosmos: The Ionian tradition of scientific philosophy. Annals of Science, 66(3):444–445, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Steele:2009:OHR

[Ste09b] John M. Steele. Ordering the heavens: Roman astronomy and cosmology in the Carolingian Renaissance. Annals of Science, 66(2): 297–299, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stein:2010:MAD

[Ste10a] Claudia Stein. Murder after death: Literature and anatomy in early modern England. Annals of Science, 67(2):273–274, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Steinicke:2010:OCN

[Ste10b] Wolfgang Steinicke. Observing and cataloguing nebulae and star clusters: from Herschel to Dreyer’s new general catalogue. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19267-6 (hardcover). xi + 648 pp. LCCN QB853 .S736 2010. URL http://assets.

cambridge.org/97805211/92675/cover/9780521192675.jpg.

Stengers:2010:C

[Ste10c] Isabelle Stengers. Cosmopolitics, volume 9–10 of Posthumanities. University of Minnesota Press, Minneapolis, MN, USA, 2010. ISBN 0-8166-5686-X (vol. 1: hc), 0-8166-5687-8 (vol. 1: pb), 0-8166-5688-6 (vol. 2). ???? pp. LCCN Q125 .S742613 2010. Translated by Robert Bononno.

Stedall:2011:CGA

[Ste11a] Jacqueline A. Stedall. From Cardano’s great art to Lagrange’s reflections: filling a gap in the history of algebra. Heritage of European mathematics. European Mathematical Society, Zürich, Switzerland, 2011. ISBN 3-03719-092-2. xi + 224 pp. LCCN QA151 .S745 2011.

Steele:2011:VAT

[Ste11b] J. M. Steele. Visual aspects of the transmission of Babylonian astronomy and its reception into Greek astronomy. Annals of Science, 68(4):453–465, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stewart:2019:RNE

[Ste19] Larry Stewart. Reading Newton in early modern Europe. Annals of Science, 76(2):219–221, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stewart:2020:CSS

[Ste20] John Stewart. Chemistry and slavery in the Scottish Enlightenment. Annals of Science, 77(2):155–168, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stimson:1937:THS

[Sti37] Dorothy Stimson, Ph.D. The teaching of the history of science in a liberal arts college for women. Annals of Science, 2(4):460–464, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Storrs:1966:LTR

[Sto66] F. C. Storrs, M.Sc., F.R.I.C., Dip.Bact. (Lond.). Lavoisier’s technical reports: 1768–1794 Part I. Annals of Science, 22(4):251–275, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Storrs:1968:LTR

[Sto68] F. C. Storrs, M.Sc., F.R.I.C., Dip.Bact. (Lond.). Lavoisier’s technical reports: 1768–1794. Part II. Annals of Science, 24(3):179–197, 1968. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stockland:2013:GIP

[Sto13] Etienne Stockland. ‘La guerre aux insectes’: Pest control and agricultural reform in the French Enlightenment. Annals of Science, 70 (4):435–460, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stranges:1993:SFP

[Str93] Anthony N. Stranges. Synthetic fuel production in prewar and World War II. Japan: a case study in technological failure. Annals of Science, 50(3):229–265, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Strauss:1994:PLW

[Str94] David Strauss. Percival Lowell, W. H. Pickering and the founding of the Lowell Observatory. Annals of Science, 51(1):37–58, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stranges:1997:UBM

[Str97] Anthony N. Stranges. The US Bureau of Mines’s synthetic fuel programme, 1920–1950s: German connections and American advances. Annals of Science, 54(1):29–68, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Strick:2014:BRB

[Str14] James Strick. Book review: Vision, Science and Literature, 1870– 1920: Ocular Horizons. Annals of Science, 71(4):588–590, 2014. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stubbs:1982:JBP

[Stu82] Mayling Stubbs. John Beale, philosophical gardener of Herefordshire. Part I. Prelude to the Royal Society (1608–1663). Annals of Science, 39(5):463–489, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Stubbs:1989:JBP

[Stu89] Mayling Stubbs. John Beale, philosophical gardener of Herefordshire. Part II. The improvement of agriculture and trade in the Royal Society (1663–1683). Annals of Science, 46(4):323–363, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sturdy:2009:SSG

[Stu09] Steve Sturdy. Suffer and survive: Gas attacks, miners’ canaries, spacesuits and the bends: The extreme life of Dr. J. S. Haldane. Annals of Science, 66(4):575–578, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Susalla:2015:BRB

[Sus15] Peter Susalla. Book review: Unravelling Starlight: William and Margaret Huggins and the Rise of the New Astronomy. Annals of Science, 72(1):124–126, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Susalla:2016:BRS

[Sus16] Peter Susalla. Book review: Selected Correspondence of William Huggins, Barbara J. Becker, Pickering and Chatto, London (2014) ISBN 1-84893-415-7 (two volumes). Annals of Science, 73(4):447– 448, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Suzuki:2009:SBN

[Suz09] Akihito Suzuki. Science and the building of a new Japan. Annals of Science, 66(2):288–290, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Svensson:2019:ECR

[Sve19] Anna Svensson. ‘And Eden from the Chaos rose’: utopian order and rebellion in the Oxford Physick Garden. Annals of Science, 76(2): 157–183, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1967:RJA

[SW67] Jessie M. Sweet, M.B.E., B.Sc. and Charles D. Waterston, B.Sc., Ph.D., F.R.S.E. Robert Jameson’s approach to the Wernerian theory of the Earth, 1796. Annals of Science, 23(2):81–95, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swann:1989:MRH

[Swa89] John P. Swann. Manuscript resources in the history of chemistry at the national library of medicine. Annals of Science, 46(3):249– 262, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1963:RJL

[Swe63] Jessie M. Sweet, M.B.E., B.Sc. Robert Jameson in London, 1793 excerpts from Robert Jameson’s Journal of a Voyage from Leith to London 1793. Annals of Science, 19(2):81–116, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1964:MGE

[Swe64] Jessie M. Sweet, M.B.E., B.Sc. Matthew Guthrie (1743–1807): An eighteenth-century gemmologist. Annals of Science, 20(4):245–302, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1967:RJI

[Swe67] Jessie M. Sweet, M.B.E., B.Sc. Robert Jameson’s Irish journal, 1797 excerpts from Robert Jameson’s Journal of my Tour in 1797. Annals of Science, 23(2):97–126, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1970:CLD

[Swe70a] Jessie M. Sweet, M.B.E., B.Sc. The collection of Louis Dufresne (1752–1832). Annals of Science, 26(1):33–71, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1970:WBC

[Swe70b] Jessie M. Sweet, M.B.E., B.Sc. William Bullock’s collection and the University of Edinburgh, 1819. Annals of Science, 26(1):23–32, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1972:ICJ

[Swe72a] Jessie M. Sweet, M.B.E., B.Sc. Instructions to collectors: John Walker (1973) and Robert Jameson (1817); with biographical notes on James Anderson (LL.D.) and James Anderson (M.D.). Annals of Science, 29(4):397–414, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1972:MWB

[Swe72b] Jessie M. Sweet, M.B.E., B.Sc. Morten Wormskiold: Botanist: (1783–1845). Annals of Science, 28(3):293–305, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sweet:1974:RJE

[Swe74] Jessie M. Sweet, M.B.E., B.Sc. Robert Jameson and the explorers: The search for the North-West Passage. Part I. 1. W. Scoresby (junior), C. L. Giesecke, M. Wormskiold and John Ross. Annals of Science, 31(1):21–47, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swerdlow:2003:ERP

[Swe03] N. M. Swerdlow. Essay review: Ptolemy’s guide to drawing a map of the Earth. J. Lennart Berggren and Alexander Jones, Ptolemy’s Geography, An Annotated Translation of the Theoretical Chapters, ISBN 0-691-09259-1 (hbk), ISBN 0-691-01042-0 (pbk). Annals of Science, 60(3):313–320, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swerdlow:2009:LTT

[Swe09] N. M. Swerdlow. The lunar theories of Tycho Brahe and Christian Longomontanus in the Progymnasmata and Astronomia Danica. Annals of Science, 66(1):5–58, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swerdlow:2021:TNP

[Swe21] N. M. Swerdlow. Essay review: Theoricae novae planetarum Georgii Peurbachii dans l’histoire de l’astronomie, Michela Malpangotto, CNRS Éditions, 2020. 751pp. 65.00 EUR. ISBN-10: 2-271-13458-7, ISBN-13: 978-2-271-13458-5. Annals of Science, 78(3):387–392, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swinburne:1962:PAT

[Swi62] R. G. Swinburne, M.A., B.Phil. The presence-and-absence theory. Annals of Science, 18(3):131–145, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swinburne:1965:GLF

[Swi65] R. G. Swinburne, M.A., B.Phil. Galton’s law — formulation and development. Annals of Science, 21(1):15–31, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Swinfin:2012:CEM

[Swi12] Stephen T. Swinfin. Chemists employed in the Manchester area, 1902–1936. Annals of Science, 69(2):239–256, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Sysak:1993:CCN

[Sys93] Janusz Sysak. Coleridge’s construction of Newton. Annals of Science, 50(1):59–81, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Szanser:1972:FGW

[Sza72] A. J. M. Szanser, M.Sc., Ph.D. F. G. W. Struve (1793–1864). Astronomer at the Pulkovo Observatory. Annals of Science, 28(4):327– 346, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tachau:1996:LGN

[Tac96] Katherine H. Tachau. Logic’s God and the natural order in late medieval Oxford: The teaching of Robert Holcot. Annals of Science, 53(3):235–267, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Takahashi:2011:MSS

[Tak11] Hidemi Takahashi. The mathematical sciences in Syriac: From Sergius of Resh-‘Aina and Severus Sebokht to Barhebraeus and Patriarch Ni‘matallah. Annals of Science, 68(4):477–491, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tanner:1977:NTB

[Tan77] R. C. H. Tanner. Nathaniel Torporley’s Congestor analyticus [Concentrated analysis] and Thomas Harriot’s De triangulis laterum rationalium [the sides of the rational triangle]. Annals of Science, 34 (4):393–428, 1977. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tanner:1980:ARM

[Tan80a] R. C. H. Tanner. The alien realm of the minus: Deviatory mathematics in Cardano’s writings. Annals of Science, 37(2):159–178, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tanner:1980:ORM

[Tan80b] R. C. H. Tanner. The ordered regiment of the minus sign: Off-beat mathematics in Harriot’s manuscripts. Annals of Science, 37(2):127– 158, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tannenbaum:2011:DDT

[Tan11] J. H. Tannenbaum. Dropsy, dialysis, transplant: a short history of failing kidneys. Annals of Science, 68(3):441–443, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tan:2014:SPN

[Tan14] Wei Yu Wayne Tan. Science, politics and networks: Shibukawa Harumi and the birth of the new almanac in Seventeenth-Century Japan. Annals of Science, 71(2):241–270, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tarrant:2020:RTA

[Tar20] Neil Tarrant. Reconstructing Thomist astrology: Robert Bellarmine and the papal bull Coeli et terrae. Annals of Science, 77(1):26– 49, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tattersall:1997:FPH

[Tat97] Robert Tattersall. Frederick Pavy (1829–1911) and his opposition to the glycogenic theory of Claude Bernard. Annals of Science, 54(4): 361–374, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tate:2017:BRE

[Tat17] Gregory Tate. Book review: The Experimental Self: Humphry Davy and the Making of a Man of Science. Annals of Science, 74(4):335– 336, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tauber:2010:CIW

[Tau10] Alfred I. Tauber. Crafting immunity. Working histories of clinical immunology. Annals of Science, 67(2):274–277, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taub:2014:BRB

[Tau14] Liba Taub. Book review: Science for the Nation: Perspectives on the History of the Science Museum, edited by P. J. T. Morris. Annals of Science, 71(1):132–135, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:1942:OT

[Tay42] F. Sherwood Taylor, M.A., Ph.D. The origin of the thermometer. Annals of Science, 5(2):129–156, 1942. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:1945:ES

[Tay45] F. Sherwood Taylor, Ph.D., M.A., B.Sc. The evolution of the still. Annals of Science, 5(3):185–202, 1945. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:1949:IH

[Tay49] F. Sherwood Taylor, Ph.D., M.A., B.Sc. The invention of the hygroscope. Annals of Science, 6(2):181–185, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:1952:CSJ

[Tay52a] F. Sherwood Taylor, M.A.B.Sc., Ph.D. The chemical studies of John Evelyn. Annals of Science, 8(4):285–292, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:1952:TSO

[Tay52b] F. Sherwood Taylor, M.A.B.Sc., Ph.D. The teaching of science at Oxford in the nineteenth century. Annals of Science, 8(1):82–112, 1952. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Taylor:2008:MDC

[Tay08] Georgette Taylor. Marking out a disciplinary common ground: The role of chemical pedagogy in establishing the doctrine of affinity at the heart of British chemistry. Annals of Science, 65(4):465–486, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1993:AAG

[TD93] Gerard L’E. Turner and Elly Dekker. An astrolabe attributed to Gerard Mercator, c. 1570. Annals of Science, 50(5):403–443, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tee:1979:ALB

[Tee79] G. J. Tee. Another link between Marx and Darwin. Annals of Science, 36(2):176, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tee:1984:ROO

[Tee84] G. J. Tee. Religious opposition to obstetric anaesthesia: Hardly a myth. Annals of Science, 41(2):179, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Teich:1975:BAP

[Tei75] Mikulás Teich. Born’s amalgamation process and the international metallurgic gathering at Skleno in 1786. Annals of Science, 32(4): 305–340, 1975. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Terry:1937:OET

[Ter37] Henry Terry. Obituary: Edward Turner, M.D., F.R.S. (1798–1837). Annals of Science, 2(2):137–152, April 15, 1937. URL http://www.

tandfonline.com/doi/abs/10.1080/00033793700200481.

Tatham:1974:AOS

[TH74] W. G. Tatham and K. A. Harwood. Astronomers and other scientists on St. Helena. Annals of Science, 31(6):489–510, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thackray:1966:FRJ

[Tha66] Arnold W. Thackray, M.A.B.Sc. Fragmentary remains of John Dalton. Part I. Letters. Annals of Science, 22(3):145–174, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Theunissen:1986:RCL

[The86] Bert Theunissen. The relevance of Cuvier’s lois zoologiques for his palaeontological work. Annals of Science, 43(6):543–556, 1986. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Theunissen:1994:CDH

[The94] Bert Theunissen. Closing the door on Hugo de Vries’ Mendelism. Annals of Science, 51(3):225–248, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thiele:1971:WDI

[Thi71] Joachim Thiele, Dr.rer.nat., Dr.phil. Zur Wirkungsgeschichte des Dopplerprinzips im Neunzehnten Jahrhundert. (German) [The historical impact of the Doppler principle in the Nineteenth Century]. Annals of Science, 27(4):393–407, December 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thiele:1973:FWK

[Thi73] Joachim Thiele. Franz Walds Kritik der theoretischen Chemie (nach Arbeiten aus den Jahren 1902–1906 und unveröffentlichten Briefen). (German) [Franz Wald’s critique of theoretical chemistry (according to works from the years 1902 to 1906 and unpublished letters)]. Annals of Science, 30(4):417–433, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thirion:2012:LKZ

[Thi12] Gabriel Thirion. L’Ensei kansho zusetsu (1823) de Yoshio Nanko: une fenêtre sur la science classique. (French) [The Ensei kansho zusetsu (1823) of Yoshio Nanko: a window into classical science]. Annals of Science, 69(1):105–126, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thomas:1947:RGI

[Tho47] H. Hamshaw Thomas, Sc.D., F.R.S. The rise of geology and its influence on contemporary thought. Annals of Science, 5(4):325– 341, 1947. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thorndike:1951:ACC

[Tho51a] Lynn Thorndike, A.M., Ph.D. Astronomical and chronological calculations at Newminster in 1428. Annals of Science, 7(3):275–283, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thornton:1951:HPS

[Tho51b] John L. Thornton, A.L.A. The history of physiology at St. Bartholomew’s Hospital, London. Annals of Science, 7(3):238–247, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thompson:1954:GGO

[Tho54] H. R. Thompson, M.A., M.Sc. The geographical and geological observations of Bernard Palissy the potter. Annals of Science, 10(2): 149–165, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thompson:1955:QCH

[Tho55] D. Thompson, M.A., B.Sc. Queenwood College, Hampshire: a mid-19th century experiment in science teaching. Annals of Science, 11 (3):246–254, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thompson:1957:JTR

[Tho57] D. Thompson, M.A., B.Sc. John Tyndall and The Royal Institution. Annals of Science, 13(1):9–22, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thornton:1967:CHW

[Tho67] John L. Thornton, F.L.A. Charles Hunnings Wilkingson (1763 or 64– 1850). Annals of Science, 23(4):277–286, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thorpe:2011:BRS

[Tho11] Charles Thorpe. Book review: Silvan S. Schweber, Einstein and Oppenheimer, ISBN-13 978-0-674-02828-9. Annals of Science, 68(4): 558–561, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thomas:2016:BRD

[Tho16] Emily Thomas. Book review: Descartes-agonistes: Physicomathematics, method and corpuscular-mechanism 1618–33. Annals of Science, 73(1):112–114, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00033790.2015.1085210.

Tirapicos:2020:CTB

[Tir20] Luís Tirapicos. On the censorship of Tycho Brahe’s books in Iberia. Annals of Science, 77(1):96–107, 2020. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tirapicos:2023:DPG

[Tir23] Luís Tirapicos. Directions of precision: George Graham’s instructions for his pendulum astronomical clocks. Annals of Science, 81(1–2): 124–138, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2023.2282782.

Tkaczyk:2014:LCS

[Tka14] Viktoria Tkaczyk. Listening in circles. spoken drama and the architects of sound, 1750–1830. Annals of Science, 71(3):299–334, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tobies:2012:IRL

[Tob12] Renate Tobies. Iris Runge: a life at the crossroads of mathematics, science and industry, volume 43 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 3-0348-0229-3 (hardcover), 3-0348-0251-X (e-book). xxii + 442 + 16 pp. LCCN QA29.R86 T63 2012.

URL http://www.springerlink.com/content/978-3-0348-0251-

2. With a foreword by Helmut Neunzert. Revised by the author and translated by Valentine A. Pakis.

Todd:1982:NFP

[Tod82] Robert B. Todd. A note on Francesco Patrizi’s use of Cleomedes. Annals of Science, 39(3):311–314, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Todd:1993:SPI

[Tod93] Jan Todd. Science at the periphery: An interpretation of Australian scientific and technological dependency and development prior to 1914. Annals of Science, 50(1):33–58, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tomic:2001:LCV

[Tom01] Sacha Tomic. L’Analyse chimique des végétaux: Le cas du quinquina. (French) [The chemical analysis of vegetables: the case of quinquina]. Annals of Science, 58(3):287–309, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tomory:2009:OGT

[Tom09] Leslie Tomory. The origins of gaslight technology in eighteenthcentury pneumatic chemistry. Annals of Science, 66(4):473–496, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tomic:2010:OCO

[Tom10] Sacha Tomic. Aux origines de la chimie organique: méthodes et pratiques des pharmaciens et des chimistes (1785–1835). (French) [At the origins of organic chemistry: methods and practices of pharmacists and chemists (1785–1835)]. Collection “Carnot”. Presses universitaires de Rennes, Rennes, France, 2010. ISBN 2-7535-1018-0. 322 pp. LCCN QD248.5.F8 T66 2010.

Tomory:2012:PEO

[Tom12] Leslie Tomory. Progressive enlightenment: the origins of the gaslight industry, 1780–1820. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01675-3 (hardcover). x + 348 pp. LCCN TP733.G6

T66 2012. URL http://muse.jhu.edu/books/9780262301244/;
http://www.jstor.org/stable/10.2307/j.ctt5vjrt6.

Tomory:2014:SAW

[Tom14] Leslie Tomory. Science and the arts in William Henry’s research into inflammable air during the Early Nineteenth Century. Annals of Science, 71(1):61–81, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Topper:1980:RMI

[Top80] David R. Topper. ‘to reason by means of images’: J. J. Thomson and the mechanical picture of nature. Annals of Science, 37(1):31– 57, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Touret:1989:CFE

[Tou89] Lydie Touret. Charles–François Exchaquet (1746–1792) et les Plans en Relief du Mont-Blanc. Annals of Science, 46(1):1–20, 1989. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1963:CRS

[Tre63] L. Trengove, B.D.M.Sc., Ph.D. Chemistry at the Royal Society of London in the eighteenth century. I. Annals of Science, 19(3):183– 237, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1964:CRS

[Tre64] L. Trengove, B.D.M.Sc., Ph.D. Chemistry at the Royal Society of London in the eighteenth century. II. Annals of Science, 20(1):1– 57, 1964. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1965:CRSb

[Tre65a] L. Trengove, B.D.M.Sc., Ph.D. Chemistry at the Royal Society of London in the eighteenth century. III(B). Metals. Annals of Science, 21(3):175–201, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1965:CRSa

[Tre65b] Leonard Trengove, B.D.M.Sc., Ph.D. Chemistry at the Royal Society of London in the eighteenth century. III(A). Metals. Annals of Science, 21(2):81–130, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1966:NTV

[Tre66] Leonard Trengove, B.D.M.Sc., Ph.D. Newton’s theological views. Annals of Science, 22(4):277–294, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1970:CRS

[Tre70] Leonard Trengove, M.A., B.D., M.Sc., Ph.D. Chemistry at the Royal Society of London in the eighteenth century. IV. Dyes. Annals of Science, 26(4):331–353, 1970. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trengove:1972:WGD

[Tre72] Leonard Trengove, M.A., B.D., M.Sc., Ph.D. William Gregor (1761– 1817) discoverer of titanium. Annals of Science, 29(4):361–395, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trenn:1974:RAT

[Tre74] Thaddeus J. Trenn, Ph.D. Rutherfords Alpha-Teilchen. (German) [Rutherford’s alpha particles]. Annals of Science, 31(1):49–72, 1974. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trenn:1978:TUE

[Tre78a] Thaddeus J. Trenn. Thoruranium (U-236) as the extinct natural parent of thorium: The premature falsification of an essentially correct theory. Annals of Science, 35(6):581–597, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trevena:1978:MBF

[Tre78b] David H. Trevena. Marcelin Berthelot’s first publication in 1850, on the subjection of liquids to tension. Annals of Science, 35(1): 45–54, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trenn:1980:PAR

[Tre80a] Thaddeus J. Trenn. The phenomenon of aggregate recoil: the premature acceptance of an essentially incorrect theory. Annals of Science, 37(1):81–100, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trevena:1980:PWF

[Tre80b] David H. Trevena. The pioneer work of François Donny on the existence of tension in liquids. Annals of Science, 37(4):379–386, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Trenn:1986:GMC

[Tre86] Thaddeus J. Trenn. The Geiger–Müller counter of 1928. Annals of Science, 43(2):111–135, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tresise:1991:CHP

[Tre91] Geoffrey Tresise. ‘Chirotherium herculis’ — problems posed by the first finds. Annals of Science, 48(6):565–576, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Treasure:2015:BRB

[Tre15] Tom Treasure. Book review: A History of Organ Transplantation: Ancient Legends to Modern Practice. Annals of Science, 72(1):138– 141, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Teichmann:2014:WHW

[TS14] Jürgen Teichmann and Arthur Stinner. From William Hyde Wollaston to Alexander von Humboldt — star spectra and celestial landscape. Annals of Science, 71(1):27–60, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Tagliaferri:1999:CPT

[TT99] Guido Tagliaferri and Pasquale Tucci. Carlini and Plana on the theory of the Moon and their dispute with Laplace. Annals of Science, 56(3):221–269, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turnbull:1953:PSF

[Tur53] G. H. Turnbull, M.A., Ph.D. Peter Stahl, the first public teacher of chemistry at Oxford. Annals of Science, 9(3):265–270, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1955:NMI

[Tur55] Joseph Turner, Ph.D. A note on Maxwell’s interpretation of some attempts at dynamical explanation. Annals of Science, 11(3):238– 245, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1967:ASE

[Tur67] G. L’E. Turner, M.A., M.Sc. The auction sales of the Earl of Bute’s instruments, 1793. Annals of Science, 23(3):213–242, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1973:MIE

[Tur73] A. J. Turner, B.A. Mathematical instruments and the education of gentlemen. Annals of Science, 30(1):51–88, 1973. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1982:AMY

[Tur82] A. J. Turner. ‘the accomplishment of many years’: Three notes towards a history of the sand-glass. Annals of Science, 39(2):161– 172, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1984:ODJ

[Tur84] G. L’e. Turner. Obituary: Derek John de Solla Price 1922–1983. Annals of Science, 41(2):105–107, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1991:ISN

[Tur91] G. L’E. Turner. Introduction: Some notes on the development of surveying and the instruments used. Annals of Science, 48(4):313– 317, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1994:TAG

[Tur94] Gerard L’E. Turner. The three astrolabes of Gerard Mercator. Annals of Science, 51(4):329–353, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:1999:NE

[Tur99] G. L’E. Turner. A new editor. Annals of Science, 56(1):1, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:2000:GEO

[Tur00] Gerard L’E. Turner. The government and the English optical glass industry, 1650–1850. Annals of Science, 57(4):399–414, 2000. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:2003:IHN

[Tur03] Gerard L’E. Turner. The Italian-hour nocturnal. Annals of Science, 60(3):249–268, 2003. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Turner:2006:NIH

[Tur06] Gerard L’E. Turner. A novel Italian-hour nocturnal by Michiel Coignet. Annals of Science, 63(2):215–219, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thornton:1956:WO

[TW56] John L. Thornton and Anna Wiles. William Odling, 1829–1921. Annals of Science, 12(4):288–295, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Usselman:2015:ATM

[UB15] Melvyn C. Usselman and Todd A. Brown. Atomic theory and multiple combining proportions: The search for whole number ratios: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):153– 169, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Uchoa:2022:SNS

[Uch22] Raphael Uchôa. From the state of nature to the state of ruins: ‘American race’ and ‘savage knowledge’ according to Carl von Martius. Annals of Science, 79(1):40–59, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ugaglia:2006:SMB

[Uga06] Monica Ugaglia. The science of magnetism before Gilbert Leonardo Garzoni’s Treatise on the Loadstone. Annals of Science, 63(1):59– 84, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Underwood:1963:ETA

[Und63] E. Ashworth Underwood, M.A., B.Sc., M.D., M.R.C.P. The early teaching of anatomy at Padua, with special reference to a model of the Padua anatomical theatre. Annals of Science, 19(1):1–26, 1963. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Underwood:1967:DEH

[Und67] E. Ashworth Underwood, M.A., B.Sc., M.D., F.R.C.P. Dumfries and the early history of surgical anaesthesia. Annals of Science, 23(1): 35–75, 1967. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Unguru:2002:BPS

[Ung02] Sabetai Unguru. Amicus Plato sed…: Fowler’s new mathematical reconstruction of the mathematics of Plato’s Academy. Annals of Science, 59(2):201–210, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Urban:1980:HGM

[Urb80] Jan Urban. On the history of geological mapping in Czechoslovakia. Annals of Science, 37(4):413–432, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Usselman:2005:RLS

[URRF05] Melvyn Usselman, Alan Rocke, Christina Reinhart, and Kelly Foulser. Restaging Liebig: a study in the replication of experiments. Annals of Science, 62(1):1–55, 2005. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Usselman:1978:WCC

[Uss78] Melvyn C. Usselman. The Wollaston/Chenevix controversy over the elemental nature of palladium: a curious episode in the history of chemistry. Annals of Science, 35(6):551–579, 1978. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Usselman:1980:WWJ

[Uss80] Melvyn C. Usselman. William Wollaston, John Johnson and Colombian alluvial platina: a study in restricted industrial enterprise. Annals of Science, 37(3):253–268, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vaccari:2000:MKE

[Vac00] Ezio Vaccari. Mining and knowledge of the Earth in eighteenthcentury Italy. Annals of Science, 57(2):163–180, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vallvey:1996:ESS

[Val96] Luis Fermín Capitán Vallvey. Export and smuggling of Spanish platina in the eighteenth century. Annals of Science, 53(5):467–487, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Valleriani:2009:TAB

[Val09] Matteo Valleriani. The transformation of Aristotle’s Mechanical Questions: a bridge between the Italian Renaissance architects and Galileo’s first new science. Annals of Science, 66(2):183–208, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

VandenBroecke:2001:DML

[Van01] Steven Vanden Broecke. Dee, Mercator, and Louvain instrument making: An undescribed astrological disc by Gerard Mercator (1551). Annals of Science, 58(3):219–240, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

VanNeste:2019:BRW

[Van19] Aaron Van Neste. Book review: Worlds of natural history, edited by H. A. Curry, N. Jardine, J. A. Secord, and E. C. Spary, Cambridge, UK, Cambridge University Press, 2018, xxv + 656 pp., 16 plts, $48.00; £36.99, ISBN 978-1-316-64971-8. Annals of Science, 76 (3–4):365–367, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vidal:2011:SSE

[VB11] Fernando Vidal and Saskia Brown. The sciences of the soul: the early modern origins of psychology. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-85586-4. xiv + 413 pp. LCCN BF101 .V5313 2011. Translated by Saskia Brown.

vanDelft:2008:ZPE

[vD08] Dirk van Delft. Zero-point energy: The case of the Leiden Low-Temperature Laboratory of Heike Kamerlingh Onnes. Annals of Science, 65(3):339–361, 2008. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

vanDongen:2010:EU

[vD10] Jeroen van Dongen. Einstein’s unification. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88346-6. x + 213 pp. LCCN QC173.6 .D66 2010. URL http://

assets.cambridge.org/97805218/83467/cover/9780521883467.
jpg;   http://www.loc.gov/catdir/enhancements/fy1009/2010017820-
b.html;   http://www.loc.gov/catdir/enhancements/fy1009/
2010017820-d.html;   http://www.loc.gov/catdir/enhancements/
fy1009/2010017820-t.html.

vanDalen:2011:BOO

[vD11] Benno van Dalen. Between Orient and Occident: Transformation of knowledge. Annals of Science, 68(4):445–451, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vuillemin-Diem:2008:MGL

[VDdM08] Gudrun Vuillemin-Diem and Guillaume de Moerbeke. Meteorologica, volume 10, 2, 1 of Aristoteles Latinus. Brepols, Bruxelles, Belgique, 2008. ISBN 2-503-53080-X, 2-503-53078-8, 2-503-53079-6. xviii + 436 pp. LCCN ????

vanderSpek:2006:STR

[vdS06] Trienke M. van der Spek. Selling a theory: The role of molecular models in J. H. van ’t Hoff’s stereochemistry theory. Annals of Science, 63(2):157–177, 2006. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Verney:1956:SAW

[Ver56] E. B. Verney, M.A., M.B., B.Ch., F.R.C.P., F.R.S. Some aspects of the work of Ernest Henry Starling. Annals of Science, 12(1):30– 47, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vernon:1994:MWM

[Ver94] Keith Vernon. Microbes at work. Micro-organisms, the D.S.I.R. and industry in Britain, 1900–1936. Annals of Science, 51(6):593–613, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vermij:2013:BRB

[Ver13] Rienk Vermij. Book review: The Disordered Police State. German Cameralism as Science and Practice, Chicago and London: University of Chicago Press, 2009. 226 pp. illus. $45.00. ISBN-13 978-0-226-87020-5. Annals of Science, 70(1):115–117, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Verburgt:2016:DFG

[Ver16] Lukas M. Verburgt. Duncan F. Gregory, William Walton and the development of British algebra: ‘algebraical geometry’, ‘geometrical algebra’, abstraction. Annals of Science, 73(1):40–67, 2016. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1025844.

Vetter:2004:SARa

[Vet04a] Jeremy Vetter. Science along the railroad: Expanding field work in the US Central West. Annals of Science, 61(2):187–211, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vetter:2004:SARb

[Vet04b] Jeremy Vetter. Science along the railroad: Expanding field work in the US Central West. Annals of Science, 61(3):271, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vetter:2010:KGE

[Vet10] Jeremy Vetter. Knowing Global Environments: New Historical Perspectives on the Field Sciences. Rutgers University Press, New Brunswick, NJ, USA, 2010. ISBN 0-8135-4875-6, 1-283-36997-4. 272 pp. LCCN Q175.

vanHelvoort:1992:CBJ

[vH92] Ton van Helvoort. The controversy between John H. Northrop and Max Delbrück on the formation of bacteriophage: Bacterial synthesis or autonomous multiplication? Annals of Science, 49(6):545–575, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

vanHelden:1994:TPA

[vH94] Anne C. van Helden. Theory and practice in air-pump construction: The cooperation between Willem Jacob ’s Gravesande and Jan van Musschenbroek. Annals of Science, 51(5):477–495, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

vanHelden:1999:LPC

[vHvG99] Anne C. van Helden and Rob H. van Gent. The lens production by Christiaan and Constantijn Huygens. Annals of Science, 56(1): 69–79, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vickers:2000:ERL

[Vic00] Brian Vickers. Essay review: Looking for the rhetoric of science. Annals of Science, 57(4):441–446, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vickers:2004:ERB

[Vic04] Brian Vickers. Essay review: ‘Neither Proper nor Useful’: Jesuit Orthodoxy and Galilean Science, Mordechai Feingold, editor, The New Science and Jesuit Science: Seventeenth-Century Perspectives Dordrecht, Boston, MA, and London: Kluwer, 2003. ix + 270 pp. EUR118, US$113, £76. ISBN 1-4020-0848-1. Annals of Science, 61 (2):213–218, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vickers:2012:ERL

[Vic12] Brian Vickers. Essay review: Library of a polymath: Urs B. Leu, Raffael Keller, and Sandra Weimann, Conrad Gessner’s Private Library, ISBN-13 978-90-04-16723-0. Annals of Science, 69(4):571–574, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vlasschaert:2010:LME

[Vla10] Anne-Marie Vlasschaert. Le Liber mahameleth: édition critique et commentaires, volume 60 of Boethius; Wissenschaftsgeschichte. Steiner, Stuttgart, Germany, 2010. ISBN 3-515-09238-2. 429 pp. LCCN QA32 .L534 2010.

vanMaanen:1986:RLQ

[vM86] Jan A. van Maanen. The refutation of Longomontanus’ quadrature by John Pell. Annals of Science, 43(4):315–352, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Voskuhl:2013:AEM

[Vos13] Adelheid Voskuhl. Androids in the Enlightenment: mechanics, artisans, and cultures of the self. University of Chicago Press, Chicago, IL, USA and London, UK, 2013. ISBN 0-226-03402-X (hardcover), 0-226-03416-X (paperback), 0-226-03433-X (e-book). xiii + 279 pp. LCCN TJ211 .V675 2013.

Vandendriessche:2016:SEP

[VPW16] Joris Vandendriessche, Evert Peeters, and Kaat Wils, editors. Scientists’ expertise as performance: between state and society, 1860– 1960, volume 6 of History and philosophy of technoscience. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2016. ISBN 1-84893-527-7, 1-315-65382-6 (PDF ebook). xv + 239 pp.

LCCN Q175.5 .S35 2016. URL http://www.tandfebooks.com/
isbn/9781315653822.

vanStrien:2015:VIL

[vS15] Marij van Strien. Vital instability: Life and free will in physics and physiology, 1860–1880. Annals of Science, 72(3):381–400, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Veracini:2017:PDN

[VT17] Cecilia Veracini and Dante Martins Teixeira. Perception and description of New World non-human primates in the travel literature of the fifteenth and sixteenth centuries: a critical review. Annals of Science, 74(1):25–63, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vieth:2022:CCJ

[VV22] Rienk Vermij and Paul Vieth. Confessionalization and comets. John Bainbridge on the comet of 1618. Annals of Science, 79(3):275–291, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2086300.

Vermij:1992:UAC

[VvM92] Rienk H. Vermij and Jan A. van Maanen. An unpublished autograph by Christiaan Huygens: His letter to David Gregory of 19 January 1694. Annals of Science, 49(6):507–523, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Vaupel:2010:DMZ

[VW10] Elisabeth Vaupel and Stefan L. Wolff. Das Deutsche Museum in der Zeit des Nationalsozialismus: eine Bestandsaufnahme. (German) [The German Museum in the time of National Socialism: an inventory], volume 27 of Deutsches Museum. Abhandlungen und Berichte. Wallstein, Göttingen, Germany, 2010. ISBN 3-8353-0596-4. 710 pp. LCCN T183.G32 M85183 2010.

Waddell:2015:BRB

[Wad15] Mark A. Waddell. Book review: Art and Science in the Early Modern Netherlands. Annals of Science, 72(1):131–133, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wakefield:2009:DPS

[Wak09] Andre Wakefield. The disordered police state: German cameralism as science and practice. University of Chicago Press, Chicago, IL, USA and London, UK, 2009. ISBN 0-226-87020-0 (hardcover). x + 226

pp. LCCN HB91 .W35 2009. URL http://www.loc.gov/catdir/
toc/ecip0826/2008036715.html.

Walker:1936:DEE

[Wal36] W. Cameron Walker, M.Sc., A.I.C. The detection and estimation of electric charges in the eighteenth century. Annals of Science, 1(1): 66–100, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walker:1937:AEB

[Wal37] W. Cameron Walker, M.Sc., A.I.C. Animal electricity before Galvani. Annals of Science, 2(1):84–113, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walker:1939:AKB

[Wal39] Oswald J. Walker, B.Sc., Ph.D. August Kekulé and the benzene problem. Annals of Science, 4(1):34–46, 1939. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wall:1976:SDC

[Wal76] Joseph Frazier Wall. Social Darwinism and constitutional law with special reference to Lochner v. New York. Annals of Science, 33(5):

465–476, 1976. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wallis:1984:GGH

[Wal84] Ruth Wallis. The glory of gravity — Halley’s Comet 1759. Annals of Science, 41(3):279–286, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wallis:1997:ECH

[Wal97] Ruth Wallis. Edward Cocker (1632?–1676) and his Arithmetick: De Morgan demolished. Annals of Science, 54(5):507–522, 1997. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wallis:2000:CCA

[Wal00] Ruth Wallis. Cross-currents in astronomy and navigation: Thomas Hornsby, FRS (1733–1810). Annals of Science, 57(3):219–240, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walton:2001:TLM

[Wal01] Steven A. Walton. Theophrastus on Lyngurium: Medieval and early modern lore from the classical lapidary tradition. Annals of Science, 58(4):357–379, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Waller:2004:BDM

[Wal04] John C. Waller. Becoming a Darwinian: the micro-politics of Sir Francis Galton’s scientific career 1859–65. Annals of Science, 61(2): 141–163, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walters:2009:MSI

[Wal09] Alice N. Walters. Making scientific instruments in the industrial revolution. Annals of Science, 66(4):563–565, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walton:2012:TCG

[Wal12] S. A. Walton. Technology and culture in Greek and Roman antiquity. Annals of Science, 69(2):295–297, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Walker:2014:BRB

[Wal14] Mark Walker. Book review: Das Deutsche Museum in der Zeit des Nationalsozialismus. Eine Bestandaufnahme, edited by Elisabeth Vaupel and Stefan Wolff. Annals of Science, 71(1):135–138, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wan:2018:NCT

[Wan18a] Zhaoyuan Wan. Newton in China: Translating the Principia into Chinese (c. 1855–2015). Annals of Science, 75(1):1–20, 2018. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wang:2018:JWN

[Wan18b] Xiaona Wang. John Wilkins (1614–1672): New essays. Annals of Science, 75(3):262–265, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Warner:1988:CCA

[War88] Deborah Jean Warner. Commodities for the classroom: Apparatus for science and education in Antebellum America. Annals of Science, 45(4):387–397, 1988. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Warner:2002:PGA

[War02] Deborah Jean Warner. Political geodesy: The army, the air force, and the world geodetic system of 1960. Annals of Science, 59(4):363– 389, 2002. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wardhaugh:2007:LEP

[War07a] Benjamin Wardhaugh. The logarithmic ear: Pietro Mengoli’s mathematics of music. Annals of Science, 64(3):327–348, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Warner:2007:HSI

[War07b] Deborah Jean Warner. How sweet it is: Sugar, science, and the state. Annals of Science, 64(2):147–170, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wardhaugh:2012:HHM

[War12] Benjamin Wardhaugh, editor. The history of the history of mathematics: case studies for the Seventeenth, Eighteenth, and Nineteenth Centuries. P. Lang, Oxford, UK, 2012. ISBN 3-0343-0708-X (paperback). 3-0353-0261-8 (e-book). vi + 187 pp. LCCN QA21 .H59

Wardhaugh:2014:BRB

[War14] Benjamin Wardhaugh. Book review: Defending Hypatia: Ramus, Savile, and the Renaissance Rediscovery of Mathematical History, by Robert Goulding. Annals of Science, 71(1):138–140, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Waterston:1982:JFM

[Wat82] Charles D. Waterston. John Farey’s mineral survey of South-East Sutherland and the age of the Brora Coalfield. Annals of Science, 39(2):173–185, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Watanabe:1992:FBP

[Wat92] Masao Watanabe. Francis Bacon: Philanthropy and the instauration of learning. Annals of Science, 49(2):163–173, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wear:2009:MNF

[Wea09] Andrew Wear. Modernizing nature. Forestry and imperial ecodevelopment 1800–1950. Annals of Science, 66(3):423–424, 2009. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weeks:2016:BRV

[Wee16] Sophie Weeks. Book review: Voyaging in Strange Seas: The Great Revolution in Science. Annals of Science, 73(1):114–116, 2016. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

URL http://www.tandfonline.com/doi/abs/10.1080/00033790.
2015.1093791.

Weil:1948:AAS

[Wei48] E. Weil, Ph.D. Andreas Albrecht, a seventeenth-century military surveyor. Annals of Science, 6(1):44–45, 1948. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weil:1950:ULD

[Wei50] E. Weil, Ph.D. An unpublished letter by Davy on the safety-lamp. Annals of Science, 6(3):306–307, 1950. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weisinger:1951:ETR

[Wei51] Herbert Weisinger, A.M., Ph.D. English treatment of the relationship between the rise of science and the Renaissance, 1740–1840. Annals of Science, 7(3):248–274, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weindling:1985:WEK

[Wei85a] Paul Weindling. Weimar eugenics: The Kaiser Wilhelm Institute for Anthropology, Human Heredity and Eugenics in social context. Annals of Science, 42(3):303–318, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weiner:1985:RMT

[Wei85b] Douglas R. Weiner. The roots of ‘Michurinism’: Transformist biology and acclimatization as currents in the Russian life sciences. Annals of Science, 42(3):243–260, 1985. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weininger:2015:BBP

[Wei15] Stephen J. Weininger. Benzene and beyond: Pursuing the core of aromaticity: Essay in honour of Alan J. Rocke. Annals of Science, 72(2):242–257, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weisser:2019:BRR

[Wei19] Olivia Weisser. Book review: Recipes and everyday knowledge: medicine, science, and the household in early modern England, by Elaine Leong, Chicago, IL, University of Chicago Press, 2018, 288 pp., 19 halftones, $90.00 (cloth); $32.50 (paper), ISBN 978-0-226-58366-2. Annals of Science, 76(3–4):369–371, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wells:1983:SLH

[Wel83] Ellen B. Wells. Scientists’ libraries: a handlist of printed sources. Annals of Science, 40(4):317–389, July 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Werrett:2010:FPA

[Wer10] Simon Werrett. Fireworks: pyrotechnic arts and sciences in European history. University of Chicago Press, Chicago, IL, USA and London, UK, 2010. ISBN 0-226-89377-4 (hardcover). viii + 359 pp. LCCN TP268 .W47 2010.

vonHumboldt:2011:AHA

[Wer11] Romy Werther. Alexander von Humboldt, August Böckh, Briefwechsel, volume 33 of Beiträge zur Alexander-von-Humboldt-Forschung.

Akademie Verlag, Berlin, Germany, 2011. ISBN 3-05-005053-5. ISSN 0232-1556. 316 pp. LCCN Q143.H9 A4 2011.

Wernimont:2014:BRB

[Wer14] Jacqueline Wernimont. Book review: Fictions of the Cosmos. Science and Literature in the Seventeenth Century. Annals of Science, 71(4): 583–585, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Westfall:1956:UBPa

[Wes56a] Richard S. Westfall, M.A., Ph.D. Unpublished Boyle papers relating to scientific method. I. Annals of Science, 12(1):63–73, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Westfall:1956:UBPb

[Wes56b] Richard S. Westfall, M.A., Ph.D. Unpublished Boyle papers relating to scientific method. II. Annals of Science, 12(2):103–117, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wess:1997:LDE

[Wes97] Jane Wess. The logic demonstrators of the 3rd Earl Stanhope (1753– 1816). Annals of Science, 54(4):375–395, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wesemael:2010:UFG

[Wes10] François Wesemael. ‘unaffected by fortune, good or bad’: Context and reception of Chandrasekhar’s mass–radius relationship for white dwarfs, 1935–1965. Annals of Science, 67(2):205–237, 2010. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wexler:2008:WWW

[Wex08] Alice Wexler. The Woman Who Walked into the Sea: Huntington’s and the Making of a Genetic Disease. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-10502-9 (hardcover), 0-300-15861-0 (paperback). xxiv + 253 pp. LCCN RC394.H85 W52 2008.

Watson:2011:WSB

[WH11] Rita Watson and Wayne Horowitz. Writing Science Before the Greeks: a Naturalistic Analysis of the Babylonian Astronomical Treatise MUL.APIN, volume 48 of Culture and history of the ancient Near East, 1566–2055. Brill, Leiden, The Netherlands, 2011. ISBN 90-04-20230-7 (hardcover). xxvi + 222 pp. LCCN DS56 .C858 v.48; QB19 .W38 2011.

Whipple:1951:JYS

[Whi51] Robert S. Whipple, M.I.E.E., F.Inst.P. John Yarwell or the story of a trade card. Annals of Science, 7(1):62–69, 1951. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Whitrow:1982:CPI

[Whi82] Magda Whitrow. Condorcet: a pioneer in information retrieval? Annals of Science, 39(6):585–592, 1982. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

White:1986:WHP

[Whi86a] John S. White. William Harvey and the primacy of the blood. Annals of Science, 43(3):239–255, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Whitehead:1986:JCS

[Whi86b] Maurice Whitehead. The Jesuit contribution to science and technical education in late-nineteenth-century Liverpool. Annals of Science, 43(4):353–368, 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Whitestone:2012:CHL

[Whi12] Sebastian Whitestone. Christian Huygens’ lost and forgotten pamphlet of his pendulum invention. Annals of Science, 69(1):91–104, 2012. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Whitmer:2013:WNP

[Whi13] Kelly Joan Whitmer. What’s in a name? Place, peoples and plants in the Danish–Halle Mission, c. 1710–1740. Annals of Science, 70(3): 337–356, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Whyte:1954:HNL

[Why54] L. L. Whyte, M.A. On the history of natural lengths. Annals of Science, 10(1):20–27, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wietzke:2019:SWG

[Wie19] Johannes Wietzke. Science writing in Greco-Roman antiquity. Annals of Science, 76(2):233–236, 2019. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wiescher:2023:GPT

[Wie23] Michael Wiescher. A German physicist’s travels in Great Britain: Julius Plücker’s visits from 1853 to 1866. Annals of Science, 80(2): 143–194, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-

505X (electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00033790.2022.2147216.

Wightman:1955:WCT

[Wig55] William P. D. Wightman, M.Sc., Ph.D. William Cullen and the teaching of chemistry. Annals of Science, 11(2):154–165, 1955. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wightman:1956:WCT

[Wig56] William P. D. Wightman, M.Sc., Ph.D. William Cullen and the teaching of chemistry. II. Annals of Science, 12(3):192–205, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wightman:1961:LC

[Wig61] William P. D. Wightman, M.Sc., Ph.D., F.R.S.E. The language of chemistry. Annals of Science, 17(4):259–267, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilson:1937:CSG

[Wil37] Forsyth J. Wilson, D.Sc., Ph.D., F.I.C. The Chemical Society of Glasgow: Minute book of 1800–1801. Annals of Science, 2(4):451– 459, 1937. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1954:SEE

[Wil54] E. Williams. Some experiments on the expansive force of freezing water. Annals of Science, 10(2):166–171, 1954. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1955:GCT

[Wil55a] J. S. Wilkie, B.Sc., Ph.D. Galton’s contribution to the theory of evolution with special reference to his use of models and metaphors. Annals of Science, 11(3):194–205, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1955:TMD

[Wil55b] E. Williams. Two metallurgical discoveries. Annals of Science, 11(1): 93–98, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williamson:1955:GTD

[Wil55c] Raymond Williamson, M.A., B.Sc., M.D. The germ theory of disease. Neglected precursors of Louis Pasteur, Richard Bradley, Benjamin Marten, Jean-Baptiste Goiffon. Annals of Science, 11(1):44–57, 1955. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1956:IEWa

[Wil56a] J. S. Wilkie, B.Sc., Ph.D. The idea of evolution in the writings of Buffon. I. Annals of Science, 12(1):48–62, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1956:IEWb

[Wil56b] J. S. Wilkie, B.Sc., Ph.D. The idea of evolution in the writings of Buffon. II. Annals of Science, 12(3):212–227, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1956:IEWc

[Wil56c] J. S. Wilkie, B.Sc., Ph.D. The idea of evolution in the writings of Buffon. III. Annals of Science, 12(4):255–266, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1956:HLC

[Wil56d] E. Williams. Hooke’s law and the concept of the elastic limit. Annals of Science, 12(1):74–83, 1956. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1957:SOL

[Wil57] E. Williams. Some observations of Leonardo, Galileo, Mariotte and others relative to size effect. Annals of Science, 13(1):23–29, 1957. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1959:SHN

[Wil59] E. Williams. Some historical notes on the ‘fusible metals’. Annals of Science, 15(1):57–62, 1959. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1960:NWFa

[Wil60a] J. S. Wilkie, B.Sc., Ph.D., F.L.S. Nägeli’s work on the fine structure of living matter. I. Annals of Science, 16(1):11–41, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1960:NWFb

[Wil60b] J. S. Wilkie, B.Sc., Ph.D., F.L.S. Nägeli’s work on the fine structure of living matter. II. The immediate reception of Nägeli’s work. Annals of Science, 16(3):171–207, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1960:NWFc

[Wil60c] J. S. Wilkie, B.Sc., Ph.D., F.L.S. Nageli’s work on the fine structure of living matter. IIIa studies tending to confirm, correct and develop Nägeli’s views I: Optical studies. Annals of Science, 16(4):209–239, 1960. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1961:NWF

[Wil61] J. S. Wilkie, B.Sc., Ph.D., F.L.S. Nägeli’s work on the fine structure of living matter. III. Studies tending to confirm, correct and develop Nägeli’s views II: X-ray studies. Annals of Science, 17(1):27–62, 1961. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilkie:1962:HLA

[Wil62] J. S. Wilkie, B.Sc., Ph.D., F.L.S. Harvey’s lectures on anatomy. Annals of Science, 18(4):255–270, 1962. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1965:TOL

[Wil65] E. Williams. Two observations of Leonardo da Vinci concerning ramifications and concavities. Annals of Science, 21(2):135–138, 1965. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Willmoth:1987:JFL

[Wil87] Frances Willmoth. John Flamsteed’s letter concerning the natural causes of earthquakes. Annals of Science, 44(1):23–70, January 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Willmoth:1991:GAA

[Wil91a] Frances Willmoth. ‘the genius of all arts’ and the use of instruments: Jonas–Moore (1617–1679) as a mathematician, surveyor, and astronomer. Annals of Science, 48(4):355–365, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilson:1991:PGT

[Wil91b] David B. Wilson. P. G. Tait and Edinburgh natural philosophy, 1860– 1901. Annals of Science, 48(3):267–287, 1991. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilson:1992:SMD

[Wil92] Philip K. Wilson. ‘out of sight, out of mind?’: The Daniel Turner– James Blondel dispute over the power of the maternal imagination. Annals of Science, 49(1):63–85, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1995:MACa

[Wil95a] J. Williams. Mathematics and the alloying of coinage 1202–1700: Part I. Annals of Science, 52(3):213–234, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:1995:MACb

[Wil95b] J. Williams. Mathematics and the alloying of coinage 1202–1700. Part II. Annals of Science, 52(3):235–263, 1995. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Willach:2001:DTO

[Wil01] Rolf Willach. The development of telescope optics in the middle of the Seventeenth Century. Annals of Science, 58(4):381–398, 2001. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wills:2007:IFE

[Wil07] Ian Wills. Instrumentalizing failure: Edison’s invention of the carbon microphone. Annals of Science, 64(3):383–409, 2007. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilson:2009:SNL

[Wil09] David B. Wilson. Seeking nature’s logic: natural philosophy in the Scottish enlightenment. Pennsylvania State University Press, University Park, PA, USA, 2009. ISBN 0-271-03525-0 (hardcover). xvi + 344 pp. LCCN QC9.G7 W55 2009.

Wilding:2011:BRI

[Wil11a] Nick Wilding. Book review: Instruments and measurements: Eileen Reeves, Galileo’s Glassworks. The Telescope and the Mirror, ISBN-13 978-0-674-02667-4; ISBN-10 0-674-02667-5. Annals of Science, 68(1):146–148, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Willis:2011:VSL

[Wil11b] Martin Willis. Vision, science and literature, 1870–1920: ocular horizons, volume 15 of Science and culture in the nineteenth century. Pickering and Chatto, London, UK, 2011. ISBN 1-84893-234-0 (hardcover), 1-84893-235-9 (e-book). xii + 295 pp. LCCN QP475 .W575 2011.

Wilson:2013:BRB

[Wil13] Philip K. Wilson. Book review: Scientists and Swindlers: Consulting on Coal and Oil in America, 1820–1890, by Paul Lucier. Annals of Science, 70(4):545–547, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Williams:2015:BRB

[Wil15] Kelsey Jackson Williams. Book review: Notebooks, English Virtuosi, and Early Modern Science. Annals of Science, 72(3):409–411, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wilson:2023:BRA

[Wil23] Catherine Wilson. Book review: Analytical essay on the faculties of the soul, by Charles Bonnet, translated with an introductory essay by Stephen Gaukroger, Oxford, Oxford University Press, 2022, pp. 234 + xxxiv, $85.00 (hardback), ISBN 978-0-19-284677-8. Annals of Science, 80(4):420–423, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00033790.2023.2200716.

Winter:1949:WGF

[Win49] H. J. J. Winter, M.Sc., Ph.D. The work of G. T. Fechner on the Galvanic circuit. Annals of Science, 6(2):197–205, 1949. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Winder:1966:BGA

[Win66] Marianne Winder, M.A., A.L.A. A bibliography of German astrological works printed between 1465 and 1600, with locations of those extant in London libraries. Annals of Science, 22(3):191–220, 1966. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Winnik:1972:RHR

[Win72] Herbert C. Winnik, B.S., M.A., Ph.D. A reconsideration of Henry A. Rowland — the man. Annals of Science, 29(1):19–34, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Winterbottom:2016:MTE

[Win16] Anna Winterbottom. Medicine, trade and empire: Garcia de Orta’s colloquies on the simples and drugs of India (1563) in context. Annals of Science, 73(3):340–??, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Winsor:2023:DDM

[Win23] Mary Pickard Winsor. Darwin’s dark matter: utter extinction. Annals of Science, 80(4):357–389, 2023. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00033790.2023.2194889.

Withers:1996:GSN

[Wit96] Charles W. J. Withers. Geography, science and national identity in early modern Britain: The case of Scotland and the work of Sir Robert Sibbald (1641–1722). Annals of Science, 53(1):29–73, 1996. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Withers:2010:GSB

[Wit10] Charles W. J. Withers. Geography and science in Britain, 1831– 1939: a study of the British Association for the Advancement of Science. Manchester University Press, Manchester, UK, 2010. ISBN 0-7190-7976-4 (hardcover). 278 pp. LCCN G.

Wittje:2013:BRD

[Wit13] Roland Wittje. Book review: David Pantalony, Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth Century Paris. Archimedes — New Studies in the History of Science and Technology, Volume 24. Dordrecht, Heidelberg, London, New York: Springer, 2009. xxxvi + 372 pp. EUR 149.75 / $189.00 / £126.00. ISBN-13 978-90-481-2815-0; e-ISBN-13 978-90-481-2816-7. Annals of Science, 70(1):131–133, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Watson:2016:CCR

[WN16] Mark F. Watson and Henry J. Noltie. Career, collections, reports and publications of Dr Francis Buchanan (later Hamilton), 1762–1829: natural history studies in Nepal, Burma (Myanmar), Bangladesh and India. Part 1. Annals of Science, 73(4):392–424, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weaire:1987:URT

[WO87] D. Weaire and S. O’Connor. Unfulfilled renown: Thomas Preston (1860–1900) and the anomalous Zeeman effect. Annals of Science, 44(6):617–644, November 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wolf:2013:BRD

[Wol13a] Jeff Wolf. Book review: David B. Wilson, Seeking Nature’s Logic: Natural Philosophy in the Scottish Enlightenment. University Park, PA: Penn State University Press, 2009. xvi + 344 pp. $55.00. ISBN-13 978-0-271-03525-3. Annals of Science, 70(1):117–119, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wolfe:2013:CSS

[Wol13b] Audra J. Wolfe. Competing with the Soviets: science, technology, and the state in Cold War America. Johns Hopkins introductory series in the history of science. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-0769-8 (hardcover), 1-4214-0771-X (paperback). viii + 166 pp. LCCN Q127.U6 W65 2013.

Woodbury:1936:THS

[Woo36] Robert S. Woodbury, S.B. The teaching of the history of science in an engineering school. Annals of Science, 1(2):226–232, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Woolf:1953:ECO

[Woo53] Harry Woolf, B.S., M.A. Eighteenth-century observations of the transits of Venus. Annals of Science, 9(2):176–190, 1953. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:1971:ISJ

[Woo71a] R. Derek Wood. The involvement of Sir John Herschel in the photographic patent case, Talbot v. Henderson, 1854. Annals of Science, 27(3):239–264, September 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:1971:JBRa

[Woo71b] R. Derek Wood. J. B. Reade, F.R.S., and the early history of photography. Part I. A re-assessment on the discovery of contemporary evidence. Annals of Science, 27(1):13–45, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:1971:JBRb

[Woo71c] R. Derek Wood. J. B. Reade, F.R.S., and the early history of photography. Part II. Gallic acid and Talbot’s calotype patent. Annals of Science, 27(1):47–83, March 1971. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:1987:BRS

[Woo87] P. B. Wood. Buffon’s reception in Scotland: the Aberdeen connection. Annals of Science, 44(2):169–190, March 1987. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:1997:SPJ

[Woo97] R. Derek Wood. A state pension for L. J. M. Daguerre for the secret of his Daguerreotype technique. Annals of Science, 54(5):489–506, 1997. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wood:2015:BRB

[Woo15] Paul Wood. Book review: An Enlightened Duke: The Life of Archibald Campbell (1682–1761), Earl of Ilay, 3rd Duke of Argyll. Annals of Science, 72(2):262–264, 2015. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wright:1992:SMS

[Wri92] Thomas Wright. Scale models, similitude and dimensions: Aspects of mid-nineteenth-century engineering science. Annals of Science, 49(3):233–254, 1992. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Wright:2016:BSP

[Wri16] Aaron Sidney Wright. A beautiful sea: P. A. M. Dirac’s epistemology and ontology of the vacuum. Annals of Science, 73(3):225–256, 2016. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Watanabe:1984:NMJ

[WT84] Masao Watanabe and Ichiro Tanaka. A Newton manuscript in Japan. Annals of Science, 41(2):159–164, 1984. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yalcinkaya:2022:KFD

[Yal22] M. Alper Yalçinkaya. Kindred fatalisms: debating science, Islam, and free will in the Darwinian era. Annals of Science, 79(3):364–385, 2022. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00033790.2022.2075938.

Yamalidou:1998:MIH

[Yam98] Maria Yamalidou. Molecular ideas in hydrodynamics. Annals of Science, 55(4):369–400, 1998. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yapp:1983:IBV

[Yap83] W. B. Yapp. The illustrations of birds in the Vatican manuscript of De arte venandi cum avibus [The art of hunting with birds] of Frederick II. Annals of Science, 40(6):597–634, November 1983. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yavetz:1993:OHS

[Yav93] Ido Yavetz. Oliver Heaviside and the significance of the British electrical debate. Annals of Science, 50(2):135–173, 1993. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yeang:2011:EEC

[Yea11] Chen-Pang Yeang. Engineering entanglement, conceptualizing quantum information. Annals of Science, 68(3):325–350, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yeo:1979:WWN

[Yeo79] Richard Yeo. William Whewell, natural theology and the philosophy of science in mid nineteenth century Britain. Annals of Science, 36 (5):493–516, 1979. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yeo:2014:NEV

[Yeo14] Richard R. Yeo. Notebooks, English virtuosi, and early modern science. University of Chicago Press, Chicago, IL, USA and London, UK, 2014. ISBN 0-226-10656-X (hardcover), 0-226-10673-X (e-book). xviii + 398 pp. LCCN Q127.G4 Y47 2014.

Yeo:2018:HCS

[Yeo18] Richard Yeo. Hippocrates’ complaint and the scientific ethos in early modern England. Annals of Science, 75(2):73–96, 2018. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yi:2013:BRR

[Yi13] Doogab Yi. Book review: Robert H. Carlson, Biology is Technology: The Promise, Peril, and New Business of Engineering Life. Cambridge, MA and London: Harvard University Press, 2010. vii + 279 pp. $39.95. ISBN-13 978-0-674-03544-7. Annals of Science, 70(1): 109–111, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yi:2014:BRB

[Yi14] Doogab Yi. Book review: Genentech: The Beginnings of Biotech. Annals of Science, 71(3):432–434, 2014. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yi:1989:IDC

[YO89] Yang Jing Yi and David Oldroyd. The introduction and development of continental drift theory and plate tectonics in China: a case study in the transference of scientific ideas from west to east. Annals of Science, 46(1):21–43, 1989. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yost:1999:PIJ

[Yos99] Robinson M. Yost. Pondering the imponderable: John Robison and magnetic theory in Britain (c.1775–1805). Annals of Science, 56(2): 143–174, 1999. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Young:1937:CBT

[You37] Frank G. Young, Ph.D. Claude Bernard and the theory of the glycogenic function of the liver. Annals of Science, 2(1):47–83, 1937. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Yoeli-Tlalim:2013:BRB

[YT13] Ronit Yoeli-Tlalim. Book review: The Dialogue of Civilizations in the Birth of Modern Science, by Arun Bala. Annals of Science, 70 (4):557–559, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zuidervaart:2016:AVL

[ZA16] Huib J. Zuidervaart and Douglas Anderson. Antony van Leeuwenhoek’s microscopes and other scientific instruments: new information from the Delft archives. Annals of Science, 73(3):257–??, 2016. CO-DEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zacharias:1972:EES

[Zac72] Peter Zacharias, Dipl.phys. Zur Entstehung des Einteilchen-Schalenmodells. (German) [The emergence of the single-particle shell model]. Annals of Science, 28(4):401–411, 1972. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zaidi:2017:BRS

[Zai17] Waqar Zaidi. Book review: Scientific Governance in Britain 1914– 79. Annals of Science, 74(2):171–172, 2017. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zweiniger-Bargielowska:2010:MBB

[ZB10] Ina Zweiniger-Bargielowska. Managing the body: beauty, health, and fitness in Britain 1880–1939. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-928052-5 (hardcover). xi + 394 pp. LCCN GV342.27 .Z94 2010.

Zevenhuizen:2000:KSS

[Zev00] Erik Zevenhuizen. Keeping and scrapping: The story of a Mendelian lecture plate of Hugo de Vries. Annals of Science, 57(4):329–352, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ziadat:1994:ERE

[Zia94] Adel A. Ziadat. Early reception of Einstein’s Relativity in the Arab periodical press. Annals of Science, 51(1):17–35, 1994. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Ziggelaar:1980:HDW

[Zig80] Augustine Ziggelaar. How did the wave theory of light take shape in the mind of Christiaan Huygens? Annals of Science, 37(2):179– 187, 1980. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Zuidervaart:2004:RPC

[Zui04] Huib J. Zuidervaart. Reflecting ‘popular culture’: The introduction, diffusion, and construction of the reflecting telescope in The Netherlands. Annals of Science, 61(4):407–452, 2004. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).